### v1

# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

Abstract.
We consider an evolving random discrete tree model called Preferential Attachment with Vertex Death, as introduced by Deijfen [13]. Initialised with an alive root labelled $1$, at each step $n\geq 1$ either a new vertex with label $n+1$ is introduced that attaches to an existing alive vertex selected preferentially according to a function $b$, or an alive vertex is selected preferentially according to a function $d$ and killed. In this article we introduce a generalised concept of persistence for evolving random graph models. Let $O_{n}$ be the smallest label among all alive vertices (the oldest alive vertex), and let $I_{n}^{m}$ be the label of the alive vertex with the $m^{\mathrm{th}}$ largest degree. We say a persistent $m$-hub exists if $I_{n}^{m}$ converges almost surely, we say that persistence occurs when $I_{n}^{1}/O_{n}$ is tight, and that lack of persistence occurs when $I_{n}^{1}/O_{n}$ tends to infinity.
We identify two regimes called the infinite lifetime and finite lifetime regimes. In the infinite lifetime regime, vertices are never killed with positive probability. Here, we provide conditions under which we prove the (non-)existence of persistent $m$-hubs for any $m\in\mathbb{N}$. This expands and generalises recent work of Iyer [25] which covers the case $d\equiv 0$ and $m=1$. In the finite lifetime regime, vertices are killed after a finite number of steps almost surely. Here we provide conditions under which we prove the occurrence of persistence, which complements recent work of Heydenreich and the author [23], where lack of persistence is studied for preferential attachment with vertex death.
1. Introduction
Random graph models based on preferential attachment are quite well-understood nowadays due to a large body of research taken place over the last three decades (see [36, 37] and the references therein for a good overview of the literature). These models have become a staple in random graph theory due to their use in modelling real-world networks. Still, a clear gap between most of these models and networks that form in the real world is the fact that preferential attachment graphs allow for growth only. Vertices and edges are sequentially attached to the graph, causing the graph to grow. However, this is not particularly realistic, as nodes and bonds in real-world networks can be removed in almost all contexts as well. As an example, people are born but also pass away; friendships are formed but can also be broken; users of social media can befriend or follow others, but also unfriend or unfollow them and even remove their accounts all together.
A limited body of work focusses on (preferential attachment type) models that allow for vertices and/or edges to also be removed from the network. Generally speaking, in such models at each step,
  1. (1)
    A new vertex is added to the graph with probability $p_{1}$, which connects to $m$ already existing vertices (preferentially or uniformly at random).
  2. (2)
    A new edge is added to the graph between existing vertices with probability $p_{2}$, where the existing vertices are selected preferentially and/or uniformly.
  3. (3)
    A vertex is selected preferentially or uniformly and is removed with probability $p_{3}$ (including the edges incident to it).
  4. (4)
    An edge is selected uniformly and is removed with probability $p_{4}$.
Examples include the models studied by Cooper et al. [12] (with $m\in\mathbb{N}$) with further work of Lindholm and Vallier [29] and Vallier [35], Chung and Lu [11] (with $m=1$), Deo and Cami [15] (with $m=1$, $p_{2}=p_{4}=0$, and vertex deletion is done anti-preferentially, i.e. favouring vertices with low degree), and Deijfen and Lindholm [14] (with $m=1$, $p_{3}=0$). Another model of Britton and Lindholm [8], with further work in [9, 28], assigns a random fitness value to each vertex, and edges are added between existing vertices with a probability proportional to the fitness values of the vertices. Work on duplication-divergence type models with edge deletion includes, among others, [34, 22, 4, 30].
The work on such models with vertex and/or edge removal is limited for a number of reasons. First, it generally only considers the degree distribution of the model, and if further properties are studied, this is under restrictive conditions and assumptions. Second, all models mentioned only study affine preferential attachment or uniform attachment for attaching edges. Third, in most models the choice to add or remove vertices and edges at each step is determined by fixed model parameters (the $p_{i}$) and does not depend on the evolution of the graph itself.
Some recent work partially addresses these concerns. Diaz, Lichev, and the author study a model with uniform attachment and uniform vertex removal ($m$ is random, $p_{2}=p_{4}=0$), where the local weak limit, existence of a giant component, and the size and location of the maximum degree are studied [17]. Work of Bellin and Kammerer et al. [5, 6, 27] studies a tree model of uniform attachment and uniform vertex removal ($m=1$, $p_{2}=p_{4}=0$), but where vertices are not removed but are ‘frozen’ and can then no longer make new connections, and where the choice to freeze vertices need not be random. This allows them to study the model in a ‘critical window’ where they obtain precise results for the height of the tree and its scaling limit.
Finally, Deijfen [13] studies a Preferential Attachment tree model with Vertex Death (PAVD). Here, death is equivalent to freezing as in the work discussed above but simply carries a different name. In this model, the probability of killing a vertex or adding a vertex, as well as to which alive vertex this new vertex connects itself, is dependent on the state of the tree. In particular, it depends on the in-degrees of the alive vertices in a general way. Deijfen studied the limiting degree distribution of the tree conditionally on survival, and shows with a number of examples how introducing death can yield novel behaviour compared to preferential attachment trees without death.
In this article we focus on the PAVD model introduced and studied by Deijfen. To this end, let us provide a definition of the model. For a tree $T$, we naturally think of its edges as being directed towards the root. We then let $\deg_{T}(v)$ denote the in-degree of a vertex $v$ in $T$. For a sequence of trees $(T_{n})_{n\in\mathbb{N}}$, we write $\deg_{n}(v)$ rather than $\deg_{T_{n}}(v)$ for ease of writing.
Definition 1.1 (Preferential Attachment with Vertex Death).
Let $b\colon\mathbb{N}_{0}\to(0,\infty)$ and $d\colon\mathbb{N}_{0}\to[0,\infty)$ be two sequences. We recursively construct a sequence of trees $(T_{n})_{n\in\mathbb{N}}$ and a sequence of sets of vertices $(\mathcal{A}_{n})_{n\in\mathbb{N}}$ as follows. We initialise $T_{1}$ as a single vertex labelled $1$ and $\mathcal{A}_{1}=\{1\}$. For $n\geq 1$, conditionally on $T_{n}$ and $\mathcal{A}_{n}$, if $\mathcal{A}_{n}\neq\varnothing$, we select a vertex $i$ from $\mathcal{A}_{n}$ with probability
$\frac{b(\deg_{n}(i))+d(\deg_{n}(i))}{\sum_{j\in\mathcal{A}_{n}}b(\deg_{n}(j))+d(\deg_{n}(j))}.$ (1.1)
Then, conditionally on $T_{n}$ and $i$, we either kill vertex $i$ with probability
$\frac{d(\deg_{n}(i))}{b(\deg_{n}(i))+d(\deg_{n}(i))},$ (1.2)
and set $T_{n+1}=T_{n}$ and $\mathcal{A}_{n+1}=\mathcal{A}_{n}\setminus\{i\}$, or otherwise construct $T_{n+1}$ from $T_{n}$ by introducing a new vertex $n+1$ which we connect by a directed edge to $i$ and set $\mathcal{A}_{n+1}=\mathcal{A}_{n}\cup\{n+1\}$.
If $\mathcal{A}_{n}=\varnothing$, we terminate the recursive construction, set $T_{i}=T_{n}$ for all $i>n$, and say that the tree has died. When $\mathcal{A}_{n}\neq\varnothing$ for all $n\in\mathbb{N}$ we say that the tree process survives.
We see from the definition that the probabilities to both select and kill a vertex depend on the evolution of the tree itself. Furthermore, in our analysis of the model, we do not restrict ourselves to the uniform case ($b$ and/or $d$ constant) or the affine case ($b(i)=b_{1}i+b_{2}$ and/or $d(i)=c_{1}i+c_{2}$). Finally, we recover the standard preferential attachment tree model when $d\equiv 0$.
Persistence of the maximum degree. The main focus of this article is persistence of the maximum degree, which in evolving random graphs is defined as the emergence of a fixed vertex that attains the largest degree in the graph for all but finitely many steps. When there is lack of persistence, such a fixed vertex does not exist and the maximum degree changes hands infinitely often (i.e. the maximum degree is attained by different vertices infinitely often). Persistence of the maximum degree is also known as degree centrality and is often leveraged in network archaeology and root-finding algorithms (see e.g. [3, 7, 10] and references therein).
The notion of persistence in evolving random graphs has been studied from different perspectives, starting with the work of Dereich and Mörters [16]. Later, conditions under which persistence and the lack thereof hold have been weakened or omitted by Galashin [20], Banerjee and Bhamidi [2], and Iyer [25]. In short (and omitting minor technical assumptions), using the PAVD formulation with $d\equiv 0$, so that there is no vertex death and we recover the classic preferential attachment model, persistence of the maximum degree occurs almost surely if and only if
$\sum_{i=0}^{\infty}\frac{1}{b(i)^{2}}<\infty.$ (1.3)
Heuristically, persistence occurs when high-degree vertices have a sufficiently strong advantage over low-degree vertices to establish new connections, so they are most likely to increase their in-degree further. This allows one vertex to attain the largest degree for all but finitely many steps, instead of the largest degree switching between vertices infinitely often. The transition between whether the advantage is strong enough or not lies exactly at the point when the series in (1.3) goes from finite to infinity.
When vertices are killed and can no longer make new attachments afterwards, as in the PAVD model, the notion of persistence as stated above is not sufficiently general. In this case, we distinguish between whether vertices can live forever with positive probability, or whether they are killed after a finite (random) number of steps almost surely. In the former case, it may be possible that a vertex attains the $m^{\mathrm{th}}$ largest degree for all but finitely many steps. We call such an individual a persistent $m$-hub. In the latter case, no individual lives forever, so that in particular no persistent $m$-hub can exist for any $m\in\mathbb{N}$. Here, a related question is whether the oldest alive individual and the the richest alive individual, that is, the alive individual with the smallest label and the alive individual with the largest degree, respectively, have labels that are ‘close’ or ‘far apart’. More precisely, whether the ratio of their labels is bounded or unbounded. When the ratio is bounded, we say that persistence occurs, whereas lack of persistence occurs when the ratio diverges to infinity.
In the case of preferential attachment without death all individuals are always alive, so that the oldest alive vertex is always the vertex with label $1$. The ratio of their labels thus equals the label of the maximum degree vertex, which implies that the existence of persistent $1$-hub is essentially equivalent to persistence. This is, however, not the case in models that incorporate vertex death.
Our contribution. In this article we study the two regimes outlined above. In the infinite lifetime regime, in which vertices can live forever with positive probability, we provide conditions, similar to (1.3), under which a persistent $m$-hub exists almost surely conditionally on the tree process $(T_{n})_{n\in\mathbb{N}}$ surviving, for each $m\in\mathbb{N}$, as well as conditions under which persistent hubs do not exist. This analysis generalises and expands recent work of Iyer [25], which focusses on the case $d\equiv 0$ and $m=1$ and provides a novel perspective used to analyse the (non-)existence of persistent hubs.
The second regime is the finite lifetime regime, in which vertices are killed after a finite number of steps almost surely. Here, we provide conditions, similar to (1.3), under which persistence occurs. This builds on recent work by the author and Heydenreich [23], in which lack of persistence is studied for the PAVD model in both the infinite and the finite lifetime regime, and extends the methodology developed by Deijfen in [13] using an embedding of the discrete tree process in a continuous-time branching process known as a Crump-Mode-Jagers branching process. We use a more precise formulation of this embedding compared to Deijfen, which allows us to simplify certain proofs of results in [13], but mainly to develop novel and precise results regarding the behaviour of the branching process embedding that were unattainable previously. Furthermore, we extend ideas used by Banerjee and Bhamidi in [2] to precisely analyse the optimal window in which the vertex with the largest degrees are found in the finite lifetime regime.
Structure of the paper. We state the main results and the necessary assumptions in Section 2. The methodology used in the analysis in presented in Section 3. Section 4 studies the existence of persistent $m$-hubs in the infinite lifetime regime, whereas Section 5 focusses on the occurrence of persistence in the finite lifetime regime. We conclude with a discussion of our work, state several problems, and discuss possibilities for future directions in Section 6.
Notation. Throughout the paper we use the following notation: we let :=$\mathbb{N}:=\{1,2,\ldots\}$ denote the natural numbers, set $\mathbb{N}_{0}:=\{0,1,\ldots\}$ to include zero and let $[t]:=\{i\in\mathbb{N}:i\leq t\}$ for any $t\geq 1$. For $x,y\in\mathbb{R}$, we let $\lceil x\rceil:=\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor:=\sup\{n\in\mathbb{Z}:n\leq x\}$, and let $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$. For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}$ such that $b_{n}$ is positive for all $n$ we say that $a_{n}=o(b_{n})$ and $a_{n}=\mathcal{O}(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$ and if there exists a constant $C>0$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$, respectively. We write $a_{n}=\Theta(b_{n})$ if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$. For random variables $X,(X_{n})_{n\in\mathbb{N}}$, and $Y$ we let $X_{n}\overset{\mathrm{d}}{\longrightarrow}X,X_{n}\overset{\mathbb{P}}{\longrightarrow}X$ and $X_{n}\overset{\mathrm{a.s.}}{\longrightarrow}X$ denote convergence in distribution, probability and almost sure convergence of $X_{n}$ to $X$, respectively. $X\preceq Y$ (resp. $X\succeq Y$) denotes that $X$ is stochastically dominated by $Y$ (resp. $X$ stochastically dominates $Y$). For a sequence $(\mathcal{E}_{t})_{t\in\mathbb{I}}$ of events, where $\mathbb{I}=\mathbb{N}$ or $\mathbb{I}=[0,\infty)$, we say that $\mathcal{E}_{t}$ holds with high probability when $\lim_{t\to\infty}\mathbb{P}\!\left(\mathcal{E}_{t}\right)=1$. Finally, for an event $\mathcal{S}$ such that $\mathbb{P}\!\left(\mathcal{S}\right)>0$ we let $\mathbb{P}_{\mathcal{S}}\left(\cdot\right):=\mathbb{P}(\cdot\,|\,\mathcal{S})$ and $\mathbb{E}_{\mathcal{S}}{}{\cdot}:=\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ denote the conditional probability measure and conditional expectation, respectively. Then, we let $\xrightarrow{\mathbb{P}_{\mathcal{S}}-\mathrm{a.s.}}$ respectively $\overset{\mathbb{P}_{\mathcal{S}}}{\longrightarrow}$ denote almost sure convergence and convergence in probability with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$.
2. Results
In this section we present the main results and the necessary assumptions that we use. Recall the PAVD model, as in Definition 1.1. We let
$\mathcal{S}:=\bigcap_{n\in\mathbb{N}}\{\mathcal{A}_{n}\neq\varnothing\}$ (2.1)
denote the event that the tree process survives, that is, the construction is never terminated. We assume throughout that $\mathbb{P}\!\left(\mathcal{S}\right)>0$. We let $\mathbb{P}_{\mathcal{S}}$ denote the probability measure $\mathbb{P}$ conditionally on $\mathcal{S}$. Upon $\mathcal{S}$, the quantities of interest are the oldest alive individual (i.e. with smallest label) and the alive individuals with the largest degrees in $T_{n}$. We thus define
$O_{n}:=\min\mathcal{A}_{n}$ (2.2)
as the ‘oldest’ alive individual (i.e. with the smallest label), and recursively define
$I_{n}^{(1)}:=\min\{i\in\mathcal{A}_{n}:\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\},$ (2.3)
and for any fixed integer $m\geq 2$,
$I_{n}^{(m)}:=\min\{i\in\mathcal{A}_{n}\setminus\{I_{n}^{(1)},\ldots,I_{n}^{(m-1)}\}:\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\setminus\{I_{n}^{(1)},\ldots,I_{n}^{(m-1)}\}\}.$ (2.4)
We can interpret $I_{n}^{(m)}$ as the oldest alive individual with the $m^{\mathrm{th}}$ largest degree in $T_{n}$ and we call $I_{n}^{(m)}$ the $m$-hub of $T_{n}$. We note that $I_{n}^{(m)}$ is well-defined only if $|\mathcal{A}_{n}|\geq m$. As we shall see, this condition is satisfied for any fixed $m\in\mathbb{N}$ and all $n$ sufficiently large almost surely conditionally on survival.
We are interested in the long-term behaviour of $I_{n}^{(m)}$ and $O_{n}$. First, we say that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent elder $\mathbb{P}_{\mathcal{S}}$-almost surely when there exists a $\mathbb{P}_{\mathcal{S}}$-almost surely finite random variable $O$ such that
[…]
We say that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, when there exists a random variable $I^{(m)}$ such that
[…]
We say that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
$\Big{(}\frac{I_{n}^{(1)}}{O_{n}}\Big{)}_{n\in\mathbb{N}}\text{ is a tight sequence of random variables with respect to $\mathbb{P}_{\mathcal{S}}$},$ (P)
and that lack of persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
$\frac{I_{n}^{(1)}}{O_{n}}\xrightarrow{\mathbb{P}_{\mathcal{S}}}\infty.$ (LP)
Clearly, when the $1$-hub persists as in (PH-$m$) with $m=1$, then the existence of a persistent elder as in (PE) and persistence as in (P) are implied, since $1\leq O_{n}\leq I_{n}^{(1)}$ and $O_{n}$ is non-decreasing. When $m$-hubs do not persist for any $m\in\mathbb{N}$, the vertex that attains the $m^{\mathrm{th}}$ largest degree changes infinitely often. Persistence (and lack thereof) quantifies the rate of change of the maximum-degree vertex. Persistence tells us that the rate of change of $I_{n}^{(1)}$ is of the same rate of $O_{n}$, so that the maximum-degree vertex has a label (or ‘age’) of the same order as $O_{n}$, whereas lack of persistence tells us that the rate of change of $I_{n}^{(1)}$ is faster than that of $O_{n}$.
In this paper, we study:
  • A class of PAVD models for which we prove when it contains a persistent $m$-hub for any $m\in\mathbb{N}$. (Theorem 2.1)
  • A class of PAVD models for which we prove that it does not contain an $m$-persistent hub for any $m\in\mathbb{N}$, but persistence does occur. (Theorem 2.6)
This complements recent work of the author and Heydenreich in [23], where a class of PAVD models is studied for which it is shown when lack of persistence occurs, i.e. when nor persistent $m$-hubs exist nor persistence occurs.
Before we state our main results, we define the following quantities. Let $(E_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent exponential random variables, where $E_{i}$ has rate $b(i)+d(i)$, and let $(B_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent Bernoulli random variables (also independent of the $E_{i}$), such that
$\mathbb{P}\!\left(B_{i}=1\right)=\frac{b(i)}{b(i)+d(i)},\qquad\text{for }i\in\mathbb{N}_{0}.$ (2.5)
Finally, define
$S_{k}:=\sum_{i=0}^{k-1}E_{i},\qquad\ D:=\inf\{i\in\mathbb{N}_{0}:B_{i}=0\},\qquad\text{and}\qquad\mathcal{R}:=\sum_{j=1}^{D}\delta_{S_{j}},$ (2.6)
where $\delta$ is a Dirac measure. Note that $\mathcal{R}$ is an empty point process when $D=0$. We let $\mu$ denote the density of the point process $\mathcal{R}$, i.e.
$\mu(t):=\lim_{\varepsilon\downarrow 0}\varepsilon^{-1}\mathbb{P}\!\left(\mathcal{R}((t,t+\varepsilon))\geq 1\right),\qquad t\geq 0.$ (2.7)
Also, let $\widehat{\mu}(\lambda)$ denote the Laplace transform of $\mu$, for $\lambda\geq 0$. That is,
$\widehat{\mu}(\lambda):=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\sum_{k=1}^{\infty}\prod_{i=0}^{k-1}\frac{b(i)}{b(i)+d(i)+\lambda}.$ (2.8)
The explicit expression for $\widehat{\mu}$ in terms of the sequences $b$ and $d$ follows from [13, Proposition $1.1$]. Then, we define for $k\in\mathbb{N}$ the sequences
$\varphi_{1}(k):=\sum_{i=0}^{k-1}\frac{1}{b(i)+d(i)},\qquad\varphi_{2}(k):=\sum_{i=0}^{k-1}\Big{(}\frac{1}{b(i)+d(i)}\Big{)}^{2},\qquad\rho_{1}(k):=\sum_{i=0}^{k-1}\frac{d(i)}{b(i)+d(i)}.$ (2.9)
We split the presentation of the results into two parts, based on the infinite lifetime regime and the finite lifetime regime. As their names suggest, vertices are able to survive forever with positive probability in the ‘infinite lifetime’ regime, whereas vertices have an almost surely finite lifetime (i.e. they are killed after a finite number of steps a.s.) in the ‘finite lifetime’ regime. Note that preferential attachment without death is a special case of the former regime, where all individuals survive forever deterministically.
In the ‘infinite lifetime’ regime we provide criteria under which persistent $m$-hubs do or do not exist, in the sense of (PH-$m$). In the ‘finite lifetime’ regime we provide criteria under which persistence occurs, in the sense of (P). These results complement the work of the author and Heydenreich in [23] regarding lack of persistence for the PAVD model, in the sense of (LP). The proof techniques for the two regimes are entirely distinct. In the ‘finite lifetime’ regime we use a similar methodology as used in [23], though we now focus on the occurrence of persistence rather than lack of persistence, which requires a finer analysis at times. In the ‘infinite lifetime’ regime we expand and generalise ideas developed by Iyer in [25]. In fact, as we shall see in Section 4, the results in this regime can be proved to hold for more general family of models.
Infinite lifetime regime. We first present our result for the model when vertices can survive forever with positive probability. To this end, we introduce the following assumptions. Recall $\widehat{\mu}$ from (2.8). We assume that
There exists λ>0 such that $\text{There exists $\lambda>0$ such that }\widehat{\mu}(\lambda)<\infty.$ (F-L)
Furthermore, we have the assumptions
$\lim_{k\to\infty}\varphi_{1}(k)$ $=\sum_{i=0}^{\infty}\frac{1}{b(i)+d(i)}=\infty,$ (N-E)
$\lim_{k\to\infty}\varphi_{2}(k)$ $=\sum_{i=0}^{\infty}\frac{1}{(b(i)+d(i))^{2}}<\infty,$ (C-V)
$\lim_{k\to\infty}\rho_{1}(k)$ $=\sum_{i=0}^{\infty}\frac{d(i)}{b(i)+d(i)}=\infty.$ (F-D)
Throughout the paper we assume that Assumption (F-L) “Finite-Laplace” holds. In fact, Assumption (F-L) combined with the continuity of the mapping $\lambda\mapsto\widehat{\mu}(\lambda)$ implies Assumption (N-E) “Non-Explosion”. It guarantees that the tree process is ‘not exploding’, in the sense that there does not exists a unique vertex that obtains an infinite degree, nor a unique path from the root of infinite length. Assumption (C-V) “Converging-Variance” implies that the variance of $S_{k}$ converges with $k$, and is a crucial assumption under which we prove that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub or that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$. Finally, Assumption (F-D) “Finite-Degree” implies that the random variable $D$, as in (2.6), is finite almost surely (see Lemma 5.1) and governs whether we are in the ‘infinite lifetime’ regime or ‘finite lifetime’ regime. We hence may or may not assume that (F-D) is satisfied.
We then have the following result, where we show under which conditions the tree process $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub or not.
Theorem 2.1 (Infinite lifetime regime).
Consider the PAVD model in Definition 1.1. Suppose that $b$ and $d$ are such that Assumption (N-E) is satisfied and let $O\sim\mathrm{Geo}(\mathbb{P}\!\left(D=\infty\right))$. Then,
$O_{n}\longrightarrow O\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (2.10)
In particular, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent elder in the sense of (PE) ((((i.e. $O<\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely)))) if and only if Assumption (F-D) is not satisfied.
Now, suppose Assumption (N-E) is satisfied and Assumptions (F-D) and (C-V) are not satisfied. Then, for any $m\in\mathbb{N}$, $(T_{n})_{n\in\mathbb{N}}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely. If, instead, Assumption (C-V) is satisfied and, additionally, Assumption (F-L) is satisfied, then for any $m\in\mathbb{N}$ there exists an almost surely finite random variable $I^{(m)}$ such that
$I_{n}^{(m)}\longrightarrow I^{(m)}\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (2.11)
That is, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub in the sense of (PH-$m$) $\mathbb{P}_{\mathcal{S}}$-almost surely. Moreover, for any $m\geq 2$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
$\deg_{n}(I_{n}^{(1)})>\deg_{n}(I_{n}^{(2)})>\cdots>\deg_{n}(I_{n}^{(m)})\quad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s}.$ (2.12)
We summarise the results of Theorem 2.1 in Table 1 and provide the proof in Section 4.
Remark 2.2.
[…]
Remark 2.3.
Assumption (F-L) is a sufficient condition for the results in (2.11) and (2.12). It can be readily verified that this assumption is satisfied for $b$ such that
$\limsup_{i\to\infty}\frac{b(i)}{i}<\infty,$ (2.13)
irrespective of $d$. We additionally need that $d$ is such that Assumption (F-D) is not satisfied to be in the infinite lifetime regime, so that either $d$ converges to zero sufficiently fast or is non-zero at a sufficiently sparse sub-sequence.
[…]
Result Assumptions
Persistent elder ((2.10)) (N-E) and not (F-D)
No persistent $m$-hub for any $m\in\mathbb{N}$ (N-E), not (F-D), and not (C-V)
Persistent $m$-hub and distinct degrees ((2.11) & (2.12)) (F-L) and (C-V)
Table 1. A summary of the results in Theorem 2.1 and the required assumptions.
Finite lifetime. We now present our result for the PAVD model when vertices die after a finite number of steps almost surely. To this end, we need some additional definitions and assumptions. First, we define the quantity
$R:=\inf_{i\in\mathbb{N}_{0}}(b(i)+d(i)),$ (2.14)
and assume that there exists $d^{*}\in[0,R)$, such that
$\lim_{i\to\infty}d(i)=d^{*},$ (2.15)
and such that Assumption (F-D) is satisfied. Here, Assumption (F-D) ensures that vertices die after a finite number of steps almost surely. Furthermore, it is crucial that the limit $d^{*}$ is strictly smaller than $R$. Indeed, the ‘finite lifetime’ regime can be divided into two further sub-regimes, called the ‘rich are old’ ‘rich die young’ and regime, based on whether $d^{*}<R$ or $d^{*}>R$ (when assuming that such a $d^{*}$ exists), respectively, see Table 2. The recent work of Heydenreich and the author [23] studies lack of persistence in these regime. They show that, under mild conditions, persistence never occurs in the ‘rich die young’ regime, see [23, Theorem $2.10$]. Our focus here thus lies on the ‘rich are old’ regime, where we show under which conditions persistence does occur.
Infinite lifetime Finite lifetime
Rich are old Rich die young
(N-E) and not (F-D) (N-E), (F-D), and (2.15), $d^{*}<R$ (N-E), (F-D), and (2.15), $d^{*}>R$
Table 2. The different regimes that can be observed in the PAVD model and the main assumptions that are required in each regime. The ‘infinite lifetime’ regime in studied in Theorem 2.1, and ‘finite lifetime’ regime, and in particular the ‘rich are old’ regime is studied in Theorem 2.6. The ‘rich die young’ regime is studied in [23].
We also define
$\alpha(k):=\rho_{1}(k)-d^{*}\varphi_{1}(k)=\sum_{i=0}^{k-1}\frac{d(i)-d^{*}}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}.$ (2.16)
Note that $\alpha\equiv\rho_{1}$ when $d^{*}=0$, that is, when $d$ converges to zero. We extend the domain of $\varphi_{1},\varphi_{2},\rho_{1}$, as in (2.9), and $\alpha$ to $\mathbb{R}_{+}$ by linear interpolation so that, for example, $\varphi_{1}(t)=\int_{0}^{t}1/(b(\lfloor x\rfloor)+d(\lfloor x\rfloor))\,\mathrm{d}x$ for $t\geq 0$. In particular, this implies that $\varphi_{1}$ and $\varphi_{2}$ are strictly increasing and thus invertible. We then define
$\mathcal{K}_{\alpha}(t):=\alpha\big{(}\varphi_{1}^{-1}(t)\big{)}\qquad\text{for }t\geq 0.$ (2.17)
Heuristically, $\mathcal{K}_{\alpha}(t)$ quantifies the difference between $\rho_{1}(\varphi_{1}^{-1}(t))$ and $d^{*}t$, and allows us to precisely quantify probabilities of the form $\mathbb{P}\!\left(D\geq\varphi_{1}^{-1}(t)\right)$ (see Lemma 5.1). We then recall $\widehat{\mu}$ from (2.8). Assumption (F-L) is not sufficient in the ‘finite lifetime’ regime. Rather, we introduce
$\underline{\lambda}:=\inf\{\lambda>0:\widehat{\mu}(\lambda)<\infty\},$ (2.18)
and assume that
$\widehat{\mu}(\lambda^{*})=1\text{ has a solution }\lambda^{*}\in(0,\infty)\qquad\text{and}\qquad\ \lambda^{*}>\underline{\lambda}.$ (Ma)
The parameter $\lambda^{*}$ is known as the Malthusian parameter, and clearly Assumption (Ma) is stronger compared to Assumption (F-L).
Finally, we have the following technical assumption for the function $\mathcal{K}_{\alpha}$, as in (2.17).
Assumption $\bm{\mathcal{K}_{\alpha}}$. For $\lambda^{*}>0,d^{*}\geq 0$, there exists a function $r\colon(0,\infty)\to\mathbb{R}_{+}$ such that
$\mathcal{K}_{\alpha}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\big{)}-\mathcal{K}_{\alpha}(r(t))=\mathcal{O}(1).$ (2.19)
Remark 2.4.
[…]
Remark 2.5.
[…]
We then have the following result, where we provide conditions under which persistence occurs.
Theorem 2.6 (Finite lifetime regime).
Consider the PAVD model, as in Definition 1.1. Suppose that $b$ and $d$ are such that Assumptions (N-E), (Ma), (F-D), and (C-V) are satisfied. Recall $R$ from (2.14) and suppose that $d$ satisfies (2.15) with $d^{*}\in[0,R)$, and let Assumption $\mathcal{K}_{\alpha}$ be satisfied. Then,
$\Big{(}\log O_{n}-\frac{d^{*}}{\lambda^{*}+d^{*}}\log n-\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{)}_{n\in\mathbb{N}},$ (2.20)
and
$\Big{(}\log I_{n}^{(1)}-\frac{d^{*}}{\lambda^{*}+d^{*}}\log n-\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{)}_{n\in\mathbb{N}},$ (2.21)
and
$\Big{(}\varphi_{1}(\max_{v\in\mathcal{A}_{n}}\deg_{n}(v))-\frac{1}{\lambda^{*}+d^{*}}\log n+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{)}_{n\in\mathbb{N}}.$ (2.22)
are tight sequences of random variables with respect to $\mathbb{P}_{\mathcal{S}}$. In particular, persistence occurs in the sense of (P).
Remark 2.7.
[…]
Remark 2.8.
[…]
The crucial assumption in both Theorem 2.1 and 2.6 is Assumption (C-V). Indeed, it follows from the first part of Theorem 2.1 that no persistent $m$-hub exists $\mathbb{P}_{\mathcal{S}}$-almost surely when Assumption (C-V) is not satisfied. Also, Theorem $2.10$ in [23] shows that lack of persistence in the sense of (LP) occurs when Assumption (C-V) is not satisfied. We note that this condition for persistence and the existence of persistent $m$-hub is similar to that for preferential attachment without vertex death, as in (1.3).
Discussion, Examples, and Related work. The results presented in this section identify the behaviour of the PAVD model in the ‘infinite lifetime’ and ‘finite lifetime’ with regards to persistence of the maximum degree (or $m^{\mathrm{th}}$-largest degree). In the former regime, we require fairly mild conditions only. In fact, we can prove some of the results in a more general set-up as well, see Section 4. In the ‘finite lifetime’ regime, stronger assumptions are required compared to the ‘infinite lifetime’ regime. This is due to the fact that we need very precise asymptotic expansions of the random variables $I_{n}$ and $O_{n}$, which in this regime no longer converge without rescaling. In particular, we require control over the growth rate of a continuous-time branching process that we analyse to prove the results in Theorem 2.6 (see Section 3 for more details). In Section 6 we discuss some open problems related to this regime.
Examples. In the ‘finite lifetime’ regime, we can further distinguish between the ‘rich are old’ and ‘rich die young’ regimes. Whether the PAVD model is in the ‘rich are old’ or ‘rich die young’ regime subtly depends on the sequences $d$ and $b$. By changing only a small number of values in either sequences, it is possible to go from one regime to the other. For example, consider the three models
[…]
In the first example we have $R=2$ and $d^{*}=\frac{3}{2}$, so that this case belongs to the ‘rich are old’ regime. In the second and third line we have changed one instance of $d$ and $b$, respectively, so that $R=\frac{5}{4}$ and $d^{*}=\frac{3}{2}$. These cases thus belong to the ‘rich die young’ regime.
Assumptions. Let us then provide some examples that satisfy the assumptions used in the two main theorems. First, we discuss Assumption $\mathcal{K}_{\alpha}$. Here, with $c\in(0,\lambda^{*}/(\lambda*+d^{*}))$, set
$\varepsilon_{t}:=\sup_{i\geq ct}|d(i)-d^{*}|.$ (2.24)
Assumption $\mathcal{K}_{\alpha}$ is then satisfied if there exists $k\in\mathbb{N}$ such that
$\varepsilon_{t}^{k}=\mathcal{O}\Big{(}\frac{1}{\mathcal{K}_{\alpha}(t)}\Big{)}.$ (2.25)
Namely, using Remark 2.5, set
$x_{1}(t):=\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t\big{)}\qquad\text{and}\qquad x_{i}(t):=\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x_{i-1}(t)\big{)}\qquad\text{for }i>1.$ (2.26)
Then, take $x(t):=x_{k}(t)$ and $r(t):=\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x(t)$. It can be shown, using (the proof of) Lemma 5.5 $(c)$ that this choice of $r$ satisfies Assumption $\mathcal{K}_{\alpha}$. We observe that $\mathcal{K}_{\alpha}(t)=o(t)$ by Lemma 5.5 $(c)$, so that $\varepsilon_{t}=t^{-\beta}$ for some $\beta>0$ satisfies (2.25) for any $k>\beta^{-1}$. That is, as long as $d$ converges to $d^{*}$ at a polynomial rate, Assumption $\mathcal{K}_{\alpha}$ is satisfied. We expect that much weaker conditions on $d$ are sufficient as well, when making more assumptions about the regularity of $b$.
It can be checked with relative ease that Assumption (Ma) is satisfied when
  • $\mathbb{E}\left[D\right]\in(1,\infty)$, or
  • $\mathbb{E}\left[D\right]=\infty$, $b$ tends to infinity, $C_{1}:=\limsup_{i\to\infty}\frac{b(i)}{i}<\infty$, and $\widehat{\mu}(C_{1})>1$.
Finally, Assumption (F-L) is implied by Assumption (Ma). It thus requires fewer conditions on the sequences $b$ and $d$. Indeed, it is satisfied when
  • $\mathbb{E}\left[D\right]\in(1,\infty)$, or
  • $\mathbb{E}\left[D\right]=\infty$, $\limsup_{i\to\infty}\frac{b(i)}{i}<\infty$.
Related work. The results in Theorem 2.1 expand and generalise recent work of Iyer in [25] in which he developed a novel methodology to study the (non-)existence of persistent hubs in evolving random trees, based on the summability of infinite sums of symmetric random variables. The proof of Theorem 2.1 is inspired by this methodology, but generalises it to include vertex death, expands the results to persistent $m$-hubs for $m>1$, and includes the result regarding the distinct degrees of $m$-hubs in (2.12). The results on persistent $m$-hubs with $m>1$, as well as the result in (2.12) are novel and also generalise earlier work of Banerjee and Bhamidi for preferential attachment trees [2, Theorem $4.9$ and Remark $4.11$].
Recent work of Heydenreich and the author in [23] also studies the PAVD, but focusses on lack of persistence, in the sense of (LP). This complements the results presented here and combined these two works provide a clear understanding of the different regimes that can be observed, see Table 2. Further extensions can still be made, however, in particular by not requiring that the death sequence $d$ converges to a constant $d^{*}\geq 0$. We discuss this in more detail in Section 6.
3. Methodology: Embedding in CMJ branching process
To prove the results presented in Section 2, we make use of a technique where one ‘embeds’ a random discrete process into a continuous-time branching process, a Crump-Mode-Jagers (CMJ) branching process to be exact. The random discrete structure is then equal in distribution to the CMJ branching process, when viewed at certain stopping times, whilst the CMJ branching process in continuous time provides more analytical advantages compared to the discrete process.
General CMJ branching processes. We first introduce some definitions and notation. We let $\mathcal{U}_{\infty}$ denote the Ulam-Harris tree. That is,
$\mathcal{U}_{\infty}:=\{\varnothing\}\cup\bigcup_{k=1}^{\infty}\mathbb{N}^{k}.$ (3.1)
For $v=v_{1}v_{2}\cdots v_{k}\in\mathcal{U}_{\infty}$, we view $v$ as the $v_{k}^{\mathrm{th}}$ child of $v_{1}\cdots v_{k-1}$ (where $v_{1}\cdots v_{k-1}$ denotes $\varnothing$ when $k=1$). In particular, we view $j\in\mathbb{N}$ as the $j^{\mathrm{th}}$ child of $\varnothing$. We assign to each individual $v\in\mathcal{U}_{\infty}$ a sequence of non-negative inter-birth time random variables $(X(vi))_{i\in\mathbb{N}}$ and an offspring random variable $D^{(v)}$, which for each $v\in\mathcal{U}_{\infty}$ are i.i.d. copies of some random variables $(X(i))_{i\in\mathbb{N}}$ and $D$, respectively. We assume throughout that $X(i)<\infty$ for all $i\in\mathbb{N}$, almost surely.
Given the inter-birth times of an individual $v$, we define its lifetime as
$L^{(v)}:=\sum_{i=1}^{D^{(v)}+1}X(vi).$ (3.2)
Furthermore, we can recursively construct all birth times $\mathcal{B}(v)$ by setting $\mathcal{B}(\varnothing):=0$ and
$\mathcal{B}(vj):=\begin{cases}\mathcal{B}(v)+\sum_{i=1}^{j}X(vi)&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{1,\ldots,D^{(v)}\},\\ \infty&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{D^{(v)}+1,\ldots\}.\end{cases}$ (3.3)
That is, given that $v$ is born in finite time, each child $v1,\ldots,vD^{(v)}$ is born finite time and children $vj$ for $j>D^{(v)}$ are never born. The individual $v$ then dies after $L^{(v)}$ time. If $v$ is never born (i.e. $\mathcal{B}(v)=\infty$) then it is clear that all descendants of $v$ are never born either.
We can then construct the branching process $(\mathrm{BP}(t))_{t\geq 0}$ by setting
$\mathrm{BP}(t):=\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t\}\qquad\text{for }t\geq 0.$ (3.4)
In words, $\mathrm{BP}(t)$ contains all individuals that are born by time $t$. It is clear that individuals $v=v_{1}\cdots v_{m}$ with $v_{1},\ldots,v_{m}\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ are born in the process $\mathrm{BP}$ if and only if $v_{i}\leq D^{(v_{1}\cdots v_{i-1})}$ for all $i\in[m]$. All other individuals are never born. We further define
$\mathcal{A}^{\mathrm{ct}}_{t}:=\{u\in\mathcal{U}_{\infty}:\sigma_{u}\leq t,\sigma_{u}+L^{(u)}>t\},\qquad t\geq 0,$ (3.5)
as the set of alive individuals at time $t$. We can view this as the continuous-time counterpart of $\mathcal{A}_{n}$, as defined in Definition 1.1. We abuse notation to write
$\mathcal{S}=\{\forall t\geq 0:\mathcal{A}_{t}^{\mathrm{ct}}\neq\varnothing\}$ (3.6)
as the event that the branching process $\mathrm{BP}$ survives.
We can view the point process $\mathcal{R}$ in (2.6), its density $\mu$ in (2.7), and the Laplace transform $\widehat{\mu}$ of $\mu$ in (2.8) from this more general set-up as well. That is, here we set
$\mathcal{R}:=\sum_{k=1}^{D}\delta_{\sum_{i=1}^{k}X(i)}.$ (3.7)
With $\mu(t)\mathrm{d}t=\mathbb{E}\left[\mathcal{R}(t)\mathrm{d}t\right]$ and using Fubini’s theorem, the Laplace transform $\widehat{\mu}$ then equals
$\widehat{\mu}(\lambda):=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\lambda\sum_{i=1}^{j}X(i)\bigg{)}\right].$ (3.8)
Persistence in BP. We can view the concepts of persistent $m$-hubs, and persistence and the lack thereof from a continuous-time perspective as well. To this end, we define
$\deg_{t}(u):=\min\bigg{\{}\sup\bigg{\{}k\in\mathbb{N}:\sum_{i=1}^{k}X(ui)\leq t-\mathcal{B}(u)\bigg{\}},D^{(u)}\bigg{\}},\qquad u\in\mathcal{U}_{\infty},t\geq 0,$ (3.9)
as the number of children $u$ has produced by time $t$. We use the convention that we set $\deg_{t}(u)=-\infty$ when $t<\mathcal{B}(u)$, i.e. when $u$ has not (yet) been born. We then define
$O_{t}^{\mathrm{ct}}:=\!\!\min_{v\in\mathcal{A}_{t}^{\mathrm{ct}}}\!\!\mathcal{B}(v)$ (3.10)
as the birth-time of the oldest alive individual at time $t$ as the continuous-time counterpart of $O_{n}$, as in (2.2). Similarly, we define
$I_{t}^{(1),\mathrm{ct}}:=\min\big{\{}\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\big{\}},$ (3.11)
and for $m\geq 2$ fixed,
$I_{t}^{(m),\mathrm{ct}}:=\min\big{\{}\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]}\big{\}}.$ (3.12)
Here $I_{t}^{(m),\mathrm{ct}}$ denotes the birth-time of the oldest alive individual that has the $m^{\mathrm{th}}$ largest number of children at time $t$ and is the continuous-time counterpart of $I_{n}^{(m)}$. For a branching process $(\mathrm{BP}(t))_{t\geq 0}$ we say that it contains a persistent $m$-hub when there exists a random variable $I^{(m),\mathrm{ct}}$ such that
[…]
When $(\mathrm{BP}(t))_{\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, we say that persistence respectively lack of persistence occurs in $(\mathrm{BP}(t))_{t\geq 0}$ when
$(I^{(1),\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t})_{t\geq 0}\text{ is a tight sequence of random variables with respect to $\mathbb{P}_{\mathcal{S}}$,}$ (3.14)
or
$I^{(1),\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}\xrightarrow{\mathbb{P}_{\mathcal{S}}}\infty.$ (3.15)
Again, persistence in $(\mathrm{BP}(t))_{t\geq 0}$ is implied when $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $1$-hub, whereas lack of persistence implies that $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$.
The continuous-time preferential attachment with vertex death branching process. We now provide a specific case for the general definition above that we coin the continuous-time preferential attachment with vertex death branching process (CTPAVD), which is directly related to the PAVD model. Recall the random variables $D$, $(E_{i})_{i\in\mathbb{N}_{0}}$, and $(S_{k})_{k\in\mathbb{N}}$ from (2.6). For each individual $v\in\mathcal{U}_{\infty}$, we let $D^{(v)}$ be an i.i.d. copy of $D$, an let each of its inter-birth times $X(vi)$ be an i.i.d. copy of $E_{i-1}$ for each $i\in\mathbb{N}$. It follows that, for each $k\in\mathbb{N}$,
$\sum_{i=1}^{k}X(vi)\overset{\mathrm{d}}{=}S_{k}\qquad\text{and, in particular,}\qquad L^{(v)}\overset{\mathrm{d}}{=}S_{D+1}.$ (3.16)
Continuous-time embedding of PAVD into CTPAVD.
Let us define the following quantities. First, we introduce the characteristics
$\chi_{\mathrm{b}}(t):=\mathbbm{1}_{\{t\geq 0\}},\qquad\text{and}\qquad\chi_{\mathrm{a}}(t):=\mathbbm{1}_{\{0\leq t<L\}},$ (3.17)
and define, for $\square\in\{\mathrm{a},\mathrm{b}\}$,
$Z_{\square}(t):=\sum_{u\in\mathcal{U}_{\infty}}\chi_{\square}(t-\mathcal{B}(u)).$ (3.18)
$Z_{\mathrm{a}}$ and $Z_{\mathrm{b}}$ ‘count’ the branching process $\mathrm{BP}$ according to the characteristics $\chi_{\mathrm{a}}$ and $\chi_{\mathrm{b}}$, respectively. Here, $Z_{\mathrm{a}}(t)$ denotes the number of alive individuals in $\mathrm{BP}(t)$ and $Z_{\mathrm{b}}(t)$ denotes the total number of individuals born in $\mathrm{BP}(t)$. We then set
$N(t):=Z^{\mathrm{b}}_{t}+(Z^{\mathrm{b}}_{t}-Z^{\mathrm{a}}_{t})\qquad\text{for }t\geq 0,$ (3.19)
as the number of births and deaths that have occurred in the branching process up to time $t$. We define the stopping times $(\tau_{n})_{n\in\mathbb{N}}$ as
$\tau_{n}:=\inf\{t\geq 0:N(t)=n\},\qquad n\in\mathbb{N}.$ (3.20)
That is, $\tau_{n}$ denotes the time at which the $n^{\mathrm{th}}$ birth or death event occurs. We then have the following correspondence between the discrete tree process $(T_{n},\mathcal{A}_{n})_{n\in\mathbb{N}}$ and the CMJ branching process $(\mathrm{BP}(\tau_{n}),\mathcal{A}_{\tau_{n}}^{\mathrm{ct}})_{n\in\mathbb{N}}$ viewed at the sequence of stopping times.
Proposition 3.1 (Embedding of PAVD in CTPAVD).
Let $(T_{n},\mathcal{A}_{n})_{n\in\mathbb{N}}$ be the sequence of trees and sets of alive vertices of a PAVD model and $(\mathrm{BP}(t),\mathcal{A}_{t}^{\mathrm{ct}})_{t\geq 0}$ be a CTPAVD model and the set of alive individuals, both constructed with the same birth and death sequences $b$ and $d$, respectively. Then,
$(T_{n},\mathcal{A}_{n})_{n\in\mathbb{N}}\overset{\mathrm{d}}{=}(\mathrm{BP}(\tau_{n}),\mathcal{A}^{\mathrm{ct}}_{\tau_{n}})_{n\in\mathbb{N}}.$ (3.21)
The proposition follows from the memoryless property of exponential random variables and properties of minima of exponential random variables. The technique of embedding evolving discrete structures was originally pioneered by Athreya and Karlin [1] and has found fruitful applications in a wide variety of discrete (combinatorial) models, such as Pólya urns, discrete randomly growing trees such as preferential attachment trees and uniform attachment trees, evolving simplicial complexes, and many more.
The description of the branching process $\mathrm{BP}$ in which we embed the PAVD tree is used by Deijfen [13] to determine an explicit description for the limiting offspring distribution of alive individuals. The description of the branching process used here is somewhat different (though equivalent), but provides an analytical advantage. In particular, defining the offspring and lifetime of an individual as an i.i.d. copy of $D$, see (2.6), and $L$, defined in (3.2), respectively, is a perspective which allows us to analyse these quantities to a greater extent.
Proposition 3.1 also implies that
$\{(O_{n},I_{n}):n\in\mathbb{N}\}\overset{\mathrm{d}}{=}\{(N(O^{\mathrm{ct}}_{\tau_{n}}),N(I^{\mathrm{ct}}_{\tau_{n}})):n\in\mathbb{N}\}.$ (3.22)
Under Assumption (Ma) it follows that $N(t)$, as defined in (3.19), satisfies
$N(t)\mathrm{e}^{-\lambda^{*}\!t}\overset{\mathrm{a.s.}}{\longrightarrow}W,$ (3.23)
where $W$ is a non-negative random variable, such that $\mathbb{P}\!\left(W>0\,|\,\mathcal{S}\right)=1$ and hence that $\tau_{n}-\frac{1}{\lambda^{*}}\log n$ converges almost surely. As a result, we can, approximately, relate the quantities in (3.22) via
$O_{n}\approx\exp\big{(}\lambda^{*}O^{\mathrm{ct}}_{\log(n)/\lambda^{*}}\big{)},\qquad\text{and},\qquad\ I_{n}\approx\exp\big{(}\lambda^{*}I^{\mathrm{ct}}_{\log(n)/\lambda^{*}}\big{)}.$ (3.24)
In particular, this relates the definition of a persistent hub and (lack of ) persistence in $(\mathrm{BP}(t))_{t\geq 0}$ to the same concepts for the PAVD discrete tree process $(T_{n})_{n\in\mathbb{N}}$.
4. Persistent hubs in general CMJ branching processes
This section is devoted to proving results regarding the (non-)existence of persistent hubs for a general family of CMJ branching processes. In particular, we verify that this family includes the CTPAVD branching process, so that via Proposition 3.1 we obtain Theorem 2.1.
4.1. General results
We first state several assumptions that are required to prove the general results. Recall the definition of the offspring $D^{(v)}$ and inter-birth times $(X(vi))_{i\in\mathbb{N}}$ of an individual $v\in\mathcal{U}_{\infty}$ after (3.1) We view these as i.i.d. copies of random variables $D$ and $(X(i))_{i\in\mathbb{N}}$. Also recall the branching process $(\mathrm{BP}(t))_{t\geq 0}$ in (3.4). We have the following assumption.
Assumption 4.1.
The random variables $D$ and $(X(i))_{i\in\mathbb{N}}$ satisfy the following:
  1. (1)
    $D$ and $(X(i))_{i\in\mathbb{N}}$ are mutually independent.
  2. (2)
    $X(i)<\infty$ for all $i\in\mathbb{N}$ almost surely.
  3. (3)
    $\sum_{k=1}^{\infty}\mathbb{P}\!\left(D\geq k\right)\prod_{i=1}^{k}\mathbb{P}\!\left(X(i)=0\right)<1$.
  4. (4)
    $\mathbb{P}\!\left(D=\infty\right)>0$.
Parts $(1)$ and $(2)$ of Assumption 4.1 are relatively straightforward. Part $(3)$ states that the expected number of children that are born instantaneously (i.e. born at the same time as their parent) is strictly smaller than one, so that in particular sub-trees of the branching process formed by individuals born instantaneously are finite almost surely. Part $(4)$ finally implies that individuals can produce an infinite offspring with positive probability, which is the crucial assumption that distinguishes between Theorems 2.1 and 2.6.
We then have the following result classifying when $(\mathrm{BP}(t))_{t\geq 0}$ does not contains a persistent $m$-hub for any $m\in\mathbb{N}$, and when $(\mathrm{BP}(t))_{t\geq 0}$ does contains a persistent $1$-hub. To this end, we recall the general form of $\widehat{\mu}$ in (3.8).
Theorem 4.2 (Existence of $1$-hubs).
Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that Assumption 4.1 is satisfied. We then have:
  1. (a)
    Suppose the series $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely. Then, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub.
  2. (b)
    Suppose there exist $\lambda>0$ and $K\in\mathbb{N}$ such that
    $\widehat{\mu}(\lambda)<1\qquad\text{and}\qquad\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]<\infty.$ (PHC)
    Then, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $1$-hub.
Remark 4.3.
We observe that the condition that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely in $(a)$, or the condition that $\widehat{\mu}(\lambda)<1$ in (PHC), implies that $\sum_{i=1}^{\infty}X(i)$ diverges almost surely. Indeed, if $\sum_{i=1}^{\infty}X(i)<\infty$ were to hold with positive probability, then
$\bigg{|}\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))\bigg{|}\leq\sum_{i=1}^{\infty}X(i)^{\prime}+\sum_{i=1}^{\infty}X(i)<\infty,$ (4.1)
with positive probability (where we use that $X(i)^{\prime}\overset{\mathrm{d}}{=}X(i)$ for all $i\in\mathbb{N}$), which contradicts the condition in $(a)$. For the first condition in (PHC),
$\widehat{\mu}(\lambda)=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\lambda\sum_{i=1}^{j}X(i)\bigg{)}\right]\geq\mathbb{E}\left[D\right]\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{\infty}X(i)\bigg{)}\right],$ (4.2)
[…]
In the case that the inter-birth times are exponential random variables, we can extend the results of Theorem 4.2 to the existence of $m$-hubs for any $m\in\mathbb{N}$.
Theorem 4.4 (Existence of $m$-hubs).
Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that $X(i)$ follows an exponential distribution for all $i\in\mathbb{N}$ and $\mathbb{P}\!\left(D=\infty\right)>0$. Suppose there exist $\lambda>0$ and $K\in\mathbb{N}$ such that (PHC) is satisfied. Then, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub.
The conditions in Theorem 4.2 $(b)$ and Theorem 4.4 ensure the almost sure emergence of a persistent $m$-hub for any $m\in\mathbb{N}$. The following result provides additional conditions under which the persistent $m$-hub is unique, and under which the offspring of the $m$-hubs are distinct.
Theorem 4.5.
Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that Assumption 4.1 and (PHC) are satisfied, and fix $m\in\mathbb{N}$. Suppose that $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub. Also suppose that either
$\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)^{\prime}-X(i)\neq 0\right)=\infty\quad\text{or}\quad\mathbb{P}\!\left(X(i)^{\prime}\neq X(i)\right)=1\quad\text{for some }i\in\mathbb{N}.$ (4.3)
Then, $\mathbb{P}_{\mathcal{S}}$-almost surely the process $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub. Additionally, suppose that for any $\varepsilon>0$ we have
$\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)>\varepsilon\right)<\infty.$ (4.4)
Then, for any $m\geq 2$ there exists $t^{\prime}>0$ such that $\mathbb{P}_{\mathcal{S}}$-almost surely for all $t>t^{\prime}$,
$\deg_{t}(I^{(1),\mathrm{ct}})>\deg_{t}(I^{(2),\mathrm{ct}})>\cdots>\deg_{t}(I^{(m),\mathrm{ct}}).$ (4.5)
Remark 4.6.
Theorem 4.5 provides general conditions under which there exists a unique persistent $m$-hub, for any $m$. This requires the existence of a persistent $m$-hub to begin with. For $m=1$ this is true under the same level of generality by Theorem 4.2 $(b)$, whilst we know it to be true for exponentially distributed inter-birth times for $m>1$ by Theorem 4.4. We believe that Theorem 4.4 can be extended to a more general class of inter-birth time distributions as well, for which Theorem 4.5 would then provide the uniqueness of the persistent $m$-hub.
Related work. The results presented in Theorems 4.2, 4.4, and 4.5 are strongly related to recent work of Iyer on persistent hubs in CMJ branching processes [25]. Iyer proves similar results as in Theorem 4.2 and 4.5 (the case $m=1$) under the additional assumption that $D=\infty$, i.e. when all individuals live forever and produce an infinite offspring. Here, based on proof techniques developed by Iyer, we expand this to the case $m>1$ and to the case that $\mathbb{P}\!\left(D=\infty\right)\in(0,1)$. Section 5 deals with Theorem 2.6, which considers the case $\mathbb{P}\!\left(D=\infty\right)=0$.
A further distinction between our results and those of Iyer are the conditions in (4.3), under which we prove that a unique persistent $m$-hub emerges. These are different to the conditions provided by Iyer, as we think that these conditions are too restrictive. Indeed, our proof requires a milder condition under which we can prove that a unique persistent $m$-hub emerges.
4.2. Application to the CTPAVD branching process: Proof of Theorem 2.1
To apply the results in Theorems 4.2, 4.4, and 4.5 to the case of the CTPAVD branching process, we first state the following lemma, which is a slight generalisation of Lemma $3.5$ in [25].
Lemma 4.7.
Recall $\widehat{\mu}$ from (3.8). Suppose that $(1)-(3)$ in Assumption 4.1 are satisfied and that there exists $\lambda>0$ such that
$\widehat{\mu}(\lambda)<\infty.$ (4.6)
Then, $|\mathrm{BP}(t)|<\infty$ for all $t\geq 0$ almost surely.
Proof.
As $\widehat{\mu}(\lambda)$ is decreasing in $\lambda$, we first observe that (4.6), combined with Assumption 4.1, implies that there exists $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$. Indeed, by using Fubini’s theorem to interchange summation and integration, we can write for any $\alpha\geq 0$,
$\widehat{\mu}(\alpha)=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]$ $=\sum_{j=1}^{\infty}\mathbb{E}\left[\mathbbm{1}_{\{D\geq j\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]$ (4.7)
$=\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right],$
where we use the independence of $D$ and the $X(i)$ in the last step. We can split the right-hand side into two terms, namely
$\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\prod_{i=1}^{j}\mathbb{P}\!\left(X(i)=0\right)+\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right].$ (4.8)
By $(3)$ in Assumption 4.1 there exists $\eta>0$ such that the first sum is at most $1-\eta$. The second sum is finite for any $\alpha\geq\lambda$, since the left-hand side of (4.7) is finite for any $\alpha\geq\lambda$ by assumption. We can thus take $J=J(\eta)\in\mathbb{N}$ sufficiently large, such that for any $\alpha\geq\lambda$,
$\sum_{j=J}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]<\frac{\eta}{2}.$ (4.9)
We can then choose $\alpha=\lambda^{\prime}\geq\lambda$ sufficiently large such that
$\sum_{j=1}^{J-1}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{j}X(i)\bigg{)}\right]<\frac{\eta}{2},$ (4.10)
by the dominated converge theorem (as each summand can be bounded from above by one). Combined, we thus arrive at $\widehat{\mu}(\lambda^{\prime})<1$.
We then prove the desired result. Let $Y$ be an exponential random variable with rate $\lambda^{\prime}$, independent of the offspring random variables $(D^{(v)})_{v\in\mathcal{U}_{\infty}}$ and inter-birth times $(X(vi))_{v\in\mathcal{U}_{\infty},i\in\mathbb{N}}$. Then, for any $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$, with $m\in\mathbb{N}$ and $v_{1},\ldots,v_{m}\in\mathbb{N}$,
$\mathbb{P}\!\left(\mathcal{B}(v)\leq Y\right)=\mathbb{P}\!\left(D^{(v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\text{, and }\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell}}X(v_{1}\cdots v_{\ell-1}i)\leq Y\right).$ (4.11)
We use the independence of the offspring random variables and the inter-birth times, together with the fact that $Y$ is exponentially distributed, to write this as
$\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$ (4.12)
By summing over all individuals $v\in\mathcal{U}_{\infty}$ (where $\mathbb{P}\!\left(\mathcal{B}(\varnothing)\leq Y\right)=1$ as $\mathcal{B}(\varnothing)=0$), we thus obtain
$\mathbb{E}\left[|\mathrm{BP}(Y)|\right]=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}\!\left(\mathcal{B}(v)\leq Y\right)=1+\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$ (4.13)
As the terms in the product depend on distinct summation indices, we can interchange summation and product to yield
$\sum_{m=0}^{\infty}\Bigg{(}\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{j}X(i)\bigg{)}\right]\Bigg{)}^{m}=\sum_{m=0}^{\infty}\widehat{\mu}(\lambda^{\prime})^{m},$ (4.14)
where we use (4.7) in the final step. As $\widehat{\mu}(\lambda^{\prime})<1$, the expected value is finite. Hence, $|\mathrm{BP}(Y)|<\infty$ almost surely. Since $Y\geq t$ with positive probability for any $t\geq 0$, it thus follows that $|\mathrm{BP}(t)|<\infty$ for any $t\geq 0$, almost surely. ∎
We then prove Theorem 2.1.
Proof of Theorem 2.1.
We first check that the CTPAVD branching process satisfies all the conditions in Theorems 4.2, 4.4, and 4.5. We then translate this back to the PAVD model.
For the CTPAVD branching process, we have
$X(i)\overset{\mathrm{d}}{=}X(i)^{\prime}\overset{\mathrm{d}}{=}\text{Exp}(b(i-1)+d(i-1))\qquad\text{for all }i\in\mathbb{N},$ (4.15)
and $D$ is as in (2.6). It directly follows that $(1)-(3)$ of Assumption 4.1 are satisfied, and $(4)$ of Assumption 4.1 is equivalent to Assumption (F-D) not being satisfied, due to Lemma 5.1 (see also [23, Lemma $5.3$]).
For $(a)$ in Theorem 4.2 it remains to check that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely. As the summands are independent and symmetric, we have
${\rm Var}\bigg{(}\sum_{i=1}^{n}(X(i)^{\prime}-X(i))\bigg{)}=2\sum_{i=1}^{n}{\rm Var}(X(i))=\sum_{i=1}^{n}\frac{2}{(b(i)+d(i))^{2}},$ (4.16)
which has a finite limit as $n$ tends to infinity if and only if Assumption (C-V) is not satisfied. Kolmogorov’s two and three series theorems (see e.g. [19, Theorems $2.5.3$ and $2.5.4$]) thus imply that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely if and only if Assumption (C-V) is not satisfied, so that $(a)$ follows. So, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub, for any $m\in\mathbb{N}$, when Assumption (C-V) is not satisfied.
For $(b)$, we directly conclude from the above that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges when Assumption (C-V) is satisfied. It remains to check the conditions in (PHC). For the second condition in (PHC), we first observe that $b+d$ tends to infinity when Assumption (C-V) is satisfied. For any $\lambda>0$ we can thus choose $K\in\mathbb{N}$ sufficiently large such that $b(i)+d(i)\geq\lambda\sqrt{2}$ for all $i\geq K$. Now,
$\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda((X(i)^{\prime}-X(i))}\right]=\prod_{i=K+1}^{\infty}\frac{(b(i)+d(i))^{2}}{(b(i)+d(i)+\lambda)(b(i)+d(i)-\lambda)}=\prod_{i=K+1}^{\infty}\frac{(b(i)+d(i))^{2}}{(b(i)+d(i))^{2}-\lambda^{2}}.$ (4.17)
Taking the logarithm of the right-hand side yields
$-\sum_{i=K+1}^{\infty}\log\Big{(}1-\frac{\lambda^{2}}{(b(i)+d(i))^{2}}\Big{)}.$ (4.18)
As $b(i)+d(i)\geq\lambda\sqrt{2}$ for all $i\geq K$ and since $\log(1-x)\geq-x-x^{2}$ for $x\in(0,1/2)$, we thus arrive at the upper bound
$\lambda^{2}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}+\lambda^{4}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{4}}\leq\frac{3}{2}\lambda^{2}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}$ (4.19)
which is finite for any $\lambda>0$ when Assumption (C-V) is satisfied. The first condition in (PHC) is implied by Assumption (F-L), i.e. that $\widehat{\mu}(\lambda)<\infty$ for some $\lambda>0$. We then, as in the first part of the proof of Lemma 4.7, conclude that there exists $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$. We note that the second condition in (PHC) is satisfied for all $\lambda>0$ and any $K\geq\lambda\sqrt{2}$, so that is satisfied for $\lambda^{\prime}$ and $K\geq\lambda^{\prime}\sqrt{2}$ in particular. So, with $\lambda^{\prime}$ and $K\geq\lambda^{\prime}\sqrt{2}$, we conclude that (PHC) is satisfied, so that part $(b)$ of Theorem 4.2 holds. Hence, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when Assumptions (C-V) and (F-L) are satisfied. Theorem 4.4 then also directly follows.
To obtain Theorem 4.5, we note that the exponential distribution has no atoms, so that there is a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, for any $m\in\mathbb{N}$. Then,
$\mathbb{P}\!\left(X(i)>\varepsilon\right)=\mathrm{e}^{-\varepsilon(b(i)+d(i))}\leq\frac{1}{(b(i)+d(i))^{2}},$ (4.20)
where the inequality follows for any $\varepsilon>0$ and all $i\geq I_{\varepsilon}$, for some fixed $I_{\varepsilon}\in\mathbb{N}$, from the fact that $b(i)+d(i)$ tends to infinity with $i$ when Assumption (C-V) holds. As a result, the second condition in (4.4) is satisfied (again applying Assumption (C-V)), so that, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely the offspring of the persistent $1$-hub, persistent $2$-hub, through the persistent $m$-hub are distinct for all $t>t^{\prime}$.
To summarise, we have the following table when the inter-birth times satisfy (4.15) and $D$ is as in (2.6).
Assumptions Result
(C-V) and (F-D) not satisfied Theorem 4.2 (a): No persistent $m$-hub
(C-V) and (F-L) satisfied, (F-D) not satisfied Theorems 4.2 (b) and 4.5: Unique persistent $m$-hub with distinct degrees
Finally, we translate this result back to the PAVD model, for which we use the embedding in Proposition 3.1. First, we claim that $\lim_{n\to\infty}\tau_{n}=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. This follows since $\mathbb{P}\!\left(\mathcal{S}\right)>0$ and since $\widehat{\mu}(\lambda)<\infty$, so that Lemma 4.7 yields that $\mathrm{BP}(t)$ is finite for all $t\geq 0$, almost surely. Hence, $\lim_{n\to\infty}\tau_{n}=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, for any $t^{\prime}>0$ there exists $n_{0}\in\mathbb{N}$ such that $\tau_{n}>\tau_{n_{0}}\geq t^{\prime}$ for all $n\geq n_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Under the conditions in Theorem 2.1 it thus follows that $(T_{n})_{n\in\mathbb{N}}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$ when Assumption (C-V) is not satisfied. When Assumption (C-V) is satisfied and Assumption (F-L) also holds, then for any $m\in\mathbb{N}$ there exists $t^{\prime}>0$ and random variables $I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}}$, such that
$(I^{(1),\mathrm{ct}}_{t},\ldots,I^{(m),\mathrm{ct}}_{t})=(I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}})$ (4.21)
for all $t>t^{\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, there exists $n_{0}\in\mathbb{N}$ and random variables $I^{(1)},\ldots,I^{(m)}$, such that
$(I_{n}^{(1)},\ldots,I_{n}^{(m)})=(I^{(1)},\ldots,I^{(m)})$ (4.22)
for all $n\geq n_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Equivalently, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, for any $m\in\mathbb{N}$, when Assumptions (C-V) and (F-L) are satisfied.
Since the exponential distribution has no atoms, it follows from the first part of Theorem 4.5 that when Assumption (C-V) is satisfied, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub. It then directly follows that $(T_{n})_{n\in\mathbb{N}}$ contains a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely as well, for any $m\in\mathbb{N}$. We remark that this also directly follows from the observation that at most one vertex increases its degree at every step in the discrete tree process $(T_{n})_{n\in\mathbb{N}}$, so that at most one vertex can attain the $m^{\mathrm{th}}$ largest degree for all but finitely many steps.
Finally, for any $m\in\mathbb{N}$ there exists $t^{\prime}>0$ and random variables $I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}}$ such that
$(I^{(1),\mathrm{ct}}_{t},\ldots,I^{(m),\mathrm{ct}}_{t})=(I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}})$ (4.23)
for all $t>t^{\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Then, there exists $t^{\prime\prime}>t^{\prime}$ such that
$\deg_{t}(I^{(1),\mathrm{ct}})>\cdots>\deg_{t}(I^{(m),\mathrm{ct}})$ (4.24)
for all $t>t^{\prime\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As $\tau_{n}$ tends to infinity with $n$ $\mathbb{P}_{\mathcal{S}}$-almost surely, and
$\{\deg_{n}(i):i\in\mathcal{A}_{n}\}\overset{\mathrm{d}}{=}\{\deg_{\tau_{n}}(v):v\in\mathcal{A}_{\tau_{n}}^{\mathrm{ct}}\}$ (4.25)
by Proposition 3.1, we thus obtain that there exists $n_{1}\in\mathbb{N}$ such that
$\deg_{n}(I^{(1)})>\cdots>\deg_{n}(I^{(m)})$ (4.26)
for all $n\geq n_{1}$ $\mathbb{P}_{\mathcal{S}}$-almost surely, which together with (4.22) yields (2.12) and concludes the proof. ∎
The remainder of this section is devoted to proving Theorems 4.2, 4.4, and 4.5.
4.3. Preliminary results
We first collect a few preliminary results that are used throughout the proof of Theorems 4.2, 4.4, and 4.5.
Lemma 4.8 (Lemma $3.4$ in [25]).
Let $(X(i))_{i\in\mathbb{N}}$, $(X(i)^{\prime})_{i\in\mathbb{N}}$, and $Y$ be mutually independent random variables, with $(X(i))\overset{\mathrm{d}}{=}(X(i)^{\prime})_{i\in\mathbb{N}}$. Suppose that there exist $\lambda>0$ and $k\in\mathbb{N}$ such that
$\prod_{i=k}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]<\infty.$ (4.27)
Then,
$\mathbb{P}\bigg{(}\exists j\in\mathbb{N}_{0}:Y+\sum_{i=1}^{k+j}X(i)\leq\sum_{i=k}^{k+j}X(i)^{\prime}\bigg{)}\leq\bigg{(}\prod_{i=k}^{\infty}\mathbb{E}\Big{[}\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\Big{]}\bigg{)}\mathbb{E}\bigg{[}\exp\bigg{(}-\lambda\bigg{(}Y+\sum_{i=1}^{k-1}X(i)\bigg{)}\bigg{)}\bigg{]}.$ (4.28)
Lemma 4.9 (Theorem $1.13$ in [24]).
Suppose that $(S_{i})_{i\in\mathbb{N}}$ is a sequence of mutually independent random variables such that $\sum_{i=1}^{\infty}S_{i}$ converges almost surely. The distribution of $\sum_{i=1}^{\infty}S_{i}$ contains an atom on $\mathbb{R}$ if and only if for some collection $(c_{i})_{i\in\mathbb{N}}\in\mathbb{R}^{\infty}$,
$\forall i\in\mathbb{N}:\mathbb{P}\!\left(S_{i}=c_{i}\right)>0\qquad\text{and}\qquad\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq c_{i}\right)<\infty.$ (4.29)
Remark 4.10.
In case the $S_{i}$ are symmetric random variables, i.e. $\mathbb{P}\!\left(S_{i}\geq x\right)=\mathbb{P}\!\left(S_{i}\leq-x\right)$ for all $x\geq 0$ and $i\in\mathbb{N}$, the statement can be simplified by restricting to collections $(c_{i})_{i\in\mathbb{N}}$ that have finitely many non-zero values only. Indeed, let $(i_{j})_{j\in\mathbb{N}}$ be a strictly increasing sub-sequence such that $c_{i_{j}}\neq 0$ for all $j\in\mathbb{N}$. The first condition in (4.29) yields $\mathbb{P}\!\left(S_{i_{j}}=-c_{i_{j}}\right)=\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)>0$. Then, the condition
$\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq c_{i}\right)<\infty$ (4.30)
in (4.29) implies that $\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)=1-o(1)$, where the $o(1)$ tends to zero with $j$. There thus exists $J\in\mathbb{N}$ such that $\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)=\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)>\frac{1}{2}$ for all $j\geq J$, which yields a contradiction. So, any collection $(c_{i})_{i\in\mathbb{N}}$ can only contain finitely many non-zero values.
Lemma 4.11 (Proposition $1.15$ in [24]).
Suppose that $(S_{i})_{j\in\mathbb{N}}$ is a sequence of mutually independent, symmetric random variables. If $\sum_{i=1}^{\infty}S_{i}$ diverges almost surely,
$\limsup_{n\to\infty}\sum_{i=1}^{n}S_{i}=\infty\qquad\text{and}\qquad\liminf_{n\to\infty}\sum_{i=1}^{n}S_{i}=-\infty,\qquad\text{almost surely.}$. (4.31)
4.4. Proving Theorems 4.2, 4.4, and 4.5
In this sub-section, we first prove part $(a)$ of Theorem 4.2, then part $(b)$ of Theorem 4.2 and Theorem 4.4, and we conclude by proving Theorem 4.5.
For $u,v\in\mathcal{U}_{\infty}$, define the event
$\text{Win}(u,v):=\{\exists T\geq 0:\forall t\geq T\ \deg_{t}(u)\geq\deg_{t}(v)\}.$ (4.32)
If an individual $u$ has not been born in $\mathrm{BP}(t)$, we use the convention that $\deg_{t}(u)=-\infty$.
We then have the following result, which is based on Lemma $3.6$ in [25].
Lemma 4.12.
Assume Assumption 4.1 is satisfied. Then,
  1. (a)
    If $\sum_{i=1}^{\infty}X(i)=\infty$ almost surely, then $\lim_{t\to\infty}\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
  2. (b)
    If $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely, then for all $u,v\in\mathcal{U}_{\infty}$,
    $\mathbb{P}_{\mathcal{S}}\left(\mathrm{Win}(u,v)\cup\mathrm{Win}(v,u)\right)=1.$ (4.33)
  3. (c)
    If $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely, there exists a function $\phi:\mathbb{N}\to\mathbb{N}$ such that $\phi(j)>j$ for all $j\in\mathbb{N}$ such that for any $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$,
    $\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,\phi(j)\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)\geq\frac{1}{2}.$ (4.34)
Proof.
$(a)$ Let $V$ be the first individual born in $(\mathrm{BP}(t))_{t\geq 0}$ such that $D^{(V)}=\infty$ (taking an arbitrary one in case multiple individuals $u_{1},\ldots,u_{M}$ with $D_{u_{1}}=\cdots=D_{u_{M}}=\infty$ are born simultaneously, for some $M\geq 2$). Conditionally on survival such an individual $V$ exists almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$. Furthermore, the birth-time of $V$ is finite almost surely. As a result, for all $t\geq\mathcal{B}(V)$,
$\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)\geq\deg_{t}(V),$ (4.35)
and the right-hand side tends to infinity with $t$ since $D^{(V)}=\infty$ and as $\sum_{i=1}^{\infty}X(Vi)=\infty$ almost surely (since $X(Vi)\overset{\mathrm{d}}{=}X(i)$ for all $i\in\mathbb{N}$), which yields the desired result.
$(b)$ The event $\text{Win}(u,v)\cup\text{Win}(v,u)$ is trivially satisfied if at most one of the two individuals $u$ and $v$ is born. Given that both vertices are born, again the event is trivially satisfied if at most one individual produces an infinite offspring. As a result, we define
$\mathcal{E}(u,v):=\{\mathcal{B}(u),\mathcal{B}(v)<\infty\}\cap\{D^{(u)},D_{v}=\infty\},$ (4.36)
and we can thus write
$\mathbb{P}\!\left(\mathrm{Win}(u,v)\cup\mathrm{Win}(v,u)\right)=\mathbb{P}\!\left((\mathrm{Win}(u,v)\cup\mathrm{Win}(v,u))\cap\mathcal{E}(u,v)\right)+\mathbb{P}\!\left(\mathcal{E}(u,v)^{c}\right).$ (4.37)
Our aim is to show that
$\mathbb{P}\!\left((\mathrm{Win}(u,v)\cup\mathrm{Win}(v,u))\cap\mathcal{E}(u,v)\right)=\mathbb{P}\!\left(\mathcal{E}(u,v)\right).$ (4.38)
Recall the assumption that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely. Then, for $u,v\in\mathcal{U}_{\infty}$, on the event $\mathcal{E}(u,v)$ we have
$\left|\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{\infty}(X(ui)-X(vi))\right|<\infty\quad\text{almost surely.}$. (4.39)
Assume, without loss of generality, that $|u|\leq|v|$. Let $N$ be chosen such that $\mathcal{B}(v)$ is independent of $(X(un))_{n\geq N}$. For example, if $u$ is a non-parent ancestor of $v$, we may choose the value $N$ such that $u(N-1)$ is the ancestor of $v$, whilst if $v=uj$ for $j\in\mathbb{N}$, we can choose $N>j$. Then, the sequences $(X(un))_{n\geq N}$ and $(X(vn))_{n\geq N}$ are independent. We now have two cases:
(I) Suppose that the sequence $(S_{i})_{i\geq N}$ defined by $S_{i}:=X(ui)-X(vi)$ does not satisfy (4.29). Then, by Lemma 4.9, the random variable
$\sum_{j=N}^{\infty}(X(uj)-X(vj))$ (4.40)
contains no atom on $\mathbb{R}$. Hence, as the summands are independent
$\mathbb{P}\bigg{(}\mathcal{E}(u,v)\cap\bigg{\{}\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{\infty}(X(uj)-X(vj))=0\bigg{\}}\bigg{)}=0.$ (4.41)
Therefore, on the event $\mathcal{E}(u,v)$ there exists $K_{0}\in\mathbb{N}$ such that either
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{k}(X(uj)-X(vj))>0\qquad\text{for all $k\geq K_{0}$,}$ (4.42)
or
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{k}(X(uj)-X(vj))<0\qquad\text{for all $k\geq K_{0}$,}$ (4.43)
is satisfied $\mathbb{P}_{\mathcal{S}}$-almost surely. In other words, we either have $\mathcal{B}(uk)>\mathcal{B}(vk)$ for all $k\geq K_{0}$, or we have $\mathcal{B}(uk)<\mathcal{B}(vk)$ for all $k\geq K_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely on the event $\mathcal{E}(u,v)$. This in turn implies that for all $k$ sufficiently large $u$ has offspring $k$ before $v$ has offspring $k$ or vice versa. Thus, we have that (4.38) is satisfied.
(II) Suppose that the sequence $(S_{i})_{i\geq N}$ defined by $S_{i}:=X(ui)-X(vi)$ does satisfy (4.29). Since $S_{i}$ has a symmetric distribution (as $X(ui)\overset{\mathrm{d}}{=}X(vi)$), any collection $(c_{i})_{i\in\mathbb{N}}$ in Lemma 4.9 can only contain finitely many non-zero entries (see Remark 4.10). As a result, this implies that $\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq 0\right)=\sum_{i=1}^{\infty}\mathbb{P}(X(ui)\neq X(vi))<\infty$. By the Borel-Cantelli lemma we obtain that $X(ui)=X(vi)$ for all but finitely many $i\in\mathbb{N}$ almost surely. Thus, for some $K_{0}\in\mathbb{N}$, it is either the case that (4.42) or (4.43) is satisfied (but not both), or for all $k\geq K_{0}$,
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{k}(X(ui)-X(vi))=0.$ (4.44)
This implies that $\mathcal{B}(uk)\geq\mathcal{B}(vk)$ for all $k\geq K_{0}$ or $\mathcal{B}(u_{k})\leq\mathcal{B}(v_{k})$ for all $k\geq K_{0}$. Again, it follows that $\text{Win}(u,v)$ or $\text{Win}(v,u)$ occurs on the event $\mathcal{E}(u,v)$, so that (4.38) holds, which concludes the proof of part $(b)$.
$(c)$ Since $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely, Lemma 4.11 implies that for any $j\in\mathbb{N}$ we have
$\limsup_{n\to\infty}\sum_{i=j+1}^{n}(X(ui)-X(uji))=\infty\quad\text{almost surely}.$ (4.45)
Since also $\sum_{i=1}^{j}X(uji)<\infty$ almost surely by Assumption 4.1, we thus obtain that
$\mathbb{P}\!\left(\exists k>j:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)=1.$ (4.46)
Then, by the monotone convergence theorem,
$\lim_{n\to\infty}\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)=1.$ (4.47)
As a result, we can define $\phi(j)$ as
$\phi(j):=\inf\bigg{\{}n>j:\mathbb{P}\bigg{(}\exists k\in\{j+1,\ldots,n\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\bigg{)}\geq\frac{1}{2}\bigg{\}},$ (4.48)
which concludes the proof. ∎
We are now ready to prove part $(a)$ of Theorem 4.2, i.e. that almost surely no persistent $m$-hub emerges $\mathbb{P}_{\mathcal{S}}$-almost surely when the series $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely.
Proof of Theorem 4.2 $(a)$.
Fix $m\in\mathbb{N}$ and $u\in\mathcal{U}_{\infty}$. We observe that, by $(a)$ of Lemma 4.12, conditionally on survival $u$ can be a persistent $m$-hub only when $\mathcal{B}(u)<\infty$ and $D^{(u)}=\infty$. That is, $u$ needs to be born in the branching process in finite time and produce an infinite number of children. Hence, we can write
[…]
$=\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\mathbb{P}\big{(}\{u\text{ is a persistent $m$-hub}\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)},$
where the last step follows from the fact that the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$ implies survival. We then use that, upon the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$, $u$ can be a persistent $m$-hub only when for some $T\geq\mathcal{B}(u)$ it has a larger offspring than all but at most $m-1$ many of its children for all time $t\geq T$. That is,
[…]
$=\bigcup_{K\in\mathbb{N}}\bigcup_{\begin{subarray}{c}C\subset\mathbb{N}\\ |C|<m\end{subarray}}\{\forall j\in\mathbb{N}\setminus C\ \forall k\geq K:\mathcal{B}(uk)\leq\mathcal{B}(ujk)\}.$
Here, the second step follows from the fact that any individual almost surely takes an infinite amount of time to produce an infinite offspring by assuming that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely (see Remark 4.3). We thus obtain
$\mathbb{P}_{\mathcal{S}}{}$ $(u\text{ is a persistent hub})$ (4.51)
[…]
We now show that the second probability on the right-hand side equals zero, by showing that
($\mathbb{P}\big{(}\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}=\mathbb{P}\big{(}\mathcal{B}(u)<\infty,D^{(u)}=\infty\big{)}.$ (4.52)
On the event $\{\mathcal{B}(u)<\infty\}$, we can write
$\mathcal{B}(uk)>\mathcal{B}(ujk)\ \ \Leftrightarrow\ \ \sum_{i=1}^{k}X(ui)>\sum_{i=1}^{j}X(ui)+\sum_{i=1}^{k}X(uji)\ \ \Leftrightarrow\ \ \sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji).$ (4.53)
We now recall the function $\phi$ in part $(c)$ of Lemma 4.12. We let $(j_{\ell})_{\ell\in\mathbb{N}}=j_{\ell}(C,K)$ be the subsequence $j_{1}=1+\max\{\max C,K\}$ and $j_{\ell+1}:=\phi(j_{\ell})$ for all $\ell\in\mathbb{N}$. As $\phi(j)>j$ for all $j\in\mathbb{N}$, it follows that the subsequence $(j_{\ell})_{\ell\in\mathbb{N}}$ is strictly increasing and thus $j_{\ell}>\max C$ and $j_{\ell}>K$ for all $\ell\in\mathbb{N}$. Then,
$\mathbb{P}{}$ ($\big{(}\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}$ (4.54)
=($=\mathbb{P}\bigg{(}\bigg{\{}\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\bigg{\}}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg{)}$
($\geq\mathbb{P}\bigg{(}\bigg{\{}\exists\ell\in\mathbb{N}\,\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}:\!\!\!\sum_{i=j_{\ell}+1}^{k}\!\!\!X(ui)>\sum_{i=1}^{k}X(uj_{\ell}i)\bigg{\}}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg{)}.$
For ease of writing, we define
$\mathcal{E}(u,\ell):=\bigg{\{}\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}:\sum_{i=j_{\ell}+1}^{k}X(ui)>\sum_{i=1}^{k}X(uj_{\ell}i)\bigg{\}}\qquad\text{for }\ell\in\mathbb{N}.$ (4.55)
We then arrive at the lower bound
($\mathbb{P}\big{(}\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}.$ (4.56)
By the choice of the sub-sequence $(j_{\ell})_{\ell\in\mathbb{N}}$, the events $(\mathcal{E}(u,\ell))_{\ell\in\mathbb{N}}$ are mutually independent, as they depend on distinct (and thus independent) inter-birth times. They are also independent of $\mathcal{B}(u)$ and $D^{(u)}$, so that we can factorize the probability into
($\mathbb{P}\big{(}\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\big{)}\mathbb{P}\big{(}\mathcal{B}(u)<\infty,D^{(u)}=\infty\big{)}.$ (4.57)
By $(c)$ of Lemma 4.12 we have that $\mathbb{P}\!\left(\mathcal{E}(u,\ell)\right)\geq\frac{1}{2}$ for all $\ell\in\mathbb{N}$. By their mutual independence and using the reverse Borel-Cantelli lemma, we thus obtain that infinitely many event $\mathcal{E}(u,\ell)$ occur almost surely. This yields (4.52). Using this in (4.51) implies that
[…]
so that a countable union bound yields
[…]
which concludes the proof. ∎
We then move to part $(b)$ of Theorem 4.2 and Theorem 4.4. To this end, we define for $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u)\leq\mathcal{B}(v)<\infty$, the event
$\{v\text{ catches up to }u\}=\{\exists j\in\mathbb{N}:\mathcal{B}(vj)\leq\mathcal{B}(uj)\}.$ (4.60)
That is, for an individual $v$ that is born at the same time as or after $u$, we say that $v$ catches up to $u$ if at some point its offspring is at least as large as that of $u$ (not including having zero children). Heuristically, one would expect that individuals that are born ‘late’ in the branching process are unlikely to catch up to the root (if the root were to produce an infinite offspring). The following result makes this heuristic precise.
Proposition 4.13.
Let the inter-birth times $(X(i))_{i\in\mathbb{N}}$ and the offspring random variable $D$ satisfy Assumption 4.1 and suppose that there exist $\lambda>0$ and $K\in\mathbb{N}$ such that (PHC) is satisfied. Then,
$\mathbb{E}_{\mathcal{S}}\left[|\{v\in\mathcal{U}_{\infty}:v_{1}\geq K,D^{(v)}=\infty,v\text{ catches up to }\varnothing\}|\mathbbm{1}_{\{D_{\varnothing}=\infty\}}\right]<\infty.$ (4.61)
Proof.
We write the size of the set in the expected value as a sum of indicators. That is, we fix $v=v_{1}v_{2}\cdots v_{m}\in\mathcal{U}_{\infty}$ for some $m\in\mathbb{N}$, $v_{1}\geq K$, and $v_{2},\ldots v_{m}\in\mathbb{N}$, and consider
$\mathbb{P}_{\mathcal{S}}{}$ $(v\text{ catches up to }\varnothing\text{ and }D^{(v)}=D^{(\varnothing)}=\infty)$ (4.62)
=($=\mathbb{P}\!\left(\{D^{(v)}=D^{(\varnothing)}=\infty\}\cap\{\mathcal{B}(v)<\infty\}\cap\{\exists j\in\mathbb{N}_{0}:\mathcal{B}(v(v_{1}+j))\leq\mathcal{B}(v_{1}+j)\}\right).$
We note that we can omit the conditioning on survival, since the event that $\varnothing$ produces an infinite offspring guarantees survival. We also need only consider children numbered $v_{1}+j$ with $j\geq 0$ in the second line, since $\varnothing$ gives birth to $v_{1}$ many children before $v$ is born. The event $\mathcal{B}(v)<\infty$ is equivalent to $D^{(v_{1}\cdots v_{\ell})}\geq v_{\ell+1}$ for all $\ell\in[m-1]$ (the event $\{D^{(\varnothing)}=\infty\}$ already implies that $v_{1}$ is born). When $D^{(v)}=\infty$, the event that $v$ gives birth to its $(v_{1}+j)^{\mathrm{th}}$ child before $\varnothing$ does is equivalent to
$\sum_{\ell=1}^{m-1}\sum_{i=1}^{v_{\ell+1}}X(v_{1}\cdots v_{\ell}i)+\sum_{i=1}^{v_{1}+j}X(vi)\leq\sum_{i=v_{1}+1}^{v_{1}+j}X(i).$ (4.63)
This yields, by the independence of the offspring and inter-birth time random variables,
$\mathbb{P}_{\mathcal{S}}{}$ $(v\text{ catches up to }\varnothing\text{ and }D^{(v)}=D^{(\varnothing)}=\infty)$ (4.64)
[…]
We bound $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$ and apply Lemma 4.8 to arrive at the upper bound
$\sum_{\begin{subarray}{c}v\in\mathcal{U}_{\infty}\\ v_{1}\geq K\end{subarray}}{}$ $\mathbb{P}_{\mathcal{S}}\left(v\text{ catches up to }\varnothing\text{ and }D_{\varnothing}=\infty\right)$ (4.65)
[…]
We extend the range $i\geq v_{1}+1$ in the first product to $i\geq K+1$ (as $v_{1}\geq K$) and then extend the range $v_{1}\geq K$ in the second sum to $v_{1}\geq 1$ to obtain an upper bound. As the different terms in the second product depend on distinct summation indices $v_{1},\ldots,v_{m}$, we can switch the summation and product to obtain
$\prod_{i=K+1}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\sum_{m=1}^{\infty}\bigg{(}\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathrm{e}^{-\lambda\sum_{i=1}^{j}X(i)}\right]\bigg{)}^{m}$ (4.66)
[…]
where we use (4.7) in the penultimate step. Under the assumptions in the proposition, it is clear that this upper bound is finite. ∎
We define
$\mathrm{Pre}(u):=\{v\in\mathcal{U}_{\infty}:\mathcal{B}(v)\leq\mathcal{B}(u)\}\qquad\text{for }u\in\mathcal{U}_{\infty}$ (4.67)
as the set of predecessors of $u$ in the branching process $\mathrm{BP}$. Fix $m\in\mathbb{N}$. We then define
$\mathcal{P}_{m}:={}$ $\bigcup_{\begin{subarray}{c}C\subseteq\mathrm{Pre}(u)\\ |C|<m\end{subarray}}\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{, }u\text{ catches up to all }v\in\mathrm{Pre}(u)\setminus C\}$ (4.68)
[…]
to be the set of all individuals $u$ whose offspring diverges to infinity and at some point in time obtain a larger offspring than all but at most $m-1$ many of its predecessors. Note that this point in time need not be the same for all these predecessors. That is, it can catch up to a predecessor $v_{1}$ at time $t_{1}$ and catch up to a predecessor $v_{2}$ at time $t_{2}\neq t_{1}$. A (potential) persistent $m$-hub $u$ needs to attain the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time. As a result, there can be at most $m-1$ many predecessors of $u$ that it does not catch up to. The main ingredient for the proof of Theorems 4.2 $(b)$ and 4.4 is to ‘extend’ the result in Proposition 4.13 to the set $\mathcal{P}_{m}$, as follows.
Proposition 4.14 ($\mathcal{P}_{m}$ is finite almost surely).
Fix $m\in\mathbb{N}$ and recall $\mathcal{P}_{m}$ from (4.68). Let the inter-birth times $(X(i))_{i\in\mathbb{N}}$ and the offspring random variable $D$ be such that Assumption 4.1 and the conditions in (PHC) are satisfied. Then, $|\mathcal{P}_{1}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. If, additionally, $X(i)$ is exponentially distributed for every $i\in\mathbb{N}$, then $|\mathcal{P}_{m}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$.
As only finitely many individuals can catch up to all but at most $m-1$ of their predecessors, and out of all pairs of these vertices one must ‘win’, as in the sense of (4.32), Theorems 4.2 $(b)$ and 4.4 then readily follow:
Proof of Theorem 4.2 $(b)$ and Theorem 4.4 subject to Proposition 4.14.
Let us first observe that combining the second condition in (PHC) with $(b)$ in Assumption 4.1 implies that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely. As a result, by Lemma 4.12 $(b)$,
$\mathbb{P}_{\mathcal{S}}\left(\bigcap_{\begin{subarray}{c}S\subset\mathcal{U}_{\infty}\\ |S|<\infty\end{subarray}}\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)\right)=1,$ (4.69)
as the intersections are over countable sets. For any finite set $S$,
$\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)=\{\exists w\in S\ \exists T\geq 0:\deg_{t}(w)=\max_{u\in S}\deg_{t}(u)\text{ for all }t\geq T\}.$ (4.70)
As a result, with $S=\mathcal{P}_{m}$ being $\mathbb{P}_{\mathcal{S}}$-almost surely finite by Proposition 4.14, it follows that
$\mathbb{P}_{\mathcal{S}}\left(\exists w\in\mathcal{P}_{m}\ \exists T\geq 0:\deg(w,\mathrm{BP}(t))=\max_{u\in\mathcal{P}_{m}}\deg_{t}(u)\text{ for all }t\geq T\right)=1.$ (4.71)
Since a persistent $m$-hub attains the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time, it must be in $\mathcal{P}_{m}$. Hence, this proves that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely. ∎
It remains to prove Proposition 4.14. We define the following (random) subset of $\mathcal{U}_{\infty}$. For $K\in\mathbb{N}$ fixed and $u\in\mathcal{U}_{\infty}$, set
$I(u):=\begin{cases}K&\mbox{if }D^{(u)}=\infty,\\ D^{(u)}&\mbox{if }D^{(u)}<\infty\end{cases}$ (4.72)
Then, we define
$\mathcal{U}_{K}^{<\infty}=\mathcal{U}_{K}^{<\infty}((D^{(u)})_{u\in\mathcal{U}_{\infty}}):=\{u\in\mathcal{U}_{\infty}:u_{j}\leq I(u_{1}\cdots u_{j-1})\text{for all }j\in[|u|]\}\cup\{\varnothing\}.$ (4.73)
That is, given $(D^{(u)})_{u\in\mathcal{U}_{\infty}}$, one constructs the set $\mathcal{U}_{K}^{<\infty}$ iteratively as follows. Initialise $\mathcal{U}^{(0)}:=\{\varnothing\}$ and, given $\mathcal{U}^{(i)}$ with $i\in\mathbb{N}_{0}$, we set
$\mathcal{U}^{(i+1)}:=\mathcal{U}^{(i)}\cup\bigcup_{u\in\mathcal{U}^{(i)}}\{u1,\ldots,uI(u)\}.$ (4.74)
Note that, if $u\in\mathcal{U}^{(i)}$ for some $i$ and $D^{(u)}=0$, then we do not add any children $uj$ to $\mathcal{U}^{(i+1)}$. This iterative process either never terminates, in which case $\mathcal{U}^{<\infty}_{K}=\cup_{i=0}^{\infty}\mathcal{U}^{(i)}$ is infinite, or it terminates at some (random) step $M\in\mathbb{N}_{0}$, so that $\mathcal{U}^{<\infty}_{K}=\mathcal{U}^{(M)}$ is finite. We observe that, conditionally on survival of the process $(\mathrm{BP}(t))_{t\geq 0}$, the set $\mathcal{U}^{<\infty}_{K}$ contains at least one individual $u$ such that $D^{(u)}=\infty$, almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$ by Assumption 4.1.
Proof of Proposition 4.14 for $m=1$.
Recall $K$ from (PHC). We define
$T=T(K):=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}|\geq 1\}$ (4.75)
to be the stopping time at which the branching process $\mathrm{BP}$ contains at least one individual that has produced $K$ many children and that satisfies that their ‘eventual’ offspring is infinite. We let $V$ denote this individual. In case there are more than one such individual who have produced $K$ children by time $T$, then choose the first one according to their lexicographical ordering. We observe that conditionally on survival such an individual $V$ exists and that $T$ is finite, almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$. We can then write
$\mathcal{P}_{1}\subseteq\mathrm{BP}(T)\cup\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\ \mathcal{B}(u)>\mathcal{B}(T),\text{ $u$ catches up to all individuals in }\mathrm{Pre}(u)\}.$ (4.76)
As $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, it follows that $|\mathrm{BP}(T)|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely by Lemma 4.7. It remains to show that the second set on the right-hand side has a finite size almost surely. To this end, we define
$\mathcal{A}^{-}:=\mathcal{U}^{<\infty}_{K}\cap\mathrm{BP}(T)\qquad\text{and}\qquad\mathcal{A}^{+}:=\bigcup_{u\in A^{-}}\{u1,\ldots,uI(u)\}\setminus\mathcal{A}^{-},$ (4.77)
where we recall $I(u)$ from (4.72). In words, $\mathcal{A}^{-}$ consists of all individuals $u$ born up to time $T$ where, for each individual $u$ such that $D^{(u)}=\infty$, its sub-tree is restricted to the first $K$ children and its descendants (if they are born by time $T$). Then, $\mathcal{A}^{+}$ consists of the individuals that are not yet born by time $T$ among the first $I(u)$ children of individuals $u\in\mathcal{A}^{-}$.
With these sets at hand, we define the following events. For $u\in\mathcal{A}^{-}$, set
$\mathcal{C}_{u}:=\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,\text{ $uv$ catches up to $u$}\},$ (4.78)
and for $u\in\mathcal{A}^{+}$ we also define
$\mathcal{D}^{(u)}:=\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,\text{ $uv$ catches up to $V$}\}.$ (4.79)
With these events at hand, we observe that
$\{u\in\mathcal{U}_{\infty}:{}$ $D^{(u)}=\infty,\ \mathcal{B}(u)>\mathcal{B}(T),\text{ $u$ catches up to all individuals in }\mathrm{Pre}(u)\}$ (4.80)
$\subseteq\bigg{(}\bigcup_{\begin{subarray}{c}u\in\mathcal{A}^{-}\\ D^{(u)}=\infty\end{subarray}}\mathcal{C}_{u}\bigg{)}\cup\bigg{(}\bigcup_{u\in\mathcal{A}^{+}}\mathcal{D}^{(u)}\bigg{)}.$
Indeed, any individual $w$ born after time $T$ satisfies that it has an ancestor in $\mathcal{A}^{+}$ or, if not, it has a ancestor $u\in\mathcal{A}^{-}$ such that $D^{(u)}=\infty$ and $w=uv$ with $v_{1}>K$. Then, any such individual $w$ that catches up to all its predecessors and that has an ancestor in $\mathcal{A}^{+}$ needs to catch up to $V$. If it does not have an ancestor in $\mathcal{A}^{+}$ but has an ancestor $u\in\mathcal{A}^{-}$ with $D^{(u)}=\infty$, it needs to catch up to $u$. We now show that each of these sets has a finite size $\mathbb{P}_{\mathcal{S}}$-almost surely.
First, it is clear that $\mathcal{A}^{-}\subseteq\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, as already mentioned above. Since for each individual in $\mathcal{A}^{-}$ we include only a finite number of children in $A^{+}$, it follows that $\mathcal{A}^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely as well. It thus suffices to show that $|\mathcal{C}_{u}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $u\in\mathcal{A}^{-}$ such that $D^{(u)}=\infty$, and that $|\mathcal{D}^{(u)}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $u\in\mathcal{A}^{+}$.
The former claim follows directly from Proposition 4.13 by replacing $D^{(\varnothing)}$ with $D^{(u)}$ and ‘$v$ catches up to $\varnothing$’ by ‘$uv$ catches up to $u$’ in the statement. As a result,
[…]
For the latter claim we use that, by its definition, any individual $u\in\mathcal{A}^{+}$ is born strictly after time $\mathcal{B}(VK)=T$. With $\mathcal{F}_{T}$ the $\sigma$-algebra generated by the process up to time $T$, we write for $u\in\mathcal{A}^{+}$,
$\mathbb{E}_{\mathcal{S}}\left[|\mathcal{D}^{(u)}|\,\Big{|}\,\mathcal{F}_{T}\right]=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,\ uv\text{ catches up to }V\,\Big{|}\,\mathcal{F}_{T}\right).$ (4.82)
As in the proof of Proposition 4.13, for each $v=v_{1}\cdots v_{m}$ with $m\in\mathbb{N}_{0},v_{1},\ldots,v_{m}\in\mathbb{N}$ we can write the probability as
$\mathbb{P}_{\mathcal{S}}\Bigg{(}{}$ $D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in[m],$ (4.83)
$\exists j\in\mathbb{N}_{0}:\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell}}X(uv_{1}\ldots v_{\ell-1}i)+\sum_{i=1}^{K+j}X(uvi)\leq\mathcal{B}(VK)+\sum_{i=K+1}^{K+j}X(Vi)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
We note that $T=\mathcal{B}(VK)$ and that $\mathcal{B}(u)>T$. The offspring random variables $D^{(uv_{1}\cdots v_{\ell-1})}$ and $D^{(uv)}$ are not contained in the $\sigma$-algebra $\mathcal{F}_{T}$. Also, as an individual $V$ that produces an infinite offspring has been born by time $T$, survival is guaranteed, so that we can omit the conditioning on the event $\mathcal{S}$. By the independence of the offspring and inter-birth random variables, and since $\mathcal{B}(u)>T=\mathcal{B}(VK)$, we obtain the upper bound
[…]
$=\bigg{(}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\bigg{)}\mathbb{P}\!\left(\exists j\in\mathbb{N}_{0}:\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell-1}}X(uv_{1}\ldots v_{\ell-1}i)+\sum_{i=1}^{K+j}X(uvi)\leq\sum_{i=K+1}^{K+j}X(Vi)\right),$
where the final step holds almost surely, since all random variables in the probability are independent of $\mathcal{F}_{T}$. As a result, by using Lemma 4.8 we arrive at
$\mathbb{E}_{\mathcal{S}}\big{[}{}$ $|\mathcal{D}^{(u)}|\,|\,\mathcal{F}_{T}\big{]}$ (4.85)
$\leq 1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\mathrm{e}^{-\lambda\sum_{i=1}^{v_{\ell}}X(i)}\right]$
$=1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\sum_{m=1}^{\infty}\mathbb{E}\left[\sum_{j=1}^{D}\mathrm{e}^{-\lambda\sum_{i=1}^{j}X(i)}\right]^{m}$
$=1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\frac{\widehat{\mu}(\lambda)}{1-\widehat{\mu}(\lambda)},$
which is finite by the conditions in (PHC). As a result, since $A^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely,
[…]
Combining all of the above, we conclude that $\mathcal{P}_{1}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. ∎
We provide a proof of Proposition 4.14 for $m=2$ and later discuss the necessary adaptations when $m\geq 3$. As these are mainly a matter of notation and careful bookkeeping, we decided to provide the proof in the case $m=2$ rather than for general $m>1$ to aid the reader. We recall that in the case $m\geq 2$ we work under the additional assumption that the inter-birth times are exponentially distributed.
Proof of Proposition 4.14 for $m=2$.
We recall $K$ from (PHC) and define
$T:=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}|\geq 2\}$ (4.87)
to be the stopping time at which the branching process contains at least two individuals that have produced $K$ many children and that satisfy that their ‘eventual’ offspring is infinite. Let $V_{1}$ and $V_{2}$ denote these two individuals, where we order them such that $\mathcal{B}(V_{1}K)<\mathcal{B}(V_{2}K)=T$. We observe that conditionally on survival such individuals exist and that $T$ is finite almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$. We now define
$\mathcal{A}(K):=\mathcal{U}_{K}^{<\infty}\cap\mathrm{BP}(T).$ (4.88)
We also define
$\mathcal{A}^{+}:=\bigg{(}\bigcup_{u\in\mathrm{BP}(T)}\{u1,\ldots,uI(u)\}\bigg{)}\setminus\mathrm{BP}(T),$ (4.89)
where we recall $I(u)$ from (4.72). In words, $\mathcal{A}^{+}$ consists of the individuals that are not yet born by time $T$ among the first $I(u)$ children of individuals $u\in\mathrm{BP}(T)$.
If $u\in\mathcal{P}_{2}$, this implies that there are two individuals $w_{1},w_{2}\in\mathrm{Pre}(u)$ such that $u$ catches up to either $w_{1}$ or to $w_{2}$. Indeed, if $u$ does not catch up to both, then $u$ can not catch up to all but at most one predecessor. Hence, for any $u\in\mathcal{P}_{2}$ we need to identify two predecessors and $u$ needs to catch up to either one of them. We let c.u.t. be short for ‘catches up to’ for ease of writing. As a result, we have
$\mathcal{P}_{2}\subseteq\mathrm{BP}(T){}$ $\cup\bigcup_{u\in\mathcal{A}^{+}}\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\}$ (4.90)
$\cup\!\!\!\!\!\!\!\!\!\!\bigcup_{\begin{subarray}{c}u\in\mathcal{A}(K)\setminus\{V_{1}\}\\ D^{(u)}=\infty\end{subarray}}\!\!\!\!\!\!\!\!\!\!\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }u\text{ or }V_{2}\}$
$\cup\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\}.$
Let us take some time to argue why this inclusion holds. First, we observe that we can partition $\mathcal{U}_{\infty}$ into all individuals born up to time $T$, i.e. $\mathrm{BP}(T)$, and all individuals born after time $T$. The three unions on the right-hand side of (4.90) consider all individuals born after time $T$. Furthermore, each set in one of the two big unions, as well as the final set on the right-hand side, identifies two predecessors to which an individual $uv$ or $V_{1}v$ has to catch up. We thus conclude that the right-hand side contains $\mathcal{P}_{2}$. To show that $\mathcal{P}_{2}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, it thus suffices to show that the right-hand side is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
To this end, we first observe that $\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, by Lemma 4.7 and the fact that $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. It then directly follows that $\mathcal{A}(K)\subseteq\mathrm{BP}(T)$ is also finite $\mathbb{P}_{\mathcal{S}}$-almost surely. Since each individual in $\mathrm{BP}(T)$ contributes only a finite number of elements to $\mathcal{A}^{+}$, it follows that $\mathcal{A}^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely as well. Furthermore, the sets $\mathrm{BP}(T)$, $\mathcal{A}^{+}$, and $\mathcal{A}(K)$ are all $\mathcal{F}_{T}$ measurable, where $\mathcal{F}_{T}$ is the $\sigma$-algebra generated by the process up to time $T$. We aim to prove that the right-hand side of (4.90) is finite conditionally on $\mathcal{F}_{T}$. As $\mathcal{A}^{+}$ and $\mathcal{A}(K)$ are $\mathcal{F}_{T}$ measurable and finite $\mathbb{P}_{\mathcal{S}}$ almost surely, we are hence left to argue that each of the sets in the unions is finite conditionally on $\mathcal{F}_{T}$. These follow from similar arguments, with minor differences only.
Let us start with the sets
$\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1}\text{ or}V_{2}\},$ (4.91)
where $u$ is in $\mathcal{A}^{+}$. We prove that for any $u\in A^{+}$ and $j\in\{1,2\}$,
[…]
We obtain this by showing that the conditional expectation of the size of the set if finite. We write
$\mathbb{E}_{\mathcal{S}}{}$ [|$\Big{[}|\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ catches up to }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.93)
$=\sum_{r=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{j}\,\Big{|}\,\mathcal{F}_{T}\right).$
We define
$K_{j}:=\max\{k\in\mathbb{N}:\mathcal{B}(V_{j}k)\leq T\}\qquad\text{for }j\in\{1,2\},$ (4.94)
as the last child $V_{j}$ gave birth to before time $T$. Note that $K_{1}\geq K$ and $K_{2}=K$ $\mathbb{P}_{\mathcal{S}}$-almost surely by definition of $V_{1}$ and $V_{2}$. We then write the conditional probability as
$\mathbb{P}_{\mathcal{S}}\Bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\mathcal{B}(u)+\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{j}+n}\!\!X(uvi)\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!\!X(V_{j}i),$ (4.95)
$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in[r]\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
We now use that $u\in\mathcal{A}^{+}$, so that $u$ is not born by time $T$. At the same time, $V_{j}K_{j}$ is born by time $T$. That is, $\mathcal{B}(u)>T\geq\mathcal{B}(V_{j}K_{j})$. As a result, for $j=2$ we obtain the upper bound
$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K+1}^{K+n}\!\!\!\!X(V_{2}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right),$ (4.96)
where we also use that $K_{2}=K$ and that the offspring random variables $D^{(uv)}$ and $D^{(uv_{1}\cdots v_{\ell-1})}$ for $\ell\in[r]$ are independent of $\mathcal{F}_{T}$, as well as independent of the inter-birth random variables. Moreover, all inter-birth times in this upper bound are independent of $\mathcal{F}_{T}$ as well, since individual $V_{2}K$ is born exactly at time $T$ by definition. As a result, we can omit the conditioning on $\mathcal{F}_{T}$ $\mathbb{P}_{\mathcal{S}}$-almost surely to arrive at
$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K+1}^{K+n}\!\!\!\!X(V_{2}i)\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$ (4.97)
When $j=1$, we instead have the upper bound
$\mathbb{P}_{\mathcal{S}}{}$ $\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{1}+n}\!\!X(uvi)\leq\mathcal{B}(V_{1}K_{1})-T+\!\!\!\!\sum_{i=K_{1}+1}^{K_{1}+n}\!\!\!\!X(V_{1}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}$ (4.98)
×$\times\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$
Here, the inter-birth time $X(V_{1}(K_{1}+1))$ (as well as $T$ and $K_{1}$) depends on $\mathcal{F}_{T}$. However, by the memoryless property of the exponential distribution, it follows that $\mathcal{B}(V_{1}K_{1})+X(V_{1}(K_{1}+1))-T\sim X(V_{1}(K_{1}+1))^{\prime}$, where $X(V_{1}(K_{1}+1))^{\prime}$ is an i.i.d. copy of $X(V_{1}(K_{1}+1))$, which is independent of $\mathcal{F}_{T}$. As a result, we arrive at the equivalent upper bound as for $j=2$ $\mathbb{P}_{\mathcal{S}}$-almost surely, namely
$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{1}+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K_{1}+1}^{K_{1}+n}\!\!\!\!X(V_{1}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$ (4.99)
Here, only the random variable $K_{1}$ depends on $\mathcal{F}_{T}$, all other random variables are independent of $\mathcal{F}_{T}$. Lemma 4.8 then yields for both $j=1$ and $j=2$ the upper bound
$\prod_{i=K_{j}+1}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)},$ (4.100)
where we recall that $K_{2}=K$ is deterministic. However, we can almost surely obtain a deterministic upper bound by using that $K_{j}\geq K$ for both $j\in\{1,2\}$, which yields
$\prod_{i=K+1}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{K}X(i)\bigg{)}\right]\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)}.$ (4.101)
Using this in (4.93), we arrive at
$\mathbb{E}_{\mathcal{S}}{}$ [|$\Big{[}|\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.102)
[…]
×$\times\sum_{r=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)}$
[…]
where we use (4.7) in the last step. The upper bound is finite by the conditions in (PHC). We thus arrive at (4.92). This implies that the first line of the right-hand side of (4.90) is $\mathbb{P}_{\mathcal{S}}$-almost surely finite conditionally on $\mathcal{F}_{T}$. We proceed with the sets on the second line. We show that for any $u\in\mathcal{A}(K)\setminus\{V_{1}\}$ such that $D^{(u)}=\infty$,
[…]
and
[…]
The first probability being equal to one $\mathbb{P}_{\mathcal{S}}$-almost surely follows directly from Proposition 4.13. Indeed, by the definition of $V_{1}$ and $V_{2}$, it follows that any individual $u\in\mathcal{A}(K)\setminus\{V_{1}\}$ has given birth to fewer than $K$ children by time $T$ if $u\neq V_{2}$ and to exactly $K$ children if $u=V_{2}$. Hence, the event in the conditional probability is independent of $\mathcal{F}_{T}$. Proposition 4.13 then yields the desired result, by replacing $\varnothing$ with $u$ and ‘$v$ catches up to $\varnothing$’ with ‘$uv$ catches up to $u$’.
The proof of (4.104) follows in an analogous way when following the proof of Proposition 4.13, since we know that $V_{2}$ has produced $K$ children by time $T$, whereas the individuals $uv_{1}$ with $v_{1}>K$ have all not been born by time $T$.
To conclude the proof, it remains to show that the set on the final line on of (4.90) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$. That is, for $j\in\{1,2\}$,
[…]
Here, individuals $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants may have been born by time $T$, so that a slightly more careful argument is required. Let us define
$\mathcal{T}_{T}:=\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)\leq T\}$ (4.106)
to be the sub-tree rooted at $V_{1}$ but restricted to the children $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants that are born by time $T$. It is clear that $\mathcal{T}_{T}\subset\mathrm{BP}(T)$, so that it is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and that it is $\mathcal{F}_{T}$ measurable. For $w\in\mathcal{T}_{T}$, we define
$K(w):=\max\{0\leq k\leq D^{(w)}:wk\in\mathcal{T}_{T}\}$ (4.107)
to be the ‘last’ child of $w$ that is included in $\mathcal{T}_{T}$. We then have the inclusion
$\{V_{1}v:{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{j}\}$ (4.108)
$\subseteq\bigcup_{w\in\mathcal{T}_{T}}\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{j}\}.$
As a result, we have
$\mathbb{E}_{\mathcal{S}}\Big{[}{}$ |$|\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.109)
[…]
Each expected value in the sum on the right-hand side can be written as
$\sum_{r=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{j}\,\Big{|}\,\mathcal{F}_{T}\right),$ (4.110)
and each such probability equals
$\mathbb{P}_{\mathcal{S}}\bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\mathcal{B}(wK(w))+\!\!\!\!\sum_{i=K(w)+1}^{v_{1}}\!\!\!\!X(wi)+\sum_{\ell=2}^{r}\sum_{i=1}^{v_{\ell}}X(wv_{1}\cdots v_{\ell-1}i)+\!\!\!\!\sum_{i=1}^{K_{1}+n}\!\!X(wvi)$ (4.111)
$\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!X(V_{j}i),D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,r\}\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
We observe that
$\mathcal{B}(wK(w))$ $<T<\mathcal{B}(w(K(w)+1))=\mathcal{B}(wK(w))+X(w(K(w)+1)),$
and
$\mathcal{B}(V_{j}K_{j})$ $\leq T<\mathcal{B}(V_{j}(K_{j}+1))=\mathcal{B}(V_{j}K_{j})+X(V_{j}(K_{j}+1)),$
with equality in the second line only for $j=2$. Moreover, again using the memoryless property of the exponential distribution,
$\mathcal{B}(wK(w))+X(w(K(w)+1))-T$ $\sim X(w(K(w)+1))^{\prime}$
and
$\mathcal{B}(V_{j}K_{j})+X(V_{j}(K_{j}+1))-T$ $\sim X(V_{j}(K_{j}+1))^{\prime},$
where $X(w(K(w)+1))^{\prime}$ and $X(V_{j}(K_{j}+1))^{\prime}$ are i.i.d. copies of $X(w(K(w)+1))$ and $X(V_{j}(K_{j}+1))$, respectively, independent of all other random variables and also independent of $\mathcal{F}_{T}$. We thus obtain that the conditional probability can be bounded from above by
$\mathbb{P}_{\mathcal{S}}\bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\sum_{i=K(w)+1}^{v_{1}}\!\!\!\!X(wi)+\sum_{\ell=2}^{r}\sum_{i=1}^{v_{\ell}}X(wv_{1}\cdots v_{\ell-1}i)+\!\!\!\!\sum_{i=1}^{K_{j}+n}\!\!X(wvi)\leq\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!X(V_{j}i),$ (4.112)
$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in\{2,\ldots,r\}\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
The dependence on $\mathcal{F}_{T}$ is only in the random variables $K(w)$ and $K_{j}$, all other random variables are independent of $\mathcal{F}_{T}$. Also using that the offspring random variables are independent of the inter-birth random variables and applying Lemma 4.8 thus yields the upper bound
$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K_{j}+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\prod_{i=K(w)+1}^{v_{1}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.113)
×$\times\prod_{\ell=2}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$
Using this in (4.110) yields
$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K_{j}+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.114)
×$\times\sum_{r=1}^{\infty}{}$ $\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\prod_{\ell=2}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$
Since $K_{j}\geq K$ by definition, we obtain the upper bound
$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.115)
×$\times\sum_{r=0}^{\infty}\Bigg{(}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v}X(i)\bigg{)}\right]\Bigg{)}^{r}$
[…]
where we use (4.7) in the final step. By the conditions in (PHC), the first three terms are finite. The final sum is finite when $D^{(w)}$ is finite. In the case that $D^{(w)}=\infty$, it equals
$\sum_{v_{1}=K(w)+1}^{\infty}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ $\leq\mathbb{P}\!\left(D=\infty\right)^{-1}\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]^{-1}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\prod_{i=1}^{v}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.116)
$=\mathbb{P}\!\left(D=\infty\right)^{-1}\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]^{-1}\widehat{\mu}(\lambda),$
where we again use (4.7) to obtain the final equality. This upper bound is $\mathbb{P}_{\mathcal{S}}$-almost surely finite for any $w\in\mathcal{T}_{T}$, since $K(w)<\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, since $\mathcal{T}_{T}$ is a set wit a $\mathbb{P}_{\mathcal{S}}$-almost surely finite size, the upper bound in (4.109) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, which yields (4.105).
Combined with (4.104), (4.103), and (4.92), we thus conclude that
$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{P}_{2}|<\infty\,|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (4.117)
Since the probability measure is bounded from above, it follows that $\mathcal{P}_{2}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, which completes the proof. ∎
Adaptations for the proof of Proposition 4.4 with $m>2$. We now discuss the adaptations of the proof for $m=2$ to extend to the general case $m\geq 2$. To show that $\mathcal{P}_{m}$ has a finite size when $m>2$, we define
$T:=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}\geq m\},$ (4.118)
and we let $V_{1},\ldots,V_{m}$ be the first $m$ individuals that produce $K$ children, ordered such that $\mathcal{B}(V_{1}K)<\mathcal{B}(V_{2}K)<\cdots<\mathcal{B}(V_{m}K)=T$. The sets $\mathcal{A}(K)$ and $\mathcal{A}^{+}$ have the same definition. We then have
$\mathcal{P}_{m}\subseteq\mathrm{BP}(T){}$ $\cup\bigcup_{u\in\mathcal{A}^{+}}\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1},V_{2},\ldots,\text{ or }V_{m}\}$ (4.119)
$\cup\bigcup_{u\in\mathcal{A}(K)\setminus\{V_{1},\ldots,V_{m-1}\}}\!\!\!\!\!\!\!\!\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{c.u.t.\ }u,\text{ or }V_{2},\ldots,V_{m}\}$
$\cup\bigcup_{j=1}^{m-1}\{V_{j}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{j}v)>T,D^{(V_{j}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{1},\ldots,\text{ or }V_{m}\}.$
Proving that each of the sets on the right-hand side is finite then follows similar arguments as in the proof for $m=2$.
4.5. Unique persistent $m$-hubs
We conclude this section by proving Theorem 4.5.
Proof of Theorem 4.5.
It follows from Theorem 4.2 that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when Assumption 4.1 and (PHC) are satisfied. From the proof of Theorem 4.2 (using Proposition 4.14 in particular), it follows that the number of persistent $m$-hubs is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. We assume that there is more than one persistent $m$-hub with positive probability and argue by contradiction. Suppose that both $u$ and $v$ are both persistent $m$-hubs. It follows that there exists $T>0$ such that their offspring is the same for all time $t\geq T$. That is, $\deg_{t}(u)=\deg_{t}(v)$ for all $t\geq T$. This is equivalent to there existing $N\in\mathbb{N}$ such that $\mathcal{B}(uk)=\mathcal{B}(vk)$ for all $k\geq N$. That is,
$\mathcal{B}(u)+\sum_{i=1}^{k}X(ui)=\mathcal{B}(v)+\sum_{i=1}^{k}X(vi)\qquad\text{or, equivalently,}\qquad\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{k}(X(ui)-X(vi))=0.$ (4.120)
Let us set $S_{i}:=X(ui)-X(vi)$. It is clear that the series $\sum_{i=1}^{\infty}S_{i}$ is finite almost surely by Assumption 4.1 and the second condition in (PHC). Since $\mathcal{B}(v)$ and $\mathcal{B}(u)$ can be represented as finite sums of inter-birth random variables, the equality on the right can only hold with positive probability when the distribution of $\sum_{i=1}^{\infty}S_{i}$ has an atom. However, the conditions in Theorem 4.5, combined with Lemma 4.9 and Remark 4.10, yield that the distribution does not have an atom. This contradiction thus implies that $u=v$ $\mathbb{P}_{\mathcal{S}}$-almost surely and there thus exists a unique persistent $m$-hub.
To conclude the proof, we show (4.5). We know that there exists a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$. Suppose $u$ is the unique persistent $m$-hub and $v$ is the unique persistent $m^{\prime}$-hub, with $1\leq m<m^{\prime}<\infty$ two integers. We now assume that
$\mathbb{P}_{\mathcal{S}}\left(\exists T>0\ \forall t\geq T:\deg_{t}(u)>\deg_{t}(v)\right)=0,$ (4.121)
and aim to show a contradiction. Since $u$ and $v$ are the unique $m$-hub and $m^{\prime}$-hub, respectively, with $m<m^{\prime}$, we have
$\mathbb{P}_{\mathcal{S}}\left(\exists T^{\prime}>0\ \forall t\geq T^{\prime}:\deg_{t}(u)\geq\deg_{t}(v)\right)$ $=1,$
and
$\mathbb{P}_{\mathcal{S}}\left(\exists T^{\prime\prime}>0\ \forall t\geq T^{\prime\prime}:\deg_{t}(u)=\deg_{t}(v)\right)$ $=0.$
Combining this with our assumption, it follows that $\mathbb{P}_{\mathcal{S}}$-almost surely,
$\mathcal{B}(uk)\leq\mathcal{B}(vk)<\mathcal{B}(u(k+1))\text{ holds for infinitely many }k\in\mathbb{N}$ (4.122)
is satisfied. That is, $u$ and $v$ have the same number of offspring in the non-empty intervals $[\mathcal{B}(vk),\mathcal{B}(u(k+1)))$, for infinitely many $k$. This implies that
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{k}(X(vi)-X(ui))<X(u(k+1))\text{holds for infinitely many }k\in\mathbb{N}.$ (4.123)
We now fix an arbitrary $\varepsilon>0$. Under the condition in (4.4), it follows that $X(u(k+1))\leq\varepsilon$ for all but finitely many $k$. Hence, the inequalities
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{k}(X(vi)-X(ui))<\varepsilon$ (4.124)
hold for infinitely many $k$. Then, since the infinite sum $\sum_{i=1}^{\infty}(X(vi)-X(ui))$ is almost surely convergent under the second condition in (PHC) and $(2)$ in Assumption 4.1,
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{\infty}(X(vi)-X(ui))\leq\varepsilon,$ (4.125)
with positive probability. As $\varepsilon$ is arbitrary, it follows that
$\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{\infty}(X(vi)-X(ui))=0$ (4.126)
holds with positive probability. However, combining the conditions in (4.3) with Lemma 4.9 yields that the distribution of $\sum_{i=1}^{\infty}(X(vi)-X(ui))$ has no atoms. Since $\mathcal{B}(v)$ and $\mathcal{B}(u)$ can be represented as finite sums of inter-birth random variables, it thus follows that (4.126) occurs with probability zero, which leads to a contradiction and concludes the proof. ∎
5. Persistence in the CTPAVD branching process
In this section we prove Theorem 2.6. Unlike in the analysis carried out in Section 4, a persistent $m$-hub, as in the sense of (PH-$m$), does not emerge $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$, since individuals live an almost surely finite time (i.e. $L<\infty$ a.s.). This requires a more direct approach where we provide precise asymptotic expansions of the random variables $O_{t}^{\mathrm{ct}}$ and $I_{t}^{(1),\mathrm{ct}}$ to show that their difference is a tight sequence of random variables (with respect to $t$), as in (3.14).
5.1. Preliminary results
We first state some preliminary results that are used throughout the remainder of this section. We recall the definition of the CTPAVD branching process in Section 3, and that we write $D$ for the offspring of an individual, $S_{k}$ for the time an individual needs to give birth to $k$ children, and $L:=S_{D+1}$ for the lifetime of an individual. Furthermore, we recall the sequences $\varphi_{1},\varphi_{2},$ and $\rho_{1}$ from (2.9) and define the sequence
$\rho_{2}(k):=\sum_{i=0}^{k-1}\Big{(}\frac{d(i)}{b(i)+d(i)}\Big{)}^{2}\qquad\text{for }k\in\mathbb{N}.$ (5.1)
We extend the domain of $\rho_{2}$ to $\mathbb{R}_{+}$ by linear interpolation. We then have the following results.
Lemma 5.1 (Offspring tail bounds, Lemmas $5.3$ and $5.4$ in [23]).
Let $D$ be as in (2.6). Then,
$\mathbb{P}\!\left(D\geq k\right)\leq\mathrm{e}^{-\rho_{1}(k)-\frac{1}{2}\rho_{2}(k)}.$ (5.2)
In particular, $\mathbb{P}\!\left(D=\infty\right)=0$ if and only if Assumption (F-D) is satisfied. Furthermore, assume that $b$ and $d$ are such that Assumption (F-D) is satisfied. When $\rho_{2}$ diverges and $d=o(b)$,
$\mathbb{P}\!\left(D\geq k\right)=\mathrm{e}^{-\rho_{1}(k)-(\frac{1}{2}+o(1))\rho_{2}(k)}.$ (5.3)
If, instead, $\lim_{k\to\infty}\rho_{2}(k)$ exists,
$\mathbb{P}\!\left(D\geq k\right)=\mathrm{e}^{-\rho_{1}(k)-\mathcal{O}(1)}.$ (5.4)
Lemma 5.2 (Probability to survive with many children, Lemma $5.9$ in [23]).
$(i)$ Suppose that there exist $x\geq 0$ and $K=K(x)\in\mathbb{N}_{0}$ such that $d(i)\geq x$ for all $i\geq K$. Then, for all $k\geq K$ and all $t^{\prime}\geq t\geq 0$,
$\mathbb{P}\!\left(D\geq k,S_{k}\leq t,S_{D+1}>t^{\prime}\right)\leq\mathrm{e}^{-x(t^{\prime}-t)}\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathbbm{1}_{\{S_{k}\leq t\}}\mathrm{e}^{x(S_{k}-t)}\right].$ (5.5)
$(ii)$ Suppose that there exist $x\geq 0$ and $K=K(x)\in\mathbb{N}$ such that $d(i)\leq x$ for all $i\geq K$. Then, for all $k\geq K$ and all $t^{\prime}\geq t\geq 0$,
$\mathbb{P}\!\left(D\geq k,S_{k}\leq t,S_{D+1}>t^{\prime}\right)\geq\mathrm{e}^{-x(t^{\prime}-t)}\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathbbm{1}_{\{S_{k}\leq t\}}\mathrm{e}^{x(S_{k}-t)}\right].$ (5.6)
We then introduce the following notation. First, recall that $\mathcal{S}$ denotes the event that the branching process survives, and recall that
$\mathbb{P}_{\mathcal{S}}\left(\cdot\right):=\mathbb{P}\!\left(\cdot\,|\,\mathcal{S}\right)\quad\text{and}\quad\mathbb{E}_{\mathcal{S}}[\cdot]:=\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ (5.7)
denote the conditional probability measure and its corresponding expected value, under the event $\mathcal{S}$. For $0\leq s<t<\infty$, we further define
$\mathcal{B}(s,t):=\{u\in\mathcal{U}_{\infty}:\sigma_{u}\in[s,t]\}$ (5.8)
as the set of individuals born in the time interval $[s,t]$. Also, recall $\mathcal{A}_{t}^{\mathrm{ct}}$ from (3.5) as the set of individuals alive at time $t$, and $\mathcal{R}$ from (2.6) as the point process that governs the production of offspring of individuals. We have the following results regarding the growth-rate of $\mathcal{A}^{\mathrm{ct}}_{t}$ and $\mathcal{B}(s,t)$. The first is similar to the well-known Kesten-Stigum theorem and is a direct consequence of results for general CMJ branching processes in [18, 32]; the second is Corollary $6.6$ in [23].
Proposition 5.3 (Growth rate of branching process).
Suppose that $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied. There exists a non-negative random variable $W$, such that
$\mathrm{e}^{-\lambda^{*}\!t}|\mathcal{A}^{\mathrm{ct}}_{t}|\overset{\mathrm{a.s.}}{\longrightarrow}W.$ (5.9)
Moreover, with
^$\widehat{\mathcal{R}}^{\lambda^{*}}(t):=\int_{0}^{t}\mathrm{e}^{-\lambda^{*}u}\,\mathcal{R}(\mathrm{d}u),$ (5.10)
the following are equivalent:
  1. (a)
    $\mathbb{E}\left[\widehat{\mathcal{R}}^{\lambda^{*}}(\infty)\log^{+}\widehat{\mathcal{R}}^{\lambda^{*}}(\infty)\right]<\infty$,
  2. (b)
    $\mathbb{E}\left[W\right]>0$,
  3. (c)
    $\mathbb{E}\left[\mathrm{e}^{-\lambda^{*}\!t}|\mathcal{A}^{\mathrm{ct}}_{t}|\right]\to\mathbb{E}\left[W\right]$ as $t\to\infty$,
  4. (d)
    $W>0$ almost surely on $\mathcal{S}$.
Proposition 5.3 is a direct result of [32, Theorem $5.4$] and [18] (see also [32, Proposition $1.1$] for a condensed version). We have the following useful corollary, which is a part of Corollary $6.6$ in [23].
Corollary 5.4 (Corollary $6.6$ in [23]).
Suppose that $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied, and let $\mathcal{S}:=\{\lim_{t\to\infty}|\mathrm{BP}(t)|=\infty\}$ be the event that the process $\mathrm{BP}$ survives. Then,
$\lim_{M\to\infty}\mathbb{P}\!\left(\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}\!t}\geq M\right)$ $=0,$ (5.11)
and
$\lim_{M\to\infty}\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}\!t}\leq 1/M\right)$ $=0.$ (5.12)
We conclude this sub-section with a few elementary results regarding the functions $\varphi_{1},\varphi_{2}$, and $\rho_{1},\rho_{2}$, as well as $\alpha$ and $\mathcal{K}_{\alpha}$, as introduced in Section 2.
Lemma 5.5 (Lemma $5.13$, [23]).
  1. (a)
    Suppose that Assumption (F-D) is satisfied and $d=o(b)$. Then, $\rho_{2}=o(\rho_{1})$.
  2. (b)
    Suppose that Assumption (N-E) is satisfied and that $\overline{d}:=\limsup_{i\to\infty}d(i)<\infty$. Then, $\limsup_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\leq\overline{d}$. In particular, when $d$ converges to zero, then $\rho_{1}=o(\varphi_{1})$. Similarly, if $\underline{d}:=\limsup_{i\to\infty}d(i)>0$, then $\liminf_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\geq\underline{d}$. In particular, if $d$ tends to infinity, then $\varphi_{1}=o(\rho_{1})$.
  3. (c)
    Suppose that Assumption (N-E) is satisfied and $\lim_{i\to\infty}d(i)=d^{*}\in[0,\infty)$. Then, we have $\mathcal{K}_{\alpha}(t)=o(t)$. Moreover, for any function $s\colon(0,\infty)\to\mathbb{R}_{+}$ such that $s(t)=o(t)$, we have $\mathcal{K}_{\alpha}(t)-\mathcal{K}_{\alpha}(t-s(t))=o(s(t))$.
5.2. The oldest alive individual
In this sub-section we study the quantity $O_{t}^{\mathrm{ct}}$; the birth-time of the oldest alive individual in $\mathrm{BP}(t)$ conditionally on survival. To this end, we first obtain precise asymptotic distributional properties of the lifetime $L=S_{D+1}$ of individuals in the branching process, to then prove when the oldest individual alive at time $t$ is born.
5.2.1. Lifetime distribution
We know from [23, Lemma $6.1$] that the lifetime $L$ of an individual is asymptotically distributed as an exponential random variable with rate $d^{*}$ when $d$ converges to $d^{*}\leq R$. That is $\mathbb{P}\!\left(L>t\right)=\exp(-(d^{*}+o(1))t)$ as $t$ tends to infinity. The following result provides a higher-order result for the tail distribution when Assumption (C-V) is also satisfied.
Lemma 5.6 (Third-order asymptotic lifetime distribution).
Suppose that $b$ and $d$ are such that Assumptions (N-E), (F-D), and (C-V) are satisfied, and recall $R$ from (2.14). Also suppose that $\lim_{i\to\infty}d(i)=d^{*}\in[0,R)$. Recall the function $\mathcal{K}_{\alpha}$ from (2.17), and suppose that Assumption $\mathcal{K}_{\alpha}$ is satisfied. Then,
$\sup_{t\geq 0}\big{|}\log(\mathbb{P}\!\left(L>t\right))+d^{*}t+\mathcal{K}_{\alpha}(t)\big{|}<\infty.$ (5.13)
Remark 5.7.
This result complements [23, Lemma $6.2$] where a different third-order rescaling is provided for $\log\mathbb{P}\!\left(L>t\right)$ in the case that Assumption (C-V) is not satisfied.
Proof.
We start by making the following observations. Since $\varphi_{2}$ converges and $d$ converges, it follows that $\rho_{2}$, as in (5.1), converges. Furthermore, as $\varphi_{2}(k)={\rm Var}(S_{k})$ converges and $\varphi_{1}(k)=\mathbb{E}\left[S_{k}\right]$, it follows that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$.
Lower bound. We set $k:=\lfloor\varphi_{1}^{-1}(t)\rfloor$. Then, using Lemma 5.1 and using the above observations,
$\mathbb{P}\!\left(L>t\right)$ $\geq\mathbb{P}\!\left(D\geq k\right)\mathbb{P}\!\left(S_{k}>t\right)$ (5.14)
$=\mathrm{e}^{-\rho_{1}(k)+\mathcal{O}(1)}\mathbb{P}\!\left(S_{k}>t\right)$
$\geq\mathrm{e}^{-\rho_{1}(\varphi_{1}^{-1}(t))+\mathcal{O}(1)}\mathbb{P}\!\left(S_{k}-\varphi_{1}(k)>0)\right)$
$=\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)+\mathcal{O}(1)}(\mathbb{P}\!\left(M_{\infty}>0\right)+o(1)),$
which yields the desired lower bound.
Upper bound. We set
$M:=\lceil(1-\varepsilon)t\rceil\qquad\text{and}\qquad k_{i}:=\lceil\varphi_{1}^{-1}(t-i)\rceil\quad\text{for }i\in\{0,\ldots,M\}.$ (5.15)
Further, we fix $\theta\in(d^{*},R)$ and take $\varepsilon\in(0,1)$ small enough such that $\theta(1-\varepsilon)>d^{*}$. Then, using Lemma 5.1 and Chernoff bounds,
$\mathbb{P}\!\left(L>t\right)$ $\leq\mathbb{P}\!\left(D\geq k_{0}\right)+\mathbb{P}\!\left(S_{k_{M}}>t\right)+\sum_{i=0}^{M-1}\mathbb{P}\!\left(D\geq k_{i+1}\right)\mathbb{P}\!\left(S_{k_{i}}>t\right)$ (5.16)
$\leq\mathrm{e}^{-\rho_{1}(k_{0})}+\mathrm{e}^{-\theta t}\prod_{i=0}^{k_{M}-1}\frac{b(i)+d(i)}{b(i)+d(i)-\theta}+\mathrm{e}^{-\theta t}\sum_{i=0}^{M-1}\mathrm{e}^{-\rho_{1}(k_{i+1})}\prod_{i=0}^{k_{i}-1}\frac{b(i)+d(i)}{b(i)+d(i)-\theta}.$
Since $\varphi_{2}$ converges, it follows that for any $i\leq M$,
$\prod_{i=0}^{k_{i}-1}\frac{b(i)+d(i)}{b(i)+d(i)-\theta}=\exp(\theta\varphi_{1}(k_{i})+\mathcal{O}(1))=\exp(\theta t-\theta i+\mathcal{O}(1)),$ (5.17)
where the $\mathcal{O}(1)$ term is independent of $t$ and $i$. By the definition of $\alpha$ in (2.16), we thus arrive at the upper bound
$\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)}+\mathrm{e}^{-\theta(1-\varepsilon)t}+\sum_{i=0}^{M-1}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}.$ (5.18)
As $\theta(1-\varepsilon)>d^{*}$ by the choice of $\theta$ and $\varepsilon$, and $\mathcal{K}_{\alpha}=o(t)$ by Lemma 5.5 $(c)$, it follows that the second term is negligible compared to the first term. To deal with the sum, we define
$\widetilde{\alpha}(k)=\sum_{i=0}^{k-1}\frac{|d(i)-d^{*}|}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}_{0},\quad\text{and}\quad\mathcal{K}_{\widetilde{\alpha}}(t):=\widetilde{\alpha}\big{(}\varphi_{1}^{-1}(t)\big{)}\qquad\text{for }t\geq 0,$ (5.19)
where we extend the domain of $\widetilde{\alpha}$ to $\mathbb{R}_{+}$ by linear interpolation. We then distinguish between two cases.
$\bm{\mathcal{K}_{\widetilde{\alpha}}}$ converges. In this case, it follows that $\mathcal{K}_{\alpha}$ is bounded, as $|\mathcal{K}_{\alpha}(t)|\leq\mathcal{K}_{\widetilde{\alpha}}(t)$ for all $t\geq 0$. Hence the term $\mathcal{K}_{\alpha}(t-(i+1))$ in the exponential summands in (5.18) can be included in the $\mathcal{O}(1)$ term for each $i\leq M-1$, to yield
$\sum_{i=0}^{M-1}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}=\sum_{i=0}^{M-1}\mathrm{e}^{-d^{*}t-(\theta-d^{*})i+\mathcal{O}(1)}\leq C\mathrm{e}^{-d^{*}t},$ (5.20)
for some constant $C>0$. Combined with the first two terms in (5.18) this implies the desired result, since $\mathcal{K}_{\alpha}$ is bounded.
$\bm{\mathcal{K}_{\widetilde{\alpha}}}$ tends to infinity. In this case, we let $C^{\prime}>2(\theta-d^{*})^{-1}$ be a constant and split the sum in (5.18) into two parts, namely
$\sum_{i=0}^{M-1}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}={}$ $\sum_{i=0}^{C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t)}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}$ (5.21)
$+\sum_{i=C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t)}^{M-1}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}.$
We note that this is possible since $\mathcal{K}_{\widetilde{\alpha}}(t)=o(t)$ (this follows in the same way as $\mathcal{K}_{\alpha}(t)=o(t)$ in Lemma 5.5 $(c)$) and $M\geq(1-\varepsilon)t$. Using the second part of Lemma 5.5 $(c)$ yields
$\mathcal{K}_{\alpha}(t-(i+1))-\mathcal{K}_{\alpha}(t)=o(i)\quad\text{for any }i\leq C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t).$ (5.22)
As a result, since $\theta>d^{*}$, the first sum on the right-hand side of (5.21) can be bounded from above for all $t$ large by
$\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)}\sum_{i=0}^{C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t)}\mathrm{e}^{-(\theta-d^{*}+o(1))i+\mathcal{O}(1)}\leq C_{2}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)},$ (5.23)
for some constant $C_{2}>0$. For the second sum on the right-hand side of (5.21), we bound for each index $i$ in the sum,
$-\mathcal{K}_{\alpha}(t-(i+1))\leq\sup_{x\in[\varepsilon t,t]}|\mathcal{K}_{\alpha}(x)|\leq\sup_{x\in[\varepsilon t,t]}\mathcal{K}_{\widetilde{\alpha}}(x)=\mathcal{K}_{\widetilde{\alpha}}(t),$ (5.24)
where the final step uses that $\mathcal{K}_{\widetilde{\alpha}}$ is increasing. As a result, we obtain
$\sum_{i=C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t)}^{M-1}\exp(-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1))\leq\exp(-d^{*}t+\mathcal{K}_{\widetilde{\alpha}}(t)-(\theta-d^{*})C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t)+\mathcal{O}(1)).$ (5.25)
By the choice of $C^{\prime}$, the upper bound is at most
$\exp(-d^{*}t-\mathcal{K}_{\widetilde{\alpha}}(t)+\mathcal{O}(1))\leq\exp(-d^{*}t-\mathcal{K}_{\alpha}(t)+\mathcal{O}(1)).$ (5.26)
Combined with the bound in (5.23) for the first sum on the right-hand side of (5.21) and the first two terms in (5.18), we thus arrive at the desired result and conclude the proof. ∎
5.2.2. Non-survival of old individuals
We conclude this sub-section by proving asymptotic results for $O_{t}^{\mathrm{ct}}$. In essence, we show that all individuals that are ‘too old’ have all died by time $t$ and that there exists ‘young enough’ individuals that survive up to time $t$ with probability close to one. This is made precise in the following result.
Proposition 5.8 (The oldest alive individual).
Suppose that the sequences $b$ and $d$ are such that Assumptions (N-E), (F-D),  (Ma), and (C-V) are satisfied. Recall $R$ from (2.14) and recall that the event $\mathcal{S}$ denotes that the branching process $\mathrm{BP}$ survives. Suppose that $\lim_{i\to\infty}d(i)=d^{*}\in[0,R)$ and that Assumption $\mathcal{K}_{\alpha}$ is satisfied. Then, for any $\varepsilon,A>0$ there exist $K,t^{\prime}>0$ such that for all $t>t^{\prime}$,
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]:\Big{|}O_{s}^{\mathrm{ct}}-\frac{d^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big{|}<K\right)>1-\varepsilon.$ (5.27)
Proof.
For $K>0$, we define
$\ell_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-K,\qquad u_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K.$ (5.28)
Then, as $O_{s}^{\mathrm{ct}}$ is increasing in $s$,
$\mathbb{P}_{\mathcal{S}}{}$ $\bigg{(}\forall s\in[t-A,t+A]:\Big{|}O^{\mathrm{ct}}_{s}-\frac{d^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big{|}<K\bigg{)}$ (5.29)
$\geq 1-\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)-\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t+A}\geq u_{t}\right).$
By conditioning on the number of individuals born up to time $\ell_{t}$ and using a union bound, we bound the first probability on the right-hand side from above by
$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)$ $=\mathbb{P}_{\mathcal{S}}\left(\exists v\in\mathcal{B}(0,\ell_{t}):L^{(v)}>t-A-\mathcal{B}(v)\right)$ (5.30)
$\leq\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\mathrm{e}^{\lambda^{*}\ell_{t}}\mathbb{P}\!\left(L>t-A-\ell_{t}\right)+\mathbb{P}_{\mathcal{S}}\left(\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}>M\right).$
The second probability on the right-hand side is at most $\varepsilon/4$ when we choose $M=M(\varepsilon)$ sufficiently large, as follows from (5.11) in Corollary 5.4. For the first probability, we use Lemma 5.6 and the definition of $\ell_{t}$ to further bound this from above by
$\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\big{(}\lambda^{*}{}$ $\ell_{t}-d^{*}(t-A-\ell_{t})-\mathcal{K}_{\alpha}(t-A-\ell_{t})+C_{1}\big{)}$ (5.31)
$=\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\Big{(}{}$ $\mathcal{K}_{\alpha}(r(t))-\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big{)}-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\Big{)},$
where $C_{1}>0$ is a sufficiently large constant. We use Lemma 5.5 $(c)$ to obtain that for any $K,A>0$ there exists $t^{\prime}>0$ such that
$\Big{|}\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big{)}-\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big{)}\Big{|}<1.$ (5.32)
That is, the left-hand side is at most $1$ for all $t\geq t^{\prime}=t^{\prime}(K,A)$. Then, Assumption $\mathcal{K}_{\alpha}$ yields that there exists a constant $C_{2}>0$ such that for all $t>0$,
$\Big{|}\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big{)}-\mathcal{K}_{\alpha}(r(t))\Big{|}\leq C_{2}.$ (5.33)
We thus arrive for all $t\geq t^{\prime}$ at the upper bound
$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)\leq\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\big{(}C_{2}+1-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\big{)}+\varepsilon/4\leq\varepsilon/2,$ (5.34)
where we choose $K$ sufficiently large so that the first term is at most $\varepsilon/4$ in the final step. We note that we can first choose such a $K$ and then take $t^{\prime}=t^{\prime}(K,A)$ sufficiently large such that (5.32) holds.
We then bound the second probability on the right-hand side of (5.29). In a similar way as for the first term, with $\widetilde{u}_{t}=u_{t}-K/2$,
$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t+A}\geq u_{t}\right)\leq\mathbb{P}_{\mathcal{S}}\left(\forall v\in\mathcal{B}(\widetilde{u}_{t},u_{t}):L^{(v)}\leq t+A-\widetilde{u}_{t}\right).$ (5.35)
For ease of writing, we define the events
$\mathcal{E}_{t}:=\{\forall v\in\mathcal{B}(\widetilde{u}_{t},u_{t}):L^{(v)}\leq t+A-\widetilde{u}_{t}\},\qquad\mathcal{D}_{t}:=\{|\mathcal{B}(\widetilde{u}_{t},u_{t})|>m\}\quad\text{for }m\in\mathbb{N},t\geq 0.$ (5.36)
We can thus bound
$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t+A}\geq u_{t}\right)\leq\mathbb{P}_{\mathcal{S}}\left(\mathcal{E}_{t}\right)=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbbm{1}_{\mathcal{S}}\mathbbm{1}_{\mathcal{E}_{t}}\right]\leq\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbbm{1}_{\mathcal{S}}\mathbbm{1}_{\mathcal{E}_{t}}\mathbbm{1}_{\mathcal{D}_{t}}\right]+\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbbm{1}_{\mathcal{S}}\mathbbm{1}_{\mathcal{D}_{t}^{c}}\right].$ (5.37)
The second term on the right-hand side equals the probability of the event $\mathcal{D}_{t}^{c}$ under the conditional probability measure $\mathbb{P}_{\mathcal{S}}$. For the first term, we condition on two things. First, we condition $\mathcal{F}_{\widetilde{u}_{t}}$, the $\sigma$-algebra generated by the branching process up to time $\widetilde{u}_{t}$. Second, we condition on $|\mathcal{B}(\widetilde{u}_{t},u_{t})|$, the number of individuals born in the branching process in the interval $(\widetilde{u}_{t},u_{t})$. We can thus write
$\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbbm{1}_{\mathcal{S}}\mathbbm{1}_{\mathcal{E}_{t}}\mathbbm{1}_{\mathcal{D}_{t}}\right]$ $=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbb{E}\left[\mathbbm{1}_{\mathcal{S}}\mathbbm{1}_{\mathcal{E}_{t}}\mathbbm{1}_{\mathcal{D}_{t}}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right]\right]$ (5.38)
$\leq\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{E}\left[\mathbbm{1}_{\mathcal{E}_{t}}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right]\right]$
$=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\mathbb{E}\left[\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{P}\!\left(\mathcal{E}_{t}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right)\right].$
Conditionally on $|\mathcal{B}(\widetilde{u}_{t},u_{t})|$ and upon the event $\mathcal{D}_{t}$, we can use the independence of the lifetimes of distinct individuals to bound the terms in the expected value from above by
$\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{P}\!\left(\mathcal{E}_{t}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right)\leq\mathbbm{1}_{\mathcal{D}_{t}}\big{(}1-\mathbb{P}\!\left(L\geq\overline{r}(t)-\widetilde{u}_{t}\right)\big{)}^{m}\leq\exp\big{(}-m\mathbb{P}\!\left(L\geq\overline{r}(t)-\widetilde{u}_{t}\right)\big{)},$ (5.39)
where we have omitted the indicator random variable and used that $1-x\leq\mathrm{e}^{-x}$ for $x\in\mathbb{R}$ in the last step. Combining all of the above we thus arrive at
$\mathbb{P}_{\mathcal{S}}\left(O_{t+A}^{\mathrm{ct}}\geq u_{t}\right)\leq\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\big{(}-m\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\big{)}+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq m\right).$ (5.40)
We choose $m=M^{-1}\exp(\lambda^{*}u_{t})$ for a large constant $M>0$ to be determined to obtain the upper bound
$\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\big{(}-M^{-1}\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\mathrm{e}^{\lambda^{*}u_{t}}\big{)}+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq M^{-1}\mathrm{e}^{\lambda^{*}u_{t}}\right).$ (5.41)
We bound the first term from above and take care of the second term later. We use Lemma 5.6 and the definition of $u_{t}$ and $\widetilde{u}_{t}$ to arrive at the upper bound
$\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\big{(}-M^{-1}\exp\big{(}-d^{*}(t{}$ $+A-\widetilde{u}_{t})-\mathcal{K}_{\alpha}(t+A-\widetilde{u}_{t})-C_{3}+\lambda^{*}u_{t}\big{)}\big{)}$ (5.42)
$=\mathbb{P}(\mathcal{S})^{-1}\exp\Big{(}-M^{-1}\exp\Big{(}{}$ $\mathcal{K}_{\alpha}(r(t))-\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-\frac{K}{2}\Big{)}$
$+(d^{*}/2+\lambda^{*})K-d^{*}A-C_{3}\Big{)}\Big{)},$
where $C_{3}>0$ is a large constant. We again use Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$ to obtain similar bounds as in (5.32) and (5.33). We thus obtain, for some constants $C_{3}>0$ and $t^{\prime}=t^{\prime}(K,A)>0$ that for all $t>t^{\prime}$ we have have the upper bound
$\mathbb{P}(\mathcal{S})^{-1}\exp\Big{(}-M^{-1}\exp\Big{(}C_{3}+(d^{*}/2+\lambda^{*})K-d^{*}A-C_{3}\Big{)}\Big{)}$ (5.43)
bound the first two terms in the exponential from below by $-C_{4}$, uniformly in $A$ and $K$ and for all $t>t^{\prime}$, where $t^{\prime}>0$ and $C_{4}>0$ are large constants. Then, we choose $K$ sufficiently large so that the entire term is at most $\varepsilon/4$. Again, we can first choose such a $K$ and then choose $t^{\prime}$.
It remains to bound the final term in (5.41) from above by $\varepsilon/4$ to obtain the desired result. To this end, we write
$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq M^{-1}\mathrm{e}^{\lambda^{*}u_{t}}\right)={}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t})|\mathrm{e}^{-\lambda^{*}u_{t}}-|\mathcal{B}(0,\widetilde{u}_{t})|\mathrm{e}^{-\lambda^{*}\widetilde{u}_{t}-\lambda^{*}K/2}\leq M^{-1}\right)$ (5.44)
[…]
We can bound the first term on the right-hand side from above by
$\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq\frac{2}{M}\right),$ (5.45)
which can be made smaller than $\varepsilon/6$ when taking $M$ sufficiently large by (5.12) in Corollary 5.4. Given such a fixed $M$, we then choose $K$ sufficiently large so that, by (5.11) in Corollary 5.3,
$\mathbb{P}_{\mathcal{S}}\left(\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\geq\frac{1}{M}\mathrm{e}^{\lambda^{*}K/2}\right)<\varepsilon/12.$ (5.46)
Combined, we thus obtain
$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq M^{-1}\mathrm{e}^{\lambda^{*}u_{t}}\right)<\varepsilon/4,$ (5.47)
as desired, which concludes the proof. ∎
5.3. The richest individual
In this sub-section we study $I^{\mathrm{ct}}_{t}$, the birth-time of the richest individual in $\mathrm{BP}(t)$. That is, the individual that is alive at time $t$ and has produced the largest offspring by time $t$. To this end, we show that individuals that are ‘too young’ produce ‘small’ offspring, whereas there exist ‘old’ individuals that produce large offspring.
5.3.1. Young individuals have small offspring and no individual has too large offspring
In this section we show that individuals that are alive in an interval of time centred at $t$ but born ‘too late’, that is, born too long after the oldest alive individual has been born, have too small offspring to attain the largest offspring among all alive individuals. Similarly, we show that there does not exist any individual that attains a ‘too large’ offspring by time $\approx t$.
Fix $C_{1},C_{1}^{\prime}>0$ and set
$k$ $:=\lceil\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big{)}\rceil,\qquad u_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t)),$ (5.48)
$k^{\prime}$ $:=\lceil\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+C_{1}^{\prime}\big{)}\rceil.$
Then, we have the following lemma.
Lemma 5.9.
Suppose $b$ and $d$ satisfy Assumptions (N-E), (F-D), (Ma), and (C-V). Recall $R$ from (2.14) and suppose that $\lim_{i\to\infty}d(i)=d^{*}\in[0,R)$. Finally, suppose that Assumption $\mathcal{K}_{\alpha}$ is satisfied. For any $\varepsilon,A,C_{1}>0$ there exist $C_{2},t^{\prime}>0$ such that for all $t>t^{\prime}$,
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)<k\right)\geq 1-\varepsilon.$ (5.49)
Similarly, for any $\varepsilon,A>0$ there exists $C_{1}^{\prime},t^{\prime}>0$ such that for all $t>t^{\prime}$,
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)<k^{\prime}\right)\geq 1-\varepsilon.$ (5.50)
Proof.
We bound the probability of the complement of the event of interest from above by $\varepsilon$. That is, we bound
$\mathbb{P}_{\mathcal{S}}{}$ $\big{(}\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k\big{)}$ (5.51)
$=\mathbb{P}_{\mathcal{S}}\left(\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t):D^{(v)}\geq k,S_{k}^{(v)}\leq s-\mathcal{B}(v),S^{(v)}_{D^{(v)}+1}>s-\mathcal{B}(v)\right).$
We now introduce the event $\{\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq M\}$, for some large but fixed $M>0$ to be determined. We then partition the set $\mathcal{B}(u_{t},t)$ into sets of individuals born in the intervals $(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, i.e. into sets $\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, with $j\in[\lceil t-u_{t}-C_{2}\rceil]$ and apply a union bound. For individuals $v\in\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, we can bound their birth-time $\mathcal{B}(v)$ from below and above by $u_{t}+C_{2}+(j-1)$ and $u_{t}+C_{2}+j$, respectively. Similarly we can bound $s$ from below and above by $t-A$ and $t+A$, respectively. As a result, we bound the above probability from above by
$\sum_{j=1}^{\lceil t-u_{t}-C_{2}\rceil}{}$ $\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\mathrm{e}^{\lambda^{*}(u_{t}+j)}\mathbb{P}\!\left(D\geq k,S_{k}\leq t+A-(u_{t}+C_{2}+(j-1)),S_{D+1}>t-A-(u_{t}+C_{2}+j)\right)$ (5.52)
$+\mathbb{P}_{\mathcal{S}}\Big{(}\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\geq M\Big{)}.$
We now choose $M$ large enough so that, by (5.11) in Corollary 5.4, the probability on the second line is at most $\varepsilon/2$. It thus remains to bound the sum from above by $\varepsilon/2$ to arrive at the desired result. For any $j\in[\lceil t-u_{t}\rceil]$, we omit the part of the event that $S_{D+1}>t-A-(u_{t}+C_{2}+j)$, use the independence of $D$ and $S_{k}$ and use a Chernoff bound to arrive at the upper bound
$\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]\exp\big{(}\theta(t+A-(u_{t}+C_{2}+(j-1)))\big{)}.$ (5.53)
Applying the inequality $\log(1+x)\geq x-x^{2}/2$ for $x>0$, we can bound
$-\log\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]=\sum_{i=0}^{k-1}\log\Big{(}1+\frac{\theta}{b(i)+d(i)}\Big{)}\geq\theta\varphi_{1}(k)-\frac{\theta^{2}}{2}\varphi_{2}(k).$ (5.54)
Since $\varphi_{2}(k)$ is increasing and converges to $\varphi_{2}(\infty):=\lim_{k\to\infty}\varphi_{2}(k)$ as $k$ tends to infinity by Assumption (C-V), and by using Lemma 5.1, we finally arrive at the upper bound
$\mathbb{P}\!\left(D\geq k\right){}$ $\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]\exp\big{(}\theta(t+A-(u_{t}+C_{2}+(j-1)))\big{)}$ (5.55)
$\leq\exp{}$ $\Big{(}-\rho_{1}(k)-\theta\varphi_{1}(k)+\frac{\theta^{2}}{2}\varphi_{2}(\infty)+\theta(t+A-(u_{t}+C_{2}+(j-1)))\Big{)}.$
Note that $\rho_{1}$ and $\varphi_{1}$ are increasing and that $\rho_{1}(x)=d^{*}\varphi_{1}(x)+\alpha(x)$ for $x\geq 0$. Then, also using the definitions of $k$ and $u_{t}$ in (5.48), we obtain the upper bound
$\exp\Big{(}{}$ $-\frac{\lambda^{*}d^{*}}{\lambda^{*}+d^{*}}t+\frac{d^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+d^{*}C_{1}-\mathcal{K}_{\alpha}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big{)}$ (5.56)
$+\theta\big{(}C_{1}-C_{2}+A-(j-1)\big{)}+\frac{\theta^{2}}{2}\varphi_{2}(\infty)\Big{)}.$
We apply Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$ and set $\theta=(C_{2}+(j-1)-(A+C_{1}))/\varphi_{2}(\infty)$ (which is positive for all $j\geq 1$ when $C_{2}$ is large), to bound this for all $t>t^{\prime}$ by
$\exp\Big{(}-\lambda^{*}\Big{(}\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big{)}+C+d^{*}C_{1}-\frac{1}{2}\frac{(C_{2}+(j-1)-(A+C_{1}))^{2}}{\varphi_{2}(\infty)}\Big{)},$ (5.57)
where $C>0$ and $t^{\prime}$ are fixed constants, where $C$ is independent of all other constants and parameters and $t^{\prime}=t^{\prime}(C_{1},d^{*},\lambda^{*})$. By substituting this upper bound in the sum in (5.52), we can bound it from above by
$\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\sum_{j=1}^{\lceil t-u_{t}-C_{2}\rceil}\exp\Big{(}\lambda^{*}C_{2}+\lambda^{*}j+C+d^{*}C_{1}-\frac{1}{2}\frac{(C_{2}+(j-1)-(A+C_{1}))^{2}}{\varphi_{2}(\infty)}\Big{)}.$ (5.58)
We now choose $C_{2}$ large enough so that $\lambda^{*}-(C_{2}-(A+C_{1}))/\varphi_{2}(\infty)<0$. This implies that the summands are exponentially decreasing in $j$ and are maximised when $j=1$. This yields, for some constant $c=c(C_{2})>0$, the upper bound
$\frac{Mc}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\Big{(}\lambda^{*}C_{2}+\lambda^{*}+C+d^{*}C_{1}-\frac{1}{2}\frac{(C_{2}-(A+C_{1}))^{2}}{\varphi_{2}(\infty)}\Big{)}.$ (5.59)
The constant $c$ is decreasing in $C_{2}$, so that we can make this upper bound smaller than $\varepsilon/2$ by again choosing $C_{2}$ sufficiently large. Using this in (5.52) thus yields the desired result.
For the second result in Lemma 5.9, an analogous proof can be used (where, at least intuitively, one would take $C_{2}<0$ such that $|C_{2}|$ is large but fixed and $C_{1}<0$ such that $|C_{1}|$ is sufficiently large and set $C_{1}^{\prime}=-C_{1}$). We leave the details to the reader. ∎
5.3.2. There exists an old individual with large offspring
The final ingredient to prove persistence is to show that there exists an ‘old’ individual that survives until time $\approx t$ and has a sufficiently large offspring. Combined with the result in Lemma 5.9, this implies that the birth-time of the individual with the largest offspring (i.e. $I_{t}^{\mathrm{ct}}$) cannot be ‘too large’ compared to $O_{t}^{\mathrm{ct}}$. We make this precise in the following result.
Lemma 5.10.
Suppose $b$ and $d$ satisfy Assumptions (N-E), (F-D), (Ma), and (C-V). Recall $R$ from (2.14) and suppose that $\lim_{i\to\infty}d(i)=d^{*}\in[0,R)$. Finally, suppose that Assumption $\mathcal{K}_{\alpha}$ is satisfied. Recall $k=k(C_{1})$ and $u_{t}$ from (5.48). For any $\varepsilon,A>0$ there exist $C_{1},t^{\prime}>0$ such that for all $t>t^{\prime}$,
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k\right)\geq 1-\varepsilon.$ (5.60)
Proof.
We first rewrite the probability as
$\mathbb{P}_{\mathcal{S}}{}$ $\big{(}\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1}):D^{(v)}\geq k,S^{(v)}_{k}\leq s-\mathcal{B}(v),S^{(v)}_{D^{(v)}+1}>s-\mathcal{B}(v)\big{)}$ (5.61)
[…]
Here we bound $s$ and $\mathcal{B}(v)$ from below and above to obtain the inequality. Similar to the steps in (5.35) through (5.40), we bound the right-hand side from below for any $m\in\mathbb{N}$ by
[…]
$-\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\Big{(}-m\mathbb{P}\!\left(D\geq k,S_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S_{D+1}>t+A-u_{t}\right)\Big{)}.$
We then bound the probability in the exponential term from below. For any $\eta>0$ there exists $t^{\prime}>0$ such that $d(i)\leq d^{*}+\eta$ for all $i\geq k$ and all $t>t^{\prime}$ (since $k$ tends to infinity with $t$). Hence, by using Lemma 5.2,
$\mathbb{P}({}$ $D\geq k,S_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S_{D+1}>t+A-u_{t})$ (5.63)
[…]
×$\times\exp\big{(}-(d^{*}+\eta)(2A+\tfrac{1}{2})C_{1}\big{)}.$
We bound the expected value from below by restricting the range of $S_{k}$ in the indicator to $S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]$, to obtain
$\mathbb{E}\big{[}{}$ $\mathbbm{1}\{S_{k}\leq t-u_{t}-A-\tfrac{1}{2}C_{1}\}\exp\big{(}-(d^{*}+\eta)(t-u_{t}-A-\tfrac{1}{2}C_{1}-S_{k})\big{)}\big{]}$ (5.64)
[…]
We then use that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$ as $k$ tends to infinity, since $\varphi_{2}(k)={\rm Var}(S_{k})$ converges by Assumption (C-V). By the choice of $k$ (as in (5.48)) and since $b+d$ tends to infinity (since $\varphi_{2}$ converges), we have
$\varphi_{1}(k)=\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}+o(1)=t-u_{t}-C_{1}+o(1).$ (5.65)
As a result, we can write
$\mathbb{P}\!\left(S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]\right)$ $=\mathbb{P}\!\left(S_{k}-\varphi_{1}(k)\in[-A+o(1),-A+\tfrac{1}{2}C_{1}+o(1)]\right)$ (5.66)
$=\mathbb{P}\!\left(M_{\infty}\in[-A,-A+\tfrac{1}{2}C_{1}]\right)+o(1).$
We observe that the right-hand side is strictly positive for any choice of $A$ and $C_{1}>0$ and can thus be bounded from below by
$c_{\infty}(A,C_{1}):=\frac{1}{2}\mathbb{P}\!\left(M_{\infty}\in[-A,-A+\tfrac{1}{2}C_{1}]\right),$ (5.67)
for all $t$ (and thus $k$) sufficiently large. Combining this with (5.64) in (5.63) yields the lower bound
$\mathbb{P}(D\geq k,S_{k}\leq t-A-u_{t},S_{D+1}>t+A-u_{t}+\tfrac{1}{2}C_{1})\geq c_{\infty}(A,C_{1})\mathbb{P}\!\left(D\geq k\right)\mathrm{e}^{-(d^{*}+\eta)(2A+C_{1})}.$ (5.68)
Then, we use Lemma 5.1 to obtain, for some small constant $\kappa>0$, the lower bound $\mathbb{P}\!\left(D\geq k\right)\geq\kappa\exp(-\rho_{1}(k))$. As in (5.65) we can write, by the definition of $\alpha$ and $\mathcal{K}_{\alpha}$ in (2.16) and (2.17), respectively,
$\rho_{1}(k)={}$ $\rho_{1}\big{(}\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big{)}\big{)}+o(1)$ (5.69)
[…]
Using Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$, we can bound the right-hand side from above for all $t>t^{\prime\prime}$ by
$\lambda^{*}\Big{[}\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big{]}-d^{*}C_{1}+C=\lambda^{*}u_{t}-d^{*}C_{1}+C,$ (5.70)
where $C>0$ is some large constant, independent of all other constants and parameters, and $t^{\prime\prime}=t^{\prime\prime}(C_{1},d^{*},\lambda^{*})$. Hence,
$\mathbb{P}\!\left(D\geq k\right)\geq\kappa\mathrm{e}^{-\lambda^{*}u_{t}+d^{*}C_{1}-C}.$ (5.71)
To conclude, we thus obtain the lower bound,
$\mathbb{P}{}$ $(D\geq k,S_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S_{D+1}>t+A-u_{t})$ (5.72)
$\geq\kappa c_{\infty}(A,C_{1})\exp\big{(}-\lambda^{*}u_{t}-\eta C_{1}-2A(d^{*}+\eta)-C\big{)}.$
We use this in (5.62) and set, for some large constant $M>0$ to be determined,
$m:=M^{-1}\exp\big{(}\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big{)},$ (5.73)
to arrive at
$\mathbb{P}_{\mathcal{S}}(\forall{}$ $s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k)$ (5.74)
$\geq 1{}$ $-\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})|\leq M^{-1}\exp\big{(}\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big{)}\right)$
$-\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\Big{(}-\frac{\kappa c_{\infty}(A,C_{1})}{M}\exp\big{(}[\tfrac{1}{2}\lambda^{*}-\eta]C_{1}-2A(d^{*}+\eta)-C\big{)}\Big{)}.$
We bound the probability on the second line by
$\mathbb{P}_{\mathcal{S}}\big{(}{}$ $|\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})|\leq M^{-1}\exp\big{(}\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big{)}$ (5.75)
[…]
$+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t})|\exp\big{(}-\lambda^{*}u_{t}\big{)}\geq M^{-1}\exp\big{(}\tfrac{\lambda^{*}}{2}C_{1}\big{)}\right)$
[…]
We first choose $M$ sufficiently large so that, by (5.12) in Corollary 5.4, the first term on the last line is smaller than $\varepsilon/3$. Next, given this $M$, we notice that the second term on the last line is decreasing in $C_{1}$. Hence, by (5.11) in Corollary 5.4, there exists $C_{1}^{\prime}>0$ such that for all $C_{1}>C_{1}^{\prime}$, the second term is at most $\varepsilon/3$. We finally turn to the term on the last line of (5.74). Here, we note that $c_{\infty}(A,C_{1})$, as defined in (5.67), is increasing in $C_{1}$, so that
$c_{\infty}(A,C_{1})\geq c_{\infty}(A,C^{\prime}_{1})>0,$ (5.76)
for all $C_{1}>C_{1}^{\prime}$. Similarly, by choosing $\eta<\lambda^{*}/2$, the inner exponential term on the last line of (5.74) is also increasing in $C_{1}$. As a result, we can take $C_{1}>C_{1}^{\prime}$ sufficiently large so that the last line of (5.74) is at most $\varepsilon/3$. This yields the desired lower bound and concludes the proof. ∎
5.4. Persistence in the PAVD model
In this sub-section we combine the result from the previous sub-sections to prove that $(I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t})_{t\geq 0}$ is a tight sequence of random variables. We then make the correspondence in (3.22) precise to translate this into the results in Theorem 2.6.
The first part is made precise in the following result.
Proposition 5.11.
Suppose $b$ and $d$ satisfy Assumptions (N-E), (F-D), (Ma), and (C-V). Recall $R$ from (2.14) and suppose that $\lim_{i\to\infty}d(i)=d^{*}\in[0,R)$. Finally, suppose that Assumption $\mathcal{K}_{\alpha}$ is satisfied. For any $\varepsilon,A>0$ there exists $M=M(\varepsilon,A),t^{\prime}>0$ such that for all $t>t^{\prime}$,
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]:I_{s}^{\mathrm{ct}}-O^{\mathrm{ct}}_{s}<M\right)\geq 1-\varepsilon.$ (5.77)
Proof.
Fix $\varepsilon,A>0$, $t^{\prime}>0$ large (according to Proposition 5.8, Lemma 5.9, and Lemma 5.10) and recall $k=k(C_{1})$ and $u_{t}$ from (5.48). Then, for any $t>t^{\prime}$:
  1. (a)
    We take $K>0$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$: $O_{s}^{\mathrm{ct}}>u_{t}-K$.
  2. (b)
    We take $C_{1}>0$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$ $\exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}$: $d_{v}(s)\geq k$.
  3. (c)
    We take $C_{2}>\tfrac{1}{2}C_{1}$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$ $\forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}$: $d_{v}(s)<k$.
$(a)$, $(b)$, and $(c)$ follow from Proposition 5.8, Lemma 5.10, and Lemma 5.9, respectively. Then, $(b)$ and $(c)$ combined imply that at $s$ the alive individual with the largest offspring has birth-time at most $u_{t}+C_{2}$, for all $s\in[t-A,t+A]$, with probability at least $1-\tfrac{2}{3}\varepsilon$. That is, $I_{s}^{\mathrm{ct}}<u_{t}+C_{2}$ for all $s\in[t-A,t+A]$ with probability at least $1-\tfrac{2}{3}\varepsilon$. Combined with $(a)$, we thus conclude that the result follows for $M=C_{2}+K$. ∎
We finally prove Theorem 2.6.
Proof of Theorem 2.6.
We fix $\varepsilon>0$ and use (3.23) to conclude that there exist $n_{0}\in\mathbb{N}$ and $A=A(\varepsilon,n_{0})>0$ sufficiently large such that
$\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\in\big{[}\tfrac{1}{\lambda^{*}}\log n-A,\tfrac{1}{\lambda^{*}}\log n+A\big{]}\right)\geq 1-\varepsilon/3,\qquad\text{for all }n\geq n_{0}.$ (5.78)
We let $F_{n,A}:=\big{[}\frac{1}{\lambda^{*}}\log n-A,\frac{1}{\lambda^{*}}\log n+A\big{]}$ for ease of writing. Then,
$\mathbb{P}_{\mathcal{S}}{}$ $\bigg{(}\Big{|}O_{\tau_{n}}^{\mathrm{ct}}-\frac{d^{*}}{\lambda^{*}(\lambda^{*}+d^{*})}\log n-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{|}<M\bigg{)}$ (5.79)
[…]
$-\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\not\in F_{n,A}\right).$
The second term on the right-hand side is at most $\varepsilon/3$ for all $n\geq n_{0}$ by (5.78). Using Proposition 5.8 with $t=\frac{1}{\lambda^{*}}\log n$ we can take $M=M(A,\varepsilon)$ and $n_{1}=n_{1}(A,\varepsilon)\geq n_{0}$ such that the first term on the right-hand side is at least $31-\varepsilon/$ 3 for all $n\geq n_{1}$ as well, so that the entire lower bound is at least $1-\frac{2}{3}\varepsilon$. We then define
$\ell_{n}:=\frac{d^{*}}{\lambda^{*}+d^{*}}\log n+\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)},$ (5.80)
and write
$\bigg{\{}\Big{|}\log O_{n}-\frac{d^{*}}{\lambda^{*}+d^{*}}\log n-\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{|}>M^{\prime}\bigg{\}}=\big{\{}O_{n}\not\in\big{[}\mathrm{e}^{-M^{\prime}+\ell_{n}},\mathrm{e}^{M^{\prime}+\ell_{n}}\big{]}\big{\}}.$ (5.81)
Now, using that $O_{n}\overset{\mathrm{d}}{=}N(O_{\tau_{n}}^{\mathrm{ct}})$, where we recall that $(N(t))_{t\geq 0}$, as in (3.19), counts the number of births and deaths in the branching process $(\mathrm{BP}(t))_{t\geq 0}$ up to time $t$, we obtain the lower bound
$\mathbb{P}_{\mathcal{S}}{}$ $\bigg{(}\Big{|}\log O_{n}-\frac{d^{*}}{\lambda^{*}+d^{*}}\log n-\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{|}<M^{\prime}\bigg{)}$ (5.82)
[…]
$-\mathbb{P}_{\mathcal{S}}\left(N\Big{(}\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\geq\mathrm{e}^{M^{\prime}+\ell_{n}}\right).$
The first probability is at least $1-\frac{2}{3}\varepsilon$ by taking $M^{\prime}=2\lambda^{*}M$ and applying (5.79) with $M$ large. Then, since $N(t)\geq|\mathcal{B}((0,t))|$ for all $t\geq 0$ deterministically,
$\mathbb{P}_{\mathcal{S}}\left(N\Big{(}-\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(\Big{|}\mathcal{B}\Big{(}0,-\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\Big{|}\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ (5.83)
$\leq\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq\mathrm{e}^{-M^{\prime}/2}\right),$
which can be made smaller than $\varepsilon/6$ when taking $M^{\prime}=2\lambda^{*}M$ large enough by Corollary 5.4. In a similar manner, using Corollary 5.4 and that $N(t)\leq 2|\mathcal{B}(0,t)|$, we can bound the final probability on the right-hand side of (5.82) from above by $\varepsilon/6$ as well when $M$ is large enough, to finally yield the lower bound $1-\varepsilon$, which concludes the proof of (2.20).
We continue by proving that $I_{n}/O_{n}$ is tight. We apply Proposition 5.11 to conclude that there exist $n_{1}\geq n_{0}$ and $M=M(\varepsilon,A,n_{1})>0$ such that for all $n\geq n_{1}$,
$\mathbb{P}_{\mathcal{S}}\left(I^{\mathrm{ct}}_{\tau_{n}}-O^{\mathrm{ct}}_{\tau_{n}}<M\right)\geq{}$ $\mathbb{P}_{\mathcal{S}}\left(\forall s\in F_{n,A}:I^{\mathrm{ct}}_{s}-O^{\mathrm{ct}}_{s}<M\right)-\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\not\in F_{n,A}\right)\geq 1-\frac{2}{3}\varepsilon.$ (5.84)
Then, we again use that $I_{n}\overset{\mathrm{d}}{=}N(I^{\mathrm{ct}}_{\tau_{n}})$ and $O_{n}\overset{\mathrm{d}}{=}N(I^{\mathrm{ct}}_{\tau_{n}})$. As a result, for some $C,M^{\prime}>0$ to be determined,
$\mathbb{P}_{\mathcal{S}}\left(\frac{I_{n}}{O_{n}}<M^{\prime}\right)=\mathbb{P}_{\mathcal{S}}\left(\frac{N(I^{\mathrm{ct}}_{\tau_{n}})}{N(O^{\mathrm{ct}}_{\tau_{n}})}<M^{\prime}\right)\geq{}$ $\mathbb{P}_{\mathcal{S}}\left(C^{2}\exp\big{(}\lambda^{*}\big{(}I^{\mathrm{ct}}_{\tau_{n}}-O^{\mathrm{ct}}_{\tau_{n}}\big{)}\big{)}<M^{\prime}\right)$ (5.85)
$-\mathbb{P}_{\mathcal{S}}\left(N(I^{\mathrm{ct}}_{\tau_{n}})\geq C\exp\big{(}\lambda^{*}I^{\mathrm{ct}}_{\tau_{n}}\big{)}\right)$
$-\mathbb{P}_{\mathcal{S}}\left(N(O^{\mathrm{ct}}_{\tau_{n}})\leq C^{-1}\exp\big{(}\lambda^{*}O^{\mathrm{ct}}_{\tau_{n}}\big{)}\right).$
We can bound the first term on the right-hand side from below by $31-2\varepsilon/$ 3 for all $n\geq n_{0}$ using (5.84) and setting $M^{\prime}:=C^{2}\mathrm{e}^{\lambda^{*}M}$. The second term can be bounded from above by using that $N(t)\leq 2|\mathcal{B}(0,t)|$ deterministically for all $t\geq 0$, so that
$\mathbb{P}_{\mathcal{S}}\left(N(I^{\mathrm{ct}}_{\tau_{n}})\geq C\exp\big{(}\lambda^{*}I^{\mathrm{ct}}_{\tau_{n}}\big{)}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,I^{\mathrm{ct}}_{\tau_{n}})|\exp\big{(}-\lambda^{*}I^{\mathrm{ct}}_{\tau_{n}}\big{)}\geq\tfrac{C}{2}\right)$ (5.86)
$\leq\mathbb{P}_{\mathcal{S}}\left(\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\geq\tfrac{C}{2}\right).$
Applying Corollary 5.4 then yields the upper bound $\varepsilon/6$ when we choose $C$ sufficiently large. In a similar way, since $N(t)\geq|\mathcal{B}(0,t)|$ deterministically for any $t\geq 0$,
$\mathbb{P}_{\mathcal{S}}\left(N(O^{\mathrm{ct}}_{\tau_{n}})\leq C^{-1}\exp\big{(}\lambda^{*}O^{\mathrm{ct}}_{\tau_{n}}\big{)}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,O^{\mathrm{ct}}_{\tau_{n}})|\exp\big{(}-\lambda^{*}O^{\mathrm{ct}}_{\tau_{n}}\big{)}\leq C^{-1}\right)$ (5.87)
$\leq\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq C^{-1}\right).$
Again, Corollary 5.4 yields that the last line is at most $\varepsilon/6$ for $C$ large enough. Combining all three bounds in (5.85) finally yields
$\mathbb{P}_{\mathcal{S}}\left(\frac{I_{n}}{O_{n}}<M^{\prime}\right)\geq 1-\varepsilon\qquad\text{for all }n\geq n_{1}.$ (5.88)
Combined with (2.20), we directly arrive at (2.21) as well.
We conclude by proving (2.22). Here, we use that, by Proposition 3.1,
$\{\varphi_{1}(\max_{v\in\mathcal{A}_{n}}\deg_{n}(v):n\in\mathbb{N}\}\overset{\mathrm{d}}{=}\{\varphi_{1}(\max_{v\in\mathcal{A}_{\tau_{n}}^{\mathrm{ct}}}\deg_{\tau_{n}}(v)):n\in\mathbb{N}\}.$ (5.89)
We thus are required to prove the result for the quantity on the right-hand side. To this end, we write
$\mathbb{P}_{\mathcal{S}}{}$ $\bigg{(}\Big{|}\varphi_{1}(\max_{v\in\mathcal{A}_{\tau_{n}}^{\mathrm{ct}}}\deg_{\tau_{n}}(v))-\frac{1}{\lambda^{*}+d^{*}}\log n+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)}\Big{|}<K\bigg{)}$ (5.90)
[…]
$-\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\not\in F_{n,A}\right).$
The final term is at most $\varepsilon/2$ for $A$ sufficiently large by (5.78). Then, the first probability on the right-hand side is also at least $21-\varepsilon/$ 2 by setting $C_{1}^{\prime}=K$ in Lemma 5.9 and $C_{1}=K$ in Lemma 5.10 and taking $K$ sufficiently large (with $t=\frac{1}{\lambda^{*}}\log n$ and $n$ sufficiently large). We thus arrive at the desired result and conclude the proof. ∎
6. Discussion and Open Problems
In this article we have analysed the existence persistent $m$-hubs and the occurrence of persistence for the preferential attachment tree with vertex death model. In both the ‘infinite lifetime’ regime and the ‘finite lifetime’ regime, we have seen that the summability condition in (C-V) is crucial for the emergence of either persistent hubs or of persistence. Complemented by the study of lack of persistence in [23], this provides a good overview of the different regimes and the behaviour therein. Still, there are several further extensions and open problems that we would like to state here.
Existence of persistent $m$-hubs for more general CMJ branching processes. In Section 4 we studied the existence of persistent $m$-hubs in CMJ branching processes. Here, we were able to provide general conditions for the non-existence of persistent $m$-hubs, but could only prove the existence of persistent $1$-hubs (i.e. a persistent maximum degree) under general conditions. For persistent $m$-hubs with $m\geq 2$ we required the inter-birth times to be exponentially distributed, as the proofs leveraged the memoryless property at a few places. This leads to the following open problem.
Open Problem 6.1.
Can Theorem 4.4 be extended to hold for more general inter-birth random variables that satisfy Assumption 4.1?
Furthermore, the results in Theorems 4.2 and 4.4 are restricted to the setting in which
$\sum_{i=1}^{\infty}X(i)=\infty\text{ almost surely and }\widehat{\mu}(\lambda)<1\text{ for some $\lambda>0$}.$ (6.1)
The existence of persistent $1$-hubs in CMJ branching processes that do not satisfy at least one conditions does exist, see e.g. [33, 26, 25, 31], but these works restrict themselves to the case $D=\infty$ almost surely. Can such results be proved in more general settings when $\mathbb{P}\!\left(D=\infty\right)\in[0,1)$ as well?
Open Problem 6.2.
Let $(\mathrm{BP}(t))_{t\geq 0}$ be a CMJ branching process whose offspring and inter-birth times satisfy Assumption 4.1. If either $\sum_{i=1}^{\infty}X(i)<\infty$ with positive probability or $\widehat{\mu}(\lambda)=\infty$ for all $\lambda>0$, under what conditions do persistent $m$-hubs exist, or does persistence occur? Additionally, can one prove an equivalent result as in Theorems 4.2 or 4.4 without the assumption that $\widehat{\mu}(\lambda)<1$ for some $\lambda>0$?
Characterising the difference between the richest and oldest individual. In the ‘rich are old’ regime, we show that the difference between the birth-time of the richest and oldest individual in $(\mathrm{BP}(t))_{t\geq 0}$, that is, $I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}$, is a tight sequence of random variables with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$. It remains unclear if one can obtain a functional convergence of this random process.
Open Problem 6.3.
Under the conditions in Theorem 2.6, characterize the random process $(I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t})_{t\geq 0}$.
Bounded non-converging death rates. The results presented in the ‘finite lifetime’ regime do not consider the case when the death rates are bounded but do not converge. This is mainly due to the difficulty to determine the lifetime distribution of individuals in complete generality. As an example, let us consider the case of converging birth rates and alternating convergent death rates. That is, take three sequences $(c_{i})_{i\in\mathbb{N}}$, $(d_{1,i})_{i\in\mathbb{N}}$ and $(d_{2,i})_{i\in\mathbb{N}}$ such that $\lim_{i\to\infty}c_{i}=c>0$ and $\lim_{i\to\infty}d_{j,i}=d_{j}$ for $j\in\{1,2\}$, with $d_{1}\neq d_{2}\geq 0$. We set
$d=(d_{1,1},d_{2,1},d_{1,2},d_{2,2},\ldots),\qquad b(i)=c_{i}.$ (6.2)
We then claim (without proof) that the lifetime distribution satisfies
$\lim_{t\to\infty}\frac{1}{t}\log\mathbb{P}\!\left(L>t\right)=-\min\bigg{\{}c+\frac{d_{1}+d_{2}}{2}-\sqrt{c^{2}+\Big{(}\frac{d_{1}-d_{2}}{2}\Big{)}^{2}},R\bigg{\}}=:-\theta.$ (6.3)
This already non-trivial result can be obtained due to the somewhat simple choice of the birth and death rates. More general birth or death rates (e.g. $b$ that diverge to infinity) make it much harder to determine the limiting value of $\theta$ that characterises the exponential decay of the lifetime distribution.
We expect that in this more general setting, persistence can be shown to occur when $\theta<R$ and when Assumption (C-V) is satisfied. We leave this as an open problem.
Open Problem 6.4.
Suppose the sequences $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied. Assume that $\theta:=\lim_{t\to\infty}-\frac{1}{t}\log\mathbb{P}\!\left(L>t\right)$ exists and recall $R$ from (2.14). If $\theta<R$ and Assumption (C-V), does persistence occur in the sense of (P)?
Non-tree graphs. The analysis carried out here considers trees only. It would be interesting to also study evolving graphs with vertex death that are not trees, e.g. when every new vertex connects to $k>1$ many alive vertices. The continuous-time embedding into a CMJ branching process is central in the analysis in this paper and naturally restricts us to the setting of trees.
In the case of affine preferential attachment graphs without vertex death, a procedure known as collapsing branching processes is used to extend the branching process analysis to non-tree graphs as well [21]. However, this is limited to the affine case only. When including vertex death, this could only have a potential when the death rates are affine as well. Even then, this introduces difficulties due to the fact that now the lifetime of distinct vertices could have dependencies, which further complicates the analysis.
Other modelling choices. In the (embedding of the) PAVD model the size of the offspring of an individual equals the number of children said individual has produced. Similarly, the rate at which an individual produces its next child or dies depends on the total number of children produced. Instead, one could count only alive children towards the offspring, or let the birth and death rates depend on the number of alive children. Either way of defining the model can be motivated from applications. However, the analysis for these alternative choices present more challenges, since now either an individual needs to produce $n$ children out of which $n-k$ should die to have an offspring of size $k$, or the rates depend not only on the parent individual but also on the lifetimes of its children. Deijfen concludes from simulations presented in [13] that the asymptotic behaviour of the degree distribution should not change significantly, and we expect this to be the case for other properties such as persistence as well. Still, it would be interesting to further investigate these models too.
Acknowledgements
Bas Lodewijks has received funding from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. $101108569$, “DynaNet”. He also thanks Tejas Iyer for a discussion regarding the proof of Proposition 4.14.

### v2

@@
 1. Introduction
-Random graph models based on preferential attachment are quite well-understood nowadays due to a large body of research taken place over the last three decades (see [36, 37] and the references therein for a good overview of the literature).
+Random graph models based on preferential attachment are quite well-understood nowadays due to a large body of research formed over the last three decades (see [34, 35] and the references therein for a good overview of the literature).
 These models have become a staple in random graph theory due to their use in modelling real-world networks.
@@
 Diaz, Lichev, and the author study a model with uniform attachment and uniform vertex removal ($m$ is random, $p_{2}=p_{4}=0$), where the local weak limit, existence of a giant component, and the size and location of the maximum degree are studied [17].
-Work of Bellin and Kammerer et al.
+Work of Bellin, Kammerer et al.
 [5, 6, 26] studies a tree model of uniform attachment and uniform vertex removal ($m=1$, $p_{2}=p_{4}=0$), but where vertices are not removed but are ‘frozen’ and can then no longer make new connections, and where the choice to freeze vertices need not be random.
@@
 The notion of persistence in evolving random graphs has been studied from different perspectives, starting with the work of Dereich and Mörters [16].
-Later, conditions under which persistence and the lack thereof hold have been weakened or omitted by Galashin [20], Banerjee and Bhamidi [2], and Iyer [25].
+Later, conditions under which persistence and the lack thereof hold have been weakened by Galashin [20], Banerjee and Bhamidi [2], and Iyer [25].
 In short (and omitting minor technical assumptions), using the PAVD formulation with $d\equiv 0$, so that there is no vertex death and we recover the classic preferential attachment model, persistence of the maximum degree occurs almost surely if and only if
 $\sum_{i=0}^{\infty}\frac{1}{b(i)^{2}}<\infty.$ (1.3)
-Heuristically, persistence occurs when high-degree vertices have a sufficiently strong advantage over low-degree vertices to establish new connections, so they are most likely to increase their in-degree further.
+Heuristically, persistence occurs when high-degree vertices have a sufficiently strong advantage over low-degree vertices when establishing new connections, so that they are most likely to increase their in-degree further.
 This allows one vertex to attain the largest degree for all but finitely many steps, instead of the largest degree switching between vertices infinitely often.
-The transition between whether the advantage is strong enough or not lies exactly at the point when the series in (1.3) goes from finite to infinity.
+The transition between whether the advantage is strong enough or not lies exactly at the point when the series in (1.3) goes from finite to infinite.
 When vertices are killed and can no longer make new attachments afterwards, as in the PAVD model, the notion of persistence as stated above is not sufficiently general.
@@
 In the infinite lifetime regime, in which vertices can live forever with positive probability, we provide conditions, similar to (1.3), under which a persistent $m$-hub exists almost surely conditionally on the tree process $(T_{n})_{n\in\mathbb{N}}$ surviving, for each $m\in\mathbb{N}$, as well as conditions under which persistent hubs do not exist.
-This analysis generalises and expands recent work of Iyer [25], which focusses on the case $d\equiv 0$ and $m=1$ and provides a novel perspective used to analyse the (non-)existence of persistent hubs.
+This analysis generalises and expands recent work of Iyer [25], which focusses on the case $d\equiv 0$ and $m=1$.
 The second regime is the finite lifetime regime, in which vertices are killed after a finite number of steps almost surely.
@@
 We use a more precise formulation of this embedding compared to Deijfen, which allows us to simplify certain proofs of results in [13], but mainly to develop novel and precise results regarding the behaviour of the branching process embedding that were unattainable previously.
-Furthermore, we extend ideas used by Banerjee and Bhamidi in [2] to precisely analyse the optimal window in which the vertex with the largest degrees are found in the finite lifetime regime.
+Furthermore, we extend ideas used by Banerjee and Bhamidi in [2] to precisely analyse the optimal window in which the vertex with the largest degrees is found in the finite lifetime regime.
 Structure of the paper.
@@
 Section 4 studies the existence of persistent $m$-hubs in the infinite lifetime regime, whereas Section 5 focusses on the occurrence of persistence in the finite lifetime regime.
-We conclude with a discussion of our work, state several problems, and discuss possibilities for future directions in Section 6.
+We conclude with a discussion of our work, state several open problems, and discuss possibilities for future directions in Section 6.
 Notation.
-Throughout the paper we use the following notation: we let :=$\mathbb{N}:=\{1,2,\ldots\}$ denote the natural numbers, set $\mathbb{N}_{0}:=\{0,1,\ldots\}$ to include zero and let $[t]:=\{i\in\mathbb{N}:i\leq t\}$ for any $t\geq 1$.
-For $x,y\in\mathbb{R}$, we let $\lceil x\rceil:=\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor:=\sup\{n\in\mathbb{Z}:n\leq x\}$, and let $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$.
+Throughout the paper we use the following notation: we let $\mathbb{N}\coloneq\{1,2,\ldots\}$ denote the natural numbers, set $\mathbb{N}_{0}\coloneq\{0,1,\ldots\}$ to include zero and let $t]\coloneq\{i\in\mathbb{N}:i\leq t\}$ for any $t\geq 1$.
+For $x,y\in\mathbb{R}$, we let $\lceil x\rceil\coloneq\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor\coloneq\sup\{n\in\mathbb{Z}:n\leq x\}$, and let $x\wedge y\coloneq\min\{x,y\}$ and $x\vee y\coloneq\max\{x,y\}$.
 For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}$ such that $b_{n}$ is positive for all $n$ we say that $a_{n}=o(b_{n})$ and $a_{n}=\mathcal{O}(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$ and if there exists a constant $C>0$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$, respectively.
@@
 For a sequence $(\mathcal{E}_{t})_{t\in\mathbb{I}}$ of events, where $\mathbb{I}=\mathbb{N}$ or $\mathbb{I}=[0,\infty)$, we say that $\mathcal{E}_{t}$ holds with high probability when $\lim_{t\to\infty}\mathbb{P}\!\left(\mathcal{E}_{t}\right)=1$.
-Finally, for an event $\mathcal{S}$ such that $\mathbb{P}\!\left(\mathcal{S}\right)>0$ we let $\mathbb{P}_{\mathcal{S}}\left(\cdot\right):=\mathbb{P}(\cdot\,|\,\mathcal{S})$ and $\mathbb{E}_{\mathcal{S}}{}{\cdot}:=\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ denote the conditional probability measure and conditional expectation, respectively.
+Finally, for an event $\mathcal{S}$ such that $\mathbb{P}\!\left(\mathcal{S}\right)>0$ we let $\mathbb{P}_{\mathcal{S}}\left(\cdot\right)\coloneq\mathbb{P}(\cdot\,|\,\mathcal{S})$ and $\mathbb{E}_{\mathcal{S}}{}{\cdot}\coloneq\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ denote the conditional probability measure and conditional expectation, respectively.
 Then, we let $\xrightarrow{\mathbb{P}_{\mathcal{S}}-\mathrm{a.s.}}$ respectively $\overset{\mathbb{P}_{\mathcal{S}}}{\longrightarrow}$ denote almost sure convergence and convergence in probability with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$.
@@
 In this section we present the main results and the necessary assumptions that we use. Recall the PAVD model, as in Definition 1.1. We let
-$\mathcal{S}:=\bigcap_{n\in\mathbb{N}}\{\mathcal{A}_{n}\neq\varnothing\}$ (2.1)
+$\mathcal{S}\coloneq\bigcap_{n\in\mathbb{N}}\{\mathcal{A}_{n}\neq\varnothing\}$ (2.1)
 denote the event that the tree process survives, that is, the construction is never terminated.
@@
 We thus define
-$O_{n}:=\min\mathcal{A}_{n}$ (2.2)
+$O_{n}\coloneq\min\mathcal{A}_{n}$ (2.2)
 as the ‘oldest’ alive individual (i.e. with the smallest label), and recursively define
-$I_{n}^{(1)}:=\min\{i\in\mathcal{A}_{n}:\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\},$ (2.3)
+$I_{n}^{(1)}\coloneq\min\{i\in\mathcal{A}_{n}\colon\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\},$ (2.3)
 and for any fixed integer $m\geq 2$,
-$I_{n}^{(m)}:=\min\{i\in\mathcal{A}_{n}\setminus\{I_{n}^{(1)},\ldots,I_{n}^{(m-1)}\}:\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\setminus\{I_{n}^{(1)},\ldots,I_{n}^{(m-1)}\}\}.$ (2.4)
+$I_{n}^{(m)}\coloneq\min\{i\in\mathcal{A}_{n}\setminus\{I_{n}^{(i)}\}_{i\in[m-1]}\colon\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\setminus\{I_{n}^{(i)}\}_{i\in[m-1]}\}.$ (2.4)
 We can interpret $I_{n}^{(m)}$ as the oldest alive individual with the $m^{\mathrm{th}}$ largest degree in $T_{n}$ and we call $I_{n}^{(m)}$ the $m$-hub of $T_{n}$.
@@
 We are interested in the long-term behaviour of $I_{n}^{(m)}$ and $O_{n}$.
-First, we say that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent elder $\mathbb{P}_{\mathcal{S}}$-almost surely when there exists a $\mathbb{P}_{\mathcal{S}}$-almost surely finite random variable $O$ such that
+Since $O_{n}$ is non-decreasing in $n$, it follows that $O_{n}$ converges $\mathbb{P}_{\mathcal{S}}$-almost surely to a limit $O$.
+We say that a persistent elder exists when
+$O\coloneq\lim_{n\to\infty}O_{n}\text{ is finite }\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (PE)
+When $O<\infty$ holds, we can think of $O$ as the first vertex that is added to the tree that will never be killed.
+It is clear that this limit exists if and only if vertices can live forever with positive probability (i.e.
+by never being both selected and killed).
+Otherwise, $O=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
+In a similar manner, we say that a persistent $m$-hub exists when there exists a $\mathbb{P}_{\mathcal{S}}$-almost surely finite random variable $I^{(m)}$ such that
+$\lim_{n\to\infty}I_{n}^{(m)}=I^{(m)}\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (PH-$m$)
+Intuitively, the existence of the limit in (PH-$m$) implies that the vertex with the $m^{\mathrm{th}}$ largest degree eventually stabilizes, in the sense that vertex $I^{(m)}$ attains the $m^{\mathrm{th}}$ largest degree in $T_{n}$ for all but finitely many $n$.
+By noting that that $I_{n}^{(m)}\geq O_{n}$ for all $m,n\in\mathbb{N}$, it follows that the limit $I^{(m)}$ in (PH-$m$) cannot exist when the limit $O$ in (PE) equals infinity.
+That is, if no persistent elder exists, then also no persistent $m$-hub exists.
+In this case, we instead focus on the growth rate of the variables $I_{n}^{(m)}$ and $O_{n}$.
+We say that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
 […]
-We say that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, when there exists a random variable $I^{(m)}$ such that
-[…]
-We say that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
-$\Big{(}\frac{I_{n}^{(1)}}{O_{n}}\Big{)}_{n\in\mathbb{N}}\text{ is a tight sequence of random variables with respect to $\mathbb{P}_{\mathcal{S}}$},$ (P)
 and that lack of persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
 $\frac{I_{n}^{(1)}}{O_{n}}\xrightarrow{\mathbb{P}_{\mathcal{S}}}\infty.$ (LP)
-Clearly, when the $1$-hub persists as in (PH-$m$) with $m=1$, then the existence of a persistent elder as in (PE) and persistence as in (P) are implied, since $1\leq O_{n}\leq I_{n}^{(1)}$ and $O_{n}$ is non-decreasing.
-When $m$-hubs do not persist for any $m\in\mathbb{N}$, the vertex that attains the $m^{\mathrm{th}}$ largest degree changes infinitely often.
-Persistence (and lack thereof) quantifies the rate of change of the maximum-degree vertex.
-Persistence tells us that the rate of change of $I_{n}^{(1)}$ is of the same rate of $O_{n}$, so that the maximum-degree vertex has a label (or ‘age’) of the same order as $O_{n}$, whereas lack of persistence tells us that the rate of change of $I_{n}^{(1)}$ is faster than that of $O_{n}$.
+We intuitively think of (P) as the vertex with the largest degree being ‘close’ to the oldest alive vertex, in terms of their label, whereas (LP) implies that they are ‘far apart‘.
+Clearly, when a persistent $1$-hub exists, as in (PH-$m$) with $m=1$, then the existence of a persistent elder, as in (PE), and the existence of a persistent $1$-hub, as in (P), are implied, since $1\leq O_{n}\leq I_{n}^{(1)}$ and $O_{n}$ is non-decreasing.
+Similarly, lack of persistence, as in (LP), implies that no persistent $1$-hub exists.
 In this paper, we study:
   • A class of PAVD models for which we prove when it contains a persistent $m$-hub for any $m\in\mathbb{N}$. (Theorem 2.1)
-  • A class of PAVD models for which we prove that it does not contain an $m$-persistent hub for any $m\in\mathbb{N}$, but persistence does occur. (Theorem 2.6)
-This complements recent work of the author and Heydenreich in [23], where a class of PAVD models is studied for which it is shown when lack of persistence occurs, i.e.
-when nor persistent $m$-hubs exist nor persistence occurs.
-Before we state our main results, we define the following quantities.
+  • A class of PAVD models for which we prove that it does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, but persistence does occur. (Theorem 2.6)
+To be able to state our main results, we need to define the following quantities.
 Let $(E_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent exponential random variables, where $E_{i}$ has rate $b(i)+d(i)$, and let $(B_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent Bernoulli random variables (also independent of the $E_{i}$), such that
 $\mathbb{P}\!\left(B_{i}=1\right)=\frac{b(i)}{b(i)+d(i)},\qquad\text{for }i\in\mathbb{N}_{0}.$ (2.5)
-Finally, define
-$S_{k}:=\sum_{i=0}^{k-1}E_{i},\qquad\ D:=\inf\{i\in\mathbb{N}_{0}:B_{i}=0\},\qquad\text{and}\qquad\mathcal{R}:=\sum_{j=1}^{D}\delta_{S_{j}},$ (2.6)
+Then, define
+$S_{k}\coloneq\sum_{i=0}^{k-1}E_{i},\qquad\ D\coloneq\inf\{i\in\mathbb{N}_{0}:B_{i}=0\},\qquad\text{and}\qquad\mathcal{R}\coloneq\sum_{j=1}^{D}\delta_{S_{j}},$ (2.6)
 where $\delta$ is a Dirac measure. Note that $\mathcal{R}$ is an empty point process when $D=0$. We let $\mu$ denote the density of the point process $\mathcal{R}$, i.e.
-$\mu(t):=\lim_{\varepsilon\downarrow 0}\varepsilon^{-1}\mathbb{P}\!\left(\mathcal{R}((t,t+\varepsilon))\geq 1\right),\qquad t\geq 0.$ (2.7)
+$\mu(t)\coloneq\lim_{\varepsilon\downarrow 0}\varepsilon^{-1}\mathbb{P}\!\left(\mathcal{R}((t,t+\varepsilon))\geq 1\right),\qquad t\geq 0.$ (2.7)
 Also, let $\widehat{\mu}(\lambda)$ denote the Laplace transform of $\mu$, for $\lambda\geq 0$. That is,
-$\widehat{\mu}(\lambda):=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\sum_{k=1}^{\infty}\prod_{i=0}^{k-1}\frac{b(i)}{b(i)+d(i)+\lambda}.$ (2.8)
-The explicit expression for $\widehat{\mu}$ in terms of the sequences $b$ and $d$ follows from [13, Proposition $1.1$]. Then, we define for $k\in\mathbb{N}$ the sequences
-$\varphi_{1}(k):=\sum_{i=0}^{k-1}\frac{1}{b(i)+d(i)},\qquad\varphi_{2}(k):=\sum_{i=0}^{k-1}\Big{(}\frac{1}{b(i)+d(i)}\Big{)}^{2},\qquad\rho_{1}(k):=\sum_{i=0}^{k-1}\frac{d(i)}{b(i)+d(i)}.$ (2.9)
+$\widehat{\mu}(\lambda)\coloneq\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\mathbb{E}\left[\sum_{k=1}^{D}\exp\bigg(-\lambda\sum_{i=0}^{k-1}E_{i}\bigg)\right]=\sum_{k=1}^{\infty}\prod_{i=0}^{k-1}\frac{b(i)}{b(i)+d(i)+\lambda}.$ (2.8)
+The explicit expression for $\widehat{\mu}$ in terms of the sequences $b$ and $d$ follows from [13, Proposition $1.1$]. Finally, we define for $k\in\mathbb{N}$ the sequences
+$\varphi_{1}(k)\coloneq\sum_{i=0}^{k-1}\frac{1}{b(i)+d(i)},\quad\varphi_{2}(k)\coloneq\sum_{i=0}^{k-1}\Big(\frac{1}{b(i)+d(i)}\Big)^{2},\quad\rho_{1}(k)\coloneq\sum_{i=0}^{k-1}\frac{d(i)}{b(i)+d(i)}.$ (2.9)
 We split the presentation of the results into two parts, based on the infinite lifetime regime and the finite lifetime regime.
@@
 Note that preferential attachment without death is a special case of the former regime, where all individuals survive forever deterministically.
-In the ‘infinite lifetime’ regime we provide criteria under which persistent $m$-hubs do or do not exist, in the sense of (PH-$m$).
-In the ‘finite lifetime’ regime we provide criteria under which persistence occurs, in the sense of (P).
-These results complement the work of the author and Heydenreich in [23] regarding lack of persistence for the PAVD model, in the sense of (LP).
-The proof techniques for the two regimes are entirely distinct.
+The proof techniques for the results related to each regime are entirely distinct.
 In the ‘finite lifetime’ regime we use a similar methodology as used in [23], though we now focus on the occurrence of persistence rather than lack of persistence, which requires a finer analysis at times.
 In the ‘infinite lifetime’ regime we expand and generalise ideas developed by Iyer in [25].
-In fact, as we shall see in Section 4, the results in this regime can be proved to hold for more general family of models.
 Infinite lifetime regime.
-We first present our result for the model when vertices can survive forever with positive probability.
+We first present our result for the PAVD model when vertices can survive forever with positive probability.
 To this end, we introduce the following assumptions.
@@
 We assume that
-There exists λ>0 such that $\text{There exists $\lambda>0$ such that }\widehat{\mu}(\lambda)<\infty.$ (F-L)
+$\text{There exists $\lambda>0$ such that }\widehat{\mu}(\lambda)<\infty.$ (F-L)
 Furthermore, we have the assumptions
@@
 Throughout the paper we assume that Assumption (F-L) “Finite-Laplace” holds.
-In fact, Assumption (F-L) combined with the continuity of the mapping $\lambda\mapsto\widehat{\mu}(\lambda)$ implies Assumption (N-E) “Non-Explosion”.
-It guarantees that the tree process is ‘not exploding’, in the sense that there does not exists a unique vertex that obtains an infinite degree, nor a unique path from the root of infinite length.
-Assumption (C-V) “Converging-Variance” implies that the variance of $S_{k}$ converges with $k$, and is a crucial assumption under which we prove that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub or that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$.
+It ensures that the branching process used in the analysis does not grow infinitely large in finite time.
+Assumption (N-E) “Non-Explosion” is of a similar nature and implies that there does not exist a unique vertex in the tree that obtains an infinite degree (as $n\to\infty$).
+Assumption (C-V) “Converging-Variance” implies that the variance of $S_{k}$ converges with $k$, and is a crucial assumption under which we prove that a persistent $m$-hub exists (as in (PH-$m$)) or that persistence occurs (as in (P)).
 Finally, Assumption (F-D) “Finite-Degree” implies that the random variable $D$, as in (2.6), is finite almost surely (see Lemma 5.1) and governs whether we are in the ‘infinite lifetime’ regime or ‘finite lifetime’ regime.
 We hence may or may not assume that (F-D) is satisfied.
-We then have the following result, where we show under which conditions the tree process $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub or not.
+Deijfen [13] shows that Assumption (F-L) implies that either Assumption (N-E) or Assumption (F-D) must be satisfied.
+We do not include an analysis of the case in which Assumption (N-E) is not satisfied but Assumption (F-D) is, which is on-going work.
+We then have the following result, where we show under which conditions a persistent $m$-hub does or does not exist.
 Theorem 2.1 (Infinite lifetime regime).
@@
 $O_{n}\longrightarrow O\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (2.10)
-In particular, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent elder in the sense of (PE) ((((i.e.
-$O<\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely)))) if and only if Assumption (F-D) is not satisfied.
-Now, suppose Assumption (N-E) is satisfied and Assumptions (F-D) and (C-V) are not satisfied.
-Then, for any $m\in\mathbb{N}$, $(T_{n})_{n\in\mathbb{N}}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely.
-If, instead, Assumption (C-V) is satisfied and, additionally, Assumption (F-L) is satisfied, then for any $m\in\mathbb{N}$ there exists an almost surely finite random variable $I^{(m)}$ such that
+In particular, a persistent elder exists, in the sense of (PE), if and only if Assumption (F-D) is not satisfied.
+Suppose Assumption (N-E) is satisfied and Assumptions (F-D) and (C-V) are not satisfied. Then, for any $m\in\mathbb{N}$, a persistent $m$-hub does not exist $\mathbb{P}_{\mathcal{S}}$-almost surely.
+If Assumptions (N-E), (C-V), and (F-L) are satisfied and Assumption (F-D) is not, then for any $m\in\mathbb{N}$ there exists an almost surely finite random variable $I^{(m)}$ such that
 $I_{n}^{(m)}\longrightarrow I^{(m)}\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (2.11)
-That is, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub in the sense of (PH-$m$) $\mathbb{P}_{\mathcal{S}}$-almost surely.
-Moreover, for any $m\geq 2$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
+That is, a persistent $m$-hub exists $\mathbb{P}_{\mathcal{S}}$-almost surely, in the sense of (PH-$m$). Moreover, for any $m\geq 2$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
 $\deg_{n}(I_{n}^{(1)})>\deg_{n}(I_{n}^{(2)})>\cdots>\deg_{n}(I_{n}^{(m)})\quad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s}.$ (2.12)
@@
 irrespective of $d$.
-We additionally need that $d$ is such that Assumption (F-D) is not satisfied to be in the infinite lifetime regime, so that either $d$ converges to zero sufficiently fast or is non-zero at a sufficiently sparse sub-sequence.
+We additionally need that $d$ is such that Assumption (F-D) is not satisfied to be in the infinite lifetime regime, so that either $d$ converges to zero sufficiently fast or is non-zero only at a sufficiently sparse sub-sequence.
 […]
@@
 No persistent $m$-hub for any $m\in\mathbb{N}$ (N-E), not (F-D), and not (C-V)
-Persistent $m$-hub and distinct degrees ((2.11) & (2.12)) (F-L) and (C-V)
+Persistent $m$-hub for any $m\in\mathbb{N}$ ((2.11) & (2.12)) (F-L), (N-E), not (F-D), and (C-V)
 Table 1. A summary of the results in Theorem 2.1 and the required assumptions.
@@
 We now present our result for the PAVD model when vertices die after a finite number of steps almost surely.
-To this end, we need some additional definitions and assumptions.
-First, we define the quantity
-$R:=\inf_{i\in\mathbb{N}_{0}}(b(i)+d(i)),$ (2.14)
+The results in Theorem 2.1 are qualitative, as they state whether certain random variables converge or diverge.
+Our result in finite lifetime, on the other hand, is quantitative.
+It shows that certain random variables grow at the same rate (in the sense of (P)).
+To this end, we need some additional definitions and assumptions to more finely control the behaviour of the model.
+First, we define
+$R\coloneq\inf_{i\in\mathbb{N}_{0}}(b(i)+d(i)),$ (2.14)
 and assume that there exists $d^{*}\in[0,R)$, such that
@@
 The recent work of Heydenreich and the author [23] studies lack of persistence in these regime.
-They show that, under mild conditions, persistence never occurs in the ‘rich die young’ regime, see [23, Theorem $2.10$].
-Our focus here thus lies on the ‘rich are old’ regime, where we show under which conditions persistence does occur.
+They show that, under mild conditions, persistence never occurs in the ‘rich die young’ regime, see [23, Theorem $2.10$], and that lack of persistence occurs in the ‘rich are old’ regime when Assumption (C-V) is not satisfied.
+Our focus here thus lies on the ‘rich are old’ regime, where we show that persistence does occur when Assumption (C-V) is satisfied (and additional technical assumptions).
 Infinite lifetime Finite lifetime
@@
 We also define
-$\alpha(k):=\rho_{1}(k)-d^{*}\varphi_{1}(k)=\sum_{i=0}^{k-1}\frac{d(i)-d^{*}}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}.$ (2.16)
+$\alpha(k)\coloneq\rho_{1}(k)-d^{*}\varphi_{1}(k)=\sum_{i=0}^{k-1}\frac{d(i)-d^{*}}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}.$ (2.16)
 Note that $\alpha\equiv\rho_{1}$ when $d^{*}=0$, that is, when $d$ converges to zero.
@@
 We then define
-$\mathcal{K}_{\alpha}(t):=\alpha\big{(}\varphi_{1}^{-1}(t)\big{)}\qquad\text{for }t\geq 0.$ (2.17)
+$\mathcal{K}_{\alpha}(t)\coloneq\alpha\big(\varphi_{1}^{-1}(t)\big)\qquad\text{for }t\geq 0.$ (2.17)
 Heuristically, $\mathcal{K}_{\alpha}(t)$ quantifies the difference between $\rho_{1}(\varphi_{1}^{-1}(t))$ and $d^{*}t$, and allows us to precisely quantify probabilities of the form $\mathbb{P}\!\left(D\geq\varphi_{1}^{-1}(t)\right)$ (see Lemma 5.1).
 We then recall $\widehat{\mu}$ from (2.8).
-Assumption (F-L) is not sufficient in the ‘finite lifetime’ regime.
+Assumption (F-L) is not sufficiently strong for the proof techniques used in the ‘finite lifetime’ regime.
 Rather, we introduce
-$\underline{\lambda}:=\inf\{\lambda>0:\widehat{\mu}(\lambda)<\infty\},$ (2.18)
+$\underline{\lambda}\coloneq\inf\{\lambda>0:\widehat{\mu}(\lambda)<\infty\},$ (2.18)
 and assume that
@@
 Finally, we have the following technical assumption for the function $\mathcal{K}_{\alpha}$, as in (2.17).
-Assumption $\bm{\mathcal{K}_{\alpha}}$. For $\lambda^{*}>0,d^{*}\geq 0$, there exists a function $r\colon(0,\infty)\to\mathbb{R}_{+}$ such that
+Assumption $\bm{\mathcal{K}_{\alpha}}$.
+For $\lambda^{*}>0,d^{*}\geq 0$, there exists a function $r\colon(0,\infty)\to\mathbb{R}_{+}$ such that
 $\mathcal{K}_{\alpha}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\big)-\mathcal{K}_{\alpha}(r(t))=\mathcal{O}(1).$ (2.19)
@@
 Discussion, Examples, and Related work.
-The results presented in this section identify the behaviour of the PAVD model in the ‘infinite lifetime’ and ‘finite lifetime’ with regards to persistence of the maximum degree (or $m^{\mathrm{th}}$-largest degree).
+The results presented in this section identify the behaviour of the PAVD model in the ‘infinite lifetime’ and ‘finite lifetime’ with regards to persistence of the maximum degree (or $m^{\mathrm{th}}$-largest degree, in general).
 In the former regime, we require fairly mild conditions only.
-In fact, we can prove some of the results in a more general set-up as well, see Section 4.
 In the ‘finite lifetime’ regime, stronger assumptions are required compared to the ‘infinite lifetime’ regime.
-This is due to the fact that we need very precise asymptotic expansions of the random variables $I_{n}$ and $O_{n}$, which in this regime no longer converge without rescaling.
+This is due to the fact that we need very precise asymptotic expansions of the random variables $I_{n}^{(1)}$ and $O_{n}$, which in this regime no longer converge without rescaling.
 In particular, we require control over the growth rate of a continuous-time branching process that we analyse to prove the results in Theorem 2.6 (see Section 3 for more details).
@@
 For example, consider the three models
-[…]
-[…]
-[…]
+$({}$ $\text{RaO})$ $b$ $=(1,2,3,\ldots)$ $\text{and }d=(1,2,\tfrac{3}{2},\tfrac{3}{2},\ldots),$ (2.23)
+$({}$ $\text{RdY}1)$ $b$ $=(1,2,3,\ldots)$ $\text{and }d=(\tfrac{1}{4},2,\tfrac{3}{2},\tfrac{3}{2},\ldots),$
+$({}$ $\text{RdY}2)$ $b$ $=(\tfrac{1}{4},2,3,\ldots)$ $\text{and }d=(1,2,\tfrac{3}{2},\tfrac{3}{2},\ldots).$
 In the first example we have $R=2$ and $d^{*}=\frac{3}{2}$, so that this case belongs to the ‘rich are old’ regime.
@@
 First, we discuss Assumption $\mathcal{K}_{\alpha}$.
-Here, with $c\in(0,\lambda^{*}/(\lambda*+d^{*}))$, set
-$\varepsilon_{t}:=\sup_{i\geq ct}|d(i)-d^{*}|.$ (2.24)
+Here, with $c\in(0,\lambda^{*}/(\lambda^{*}+d^{*}))$, set
+$\varepsilon_{t}\coloneq\sup_{i\geq ct}|d(i)-d^{*}|.$ (2.24)
 Assumption $\mathcal{K}_{\alpha}$ is then satisfied if there exists $k\in\mathbb{N}$ such that
@@
 Namely, using Remark 2.5, set
-$x_{1}(t):=\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t\big{)}\qquad\text{and}\qquad x_{i}(t):=\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x_{i-1}(t)\big{)}\qquad\text{for }i>1.$ (2.26)
-Then, take $x(t):=x_{k}(t)$ and $r(t):=\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x(t)$.
-It can be shown, using (the proof of) Lemma 5.5 $(c)$ that this choice of $r$ satisfies Assumption $\mathcal{K}_{\alpha}$.
-We observe that $\mathcal{K}_{\alpha}(t)=o(t)$ by Lemma 5.5 $(c)$, so that $\varepsilon_{t}=t^{-\beta}$ for some $\beta>0$ satisfies (2.25) for any $k>\beta^{-1}$.
+$x_{1}(t)\coloneq\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t\big)\qquad\text{and}\qquad x_{i}(t)\coloneq\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x_{i-1}(t)\big)\qquad\text{for }i>1.$ (2.26)
+Then, take $x(t)\coloneq x_{k}(t)$ and $r(t)\coloneq\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x(t)$.
+It can be shown, using (the proof of) Lemma 5.5(c)(c) that this choice of $r$ satisfies Assumption $\mathcal{K}_{\alpha}$.
+We observe that $\mathcal{K}_{\alpha}(t)=o(t)$ by Lemma 5.5(c)(c), so that $\varepsilon_{t}=t^{-\beta}$ for some $\beta>0$ satisfies (2.25) for any $k>\beta^{-1}$.
 That is, as long as $d$ converges to $d^{*}$ at a polynomial rate, Assumption $\mathcal{K}_{\alpha}$ is satisfied.
@@
   • $\mathbb{E}\left[D\right]\in(1,\infty)$, or
-  • $\mathbb{E}\left[D\right]=\infty$, $b$ tends to infinity, $C_{1}:=\limsup_{i\to\infty}\frac{b(i)}{i}<\infty$, and $\widehat{\mu}(C_{1})>1$.
+  • $\mathbb{E}\left[D\right]=\infty$, $b$ tends to infinity, $C_{1}\coloneq\limsup_{i\to\infty}\frac{b(i)}{i}<\infty$, and $\widehat{\mu}(C_{1})>1$.
 Finally, Assumption (F-L) is implied by Assumption (Ma). It thus requires fewer conditions on the sequences $b$ and $d$. Indeed, it is satisfied when
@@
 The results on persistent $m$-hubs with $m>1$, as well as the result in (2.12) are novel and also generalise earlier work of Banerjee and Bhamidi for preferential attachment trees [2, Theorem $4.9$ and Remark $4.11$].
-Recent work of Heydenreich and the author in [23] also studies the PAVD, but focusses on lack of persistence, in the sense of (LP).
+Recent work of Heydenreich and the author in [23] also studies the PAVD model, but focusses on lack of persistence, in the sense of (LP).
 This complements the results presented here and combined these two works provide a clear understanding of the different regimes that can be observed, see Table 2.
@@
 3. Methodology: Embedding in CMJ branching process
-To prove the results presented in Section 2, we make use of a technique where one ‘embeds’ a random discrete process into a continuous-time branching process, a Crump-Mode-Jagers (CMJ) branching process to be exact.
+To prove the results presented in Section 2, we make use of a technique where one ‘embeds’ a random discrete process into a Crump–Mode-Jagers (CMJ) continuous-time branching process.
 The random discrete structure is then equal in distribution to the CMJ branching process, when viewed at certain stopping times, whilst the CMJ branching process in continuous time provides more analytical advantages compared to the discrete process.
-General CMJ branching processes. We first introduce some definitions and notation. We let $\mathcal{U}_{\infty}$ denote the Ulam-Harris tree. That is,
-$\mathcal{U}_{\infty}:=\{\varnothing\}\cup\bigcup_{k=1}^{\infty}\mathbb{N}^{k}.$ (3.1)
+General CMJ branching processes.
+We first introduce some definitions and notation to facilitate the introduction of the CMJ branching process that relates to the discrete PAVD model.
+We let $\mathcal{U}_{\infty}$ denote the Ulam-Harris tree.
+That is,
+$\mathcal{U}_{\infty}\coloneq\{\varnothing\}\cup\bigcup_{k=1}^{\infty}\mathbb{N}^{k}.$ (3.1)
 For $v=v_{1}v_{2}\cdots v_{k}\in\mathcal{U}_{\infty}$, we view $v$ as the $v_{k}^{\mathrm{th}}$ child of $v_{1}\cdots v_{k-1}$ (where $v_{1}\cdots v_{k-1}$ denotes $\varnothing$ when $k=1$).
@@
 Given the inter-birth times of an individual $v$, we define its lifetime as
-$L^{(v)}:=\sum_{i=1}^{D^{(v)}+1}X(vi).$ (3.2)
-Furthermore, we can recursively construct all birth times $\mathcal{B}(v)$ by setting $\mathcal{B}(\varnothing):=0$ and
-$\mathcal{B}(vj):=\begin{cases}\mathcal{B}(v)+\sum_{i=1}^{j}X(vi)&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{1,\ldots,D^{(v)}\},\\ \infty&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{D^{(v)}+1,\ldots\}.\end{cases}$ (3.3)
-That is, given that $v$ is born in finite time, each child $v1,\ldots,vD^{(v)}$ is born finite time and children $vj$ for $j>D^{(v)}$ are never born.
+$L^{(v)}\coloneq\sum_{i=1}^{D^{(v)}+1}X(vi).$ (3.2)
+Furthermore, we can recursively construct all birth times $\mathcal{B}(v)$ by setting $\mathcal{B}(\varnothing)\coloneq 0$ and
+$\mathcal{B}(vj)\coloneq\begin{cases}\mathcal{B}(v)+\sum_{i=1}^{j}X(vi)&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{1,\ldots,D^{(v)}\},\\ \infty&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{D^{(v)}+1,\ldots\}.\end{cases}$ (3.3)
+Given that $v$ is born in finite time, each child $v1,\ldots,vD^{(v)}$ is born finite time and children $vj$ for $j>D^{(v)}$ are never born (as their birth-time is infinite).
 The individual $v$ then dies after $L^{(v)}$ time.
@@
 $\mathcal{B}(v)=\infty$) then it is clear that all descendants of $v$ are never born either.
-We can then construct the branching process $(\mathrm{BP}(t))_{t\geq 0}$ by setting
-$\mathrm{BP}(t):=\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t\}\qquad\text{for }t\geq 0.$ (3.4)
-In words, $\mathrm{BP}(t)$ contains all individuals that are born by time $t$.
+We can then construct a branching process $(\mathrm{BP}(t))_{t\geq 0}$ by setting
+$\mathrm{BP}(t)\coloneq\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t\}\qquad\text{for }t\geq 0.$ (3.4)
 It is clear that individuals $v=v_{1}\cdots v_{m}$ with $v_{1},\ldots,v_{m}\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ are born in the process $\mathrm{BP}$ if and only if $v_{i}\leq D^{(v_{1}\cdots v_{i-1})}$ for all $i\in[m]$.
@@
 We further define
-$\mathcal{A}^{\mathrm{ct}}_{t}:=\{u\in\mathcal{U}_{\infty}:\sigma_{u}\leq t,\sigma_{u}+L^{(u)}>t\},\qquad t\geq 0,$ (3.5)
+$\mathcal{A}^{\mathrm{ct}}_{t}\coloneq\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t<\mathcal{B}(u)+L^{(u)}\},\qquad\text{ for }t\geq 0,$ (3.5)
 as the set of alive individuals at time $t$. We can view this as the continuous-time counterpart of $\mathcal{A}_{n}$, as defined in Definition 1.1. We abuse notation to write
@@
 as the event that the branching process $\mathrm{BP}$ survives.
-We can view the point process $\mathcal{R}$ in (2.6), its density $\mu$ in (2.7), and the Laplace transform $\widehat{\mu}$ of $\mu$ in (2.8) from this more general set-up as well.
-That is, here we set
-$\mathcal{R}:=\sum_{k=1}^{D}\delta_{\sum_{i=1}^{k}X(i)}.$ (3.7)
-With $\mu(t)\mathrm{d}t=\mathbb{E}\left[\mathcal{R}(t)\mathrm{d}t\right]$ and using Fubini’s theorem, the Laplace transform $\widehat{\mu}$ then equals
-$\widehat{\mu}(\lambda):=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\lambda\sum_{i=1}^{j}X(i)\bigg{)}\right].$ (3.8)
-Persistence in BP. We can view the concepts of persistent $m$-hubs, and persistence and the lack thereof from a continuous-time perspective as well. To this end, we define
-$\deg_{t}(u):=\min\bigg{\{}\sup\bigg{\{}k\in\mathbb{N}:\sum_{i=1}^{k}X(ui)\leq t-\mathcal{B}(u)\bigg{\}},D^{(u)}\bigg{\}},\qquad u\in\mathcal{U}_{\infty},t\geq 0,$ (3.9)
-as the number of children $u$ has produced by time $t$. We use the convention that we set $\deg_{t}(u)=-\infty$ when $t<\mathcal{B}(u)$, i.e. when $u$ has not (yet) been born. We then define
-$O_{t}^{\mathrm{ct}}:=\!\!\min_{v\in\mathcal{A}_{t}^{\mathrm{ct}}}\!\!\mathcal{B}(v)$ (3.10)
-as the birth-time of the oldest alive individual at time $t$ as the continuous-time counterpart of $O_{n}$, as in (2.2). Similarly, we define
-$I_{t}^{(1),\mathrm{ct}}:=\min\big{\{}\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\big{\}},$ (3.11)
-and for $m\geq 2$ fixed,
-$I_{t}^{(m),\mathrm{ct}}:=\min\big{\{}\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]}\big{\}}.$ (3.12)
-Here $I_{t}^{(m),\mathrm{ct}}$ denotes the birth-time of the oldest alive individual that has the $m^{\mathrm{th}}$ largest number of children at time $t$ and is the continuous-time counterpart of $I_{n}^{(m)}$.
-For a branching process $(\mathrm{BP}(t))_{t\geq 0}$ we say that it contains a persistent $m$-hub when there exists a random variable $I^{(m),\mathrm{ct}}$ such that
-[…]
-When $(\mathrm{BP}(t))_{\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, we say that persistence respectively lack of persistence occurs in $(\mathrm{BP}(t))_{t\geq 0}$ when
-$(I^{(1),\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t})_{t\geq 0}\text{ is a tight sequence of random variables with respect to $\mathbb{P}_{\mathcal{S}}$,}$ (3.14)
-or
-$I^{(1),\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}\xrightarrow{\mathbb{P}_{\mathcal{S}}}\infty.$ (3.15)
-Again, persistence in $(\mathrm{BP}(t))_{t\geq 0}$ is implied when $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $1$-hub, whereas lack of persistence implies that $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$.
-The continuous-time preferential attachment with vertex death branching process.
+Continuous-time embedding of PAVD.
 We now provide a specific case for the general definition above that we coin the continuous-time preferential attachment with vertex death branching process (CTPAVD), which is directly related to the PAVD model.
@@
 For each individual $v\in\mathcal{U}_{\infty}$, we let $D^{(v)}$ be an i.i.d.
-copy of $D$, an let each of its inter-birth times $X(vi)$ be an i.i.d.
+copy of $D$, and set $X(vi)=E_{i-1}^{(v)}$, where $E^{(v)}_{i-1}$ is an i.i.d.
 copy of $E_{i-1}$ for each $i\in\mathbb{N}$.
-It follows that, for each $k\in\mathbb{N}$,
-$\sum_{i=1}^{k}X(vi)\overset{\mathrm{d}}{=}S_{k}\qquad\text{and, in particular,}\qquad L^{(v)}\overset{\mathrm{d}}{=}S_{D+1}.$ (3.16)
-Continuous-time embedding of PAVD into CTPAVD.
-Let us define the following quantities. First, we introduce the characteristics
-$\chi_{\mathrm{b}}(t):=\mathbbm{1}_{\{t\geq 0\}},\qquad\text{and}\qquad\chi_{\mathrm{a}}(t):=\mathbbm{1}_{\{0\leq t<L\}},$ (3.17)
+With $S_{k}^{(v)}\coloneq\sum_{i=0}^{k-1}E^{(v)}_{i}$, we can write for each $v\in\mathcal{U}_{\infty}$ and $k\in\mathbb{N}$,
+$\sum_{i=1}^{k}X(vi)=S_{k}^{(v)}\qquad\text{and, in particular,}\qquad L^{(v)}\overset{\mathrm{d}}{=}S_{D+1}^{(v)}.$ (3.7)
+Let us introduce for $v\in\mathcal{U}_{\infty}$ the characteristics
+$\chi^{(v)}_{\mathrm{b}}(t)\coloneq\mathbbm{1}_{\{t\geq 0\}},\qquad\text{and}\qquad\chi^{(v)}_{\mathrm{a}}(t)\coloneq\mathbbm{1}_{\{0\leq t<L^{(v)}\}},$ (3.8)
 and define, for $\square\in\{\mathrm{a},\mathrm{b}\}$,
-$Z_{\square}(t):=\sum_{u\in\mathcal{U}_{\infty}}\chi_{\square}(t-\mathcal{B}(u)).$ (3.18)
+$Z_{\square}(t)\coloneq\sum_{v\in\mathcal{U}_{\infty}}\chi^{(v)}_{\square}(t-\mathcal{B}(v)).$ (3.9)
 $Z_{\mathrm{a}}$ and $Z_{\mathrm{b}}$ ‘count’ the branching process $\mathrm{BP}$ according to the characteristics $\chi_{\mathrm{a}}$ and $\chi_{\mathrm{b}}$, respectively.
@@
 We then set
-$N(t):=Z^{\mathrm{b}}_{t}+(Z^{\mathrm{b}}_{t}-Z^{\mathrm{a}}_{t})\qquad\text{for }t\geq 0,$ (3.19)
-as the number of births and deaths that have occurred in the branching process up to time $t$. We define the stopping times $(\tau_{n})_{n\in\mathbb{N}}$ as
-$\tau_{n}:=\inf\{t\geq 0:N(t)=n\},\qquad n\in\mathbb{N}.$ (3.20)
+$N(t)\coloneq Z^{\mathrm{b}}_{t}+(Z^{\mathrm{b}}_{t}-Z^{\mathrm{a}}_{t})\qquad\text{for }t\geq 0,$ (3.10)
+as the number of births and deaths that have occurred in the branching process up to time $t$. We let $(\tau_{n})_{n\in\mathbb{N}}$ be stopping times, defined as
+$\tau_{n}\coloneq\inf\{t\geq 0:N(t)=n\},\qquad n\in\mathbb{N}.$ (3.11)
 That is, $\tau_{n}$ denotes the time at which the $n^{\mathrm{th}}$ birth or death event occurs.
@@
 The proposition follows from the memoryless property of exponential random variables and properties of minima of exponential random variables.
-The technique of embedding evolving discrete structures was originally pioneered by Athreya and Karlin [1] and has found fruitful applications in a wide variety of discrete (combinatorial) models, such as Pólya urns, discrete randomly growing trees such as preferential attachment trees and uniform attachment trees, evolving simplicial complexes, and many more.
+The technique of embedding evolving discrete structures was originally pioneered by Athreya and Karlin [1] and has found fruitful applications in a wide variety of discrete (combinatorial) models, such as Pólya urns, discrete randomly growing trees such as preferential attachment trees and uniform attachment trees (as first observed by Pittel [31]), evolving simplicial complexes, and many more.
 The description of the branching process $\mathrm{BP}$ in which we embed the PAVD tree is used by Deijfen [13] to determine an explicit description for the limiting offspring distribution of alive individuals.
-The description of the branching process used here is somewhat different (though equivalent), but provides an analytical advantage.
-In particular, defining the offspring and lifetime of an individual as an i.i.d.
-copy of $D$, see (2.6), and $L$, defined in (3.2), respectively, is a perspective which allows us to analyse these quantities to a greater extent.
+The description of the branching process used here is somewhat different (though equivalent).
+Namely, the explicit construction in terms of the sum of inter-birth times $S_{k}$, the offspring random variable $D$, and the lifetime $L$, was not used by Deijfen.
+This viewpoint provides an analytical advantage.
+In particular, it allows us to analyse the lifetime distribution of individuals and the number of children an individual produces over time to a greater extent.
+Persistence in the CTPAVD branching process.
+We can view the concepts of persistent $m$-hubs, and persistence and the lack thereof from a continuous-time perspective as well. To this end, we define
+$\deg_{t}(u)\coloneq\min\big\{\sup\big\{k\in\mathbb{N}_{0}:S_{k}^{(u)}\leq t-\mathcal{B}(u)\big\},D^{(u)}\big\},\qquad u\in\mathcal{U}_{\infty},t\geq 0,$ (3.13)
+as the number of children $u$ has produced by time $t$.
+We use the convention that $S_{0}^{(v)}=0$ for all $v$ and that we set $\deg_{t}(u)=$-$\infty$ when $t<\mathcal{B}(u)$, i.e.
+when $u$ has not (yet) been born.
+We also define
+$O_{t}^{\mathrm{ct}}\coloneq\!\!\min_{v\in\mathcal{A}_{t}^{\mathrm{ct}}}\!\mathcal{B}(v)$ (3.14)
+as the birth-time of the oldest alive individual at time $t$, which is the continuous-time counterpart of $O_{n}$, as in (2.2). Similarly, we define
+$I_{t}^{(1),\mathrm{ct}}\coloneq\min\big\{\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\big\},$ (3.15)
+and for $m\geq 2$ fixed,
+$I_{t}^{(m),\mathrm{ct}}\coloneq\min\big\{\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]}\big\}.$ (3.16)
+Here $I_{t}^{(m),\mathrm{ct}}$ denotes the birth-time of the oldest alive individual that has the $m^{\mathrm{th}}$ largest number of children at time $t$ and is the continuous-time counterpart of $I_{n}^{(m)}$.
+We say a branching process $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent elder when there exists a random variable $O^{\mathrm{ct}}$ that is finite almost surely, such that
+$O^{\mathrm{ct}}_{t}\to O^{\mathrm{ct}}\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely as $t\to\infty$}.$ (3.17)
+We say that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when there exists an almost surely finite random variable $I^{(m),\mathrm{ct}}$ such that
+$I_{t}^{(m),\mathrm{ct}}\to I^{(m),\mathrm{ct}}\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely as $t\to\infty$}.$ (3.18)
+When $(\mathrm{BP}(t))_{\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, we say that persistence occurs in $(\mathrm{BP}(t))_{t\geq 0}$ when
+[…]
+and that lack of persistence occurs when
+$I^{(1),\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}\xrightarrow{\mathbb{P}_{\mathcal{S}}}\infty.$ (3.20)
+As in the discrete setting, $(\mathrm{BP}(t))_{t\geq 0}$ containing a persistent $1$-hub implies that it contains a persistent elder, and it implies persistence in $(\mathrm{BP}(t))_{t\geq 0}$.
+On the other, hand, lack of persistence implies that $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $1$-hub.
 Proposition 3.1 also implies that
@@
 $N(t)\mathrm{e}^{-\lambda^{*}\!t}\overset{\mathrm{a.s.}}{\longrightarrow}W,$ (3.22)
-where $W$ is a non-negative random variable, such that $\mathbb{P}\!\left(W>0\,|\,\mathcal{S}\right)=1$ and hence that $\tau_{n}-\frac{1}{\lambda^{*}}\log n$ converges almost surely.
+where $W$ is a non-negative random variable, such that $\mathbb{P}\!\left(W>0\,|\,\mathcal{S}\right)=1$ and hence that $\tau_{n}-\frac{1}{\lambda^{*}}\log n$ converges almost surely (see also Proposition 5.3).
 As a result, we can, approximately, relate the quantities in (3.21) via
@@
 In particular, this relates the definition of a persistent hub and (lack of ) persistence in $(\mathrm{BP}(t))_{t\geq 0}$ to the same concepts for the PAVD discrete tree process $(T_{n})_{n\in\mathbb{N}}$.
-4. Persistent hubs in general CMJ branching processes
-This section is devoted to proving results regarding the (non-)existence of persistent hubs for a general family of CMJ branching processes.
-In particular, we verify that this family includes the CTPAVD branching process, so that via Proposition 3.1 we obtain Theorem 2.1.
-4.1. General results
-We first state several assumptions that are required to prove the general results.
-Recall the definition of the offspring $D^{(v)}$ and inter-birth times $(X(vi))_{i\in\mathbb{N}}$ of an individual $v\in\mathcal{U}_{\infty}$ after (3.1) We view these as i.i.d.
-copies of random variables $D$ and $(X(i))_{i\in\mathbb{N}}$.
-Also recall the branching process $(\mathrm{BP}(t))_{t\geq 0}$ in (3.4).
-We have the following assumption.
-Assumption 4.1.
-The random variables $D$ and $(X(i))_{i\in\mathbb{N}}$ satisfy the following:
-  1. (1)
-    $D$ and $(X(i))_{i\in\mathbb{N}}$ are mutually independent.
-  2. (2)
-    $X(i)<\infty$ for all $i\in\mathbb{N}$ almost surely.
-  3. (3)
-    $\sum_{k=1}^{\infty}\mathbb{P}\!\left(D\geq k\right)\prod_{i=1}^{k}\mathbb{P}\!\left(X(i)=0\right)<1$.
-  4. (4)
-    $\mathbb{P}\!\left(D=\infty\right)>0$.
-Parts $(1)$ and $(2)$ of Assumption 4.1 are relatively straightforward.
-Part $(3)$ states that the expected number of children that are born instantaneously (i.e.
-born at the same time as their parent) is strictly smaller than one, so that in particular sub-trees of the branching process formed by individuals born instantaneously are finite almost surely.
-Part $(4)$ finally implies that individuals can produce an infinite offspring with positive probability, which is the crucial assumption that distinguishes between Theorems 2.1 and 2.6.
-We then have the following result classifying when $(\mathrm{BP}(t))_{t\geq 0}$ does not contains a persistent $m$-hub for any $m\in\mathbb{N}$, and when $(\mathrm{BP}(t))_{t\geq 0}$ does contains a persistent $1$-hub.
-To this end, we recall the general form of $\widehat{\mu}$ in (3.8).
-Theorem 4.2 (Existence of $1$-hubs).
-Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that Assumption 4.1 is satisfied. We then have:
+In Sections 4 and 5 we analyse the CMJ branching process defined in this section and, in particular, prove results under which conditions the behaviour in (3.18) and (3.19) occurs.
+In each of these sections, we use these results when proving the main theorems by leveraging Proposition 3.1.
+4. (Non)-existence of persistent hubs
+This section is devoted to proving results regarding the (non-)existence of persistent hubs for the CTPAVD branching process introduced in Section 3. Via Proposition 3.1 we then obtain Theorem 2.1.
+We first state the following result classifying when $(\mathrm{BP}(t))_{t\geq 0}$ does (not) contain a persistent $m$-hub for any $m\in\mathbb{N}$
+Theorem 4.1 ((Non-)existence of $m$-hubs).
+Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7).
+Assume that Assumption (N-E) is satisfied and Assumption (F-D) is not satisfied.
   1. (a)
-    Suppose the series $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely.
-Then, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub.
+    If Assumption (C-V) is not satisfied, then for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub.
   2. (b)
-    Suppose there exist $\lambda>0$ and $K\in\mathbb{N}$ such that
-    $\widehat{\mu}(\lambda)<1\qquad\text{and}\qquad\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]<\infty.$ (PHC)
-    Then, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $1$-hub.
-Remark 4.3.
-We observe that the condition that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely in $(a)$, or the condition that $\widehat{\mu}(\lambda)<1$ in (PHC), implies that $\sum_{i=1}^{\infty}X(i)$ diverges almost surely.
-Indeed, if $\sum_{i=1}^{\infty}X(i)<\infty$ were to hold with positive probability, then
-$\bigg{|}\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))\bigg{|}\leq\sum_{i=1}^{\infty}X(i)^{\prime}+\sum_{i=1}^{\infty}X(i)<\infty,$ (4.1)
-with positive probability (where we use that $X(i)^{\prime}\overset{\mathrm{d}}{=}X(i)$ for all $i\in\mathbb{N}$), which contradicts the condition in $(a)$. For the first condition in (PHC),
-$\widehat{\mu}(\lambda)=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\lambda\sum_{i=1}^{j}X(i)\bigg{)}\right]\geq\mathbb{E}\left[D\right]\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{\infty}X(i)\bigg{)}\right],$ (4.2)
-[…]
-In the case that the inter-birth times are exponential random variables, we can extend the results of Theorem 4.2 to the existence of $m$-hubs for any $m\in\mathbb{N}$.
-Theorem 4.4 (Existence of $m$-hubs).
-Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that $X(i)$ follows an exponential distribution for all $i\in\mathbb{N}$ and $\mathbb{P}\!\left(D=\infty\right)>0$.
-Suppose there exist $\lambda>0$ and $K\in\mathbb{N}$ such that (PHC) is satisfied.
-Then, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub.
-The conditions in Theorem 4.2 $(b)$ and Theorem 4.4 ensure the almost sure emergence of a persistent $m$-hub for any $m\in\mathbb{N}$.
-The following result provides additional conditions under which the persistent $m$-hub is unique, and under which the offspring of the $m$-hubs are distinct.
-Theorem 4.5.
-Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that Assumption 4.1 and (PHC) are satisfied, and fix $m\in\mathbb{N}$.
-Suppose that $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub.
-Also suppose that either
-$\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)^{\prime}-X(i)\neq 0\right)=\infty\quad\text{or}\quad\mathbb{P}\!\left(X(i)^{\prime}\neq X(i)\right)=1\quad\text{for some }i\in\mathbb{N}.$ (4.3)
-Then, $\mathbb{P}_{\mathcal{S}}$-almost surely the process $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub. Additionally, suppose that for any $\varepsilon>0$ we have
-$\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)>\varepsilon\right)<\infty.$ (4.4)
-Then, for any $m\geq 2$ there exists $t^{\prime}>0$ such that $\mathbb{P}_{\mathcal{S}}$-almost surely for all $t>t^{\prime}$,
+    If Assumption (C-V) is satisfied and Assumption (F-L) is satisfied, then for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub, in the sense of (3.18).
+For any $m\geq 2$ there exists $t^{\prime}>0$ such that $\mathbb{P}_{\mathcal{S}}$-almost surely for all $t>t^{\prime}$,
     $\deg_{t}(I^{(1),\mathrm{ct}})>\deg_{t}(I^{(2),\mathrm{ct}})>\cdots>\deg_{t}(I^{(m),\mathrm{ct}}).$ (4.1)
-Remark 4.6.
-Theorem 4.5 provides general conditions under which there exists a unique persistent $m$-hub, for any $m$.
-This requires the existence of a persistent $m$-hub to begin with.
-For $m=1$ this is true under the same level of generality by Theorem 4.2 $(b)$, whilst we know it to be true for exponentially distributed inter-birth times for $m>1$ by Theorem 4.4.
-We believe that Theorem 4.4 can be extended to a more general class of inter-birth time distributions as well, for which Theorem 4.5 would then provide the uniqueness of the persistent $m$-hub.
 Related work.
-The results presented in Theorems 4.2, 4.4, and 4.5 are strongly related to recent work of Iyer on persistent hubs in CMJ branching processes [25].
-Iyer proves similar results as in Theorem 4.2 and 4.5 (the case $m=1$) under the additional assumption that $D=\infty$, i.e.
-when all individuals live forever and produce an infinite offspring.
-Here, based on proof techniques developed by Iyer, we expand this to the case $m>1$ and to the case that $\mathbb{P}\!\left(D=\infty\right)\in(0,1)$.
-Section 5 deals with Theorem 2.6, which considers the case $\mathbb{P}\!\left(D=\infty\right)=0$.
-A further distinction between our results and those of Iyer are the conditions in (4.3), under which we prove that a unique persistent $m$-hub emerges.
-These are different to the conditions provided by Iyer, as we think that these conditions are too restrictive.
-Indeed, our proof requires a milder condition under which we can prove that a unique persistent $m$-hub emerges.
-4.2. Application to the CTPAVD branching process: Proof of Theorem 2.1
-To apply the results in Theorems 4.2, 4.4, and 4.5 to the case of the CTPAVD branching process, we first state the following lemma, which is a slight generalisation of Lemma $3.5$ in [25].
+The results presented in Theorem 4.1 are related to recent work of Iyer on persistent hubs in general CMJ branching processes [25].
+The main difference between the results of Iyer and the results presented here are:
+  • Iyer focusses on general branching processes, where the inter-birth times $X(i)$ are not necessarily exponentially distributed.
+  • Iyer focusses the case where the offspring $D^{(v)}$ is infinite for each individual $v$.
+  • The results of Iyer are restricted to the case $m=1$ and do not include (4.1).
+In short, based on proof techniques developed by Iyer, we extend his results to the case $m>1$ and to the case that $\mathbb{P}\!\left(D=\infty\right)\in(0,1]$ (which is the case when Assumption (F-D) is not satisfied) when the inter-birth times are exponentially distributed.
+These extensions can be proved for more general CMJ branching processes, where we do not assume the inter-birth times to be exponentially distributed, as well.
+As the focus of this work lies on the PAVD model, we do not prove this explicitly, but rather provide a discussion on how to extend (part of) the results in Theorem 4.1 to general CMJ branching processes can be found Appendix A.
+4.1. Proof of Theorem 2.1
+To translate the results in Theorem 4.1 to the results of Theorem 2.1, we first state the following lemma, which is a slight generalisation of Lemma $3.5$ in [25].
 Lemma 4.2.
-Recall $\widehat{\mu}$ from (3.8). Suppose that $(1)-(3)$ in Assumption 4.1 are satisfied and that there exists $\lambda>0$ such that
-$\widehat{\mu}(\lambda)<\infty.$ (4.6)
+Recall the Laplace transform $\widehat{\mu}$ from (2.8) and suppose that Assumption (F-L) is satisfied.
+That is, suppose that there exists $\lambda>0$ such that $\widehat{\mu}(\lambda)<\infty$.
 Then, $|\mathrm{BP}(t)|<\infty$ for all $t\geq 0$ almost surely.
 Proof.
-As $\widehat{\mu}(\lambda)$ is decreasing in $\lambda$, we first observe that (4.6), combined with Assumption 4.1, implies that there exists $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$.
-Indeed, by using Fubini’s theorem to interchange summation and integration, we can write for any $\alpha\geq 0$,
-$\widehat{\mu}(\alpha)=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]$ $=\sum_{j=1}^{\infty}\mathbb{E}\left[\mathbbm{1}_{\{D\geq j\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]$ (4.7)
-$=\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right],$
-where we use the independence of $D$ and the $X(i)$ in the last step. We can split the right-hand side into two terms, namely
-$\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\prod_{i=1}^{j}\mathbb{P}\!\left(X(i)=0\right)+\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right].$ (4.8)
-By $(3)$ in Assumption 4.1 there exists $\eta>0$ such that the first sum is at most $1-\eta$.
-The second sum is finite for any $\alpha\geq\lambda$, since the left-hand side of (4.7) is finite for any $\alpha\geq\lambda$ by assumption.
-We can thus take $J=J(\eta)\in\mathbb{N}$ sufficiently large, such that for any $\alpha\geq\lambda$,
-$\sum_{j=J}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]<\frac{\eta}{2}.$ (4.9)
-We can then choose $\alpha=\lambda^{\prime}\geq\lambda$ sufficiently large such that
-$\sum_{j=1}^{J-1}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{j}X(i)\bigg{)}\right]<\frac{\eta}{2},$ (4.10)
-by the dominated converge theorem (as each summand can be bounded from above by one). Combined, we thus arrive at $\widehat{\mu}(\lambda^{\prime})<1$.
-We then prove the desired result.
-Let $Y$ be an exponential random variable with rate $\lambda^{\prime}$, independent of the offspring random variables $(D^{(v)})_{v\in\mathcal{U}_{\infty}}$ and inter-birth times $(X(vi))_{v\in\mathcal{U}_{\infty},i\in\mathbb{N}}$.
-Then, for any $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$, with $m\in\mathbb{N}$ and $v_{1},\ldots,v_{m}\in\mathbb{N}$,
-$\mathbb{P}\!\left(\mathcal{B}(v)\leq Y\right)=\mathbb{P}\!\left(D^{(v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\text{, and }\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell}}X(v_{1}\cdots v_{\ell-1}i)\leq Y\right).$ (4.11)
-We use the independence of the offspring random variables and the inter-birth times, together with the fact that $Y$ is exponentially distributed, to write this as
-$\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$ (4.12)
-By summing over all individuals $v\in\mathcal{U}_{\infty}$ (where $\mathbb{P}\!\left(\mathcal{B}(\varnothing)\leq Y\right)=1$ as $\mathcal{B}(\varnothing)=0$), we thus obtain
-$\mathbb{E}\left[|\mathrm{BP}(Y)|\right]=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}\!\left(\mathcal{B}(v)\leq Y\right)=1+\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$ (4.13)
+We start by observing that Assumption (F-L) is equivalent to there existing $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$.
+Indeed, since $\widehat{\mu}(\lambda)<\infty$ and $\widehat{\mu}(\alpha)$ is decreasing in $\alpha$, it follows from the dominated convergence theorem that
+$\lim_{\alpha\to\infty}\widehat{\mu}(\alpha)=\mu(\{0\})=0\qquad\text{and so there exists $\lambda^{\prime}\in[\lambda,\infty)$ such that }\widehat{\mu}(\lambda^{\prime})<1.$ (4.2)
+We then prove the result. For any $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$, with $m\in\mathbb{N}$ and $v_{1},\ldots,v_{m}\in\mathbb{N}$,
+$\mathbb{P}\!\left(\mathcal{B}(v)\leq t\right)=\mathbb{P}\!\left(D^{(v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\text{, and }\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(v_{1}\cdots v_{\ell-1})}_{i}\leq t\right).$ (4.3)
+We use the independence of the offspring random variables and the inter-birth times, together with a Chernoff bound with $\lambda^{\prime}$, to bound the right-hand side from above by
+$\mathrm{e}^{\lambda^{\prime}t}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$ (4.4)
+By summing over all individuals $v\in\mathcal{U}_{\infty}$ (where $\mathbb{P}\!\left(\mathcal{B}(\varnothing)\leq t\right)=1$ as $\mathcal{B}(\varnothing)=0$), we obtain
+$\mathbb{E}\left[|\mathrm{BP}(t)|\right]$ $=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}\!\left(\mathcal{B}(v)\leq t\right)$ (4.5)
+$\leq 1+\mathrm{e}^{\lambda^{\prime}t}\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
 As the terms in the product depend on distinct summation indices, we can interchange summation and product to yield
-$\sum_{m=0}^{\infty}\Bigg{(}\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{j}X(i)\bigg{)}\right]\Bigg{)}^{m}=\sum_{m=0}^{\infty}\widehat{\mu}(\lambda^{\prime})^{m},$ (4.14)
-where we use (4.7) in the final step.
-As $\widehat{\mu}(\lambda^{\prime})<1$, the expected value is finite.
-Hence, $|\mathrm{BP}(Y)|<\infty$ almost surely.
-Since $Y\geq t$ with positive probability for any $t\geq 0$, it thus follows that $|\mathrm{BP}(t)|<\infty$ for any $t\geq 0$, almost surely.
+$\sum_{m=0}^{\infty}\Bigg(\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{j-1}E_{i}\bigg)\right]\Bigg)^{m}.$ (4.6)
+By rewriting the inner sum and using that $D$ is independent of $(E_{i})_{i\in\mathbb{N}}$, we obtain
+$\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{j-1}E_{i}\bigg)\right]=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{j-1}E_{i}\bigg)\right]=\widehat{\mu}(\lambda^{\prime}),$ (4.7)
+where we recall $\widehat{\mu}$ from (2.8).
+We finally obtain that $\mathbb{E}\left[|\mathrm{BP}(t)|\right]\leq\mathrm{e}^{\lambda^{\prime}t}\sum_{m=0}^{\infty}\widehat{\mu}(\lambda^{\prime})^{m}$, which is finite since $\widehat{\mu}(\lambda^{\prime})<1$.
+Hence, $|\mathrm{BP}(t)|<\infty$ almost surely, as desired.
 ∎
-We then prove Theorem 2.1.
+Theorem 2.1 then readily follows.
 Proof of Theorem 2.1.
-We first check that the CTPAVD branching process satisfies all the conditions in Theorems 4.2, 4.4, and 4.5. We then translate this back to the PAVD model.
-For the CTPAVD branching process, we have
-$X(i)\overset{\mathrm{d}}{=}X(i)^{\prime}\overset{\mathrm{d}}{=}\text{Exp}(b(i-1)+d(i-1))\qquad\text{for all }i\in\mathbb{N},$ (4.15)
-and $D$ is as in (2.6).
-It directly follows that $(1)-(3)$ of Assumption 4.1 are satisfied, and $(4)$ of Assumption 4.1 is equivalent to Assumption (F-D) not being satisfied, due to Lemma 5.1 (see also [23, Lemma $5.3$]).
-For $(a)$ in Theorem 4.2 it remains to check that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely. As the summands are independent and symmetric, we have
-${\rm Var}\bigg{(}\sum_{i=1}^{n}(X(i)^{\prime}-X(i))\bigg{)}=2\sum_{i=1}^{n}{\rm Var}(X(i))=\sum_{i=1}^{n}\frac{2}{(b(i)+d(i))^{2}},$ (4.16)
-which has a finite limit as $n$ tends to infinity if and only if Assumption (C-V) is not satisfied.
-Kolmogorov’s two and three series theorems (see e.g.
-[19, Theorems $2.5.3$ and $2.5.4$]) thus imply that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely if and only if Assumption (C-V) is not satisfied, so that $(a)$ follows.
-So, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub, for any $m\in\mathbb{N}$, when Assumption (C-V) is not satisfied.
-For $(b)$, we directly conclude from the above that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges when Assumption (C-V) is satisfied.
-It remains to check the conditions in (PHC).
-For the second condition in (PHC), we first observe that $b+d$ tends to infinity when Assumption (C-V) is satisfied.
-For any $\lambda>0$ we can thus choose $K\in\mathbb{N}$ sufficiently large such that $b(i)+d(i)\geq\lambda\sqrt{2}$ for all $i\geq K$.
-Now,
-$\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda((X(i)^{\prime}-X(i))}\right]=\prod_{i=K+1}^{\infty}\frac{(b(i)+d(i))^{2}}{(b(i)+d(i)+\lambda)(b(i)+d(i)-\lambda)}=\prod_{i=K+1}^{\infty}\frac{(b(i)+d(i))^{2}}{(b(i)+d(i))^{2}-\lambda^{2}}.$ (4.17)
-Taking the logarithm of the right-hand side yields
-$-\sum_{i=K+1}^{\infty}\log\Big{(}1-\frac{\lambda^{2}}{(b(i)+d(i))^{2}}\Big{)}.$ (4.18)
-As $b(i)+d(i)\geq\lambda\sqrt{2}$ for all $i\geq K$ and since $\log(1-x)\geq-x-x^{2}$ for $x\in(0,1/2)$, we thus arrive at the upper bound
-$\lambda^{2}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}+\lambda^{4}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{4}}\leq\frac{3}{2}\lambda^{2}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}$ (4.19)
-which is finite for any $\lambda>0$ when Assumption (C-V) is satisfied.
-The first condition in (PHC) is implied by Assumption (F-L), i.e.
-that $\widehat{\mu}(\lambda)<\infty$ for some $\lambda>0$.
-We then, as in the first part of the proof of Lemma 4.7, conclude that there exists $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$.
-We note that the second condition in (PHC) is satisfied for all $\lambda>0$ and any $K\geq\lambda\sqrt{2}$, so that is satisfied for $\lambda^{\prime}$ and $K\geq\lambda^{\prime}\sqrt{2}$ in particular.
-So, with $\lambda^{\prime}$ and $K\geq\lambda^{\prime}\sqrt{2}$, we conclude that (PHC) is satisfied, so that part $(b)$ of Theorem 4.2 holds.
-Hence, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when Assumptions (C-V) and (F-L) are satisfied.
-Theorem 4.4 then also directly follows.
-To obtain Theorem 4.5, we note that the exponential distribution has no atoms, so that there is a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, for any $m\in\mathbb{N}$. Then,
-$\mathbb{P}\!\left(X(i)>\varepsilon\right)=\mathrm{e}^{-\varepsilon(b(i)+d(i))}\leq\frac{1}{(b(i)+d(i))^{2}},$ (4.20)
-where the inequality follows for any $\varepsilon>0$ and all $i\geq I_{\varepsilon}$, for some fixed $I_{\varepsilon}\in\mathbb{N}$, from the fact that $b(i)+d(i)$ tends to infinity with $i$ when Assumption (C-V) holds.
-As a result, the second condition in (4.4) is satisfied (again applying Assumption (C-V)), so that, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely the offspring of the persistent $1$-hub, persistent $2$-hub, through the persistent $m$-hub are distinct for all $t>t^{\prime}$.
-To summarise, we have the following table when the inter-birth times satisfy (4.15) and $D$ is as in (2.6).
-Assumptions Result
-(C-V) and (F-D) not satisfied Theorem 4.2 (a): No persistent $m$-hub
-(C-V) and (F-L) satisfied, (F-D) not satisfied Theorems 4.2 (b) and 4.5: Unique persistent $m$-hub with distinct degrees
-Finally, we translate this result back to the PAVD model, for which we use the embedding in Proposition 3.1.
-First, we claim that $\lim_{n\to\infty}\tau_{n}=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
-This follows since $\mathbb{P}\!\left(\mathcal{S}\right)>0$ and since $\widehat{\mu}(\lambda)<\infty$, so that Lemma 4.7 yields that $\mathrm{BP}(t)$ is finite for all $t\geq 0$, almost surely.
-Hence, $\lim_{n\to\infty}\tau_{n}=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
-As a result, for any $t^{\prime}>0$ there exists $n_{0}\in\mathbb{N}$ such that $\tau_{n}>\tau_{n_{0}}\geq t^{\prime}$ for all $n\geq n_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
-Under the conditions in Theorem 2.1 it thus follows that $(T_{n})_{n\in\mathbb{N}}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$ when Assumption (C-V) is not satisfied.
-When Assumption (C-V) is satisfied and Assumption (F-L) also holds, then for any $m\in\mathbb{N}$ there exists $t^{\prime}>0$ and random variables $I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}}$, such that
-$(I^{(1),\mathrm{ct}}_{t},\ldots,I^{(m),\mathrm{ct}}_{t})=(I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}})$ (4.21)
-for all $t>t^{\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, there exists $n_{0}\in\mathbb{N}$ and random variables $I^{(1)},\ldots,I^{(m)}$, such that
-$(I_{n}^{(1)},\ldots,I_{n}^{(m)})=(I^{(1)},\ldots,I^{(m)})$ (4.22)
-for all $n\geq n_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
-Equivalently, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, for any $m\in\mathbb{N}$, when Assumptions (C-V) and (F-L) are satisfied.
-Since the exponential distribution has no atoms, it follows from the first part of Theorem 4.5 that when Assumption (C-V) is satisfied, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub.
-It then directly follows that $(T_{n})_{n\in\mathbb{N}}$ contains a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely as well, for any $m\in\mathbb{N}$.
-We remark that this also directly follows from the observation that at most one vertex increases its degree at every step in the discrete tree process $(T_{n})_{n\in\mathbb{N}}$, so that at most one vertex can attain the $m^{\mathrm{th}}$ largest degree for all but finitely many steps.
-Finally, for any $m\in\mathbb{N}$ there exists $t^{\prime}>0$ and random variables $I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}}$ such that
-$(I^{(1),\mathrm{ct}}_{t},\ldots,I^{(m),\mathrm{ct}}_{t})=(I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}})$ (4.23)
-for all $t>t^{\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Then, there exists $t^{\prime\prime}>t^{\prime}$ such that
-$\deg_{t}(I^{(1),\mathrm{ct}})>\cdots>\deg_{t}(I^{(m),\mathrm{ct}})$ (4.24)
-for all $t>t^{\prime\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As $\tau_{n}$ tends to infinity with $n$ $\mathbb{P}_{\mathcal{S}}$-almost surely, and
-$\{\deg_{n}(i):i\in\mathcal{A}_{n}\}\overset{\mathrm{d}}{=}\{\deg_{\tau_{n}}(v):v\in\mathcal{A}_{\tau_{n}}^{\mathrm{ct}}\}$ (4.25)
-by Proposition 3.1, we thus obtain that there exists $n_{1}\in\mathbb{N}$ such that
-$\deg_{n}(I^{(1)})>\cdots>\deg_{n}(I^{(m)})$ (4.26)
-for all $n\geq n_{1}$ $\mathbb{P}_{\mathcal{S}}$-almost surely, which together with (4.22) yields (2.12) and concludes the proof. ∎
-The remainder of this section is devoted to proving Theorems 4.2, 4.4, and 4.5.
+Recall the stopping times $(\tau_{n})_{n\in\mathbb{N}}$ from (3.11).
+By using Lemma 4.2 we observe that $\tau_{n}\to\infty$ almost surely conditionally on survival.
+As a result, by using Proposition 3.1 we obtain that Theorem 4.1 directly implies Theorem 2.1.
+∎
+The remainder of this section is devoted to proving Theorem 4.1.
 4.2. Preliminary results
-We first collect a few preliminary results that are used throughout the proof of Theorems 4.2, 4.4, and 4.5.
+We first collect a few preliminary results that are used throughout the proof of Theorem 4.1.
 Lemma 4.3 (Lemma $3.4$ in [25]).
@@
 Then,
-$\mathbb{P}\bigg{(}\exists j\in\mathbb{N}_{0}:Y+\sum_{i=1}^{k+j}X(i)\leq\sum_{i=k}^{k+j}X(i)^{\prime}\bigg{)}\leq\bigg{(}\prod_{i=k}^{\infty}\mathbb{E}\Big{[}\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\Big{]}\bigg{)}\mathbb{E}\bigg{[}\exp\bigg{(}-\lambda\bigg{(}Y+\sum_{i=1}^{k-1}X(i)\bigg{)}\bigg{)}\bigg{]}.$ (4.28)
-Lemma 4.9 (Theorem $1.13$ in [24]).
-Suppose that $(S_{i})_{i\in\mathbb{N}}$ is a sequence of mutually independent random variables such that $\sum_{i=1}^{\infty}S_{i}$ converges almost surely.
-The distribution of $\sum_{i=1}^{\infty}S_{i}$ contains an atom on $\mathbb{R}$ if and only if for some collection $(c_{i})_{i\in\mathbb{N}}\in\mathbb{R}^{\infty}$,
-$\forall i\in\mathbb{N}:\mathbb{P}\!\left(S_{i}=c_{i}\right)>0\qquad\text{and}\qquad\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq c_{i}\right)<\infty.$ (4.29)
+$\mathbb{P}\bigg(\exists j\in\mathbb{N}_{0}\colon Y+\sum_{i=1}^{k+j}X(i)\leq\sum_{i=k}^{k+j}X(i)^{\prime}\bigg)\leq\bigg(\prod_{i=k}^{\infty}\mathbb{E}\Big[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\Big]\bigg)\mathbb{E}\bigg[\exp\bigg(-\lambda\bigg(Y+\sum_{i=1}^{k-1}X(i)\bigg)\bigg)\bigg].$ (4.9)
+In short, the proof of Lemma 4.3 is based on the fact that the random variable $\exp(\lambda\sum_{i=k}^{k+j}(X(i)-X(i)^{\prime}))$ is a sub-martingale combined with Chernoff’s inequality.
+Lemma 4.4 (Proposition $1.15$ in [24]).
+Suppose that $(S_{i})_{j\in\mathbb{N}}$ is a sequence of mutually independent, symmetric random variables. If $\sum_{i=1}^{\infty}S_{i}$ diverges almost surely, then
+$\limsup_{n\to\infty}\sum_{i=1}^{n}S_{i}=\infty\qquad\text{and}\qquad\liminf_{n\to\infty}\sum_{i=1}^{n}S_{i}=-\infty,\qquad\text{almost surely.}$ (4.10)
+The proof of Lemma 4.4 is based on the fact that $\sum_{i=1}^{n}S_{i}$ is a martingale, combined with the fact that divergent martingales are not bounded almost surely.
+Lemma 4.5.
+Let $(E_{i})_{i\in\mathbb{N}_{0}}$ and $(E_{i}^{\prime})_{i\in\mathbb{N}_{0}}$ be independent exponential random variables, where $E_{i}$ and $E_{i}^{\prime}$ have rate $\lambda_{i}>0$ for each $i\in\mathbb{N}_{0}$.
+Then,
+$\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})\text{ converges almost surely}\quad\Leftrightarrow\quad\sum_{i=0}^{\infty}\frac{1}{\lambda_{i}^{2}}<\infty.$ (4.11)
 Remark 4.6.
-In case the $S_{i}$ are symmetric random variables, i.e.
-$\mathbb{P}\!\left(S_{i}\geq x\right)=\mathbb{P}\!\left(S_{i}\leq-x\right)$ for all $x\geq 0$ and $i\in\mathbb{N}$, the statement can be simplified by restricting to collections $(c_{i})_{i\in\mathbb{N}}$ that have finitely many non-zero values only.
-Indeed, let $(i_{j})_{j\in\mathbb{N}}$ be a strictly increasing sub-sequence such that $c_{i_{j}}\neq 0$ for all $j\in\mathbb{N}$.
-The first condition in (4.29) yields $\mathbb{P}\!\left(S_{i_{j}}=-c_{i_{j}}\right)=\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)>0$.
-Then, the condition
-$\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq c_{i}\right)<\infty$ (4.30)
-in (4.29) implies that $\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)=1-o(1)$, where the $o(1)$ tends to zero with $j$.
-There thus exists $J\in\mathbb{N}$ such that $\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)=\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)>\frac{1}{2}$ for all $j\geq J$, which yields a contradiction.
-So, any collection $(c_{i})_{i\in\mathbb{N}}$ can only contain finitely many non-zero values.
-Lemma 4.11 (Proposition $1.15$ in [24]).
-Suppose that $(S_{i})_{j\in\mathbb{N}}$ is a sequence of mutually independent, symmetric random variables. If $\sum_{i=1}^{\infty}S_{i}$ diverges almost surely,
-$\limsup_{n\to\infty}\sum_{i=1}^{n}S_{i}=\infty\qquad\text{and}\qquad\liminf_{n\to\infty}\sum_{i=1}^{n}S_{i}=-\infty,\qquad\text{almost surely.}$. (4.31)
-4.4. Proving Theorems 4.2, 4.4, and 4.5
-In this sub-section, we first prove part $(a)$ of Theorem 4.2, then part $(b)$ of Theorem 4.2 and Theorem 4.4, and we conclude by proving Theorem 4.5.
+$(i)$ If we set $X(i)=E_{i}$ and $X(i)^{\prime}=E_{i}^{\prime}$ in Lemma 4.3, it also follows that for any $\lambda>0$ there exist $k\in\mathbb{N}$ such that (4.8) is satisfied when
+$\sum_{i=0}^{\infty}\frac{1}{\lambda_{i}^{2}}<\infty.$ (4.12)
+Indeed, as the sum is finite it follows that $\lambda_{i}\to\infty$ with $i$, so that for any $\lambda>0$ we can take $k$ large enough such that $\lambda_{i}>\lambda$ for all $i\geq k$, so that
+$\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]=\frac{\lambda_{i}^{2}}{(\lambda_{i}-\lambda)(\lambda_{i}+\lambda)}=\frac{1}{1-\lambda^{2}/\lambda_{i}^{2}}.$ (4.13)
+The product in (4.8) hence is finite when the sum in (4.12) is finite.
+[…]
+Proof.
+To prove the convergence of the series of exponential random variables, we check the conditions of Kolmogorov’s three series theorem (see e.g.
+[19, Theorem $2.5.4$]).
+This theorem states that a random series $\sum_{i=0}^{\infty}X_{i}$ of independent random variables $(X_{i})_{i\in\mathbb{N}_{0}}$ converges almost surely if the following conditions holds for some $A>0$, and only if the following conditions hold for any $A>0$:
+[…]
+We notice that for symmetric random variables, the second condition is always satisfied. It thus suffices to focus on conditions $(i)$ and $(iii)$, for $X_{i}=E_{i}-E_{i}^{\prime}$.
+Let us first prove the ‘if’ direction. We suppose that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$. Then,
+$\sum_{i=0}^{\infty}{\rm Var}(X_{i}\mathbbm{1}_{\{|X_{i}|\leq A\}})=\sum_{i=0}^{\infty}\mathbb{E}\left[X_{i}^{2}\mathbbm{1}_{\{|X_{i}|\leq A\}}\right]\leq\sum_{i=0}^{\infty}\mathbb{E}\left[X_{i}^{2}\right]=\sum_{i=0}^{\infty}2{\rm Var}(E_{i})=\sum_{i=0}^{\infty}\frac{2}{\lambda_{i}^{2}},$ (4.15)
+which is finite by assumption. Condition $(iii)$ is hence satisfied. Also,
+$\sum_{i=0}^{\infty}\mathbb{P}\!\left(|X_{i}|\geq A\right)\leq\sum_{i=0}^{\infty}\mathbb{P}\!\left(E_{i}+E_{i}^{\prime}\geq A\right)\leq 2\sum_{i=0}^{\infty}\mathbb{P}\!\left(E_{i}\geq\frac{A}{2}\right)=2\sum_{i=0}^{\infty}\mathrm{e}^{-\lambda_{i}A/2}$ (4.16)
+By assuming that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$, it follows that $\lambda_{i}\to\infty$ with $i$.
+Hence, there exists $I\in\mathbb{N}$ such that $\mathrm{e}^{-\lambda_{i}A/2}\leq 1/\lambda_{i}^{2}$, from which Condition $(i)$ follows.
+For the ‘only if’ direction, we argue by contradiction.
+We assume that there exists $A>0$ such that either Condition $(i)$ or $(iii)$ does not hold, and additionally assume that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$.
+We show that this leads to a contradiction.
+Suppose first that Condition $(i)$ is not satisfied for some $A>0$.
+By the bound in (4.16) and the assumption that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$, this directly leads to a contradiction.
+Hence, now assume that Condition $(i)$ is satisfied for all $A>0$, but that Condition $(iii)$ is not satisfied for some $A>0$.
+Then, by the second Borel-Cantelli lemma, $|X_{i}|\geq A$ only finitely often almost surely, so that
+$\sum_{i=0}^{\infty}{\rm Var}(X_{i}\mathbbm{1}_{\{|X_{i}|\leq A\}})=\infty\quad\Leftrightarrow\quad\sum_{i=0}^{\infty}{\rm Var}(X_{i})=\infty.$ (4.17)
+However, the equality on the right-hand side contradicts the assumption that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$, since ${\rm Var}(X_{i})=\mathbb{E}\left[X_{i}^{2}\right]=2{\rm Var}{E_{i}}=2\lambda_{i}^{-2}$.
+∎
+4.3. Proving Theorem 4.1(a)
 For $u,v\in\mathcal{U}_{\infty}$, define the event
-$\text{Win}(u,v):=\{\exists T\geq 0:\forall t\geq T\ \deg_{t}(u)\geq\deg_{t}(v)\}.$ (4.32)
+$\text{Win}(u,v)\coloneq\{\exists T\geq 0:\forall t\geq T\ \deg_{t}(u)\geq\deg_{t}(v)\}.$ (4.18)
 If an individual $u$ has not been born in $\mathrm{BP}(t)$, we use the convention that $\deg_{t}(u)=$-$\infty$.
@@
 Lemma 4.7.
-Assume Assumption 4.1 is satisfied. Then,
+Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7).
+Assume that Assumption (N-E) is satisfied and Assumption (F-D) is not satisfied.
+The following statements are true:
   1. (a)
-    If $\sum_{i=1}^{\infty}X(i)=\infty$ almost surely, then $\lim_{t\to\infty}\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
+    $\lim_{t\to\infty}\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
   2. (b)
-    If $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely, then for all $u,v\in\mathcal{U}_{\infty}$,
+    If Assumption (C-V) is satisfied, then for all $u,v\in\mathcal{U}_{\infty}$,
     $\mathbb{P}_{\mathcal{S}}\left(\mathrm{Win}(u,v)\cup\mathrm{Win}(v,u)\right)=1.$ (4.19)
   3. (c)
-    If $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely, there exists a function $\phi:\mathbb{N}\to\mathbb{N}$ such that $\phi(j)>j$ for all $j\in\mathbb{N}$ such that for any $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$,
-    $\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,\phi(j)\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)\geq\frac{1}{2}.$ (4.34)
+    If Assumption (C-V) is not satisfied, there exists a function $\phi:\mathbb{N}\to\mathbb{N}$ such that $\phi(j)>j$ for all $j\in\mathbb{N}$ such that for any $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$,
+    $\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,\phi(j)\}:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\right)\geq\frac{1}{2}.$ (4.20)
 Proof.
-$(a)$ Let $V$ be the first individual born in $(\mathrm{BP}(t))_{t\geq 0}$ such that $D^{(V)}=\infty$ (taking an arbitrary one in case multiple individuals $u_{1},\ldots,u_{M}$ with $D_{u_{1}}=\cdots=D_{u_{M}}=\infty$ are born simultaneously, for some $M\geq 2$).
-Conditionally on survival such an individual $V$ exists almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$.
+$(a)$ As Assumption (F-D) is not satisfied, it follows from [23, Lemma 5.3 and Corollary 6.4] that offspring random variable $D$ is such that $\mathbb{P}\!\left(D=\infty\right)>0$.
+We can hence define $V$ be the first individual born in $(\mathrm{BP}(t))_{t\geq 0}$ such that $D^{(V)}=\infty$.
+Conditionally on survival such an individual $V$ exists almost surely.
 Furthermore, the birth-time of $V$ is finite almost surely.
@@
 $\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)\geq\deg_{t}(V),$ (4.21)
-and the right-hand side tends to infinity with $t$ since $D^{(V)}=\infty$ and as $\sum_{i=1}^{\infty}X(Vi)=\infty$ almost surely (since $X(Vi)\overset{\mathrm{d}}{=}X(i)$ for all $i\in\mathbb{N}$), which yields the desired result.
+and the right-hand side tends to infinity with $t$ since $D^{(V)}=\infty$.
 $(b)$ The event $\text{Win}(u,v)\cup\text{Win}(v,u)$ is trivially satisfied if at most one of the two individuals $u$ and $v$ is born.
-Given that both vertices are born, again the event is trivially satisfied if at most one individual produces an infinite offspring.
+Given that both vertices are born, again the event is trivially satisfied if at most one individual produces infinite offspring.
 As a result, we define
-$\mathcal{E}(u,v):=\{\mathcal{B}(u),\mathcal{B}(v)<\infty\}\cap\{D^{(u)},D_{v}=\infty\},$ (4.36)
+$\mathcal{E}(u,v)\coloneq\{\mathcal{B}(u),\mathcal{B}(v)<\infty\}\cap\{D^{(u)},D_{v}=\infty\},$ (4.22)
 and we can thus write
@@
 $\mathbb{P}\!\left((\mathrm{Win}(u,v)\cup\mathrm{Win}(v,u))\cap\mathcal{E}(u,v)\right)=\mathbb{P}\!\left(\mathcal{E}(u,v)\right).$ (4.24)
-Recall the assumption that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely. Then, for $u,v\in\mathcal{U}_{\infty}$, on the event $\mathcal{E}(u,v)$ we have
-$\left|\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{\infty}(X(ui)-X(vi))\right|<\infty\quad\text{almost surely.}$. (4.39)
+Lemma 4.5 tells us that Assumption (C-V) being satisfied is equivalent to the infinite series $\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})$ converging almost surely.
+Then, for $u,v\in\mathcal{U}_{\infty}$, on the event $\mathcal{E}(u,v)$ we have
+$\left|\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=0}^{\infty}(E^{(u)}_{i}-E^{(v)}_{i})\right|<\infty\quad\text{almost surely,}$ (4.25)
+since $E^{(u)}_{i}\overset{\mathrm{d}}{=}E_{i}$ and $E^{(v)}_{i}\overset{\mathrm{d}}{=}E_{i}^{\prime}$.
 Assume, without loss of generality, that $|u|\leq|v|$.
-Let $N$ be chosen such that $\mathcal{B}(v)$ is independent of $(X(un))_{n\geq N}$.
-For example, if $u$ is a non-parent ancestor of $v$, we may choose the value $N$ such that $u(N-1)$ is the ancestor of $v$, whilst if $v=uj$ for $j\in\mathbb{N}$, we can choose $N>j$.
-Then, the sequences $(X(un))_{n\geq N}$ and $(X(vn))_{n\geq N}$ are independent.
-We now have two cases:
-(I) Suppose that the sequence $(S_{i})_{i\geq N}$ defined by $S_{i}:=X(ui)-X(vi)$ does not satisfy (4.29). Then, by Lemma 4.9, the random variable
-$\sum_{j=N}^{\infty}(X(uj)-X(vj))$ (4.40)
-contains no atom on $\mathbb{R}$. Hence, as the summands are independent
-$\mathbb{P}\bigg{(}\mathcal{E}(u,v)\cap\bigg{\{}\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{\infty}(X(uj)-X(vj))=0\bigg{\}}\bigg{)}=0.$ (4.41)
+Let $N$ be chosen such that $\mathcal{B}(v)$ is independent of $(E^{(u)}_{n})_{n\geq N}$.
+For example, if $u$ is a non-parent ancestor of $v$, we may choose the value $N$ such that u(N1)u(N-1) is the ancestor of $v$, whilst if $v=uj$ for $j\in\mathbb{N}$, we can choose $N>j$.
+Then, the sequences $(E^{(u)}_{n})_{n\geq N}$ and $(E^{(v)}_{n})_{n\geq N}$ are independent.
 Therefore, on the event $\mathcal{E}(u,v)$ there exists $K_{0}\in\mathbb{N}$ such that either
-$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{k}(X(uj)-X(vj))>0\qquad\text{for all $k\geq K_{0}$,}$ (4.42)
+$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=0}^{k-1}(E^{(u)}_{j}-E^{(v)}_{j})>0\qquad\text{for all $k\geq K_{0}$,}$ (4.26)
 or
-$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{k}(X(uj)-X(vj))<0\qquad\text{for all $k\geq K_{0}$,}$ (4.43)
+$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=0}^{k-1}(E^{(u)}_{j}-E^{(v)}_{j})<0\qquad\text{for all $k\geq K_{0}$,}$ (4.27)
 is satisfied $\mathbb{P}_{\mathcal{S}}$-almost surely.
@@
 This in turn implies that for all $k$ sufficiently large $u$ has offspring $k$ before $v$ has offspring $k$ or vice versa.
-Thus, we have that (4.38) is satisfied.
-(II) Suppose that the sequence $(S_{i})_{i\geq N}$ defined by $S_{i}:=X(ui)-X(vi)$ does satisfy (4.29).
-Since $S_{i}$ has a symmetric distribution (as $X(ui)\overset{\mathrm{d}}{=}X(vi)$), any collection $(c_{i})_{i\in\mathbb{N}}$ in Lemma 4.9 can only contain finitely many non-zero entries (see Remark 4.10).
-As a result, this implies that $\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq 0\right)=\sum_{i=1}^{\infty}\mathbb{P}(X(ui)\neq X(vi))<\infty$.
-By the Borel-Cantelli lemma we obtain that $X(ui)=X(vi)$ for all but finitely many $i\in\mathbb{N}$ almost surely.
-Thus, for some $K_{0}\in\mathbb{N}$, it is either the case that (4.42) or (4.43) is satisfied (but not both), or for all $k\geq K_{0}$,
-$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{k}(X(ui)-X(vi))=0.$ (4.44)
-This implies that $\mathcal{B}(uk)\geq\mathcal{B}(vk)$ for all $k\geq K_{0}$ or $\mathcal{B}(u_{k})\leq\mathcal{B}(v_{k})$ for all $k\geq K_{0}$.
-Again, it follows that $\text{Win}(u,v)$ or $\text{Win}(v,u)$ occurs on the event $\mathcal{E}(u,v)$, so that (4.38) holds, which concludes the proof of part $(b)$.
-$(c)$ Since $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely, Lemma 4.11 implies that for any $j\in\mathbb{N}$ we have
-$\limsup_{n\to\infty}\sum_{i=j+1}^{n}(X(ui)-X(uji))=\infty\quad\text{almost surely}.$ (4.45)
-Since also $\sum_{i=1}^{j}X(uji)<\infty$ almost surely by Assumption 4.1, we thus obtain that
-$\mathbb{P}\!\left(\exists k>j:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)=1.$ (4.46)
+Thus, we have that (4.24) is satisfied, which concludes the proof of part $(b)$.
+$(c)$ We again use Lemma 4.5 to conclude that Assumption (C-V) not being satisfied is equivalent to the infinite series $\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})$ diverging almost surely, and use again that $E^{(u)}_{i}\overset{\mathrm{d}}{=}E_{i}$ and $E^{(uj)}_{i}\overset{\mathrm{d}}{=}E_{i}^{\prime}$ for any $j\in\mathbb{N}$ and any $u\in\mathcal{U}_{\infty}$.
+We then use Lemma 4.4 to obtain that for any $j\in\mathbb{N}$ we have
+$\limsup_{n\to\infty}\sum_{i=j}^{n}(E^{(u)}_{i}-E^{(uj)}_{i})=\infty\quad\text{almost surely}.$ (4.28)
+Since also $\sum_{i=0}^{j-1}E^{(u)}_{i}$$\infty$ almost surely, we thus obtain that
+$\mathbb{P}\!\left(\exists k>j:\sum_{i=j}^{k}E^{(u)}_{i}>\sum_{i=0}^{k}E^{(uj)}_{i}\right)=1.$ (4.29)
 Then, by the monotone convergence theorem,
-$\lim_{n\to\infty}\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)=1.$ (4.47)
+$\lim_{n\to\infty}\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\right)=1.$ (4.30)
 As a result, we can define $\phi(j)$ as
-$\phi(j):=\inf\bigg{\{}n>j:\mathbb{P}\bigg{(}\exists k\in\{j+1,\ldots,n\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\bigg{)}\geq\frac{1}{2}\bigg{\}},$ (4.48)
+$\phi(j)\coloneq\inf\bigg\{n>j:\mathbb{P}\bigg(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\bigg)\geq\frac{1}{2}\bigg\},$ (4.31)
 which concludes the proof. ∎
 We are now ready to prove part $(a)$ of Theorem 4.1, i.e.
-that almost surely no persistent $m$-hub emerges $\mathbb{P}_{\mathcal{S}}$-almost surely when the series $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely.
-Proof of Theorem 4.2 $(a)$.
+that almost surely no persistent $m$-hub emerges $\mathbb{P}_{\mathcal{S}}$-almost surely when Assumption (C-V) is not satisfied (i.e.
+when the series $\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})$ diverges almost surely.
+Proof of Theorem 4.1(a)(a).
 Fix $m\in\mathbb{N}$ and $u\in\mathcal{U}_{\infty}$.
@@
 Hence, we can write
-[…]
+$\mathbb{P}_{\mathcal{S}}\left(u\text{ is a persistent $m$-hub}\right)$ $=\mathbb{P}_{\mathcal{S}}\big(\{u\text{ is a persistent $m$-hub}\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big)$ (4.32)
 $=\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\mathbb{P}\big(\{u\text{ is a persistent $m$-hub}\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big),$
 where the last step follows from the fact that the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$ implies survival.
-We then use that, upon the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$, $u$ can be a persistent $m$-hub only when for some $T\geq\mathcal{B}(u)$ it has a larger offspring than all but at most $m-1$ many of its children for all time $t\geq T$.
+We then use that, upon the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$, $u$ can be a persistent $m$-hub only when for some $T\geq\mathcal{B}(u)$ it has a larger offspring than all but at most m1m-1 many of its children for all time $t\geq T$.
 That is,
-[…]
+$\{u\text{ is a persistent $m$-hub}\}$ $\subseteq\bigcup_{T\geq\mathcal{B}(u)}\bigcup_{\begin{subarray}{c}C\subset\mathbb{N}\\ |C|<m\end{subarray}}\{\forall j\in\mathbb{N}\setminus C\ \forall t\geq T:\deg_{t}(u)\geq\deg_{t}(uj)\}$ (4.33)
 $=\bigcup_{K\in\mathbb{N}}\bigcup_{\begin{subarray}{c}C\subset\mathbb{N}\\ |C|<m\end{subarray}}\{\forall j\in\mathbb{N}\setminus C\ \forall k\geq K:\mathcal{B}(uk)\leq\mathcal{B}(ujk)\}.$
-Here, the second step follows from the fact that any individual almost surely takes an infinite amount of time to produce an infinite offspring by assuming that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely (see Remark 4.3).
-We thus obtain
+Here, the second step follows from the fact that any individual almost surely takes an infinite amount of time to produce infinite offspring by Assumption (N-E). We thus obtain
 $\mathbb{P}_{\mathcal{S}}{}$ $(u\text{ is a persistent hub})$ (4.34)
-[…]
-We now show that the second probability on the right-hand side equals zero, by showing that
-($\mathbb{P}\big{(}\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}=\mathbb{P}\big{(}\mathcal{B}(u)<\infty,D^{(u)}=\infty\big{)}.$ (4.52)
+$\leq{}$ $\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\sum_{K=1}^{\infty}\sum_{\begin{subarray}{c}C\subset\mathbb{N}\\ |C|<m\end{subarray}}\mathbb{P}\big(\{\forall j\in\mathbb{N}\setminus C\ \forall k\geq K:\mathcal{B}(uk)\leq\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big).$
+We now show that each probability on the right-hand side equals zero, by showing that
+$\mathbb{P}\big(\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big)=\mathbb{P}\big(\mathcal{B}(u)<\infty,D^{(u)}=\infty\big).$ (4.35)
 On the event $\{\mathcal{B}(u)<\infty\}$, we can write
-$\mathcal{B}(uk)>\mathcal{B}(ujk)\ \ \Leftrightarrow\ \ \sum_{i=1}^{k}X(ui)>\sum_{i=1}^{j}X(ui)+\sum_{i=1}^{k}X(uji)\ \ \Leftrightarrow\ \ \sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji).$ (4.53)
+$\mathcal{B}(uk)>\mathcal{B}(ujk)\ \ \Leftrightarrow\ \ \sum_{i=0}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{j-1}E^{(u)}_{i}+\sum_{i=0}^{k-1}E^{(uj)}_{i}\ \ \Leftrightarrow\ \ \sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}$.
+(4.36)
 We now recall the function $\phi$ in part $(c)$ of Lemma 4.7.
-We let $(j_{\ell})_{\ell\in\mathbb{N}}=j_{\ell}(C,K)$ be the subsequence $j_{1}=1+\max\{\max C,K\}$ and $j_{\ell+1}:=\phi(j_{\ell})$ for all $\ell\in\mathbb{N}$.
+We let $(j_{\ell})_{\ell\in\mathbb{N}}=j_{\ell}(C,K)$ be the subsequence $j_{1}=1+\max\{\max C,K\}$ and $j_{\ell+1}\coloneq\phi(j_{\ell})$ for all $\ell\in\mathbb{N}$.
 As $\phi(j)>j$ for all $j\in\mathbb{N}$, it follows that the subsequence $(j_{\ell})_{\ell\in\mathbb{N}}$ is strictly increasing and thus $j_{\ell}>\max C$ and $j_{\ell}>K$ for all $\ell\in\mathbb{N}$.
 Then,
-$\mathbb{P}{}$ ($\big{(}\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}$ (4.54)
-=($=\mathbb{P}\bigg{(}\bigg{\{}\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\bigg{\}}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg{)}$
-($\geq\mathbb{P}\bigg{(}\bigg{\{}\exists\ell\in\mathbb{N}\,\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}:\!\!\!\sum_{i=j_{\ell}+1}^{k}\!\!\!X(ui)>\sum_{i=1}^{k}X(uj_{\ell}i)\bigg{\}}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg{)}.$
+$\mathbb{P}{}$ $\big(\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big)$ (4.37)
+$=\mathbb{P}\bigg(\bigg\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\bigg\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg)$
+$\geq\mathbb{P}\bigg(\bigg\{\exists\ell\in\mathbb{N}\,\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}\colon\sum_{i=j_{\ell}}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj_{\ell})}_{i}\bigg\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg).$
 For ease of writing, we define
-$\mathcal{E}(u,\ell):=\bigg{\{}\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}:\sum_{i=j_{\ell}+1}^{k}X(ui)>\sum_{i=1}^{k}X(uj_{\ell}i)\bigg{\}}\qquad\text{for }\ell\in\mathbb{N}.$ (4.55)
+[…]
 We then arrive at the lower bound
-($\mathbb{P}\big{(}\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}.$ (4.56)
+$\mathbb{P}\big(\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big).$ (4.39)
 By the choice of the sub-sequence $(j_{\ell})_{\ell\in\mathbb{N}}$, the events $(\mathcal{E}(u,\ell))_{\ell\in\mathbb{N}}$ are mutually independent, as they depend on distinct (and thus independent) inter-birth times.
 They are also independent of $\mathcal{B}(u)$ and $D^{(u)}$, so that we can factorize the probability into
-($\mathbb{P}\big{(}\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\big{)}\mathbb{P}\big{(}\mathcal{B}(u)<\infty,D^{(u)}=\infty\big{)}.$ (4.57)
+$\mathbb{P}\big(\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\big)\mathbb{P}\big(\mathcal{B}(u)<\infty,D^{(u)}=\infty\big).$ (4.40)
 By $(c)$ of Lemma 4.7 we have that $\mathbb{P}\!\left(\mathcal{E}(u,\ell)\right)\geq\frac{1}{2}$ for all $\ell\in\mathbb{N}$.
@@
 Using this in (4.34) implies that
-[…]
+$\mathbb{P}_{\mathcal{S}}\left(u\text{ is a persistent $m$-hub}\right)=0,$ (4.41)
 so that a countable union bound yields
-[…]
+$\mathbb{P}_{\mathcal{S}}\left(\exists u\in\mathcal{U}_{\infty}:u\text{ is a persistent $m$-hub}\right)=0,$ (4.42)
 which concludes the proof. ∎
-We then move to part $(b)$ of Theorem 4.2 and Theorem 4.4. To this end, we define for $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u)\leq\mathcal{B}(v)<\infty$, the event
+4.4. Proving Theorem 4.1(b)
+We then move to part $(b)$ of Theorem 4.1. To this end, we define for $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u)\leq\mathcal{B}(v)<\infty$, the event
 $\{v\text{ catches up to }u\}=\{\exists j\in\mathbb{N}:\mathcal{B}(vj)\leq\mathcal{B}(uj)\}.$ (4.43)
 That is, for an individual $v$ that is born at the same time as or after $u$, we say that $v$ catches up to $u$ if at some point its offspring is at least as large as that of $u$ (not including having zero children).
-Heuristically, one would expect that individuals that are born ‘late’ in the branching process are unlikely to catch up to the root (if the root were to produce an infinite offspring).
+Heuristically, one would expect that individuals that are born ‘late’ in the branching process are unlikely to catch up to the root (if the root were to produce infinite offspring).
 The following result makes this heuristic precise.
 Proposition 4.8.
-Let the inter-birth times $(X(i))_{i\in\mathbb{N}}$ and the offspring random variable $D$ satisfy Assumption 4.1 and suppose that there exist $\lambda>0$ and $K\in\mathbb{N}$ such that (PHC) is satisfied.
-Then,
+Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7).
+Assume that Assumptions (N-E), (F-L), and (C-V) are satisfied and Assumption (F-D) is not satisfied.
+Then, there exists $K\in\mathbb{N}$ such that
 $\mathbb{E}_{\mathcal{S}}\left[|\{v\in\mathcal{U}_{\infty}:v_{1}\geq K,D^{(v)}=\infty,v\text{ catches up to }\varnothing\}|\mathbbm{1}_{\{D_{\varnothing}=\infty\}}\right]<\infty.$ (4.44)
@@
 That is, we fix $v=v_{1}v_{2}\cdots v_{m}\in\mathcal{U}_{\infty}$ for some $m\in\mathbb{N}$, $v_{1}\geq K$, and $v_{2},\ldots v_{m}\in\mathbb{N}$, and consider
-$\mathbb{P}_{\mathcal{S}}{}$ $(v\text{ catches up to }\varnothing\text{ and }D^{(v)}=D^{(\varnothing)}=\infty)$ (4.62)
-=($=\mathbb{P}\!\left(\{D^{(v)}=D^{(\varnothing)}=\infty\}\cap\{\mathcal{B}(v)<\infty\}\cap\{\exists j\in\mathbb{N}_{0}:\mathcal{B}(v(v_{1}+j))\leq\mathcal{B}(v_{1}+j)\}\right).$
-We note that we can omit the conditioning on survival, since the event that $\varnothing$ produces an infinite offspring guarantees survival.
+[…]
+$=\mathbb{P}\!\left(\{D^{(v)}=D^{(\varnothing)}=\infty\}\cap\{\mathcal{B}(v)<\infty\}\cap\{\exists j\in\mathbb{N}_{0}:\mathcal{B}(v(v_{1}+j))\leq\mathcal{B}(v_{1}+j)\}\right).$
+We note that we can omit the conditioning on survival, since the event that $\varnothing$ produces infinite offspring guarantees survival.
 We also need only consider children numbered $v_{1}+j$ with $j\geq 0$ in the second line, since $\varnothing$ gives birth to $v_{1}$ many children before $v$ is born.
 The event $\mathcal{B}(v)<\infty$ is equivalent to $D^{(v_{1}\cdots v_{\ell})}\geq v_{\ell+1}$ for all $\ell\in[m-1]$ (the event $\{D^{(\varnothing)}=\infty\}$ already implies that $v_{1}$ is born).
-When $D^{(v)}=\infty$, the event that $v$ gives birth to its $(v_{1}+j)^{\mathrm{th}}$ child before $\varnothing$ does is equivalent to
-$\sum_{\ell=1}^{m-1}\sum_{i=1}^{v_{\ell+1}}X(v_{1}\cdots v_{\ell}i)+\sum_{i=1}^{v_{1}+j}X(vi)\leq\sum_{i=v_{1}+1}^{v_{1}+j}X(i).$ (4.63)
+When $D^{(v)}=\infty$, for $j\in\mathbb{N}_{0}$ the event that $v$ gives birth to its $(v_{1}+1+j)^{\mathrm{th}}$ child before $\varnothing$ does is equivalent to
+$\sum_{\ell=1}^{m-1}\sum_{i=0}^{v_{\ell+1}-1}E^{(v_{1}\cdots v_{\ell})}_{i}+\sum_{i=0}^{v_{1}+j}E^{(v)}_{i}\leq\sum_{i=v_{1}}^{v_{1}+j}E^{(\varnothing)}_{i}$. (4.46)
 This yields, by the independence of the offspring and inter-birth time random variables,
-$\mathbb{P}_{\mathcal{S}}{}$ $(v\text{ catches up to }\varnothing\text{ and }D^{(v)}=D^{(\varnothing)}=\infty)$ (4.64)
 […]
-We bound $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$ and apply Lemma 4.8 to arrive at the upper bound
+$={}$ $\mathbb{P}\!\left(\exists j\in\mathbb{N}_{0}\colon\sum_{\ell=1}^{m-1}\sum_{i=0}^{v_{\ell+1}-1}\!\!\!E^{(v_{1}\cdots v_{\ell})}_{i}+\!\sum_{i=0}^{v_{1}+j}\!E^{(v)}_{i}\leq\sum_{i=v_{1}}^{v_{1}+j}\!E^{(\varnothing)}_{i}\right)\mathbb{P}\!\left(D=\infty\right)^{2}\prod_{\ell=1}^{m-1}\mathbb{P}\!\left(D\geq v_{\ell+1}\right).$
+We bound $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$ and apply Lemma 4.3 (the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K=K(\lambda)$ large, see Remark 4.6(i)(i) and $(ii)$) with
+$Y=\sum_{\ell=1}^{m-1}\sum_{i=0}^{v_{\ell+1}-1}\!\!\!E^{(v_{1}\cdots v_{\ell})}_{i},\qquad X(i)=E^{(v)}_{i-1},\qquad X(i)^{\prime}=E^{(\varnothing)}_{i-1},\qquad k=v_{1}+1,$ (4.48)
+to arrive at the upper bound
 $\sum_{\begin{subarray}{c}v\in\mathcal{U}_{\infty}\\ v_{1}\geq K\end{subarray}}{}$ $\mathbb{P}_{\mathcal{S}}\left(v\text{ catches up to }\varnothing\text{ and }D_{\varnothing}=\infty\right)$ (4.49)
-[…]
-We extend the range $i\geq v_{1}+1$ in the first product to $i\geq K+1$ (as $v_{1}\geq K$) and then extend the range $v_{1}\geq K$ in the second sum to $v_{1}\geq 1$ to obtain an upper bound.
+$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K}^{\infty}\sum_{v_{2},\ldots v_{m}=1}^{\infty}\prod_{i=v_{1}}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
+We observe that $\mathbb{E}\left[\exp(\lambda(E_{i}^{\prime}-E_{i}))\right]\geq 1$ by Jensen’s inequality).
+Hence, as $v_{1}\geq K$ we can extend the range $i\geq v_{1}$ in the first product to $i\geq K$, and also extend the range $v_{1}\geq K$ in the second sum to $v_{1}\geq 1$, to obtain an upper bound.
 As the different terms in the second product depend on distinct summation indices $v_{1},\ldots,v_{m}$, we can switch the summation and product to obtain
-$\prod_{i=K+1}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\sum_{m=1}^{\infty}\bigg{(}\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathrm{e}^{-\lambda\sum_{i=1}^{j}X(i)}\right]\bigg{)}^{m}$ (4.66)
-[…]
-where we use (4.7) in the penultimate step. Under the assumptions in the proposition, it is clear that this upper bound is finite. ∎
+$\prod_{i=K}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\bigg(\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{j-1}E_{i}\bigg)\right]\bigg)^{m}$ (4.50)
+$={}$ $\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg(-\lambda\sum_{i=1}^{j}E_{i}\bigg)\right]^{m}=\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\widehat{\mu}(\lambda)^{m},$
+where we use the definition of $\widehat{\mu}$, as in (2.8), in the final step.
+We now choose $\lambda$ large enough so that $\widehat{\mu}(\lambda)<1$, which is possible by Assumption (F-L) (as follows from (4.2)).
+It is then clear that the upper bound is finite, which concludes the proof.
+∎
 We define
-$\mathrm{Pre}(u):=\{v\in\mathcal{U}_{\infty}:\mathcal{B}(v)\leq\mathcal{B}(u)\}\qquad\text{for }u\in\mathcal{U}_{\infty}$ (4.67)
+$\mathrm{Pre}(u)\coloneq\{v\in\mathcal{U}_{\infty}:\mathcal{B}(v)\leq\mathcal{B}(u)\}\qquad\text{for }u\in\mathcal{U}_{\infty}$ (4.51)
 as the set of predecessors of $u$ in the branching process $\mathrm{BP}$. Fix $m\in\mathbb{N}$. We then define
-$\mathcal{P}_{m}:={}$ $\bigcup_{\begin{subarray}{c}C\subseteq\mathrm{Pre}(u)\\ |C|<m\end{subarray}}\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{, }u\text{ catches up to all }v\in\mathrm{Pre}(u)\setminus C\}$ (4.68)
+$\mathcal{P}_{m}\coloneq{}$ $\bigcup_{\begin{subarray}{c}C\subseteq\mathrm{Pre}(u)\\ |C|<m\end{subarray}}\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{, }u\text{ catches up to all }v\in\mathrm{Pre}(u)\setminus C\}$ (4.52)
 […]
 […]
-to be the set of all individuals $u$ whose offspring diverges to infinity and at some point in time obtain a larger offspring than all but at most $m-1$ many of its predecessors.
+to be the set of all individuals $u$ whose offspring diverges to infinity and at some point in time obtain a larger offspring than all but at most m1m-1 many of its predecessors.
 Note that this point in time need not be the same for all these predecessors.
@@
 A (potential) persistent $m$-hub $u$ needs to attain the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time.
-As a result, there can be at most $m-1$ many predecessors of $u$ that it does not catch up to.
-The main ingredient for the proof of Theorems 4.2 $(b)$ and 4.4 is to ‘extend’ the result in Proposition 4.13 to the set $\mathcal{P}_{m}$, as follows.
+As a result, there can be at most m1m-1 many predecessors of $u$ that it does not catch up to.
+The main ingredient for the proof of Theorem 4.1(b)(b) is to ‘extend’ the result in Proposition 4.8 to the set $\mathcal{P}_{m}$, as follows.
 Proposition 4.9 ($\mathcal{P}_{m}$ is finite almost surely).
+Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7).
+Assume that Assumptions (N-E), (F-L), and (C-V) are satisfied and Assumption (F-D) is not satisfied.
 Fix $m\in\mathbb{N}$ and recall $\mathcal{P}_{m}$ from (4.52).
-Let the inter-birth times $(X(i))_{i\in\mathbb{N}}$ and the offspring random variable $D$ be such that Assumption 4.1 and the conditions in (PHC) are satisfied.
-Then, $|\mathcal{P}_{1}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
-If, additionally, $X(i)$ is exponentially distributed for every $i\in\mathbb{N}$, then $|\mathcal{P}_{m}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$.
-As only finitely many individuals can catch up to all but at most $m-1$ of their predecessors, and out of all pairs of these vertices one must ‘win’, as in the sense of (4.32), Theorems 4.2 $(b)$ and 4.4 then readily follow:
-Proof of Theorem 4.2 $(b)$ and Theorem 4.4 subject to Proposition 4.14.
-Let us first observe that combining the second condition in (PHC) with $(b)$ in Assumption 4.1 implies that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely.
-As a result, by Lemma 4.12 $(b)$,
+Then, $|\mathcal{P}_{m}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
+As only finitely many individuals can catch up to all but at most m1m-1 of their predecessors, and out of all pairs of these vertices one must ‘win’, as in the sense of (4.18), Theorem 4.1(b)(b) readily follows:
+Proof of Theorem 4.1(b)(b) subject to Proposition 4.9.
+Fix $m\in\mathbb{N}$. By Lemma 4.7(b)(b),
 $\mathbb{P}_{\mathcal{S}}\left(\bigcap_{\begin{subarray}{c}S\subset\mathcal{U}_{\infty}\\ |S|<\infty\end{subarray}}\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)\right)=1,$ (4.53)
 as the intersections are over countable sets. For any finite set $S$,
-$\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)=\{\exists w\in S\ \exists T\geq 0:\deg_{t}(w)=\max_{u\in S}\deg_{t}(u)\text{ for all }t\geq T\}.$ (4.70)
-As a result, with $S=\mathcal{P}_{m}$ being $\mathbb{P}_{\mathcal{S}}$-almost surely finite by Proposition 4.14, it follows that
-$\mathbb{P}_{\mathcal{S}}\left(\exists w\in\mathcal{P}_{m}\ \exists T\geq 0:\deg(w,\mathrm{BP}(t))=\max_{u\in\mathcal{P}_{m}}\deg_{t}(u)\text{ for all }t\geq T\right)=1.$ (4.71)
-Since a persistent $m$-hub attains the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time, it must be in $\mathcal{P}_{m}$.
-Hence, this proves that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely.
+$\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)\subseteq\{\exists w\in S\ \exists T\geq 0:\deg_{t}(w)=\max_{u\in S}\deg_{t}(u)\text{ for all }t\geq T\}.$ (4.54)
+As a result, with $S=\mathcal{P}_{1}$ being $\mathbb{P}_{\mathcal{S}}$-almost surely finite by Proposition 4.9, it follows that
+$\mathbb{P}_{\mathcal{S}}\left(\exists w_{1}\in\mathcal{P}_{1}\ \exists T_{1}\geq 0:\deg_{t}(w_{1})=\max_{u\in\mathcal{P}_{1}}\deg_{t}(u)\text{ for all }t\geq T_{1}\right)=1.$ (4.55)
+Let us call this individual the ‘winner’ of $\mathcal{P}_{1}$.
+Note that this individual $w_{1}$ is unique, since otherwise, if there were winners $w_{1}$ and $w_{1}^{\prime}$, then $\deg_{t}(w_{1})=\deg_{t}(w_{1}^{\prime})$ for all $t\geq\widetilde{T}$ and some time $\widetilde{T}$.
+But this would imply that infinitely many of their inter-birth times are of the same length, which occurs with probability zero.
+We can then also recursively obtain that, for any $m\geq 2$ and with $\mathcal{W}_{m}\coloneq\{w_{1},\ldots,w_{m-1}\}$,
+$\mathbb{P}_{\mathcal{S}}\left(\exists w_{m}\in\mathcal{P}_{m}\setminus\mathcal{W}_{m}\ \exists T_{m}\geq 0:\deg_{t}(w_{m})=\max_{u\in\mathcal{P}_{m}\setminus\mathcal{W}_{m}}\deg_{t}(u)\text{ for all }t\geq T_{m}\right)=1.$ (4.56)
+Now, for any $m\in\mathbb{N}$ and any $v\not\in\mathcal{P}_{m}$ there exist at least $m$ predecessors $v^{\prime}_{1},\ldots,v^{\prime}_{m}$ of $v$ such that $\deg_{t}(v^{\prime}_{i})>\deg_{t}(v)$ for all $t\geq\mathcal{B}(v^{\prime}_{i}1)$ and all $i\in[m]$ (that is, we do not include the time when both $v$ and $v_{i}^{\prime}$ have no children).
+Either $v_{i}^{\prime}\in\mathcal{P}_{m}$ for all $i\in[m]$ or there exists yet another $m$ predecessors $v^{\prime\prime}_{1},\ldots,v^{\prime\prime}_{m}$ of $v_{I}^{\prime}$ for some $I\in[m]$ such that $\deg_{t}(v_{i}^{\prime\prime})>\deg_{t}(v_{I}^{\prime})>\deg_{t}(v)$ for all $t\geq\mathcal{B}(v^{\prime\prime}_{i}1)$ and all $i\in[m]$.
+This procedure terminates, since the number of predecessors of $v$ is finite, so that there exist $u_{1},\ldots,u_{m}\in\mathcal{P}_{m}$ such that $\deg_{t}(u_{i})>\deg_{t}(v)$ for al $t\geq\mathcal{B}(u_{i}1)$ and for all $i\in[m]$.
+As a result, for any $v\not\in\mathcal{P}_{m}$ there does not exist a time $t$ such that $v$ has the $m^{\mathrm{th}}$ largest offspring among all individuals in $\mathrm{BP}(t)$.
+We now show that a persistent $m$-hub exists $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$ by induction.
+Fix first $m=1$.
+The unique winner $w_{1}$ of $\mathcal{P}_{1}=\mathcal{P}_{1}\setminus\mathcal{W}_{1}$ (with $\mathcal{W}_{1}$ the empty set) has the largest offspring among all in $\mathcal{P}_{1}$ for all $t\geq T_{1}$ by (4.55).
+Since any $v\not\in\mathcal{P}_{1}$ does not attain the largest offspring for any time $t$, it follows that $w_{1}$ has the largest offspring among all individuals for all time $t\geq T_{1}$.
+This implies that a unique persistent $1$-hub $I^{(1),\mathrm{ct}}=w_{1}$ exists $\mathbb{P}_{\mathcal{S}}$-almost surely.
+[…]
+Finally, we prove (4.1).
+Fix two individuals $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u),\mathcal{B}(v)<\infty$ and $D^{(u)}=D^{(v)}=\infty$ and consider their number of children $\deg_{t}(u)$ and $\deg_{t}(v)$.
+Suppose that with positive probability there exists an increasing and diverging sequence $(t_{k})_{k\in\mathbb{N}}$ of times, such that $\deg_{t_{k}}(u)=\deg_{t_{k}}(v)$ for all $k\in\mathbb{N}$.
+As the number of children of $u$ and $v$ tend to infinity with $k$ almost surely conditionally on survival (and that both are born in the branching processes), this equality is equivalent to the existence of a sequence $(n_{\ell})_{\ell\in\mathbb{N}}$ such that
+[…]
+We can rewrite this as
+$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=0}^{n_{\ell}-1}E^{(v)}_{i}-E^{(u)}_{i}\leq E^{(u)}_{n_{\ell}}\qquad\text{for all }\ell\in\mathbb{N}.$ (4.58)
+By Lemma 4.5 and Remark 4.6 the series $\sum_{i=0}^{\infty}(E^{(v)}_{i}-E^{(u)}_{i})$ converges almost surely.
+Hence, the distribution of the random variable $\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=0}^{\infty}(E^{(v)}_{i}-E^{(u)}_{i})$ does not contain an atom on $\mathbb{R}$ by [25, Lemma 3.2].
+However, for any $\varepsilon>0$, the inequality $E^{(u)}_{n_{\ell}}<\varepsilon$ will hold for infinitely many $\ell$ almost surely by the Borel-Cantelli lemma, since
+$\sum_{i=0}^{\infty}\mathbb{P}\!\left(E_{i}\geq\varepsilon\right)\leq\sum_{i=0}^{\infty}\mathrm{e}^{-(b(i)+d(i))\varepsilon}\leq\sum_{i=0}^{K}\mathrm{e}^{-(b(i)+d(i))\varepsilon}+\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}<\infty.$ (4.59)
+Here, the final sum is finite and $K\in\mathbb{N}$ is such that $\mathrm{e}^{-(b(i)+d(i))\varepsilon}\leq(b(i)+d(i))^{-2}$ for all $i\geq K$, which is possible since $b(i)+d(i)$ tends to infinity, both due to Assumption (C-V).
+The inequalities in (4.58) thus lead to a contradiction, so that any pair of vertices satisfies that their degrees are equal for only a finite amount of time.
+This directly implies (4.1), which concludes the proof.
 ∎
 It remains to prove Proposition 4.9. We define the following (random) subset of $\mathcal{U}_{\infty}$. For $K\in\mathbb{N}$ fixed and $u\in\mathcal{U}_{\infty}$, set
-$I(u):=\begin{cases}K&\mbox{if }D^{(u)}=\infty,\\ D^{(u)}&\mbox{if }D^{(u)}<\infty\end{cases}$ (4.72)
+$I(u)\coloneq\begin{cases}K&\mbox{if }D^{(u)}=\infty,\\ D^{(u)}&\mbox{if }D^{(u)}<\infty\end{cases}$ (4.60)
 Then, we define
-$\mathcal{U}_{K}^{<\infty}=\mathcal{U}_{K}^{<\infty}((D^{(u)})_{u\in\mathcal{U}_{\infty}}):=\{u\in\mathcal{U}_{\infty}:u_{j}\leq I(u_{1}\cdots u_{j-1})\text{for all }j\in[|u|]\}\cup\{\varnothing\}.$ (4.73)
+$\mathcal{U}_{K}^{<\infty}=\mathcal{U}_{K}^{<\infty}((D^{(u)})_{u\in\mathcal{U}_{\infty}})\coloneq\{u\in\mathcal{U}_{\infty}:u_{j}\leq I(u_{1}\cdots u_{j-1})\text{ for all }j\in[|u|]\}\cup\{\varnothing\}.$ (4.61)
 That is, given $(D^{(u)})_{u\in\mathcal{U}_{\infty}}$, one constructs the set $\mathcal{U}_{K}^{<\infty}$ iteratively as follows.
-Initialise $\mathcal{U}^{(0)}:=\{\varnothing\}$ and, given $\mathcal{U}^{(i)}$ with $i\in\mathbb{N}_{0}$, we set
-$\mathcal{U}^{(i+1)}:=\mathcal{U}^{(i)}\cup\bigcup_{u\in\mathcal{U}^{(i)}}\{u1,\ldots,uI(u)\}.$ (4.74)
+Initialise $\mathcal{U}^{(0)}\coloneq\{\varnothing\}$ and, given $\mathcal{U}^{(i)}$ with $i\in\mathbb{N}_{0}$, we set
+$\mathcal{U}^{(i+1)}\coloneq\mathcal{U}^{(i)}\cup\bigcup_{u\in\mathcal{U}^{(i)}}\{u1,\ldots,uI(u)\}.$ (4.62)
 Note that, if $u\in\mathcal{U}^{(i)}$ for some $i$ and $D^{(u)}=0$, then we do not add any children $uj$ to $\mathcal{U}^{(i+1)}$.
 This iterative process either never terminates, in which case $\mathcal{U}^{<\infty}_{K}=\cup_{i=0}^{\infty}\mathcal{U}^{(i)}$ is infinite, or it terminates at some (random) step $M\in\mathbb{N}_{0}$, so that $\mathcal{U}^{<\infty}_{K}=\mathcal{U}^{(M)}$ is finite.
-We observe that, conditionally on survival of the process $(\mathrm{BP}(t))_{t\geq 0}$, the set $\mathcal{U}^{<\infty}_{K}$ contains at least one individual $u$ such that $D^{(u)}=\infty$, almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$ by Assumption 4.1.
-Proof of Proposition 4.14 for $m=1$.
-Recall $K$ from (PHC). We define
-$T=T(K):=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}|\geq 1\}$ (4.75)
-to be the stopping time at which the branching process $\mathrm{BP}$ contains at least one individual that has produced $K$ many children and that satisfies that their ‘eventual’ offspring is infinite.
-We let $V$ denote this individual.
-In case there are more than one such individual who have produced $K$ children by time $T$, then choose the first one according to their lexicographical ordering.
-We observe that conditionally on survival such an individual $V$ exists and that $T$ is finite, almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$.
-We can then write
-$\mathcal{P}_{1}\subseteq\mathrm{BP}(T)\cup\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\ \mathcal{B}(u)>\mathcal{B}(T),\text{ $u$ catches up to all individuals in }\mathrm{Pre}(u)\}.$ (4.76)
-As $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, it follows that $|\mathrm{BP}(T)|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely by Lemma 4.7.
-It remains to show that the second set on the right-hand side has a finite size almost surely.
-To this end, we define
-$\mathcal{A}^{-}:=\mathcal{U}^{<\infty}_{K}\cap\mathrm{BP}(T)\qquad\text{and}\qquad\mathcal{A}^{+}:=\bigcup_{u\in A^{-}}\{u1,\ldots,uI(u)\}\setminus\mathcal{A}^{-},$ (4.77)
-where we recall $I(u)$ from (4.72).
-In words, $\mathcal{A}^{-}$ consists of all individuals $u$ born up to time $T$ where, for each individual $u$ such that $D^{(u)}=\infty$, its sub-tree is restricted to the first $K$ children and its descendants (if they are born by time $T$).
-Then, $\mathcal{A}^{+}$ consists of the individuals that are not yet born by time $T$ among the first $I(u)$ children of individuals $u\in\mathcal{A}^{-}$.
-With these sets at hand, we define the following events. For $u\in\mathcal{A}^{-}$, set
-$\mathcal{C}_{u}:=\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,\text{ $uv$ catches up to $u$}\},$ (4.78)
-and for $u\in\mathcal{A}^{+}$ we also define
-$\mathcal{D}^{(u)}:=\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,\text{ $uv$ catches up to $V$}\}.$ (4.79)
-With these events at hand, we observe that
-$\{u\in\mathcal{U}_{\infty}:{}$ $D^{(u)}=\infty,\ \mathcal{B}(u)>\mathcal{B}(T),\text{ $u$ catches up to all individuals in }\mathrm{Pre}(u)\}$ (4.80)
-$\subseteq\bigg{(}\bigcup_{\begin{subarray}{c}u\in\mathcal{A}^{-}\\ D^{(u)}=\infty\end{subarray}}\mathcal{C}_{u}\bigg{)}\cup\bigg{(}\bigcup_{u\in\mathcal{A}^{+}}\mathcal{D}^{(u)}\bigg{)}.$
-Indeed, any individual $w$ born after time $T$ satisfies that it has an ancestor in $\mathcal{A}^{+}$ or, if not, it has a ancestor $u\in\mathcal{A}^{-}$ such that $D^{(u)}=\infty$ and $w=uv$ with $v_{1}>K$.
-Then, any such individual $w$ that catches up to all its predecessors and that has an ancestor in $\mathcal{A}^{+}$ needs to catch up to $V$.
-If it does not have an ancestor in $\mathcal{A}^{+}$ but has an ancestor $u\in\mathcal{A}^{-}$ with $D^{(u)}=\infty$, it needs to catch up to $u$.
-We now show that each of these sets has a finite size $\mathbb{P}_{\mathcal{S}}$-almost surely.
-First, it is clear that $\mathcal{A}^{-}\subseteq\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, as already mentioned above.
-Since for each individual in $\mathcal{A}^{-}$ we include only a finite number of children in $A^{+}$, it follows that $\mathcal{A}^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely as well.
-It thus suffices to show that $|\mathcal{C}_{u}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $u\in\mathcal{A}^{-}$ such that $D^{(u)}=\infty$, and that $|\mathcal{D}^{(u)}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $u\in\mathcal{A}^{+}$.
-The former claim follows directly from Proposition 4.13 by replacing $D^{(\varnothing)}$ with $D^{(u)}$ and ‘$v$ catches up to $\varnothing$’ by ‘$uv$ catches up to $u$’ in the statement. As a result,
+We observe that, conditionally on survival of the process $(\mathrm{BP}(t))_{t\geq 0}$, the set $\mathcal{U}^{<\infty}_{K}$ contains at least one individual $u$ such that $D^{(u)}=\infty$, almost surely, when $\mathbb{P}\!\left(D=\infty\right)>0$ (which is the case when Assumption (F-D) is satisfied).
+We also define
+$\mathcal{U}^{<\infty}\coloneq\mathcal{U}_{\infty}^{<\infty}=\{u\in\mathcal{U}_{\infty}\colon u_{j}\leq D^{(u)}\text{ for all }j\in[|u|]\}\cup\{\varnothing\}$ (4.63)
+as the set of all individuals that are eventually born in $\mathrm{BP}$.
+We now prove Proposition 4.8 for $m=2$.
+The case $m=1$ follows similarly, but requires fewer steps, whilst the difference for $m>2$ mainly lies in more notation and careful bookkeeping.
+As such, we decided to prove only the case $m=2$ to aid the reader.
+We provide a brief discussion after the proof where changes are necessary when $m=1$ or $m>2$.
+Proof of Proposition 4.9 for $m=2$.
+We first note that $\mathcal{P}_{2}\subseteq\mathcal{U}^{<\infty}$, since any individual that catches up to all its predecessors clearly needs to be born.
+We first construct a superset of $\mathcal{U}^{<\infty}$, which then informs a superset for $\mathcal{P}_{2}$.
+Recall $K$ from Proposition 4.8.
+We define
+$T=T(K)\coloneq\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{ and }\mathcal{B}(uK)\leq t\}|\geq 2\}$ (4.64)
+to be the stopping time at which the branching process contains at least two individuals that have produced (at least) $K$ many children and whose offspring $D$ is infinite.
+Let $V_{1}$ and $V_{2}$ denote these two individuals, where we assume without loss of generality that $\mathcal{B}(V_{1}K)\leq\mathcal{B}(V_{2}K)=T$.
+Note that, conditionally on survival, such individuals exist and that $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$ by Assumption (F-D) (see [23, Lemma 5.3 and Corollary 6.4]).
+We then define
+$\mathcal{A}_{-}\coloneq\mathcal{U}_{K}^{<\infty}\cap\mathrm{BP}(T)\quad\text{and}\quad\mathcal{A}_{+}\coloneq\bigg(\bigcup_{u\in\mathcal{A}_{-}}\{u1,\ldots,uI(u)\}\bigg)\setminus\mathcal{A}_{-},$ (4.65)
+where we recall $I(u)$ from (4.60) and $\mathcal{U}_{K}^{<\infty}$ from (4.61).
+We think of $\mathcal{A}_{+}$ as the ‘boundary’ of $\mathcal{A}_{-}$ restricted to $\mathcal{U}_{K}^{<\infty}$.
+That is, $\mathcal{A}_{+}$ consists of all individuals in $\mathcal{U}_{K}^{<\infty}$ that have not been born by time $T$ but whose parent has been born by time $T$.
+We then have
+$\mathcal{U}^{<\infty}\subseteq\mathcal{A}_{-}\cup\bigcup_{\begin{subarray}{c}u\in\mathcal{A}_{-}\\ D^{(u)}=\infty\end{subarray}}\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}\cup\bigcup_{u\in\mathcal{A}_{+}}\{uv\colon v\in\mathcal{U}_{\infty}\}.$ (4.66)
+Indeed, let $w\in\mathcal{U}^{<\infty}$.
+We show that $wj$ is in one of the sets on the right-hand side for each $j\in[D^{(w)}]$ when $D^{(w)}<\infty$ and for each $j\in\mathbb{N}$ when $D^{(w)}=\infty$.
+Since $\varnothing\in\mathcal{A}_{-}$, this recursively implies the inclusion.
+First, suppose that $w\in\mathcal{A}_{-}$.
+If $D^{(w)}<\infty$, then for any $j\in[D^{(w)}]$ the individual $wj\in\mathcal{A}_{-}$ if $wj$ is born by time $T$, or it is in $\mathcal{A}_{+}$ if it is not born by time $T$ (and so it is included in the set $\{wjv\colon v\in\mathcal{U}_{\infty}\}$).
+If $D^{(w)}=\infty$, then for any $j\in[K]$ we can repeat the same argument as in the previous case, and for any $j>K$ the individual $wj$ is in $\{wv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$.
+Second, suppose that $w\in\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ for some $u\in\mathcal{A}_{-}$ with $D^{(u)}=\infty$.
+It then directly follows that $wj\in\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ for any $j\in\mathbb{N}$ (irrespective of the value of $D^{(w)}$).
+Third, suppose that $w\in\{uv\colon v\in\mathcal{U}_{\infty}\}$ for some $u\in\mathcal{A}_{+}$.
+Again, it directly follows that $wj\in\{uv\colon v\in\mathcal{U}_{\infty}\}$ for any $j\in\mathbb{N}$ (irrespective of the value of $D^{(w)}$).
+Using this superset, we then construct a superset for $\mathcal{P}_{2}$.
+To this end, we observe that $u\in\mathcal{P}_{2}$ implies that $D^{(u)}=\infty$ and that, for any $w_{1}\neq w_{2}\in\text{Pre}(u)$, either $u$ catches up to $w_{1}$ or $u$ catches up to $w_{2}$, as $u$ catches up to all but at most one predecessor.
+Abbreviating ‘catches up to’ by c.u.t.
+and using (4.66), we thus have the inclusion
+$\mathcal{P}_{2}\subseteq\mathcal{A}_{-}{}$ $\cup\bigcup_{\begin{subarray}{c}u\in\mathcal{A}_{-}\setminus\{V_{2}\}\\ D^{(u)}=\infty\end{subarray}}\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.
+}u\text{ or }V_{2}\}$ (4.67)
 […]
-For the latter claim we use that, by its definition, any individual $u\in\mathcal{A}^{+}$ is born strictly after time $\mathcal{B}(VK)=T$.
-With $\mathcal{F}_{T}$ the $\sigma$-algebra generated by the process up to time $T$, we write for $u\in\mathcal{A}^{+}$,
-$\mathbb{E}_{\mathcal{S}}\left[|\mathcal{D}^{(u)}|\,\Big{|}\,\mathcal{F}_{T}\right]=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,\ uv\text{ catches up to }V\,\Big{|}\,\mathcal{F}_{T}\right).$ (4.82)
-As in the proof of Proposition 4.13, for each $v=v_{1}\cdots v_{m}$ with $m\in\mathbb{N}_{0},v_{1},\ldots,v_{m}\in\mathbb{N}$ we can write the probability as
+$\cup\bigcup_{u\in\mathcal{A}_{+}}\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{1}\text{ or }V_{2}\}.$
+Again, this inclusion is clear.
+By (4.66), we take all individuals in $u\in\mathcal{A}_{-}\setminus\{V_{2}\}$ such that $D^{(u)}=\infty$ and require an individual in the set $\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ to catch up to either $u$ or to $V_{2}$.
+Similarly, we require any individual in the set $\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ to catch up to either $V_{1}$ or $V_{2}$.
+Finally, for any individual in a set $\{uv\colon v\in\mathcal{U}_{\infty}\}$ for some $u\in\ \mathcal{A}_{+}$, we require they catch up to either $V_{1}$ or $V_{2}$.
+It remains to show that each of the sets on the right-hand side of the inclusion are finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
+First, since $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely it follows from Lemma 4.2 that $\mathcal{A}_{-}\subseteq\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
+As $I(u)$ is finite almost surely for each $u\in\mathcal{U}_{\infty}$, it also follows that $\mathcal{A}_{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
+We also have the inclusions
+$\{uv{}\colon{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }u\text{ or }V_{2}\}$ (4.68)
+[…]
+$\cup\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }V_{2}\},$
+and
+$\{V_{2}v\colon{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t. }V_{1}\text{ or }V_{2}\}$ (4.69)
+$\subseteq{}$ $\bigcup_{j=1}^{2}\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t. }V_{j}\},$
+and
+$\{uv{}\colon{}$ $v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{1}\text{ or }V_{2}\}$ (4.70)
+$\subseteq{}$ $\bigcup_{j=1}^{2}\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\}.$
+We thus obtain the desired result if we show that, for any $u\in\mathcal{A}_{-}$ such that $D^{(u)}=\infty$, the sets on the right-hand side of (4.68) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely, the sets on the right-hand side of (4.69) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and, for any $u\in\mathcal{A}_{+}$, the sets on the right-hand side of (4.70) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
+We show that this is indeed the case, distinguishing per equation.
+The proof for each case is similar to the proof of Proposition 4.8, but with slight, case specific, differences.
+[…]
+$\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }V_{2}\}\qquad\text{for }u\in\mathcal{A}_{-}\setminus\{V_{2}\}\text{ such that }D^{(u)}=\infty.$ (4.71)
+Let $\mathcal{F}_{T}$ be the $\sigma$-algebra generated by $(\mathrm{BP}(t);t\leq T)$.
+We note that $T$, $\mathcal{A}_{-}$, $\mathcal{A}_{+}$, $V_{1}$, and $V_{2}$ are all measurable with respect to $\mathcal{F}_{T}$.
+Take $u\in\mathcal{A}_{-}\setminus\{V_{2}\}$ with $D^{(u)}=\infty$.
+We distinguish between two cases.
+Case (i) $u\neq V_{1}$.
+Similar to the proof of Proposition 4.8, we can write for any $uv$ with $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$ and with $m\in\mathbb{N}_{0}$, $v_{1}>K$, and $v_{2},\ldots,v_{m}\in\mathbb{N}$,
+$\mathbb{P}_{\mathcal{S}}\big(uv{}$ $\text{ catches up to }V_{2}\text{ and }D^{(uv)}=\infty\,\big|\,\mathcal{F}_{T}\big)$ (4.72)
+$=\mathbb{P}_{\mathcal{S}}\Bigg({}$ $D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m],$
+$\exists j\in\mathbb{N}_{0}\colon\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\sum_{i=0}^{K+j}E^{(uv)}_{i}\leq\mathcal{B}(V_{2}K)+\!\!\!\sum_{i=K}^{K+j}E^{(V_{2})}_{i}\,\Bigg|\,\mathcal{F}_{T}\Bigg).$
+Since $V_{2}$ produces its $K^{\mathrm{th}}$ child at time $T$ by the definition of $T$ and $V_{2}$, it follows that $\mathcal{B}(V_{2}K)=T$.
+Furthermore, since we suppose that $u\neq V_{1}$, we know that $u$ has produced at most $K$ children (in fact, fewer than $K$ children) by time $T$, so that $\mathcal{B}(uK)\leq T$.
+We can thus omit $\mathcal{B}(u)+\sum_{i=0}^{K-1}E^{(u)}_{i}=\mathcal{B}(uK)$ and $\mathcal{B}(V_{2}K)$ from the event to obtain the upper bound
 $\mathbb{P}_{\mathcal{S}}\Bigg({}$ $D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m],$ (4.73)
-$\exists j\in\mathbb{N}_{0}:\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell}}X(uv_{1}\ldots v_{\ell-1}i)+\sum_{i=1}^{K+j}X(uvi)\leq\mathcal{B}(VK)+\sum_{i=K+1}^{K+j}X(Vi)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
-We note that $T=\mathcal{B}(VK)$ and that $\mathcal{B}(u)>T$.
-The offspring random variables $D^{(uv_{1}\cdots v_{\ell-1})}$ and $D^{(uv)}$ are not contained in the $\sigma$-algebra $\mathcal{F}_{T}$.
-Also, as an individual $V$ that produces an infinite offspring has been born by time $T$, survival is guaranteed, so that we can omit the conditioning on the event $\mathcal{S}$.
-By the independence of the offspring and inter-birth random variables, and since $\mathcal{B}(u)>T=\mathcal{B}(VK)$, we obtain the upper bound
+$\exists j\in\mathbb{N}_{0}\colon\sum_{i=K}^{v_{1}-1}E^{(u)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\sum_{i=0}^{K+j}E^{(uv)}_{i}\leq\sum_{i=K}^{K+j}E^{(V_{2})}_{i}\,\Bigg|\,\mathcal{F}_{T}\Bigg).$
+We now observe that all random variables in the conditional probability, with the exception of $D^{(u)}$, are independent of $\mathcal{F}_{T}$.
+Since we know, however, that $D^{(u)}=\infty$ and thus $D^{(u)}\geq v_{1}$ is satisfied, the probability almost surely equals
+$\mathbb{P}_{\mathcal{S}}\Bigg({}$ $D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\},$ (4.74)
+$\exists j\in\mathbb{N}_{0}\colon\sum_{i=K}^{v_{1}-1}E^{(u)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\sum_{i=0}^{K+j}E^{(uv)}_{i}\leq\sum_{i=K}^{K+j}E^{(V_{2})}_{i}\Bigg)$.
+By bounding $\mathbb{P}_{\mathcal{S}}\left(\mathcal{E}\right)\leq\mathbb{P}\!\left(\mathcal{E}\right)/\mathbb{P}\!\left(\mathcal{S}\right)$ for any event $\mathcal{E}$, using the independence of the offspring and inter-birth random variables, and applying Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6), we can bound the probability from above by
+$\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}{}$ $\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=K}^{v_{1}-1}E_{i}\bigg)\right]$ (4.75)
+$\times\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
+Since the inter-birth random variables in the final two terms on the first line are independent, we can combine them into a single expected value.
+By also bounding $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$, we thus obtain the upper bound
+$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$ (4.76)
+We set $c=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}$. By summing over all $uv$ with $v\in\mathcal{U}_{\infty}$ such that $v_{1}>K$, we thus obtain
+$\mathbb{E}_{\mathcal{S}}\Big[{}$ $|\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,uv\text{ catches up to }V_{2}\text{ and }D^{(uv)}=\infty\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.77)
+$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K+1}^{\infty}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=1}^{v_{\ell}}E_{i}\bigg)\right]$
+$\leq c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m},$
+where in the last step we interchanged summation and products and extended the range of the sum over $v_{1}$ to all $v_{1}\in\mathbb{N}$.
+We observe that the inner sum on the last line equals $\widehat{\mu}(\lambda)$, so that by Assumption (F-L) we can take $\lambda>0$ large enough such that $\widehat{\mu}(\lambda)<1$, as follows from (4.2).
+The upper bound is thus finite, from which it follows that
+$\mathbb{P}_{\mathcal{S}}\left(\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,uv\text{ catches up to }V_{2}\text{ and }D^{(uv)}=\infty\}|<1\,\Big|\,\mathcal{F}_{T}\right)=1,\quad\mathbb{P}_{\mathcal{S}}\text{-a.s.}$ (4.78)
+Case (ii) $u=V_{1}$. In this case, let
+$\mathcal{T}_{T}(V_{1})\coloneq\{V_{1}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)\leq T\}$ (4.79)
+be the sub-tree rooted at $V_{1}$, restricted to the children $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants that are born by time $T$.
+It is clear that $\mathcal{T}_{T}(V_{1})\subset\mathrm{BP}(T)$, so that it is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and that it is $\mathcal{F}_{T}$ measurable.
+For $w\in\mathcal{T}_{T}(V_{1})$, we define
+$K(w)\coloneq\max\{0\leq k\leq D^{(w)}:wk\in\mathcal{T}_{T}(V_{1})\}$ (4.80)
+to be the ‘last’ child of $w$ that is included in $\mathcal{T}_{T}(V_{1})$. We then have the inclusion
+$\{V_{1}v:{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{2}\}$ (4.81)
+$\subseteq\mathcal{T}_{T}(V_{1})\cup\bigcup_{w\in\mathcal{T}_{T}(V_{1})}\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\}.$
+As a result, we have
+$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{2}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.82)
+$\leq{}$ $|\mathcal{T}_{T}(V_{1})|+\!\!\!\!\sum_{w\in\mathcal{T}_{T}(V_{1})}\!\!\!\!\!\!\mathbb{E}_{\mathcal{S}}\left[|\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\}|\,\Big|\,\mathcal{F}_{T}\right].$
+As already concluded, the first term on the right-hand side is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. Each expected value in the sum on the right-hand side can be written as
+$\sum_{m=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\,\Big|\,\mathcal{F}_{T}\right),$ (4.83)
+and each such probability equals
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\mathcal{B}(wK(w))+\!\!\!\!\!\sum_{i=K(w)}^{v_{1}-1}\!\!\!\!E^{(w)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(wv_{1}\cdots v_{\ell-1})}_{i}+\!\sum_{i=0}^{K+j}\!E^{(wv)}_{i}\!\leq\mathcal{B}(V_{2}K)+\!\sum_{i=K}^{K+j}\!E^{(V_{2})}_{i}$, (4.84)
+$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
+We observe that $\mathcal{B}(w(K(w)+1))=\mathcal{B}(wK(w))+E^{(w)}_{K(w)}>T=\mathcal{B}(V_{2}K)$ by the definition of $K(w)$, so that we obtain the upper bound
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}:\!\!\!\!\sum_{i=K(w)+1}^{v_{1}-1}\!\!\!\!E^{(w)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(wv_{1}\cdots v_{\ell-1})}_{i}+\!\sum_{i=0}^{K+j}\!E^{(wv)}_{i}\!\leq\sum_{i=K}^{K+j}\!E^{(V_{2})}_{i}$, (4.85)
+$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).\textbf{}$
+Only the random variables $D^{(w)}$, $K(w)$, and $V_{2}$ depend on $\mathcal{F}_{T}$.
+Also, the random variable $K(w)$ is independent of the inter-birth times $E^{(w)}_{i}$ for $i\geq K(w)+1$.
+As a result, we can apply Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6) to obtain the upper bound
+$\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}{}$ $\mathbbm{1}_{\{D^{(w)}\geq v_{1}\}}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]\prod_{i=K(w)+1}^{v_{1}-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$ (4.86)
+$\times\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
+Substituting this upper bound in (4.83), we thus arrive at the upper bound
+$\mathbb{E}_{\mathcal{S}}\Big[{}$ $|\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.87)
+$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]$
+$\times\prod_{i=K(w)+1}^{v_{1}-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
+Switching summation and products yields
+$\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]$ (4.88)
+$\times\sum_{v_{1}=K(w)+1}^{D^{(w)}}{}$ $\prod_{i=K(w)+1}^{v_{1}-1}\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m-1}.$
+As in Case (i), we can write the inner sum as $\widehat{\mu}(\lambda)$ and by Assumption (F-L) we can choose $\lambda>0$ such that $\widehat{\mu}(\lambda)<1$ (see the proof of Lemma 4.2).
+We then observe that
+$\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}-1}\!\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$ $\leq\sum_{v_{1}=K(w)+1}^{\infty}\prod_{i=K(w)+1}^{v_{1}-1}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$ (4.89)
+$\leq\mathbb{P}\!\left(D=\infty\right)^{-1}\!\!\!\prod_{i=0}^{K(w)}\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]^{-1}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\prod_{i=1}^{v-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$
+$=\mathbb{P}\!\left(D=\infty\right)^{-1}\widehat{\mu}(\lambda)\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]^{-1}.$
+This upper bound is $\mathbb{P}_{\mathcal{S}}$-almost surely finite for any $w\in\mathcal{T}_{T}$, since $K(w)\leq K$ by the definition of $T$ and $V_{1},V_{2}$ and since we can take $\lambda>0$ such that $\widehat{\mu}(\lambda)<\infty$ by Assumption (F-L).
+As a result, since $|\mathcal{T}_{T}(V_{1})|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, the upper bound in (4.82) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
+We thus conclude that sets on the left-hand side of (4.68) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$, for any $u\in\mathcal{A}_{-}\setminus\{V_{2}\}$ .
 […]
-$=\bigg{(}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\bigg{)}\mathbb{P}\!\left(\exists j\in\mathbb{N}_{0}:\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell-1}}X(uv_{1}\ldots v_{\ell-1}i)+\sum_{i=1}^{K+j}X(uvi)\leq\sum_{i=K+1}^{K+j}X(Vi)\right),$
-where the final step holds almost surely, since all random variables in the probability are independent of $\mathcal{F}_{T}$. As a result, by using Lemma 4.8 we arrive at
-$\mathbb{E}_{\mathcal{S}}\big{[}{}$ $|\mathcal{D}^{(u)}|\,|\,\mathcal{F}_{T}\big{]}$ (4.85)
-$\leq 1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\mathrm{e}^{-\lambda\sum_{i=1}^{v_{\ell}}X(i)}\right]$
-$=1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\sum_{m=1}^{\infty}\mathbb{E}\left[\sum_{j=1}^{D}\mathrm{e}^{-\lambda\sum_{i=1}^{j}X(i)}\right]^{m}$
-$=1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\frac{\widehat{\mu}(\lambda)}{1-\widehat{\mu}(\lambda)},$
-which is finite by the conditions in (PHC). As a result, since $A^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely,
-[…]
-Combining all of the above, we conclude that $\mathcal{P}_{1}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. ∎
-We provide a proof of Proposition 4.14 for $m=2$ and later discuss the necessary adaptations when $m\geq 3$.
-As these are mainly a matter of notation and careful bookkeeping, we decided to provide the proof in the case $m=2$ rather than for general $m>1$ to aid the reader.
-We recall that in the case $m\geq 2$ we work under the additional assumption that the inter-birth times are exponentially distributed.
-Proof of Proposition 4.14 for $m=2$.
-We recall $K$ from (PHC) and define
-$T:=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}|\geq 2\}$ (4.87)
-to be the stopping time at which the branching process contains at least two individuals that have produced $K$ many children and that satisfy that their ‘eventual’ offspring is infinite.
-Let $V_{1}$ and $V_{2}$ denote these two individuals, where we order them such that $\mathcal{B}(V_{1}K)<\mathcal{B}(V_{2}K)=T$.
-We observe that conditionally on survival such individuals exist and that $T$ is finite almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$.
-We now define
-$\mathcal{A}(K):=\mathcal{U}_{K}^{<\infty}\cap\mathrm{BP}(T).$ (4.88)
-We also define
-$\mathcal{A}^{+}:=\bigg{(}\bigcup_{u\in\mathrm{BP}(T)}\{u1,\ldots,uI(u)\}\bigg{)}\setminus\mathrm{BP}(T),$ (4.89)
-where we recall $I(u)$ from (4.72). In words, $\mathcal{A}^{+}$ consists of the individuals that are not yet born by time $T$ among the first $I(u)$ children of individuals $u\in\mathrm{BP}(T)$.
-If $u\in\mathcal{P}_{2}$, this implies that there are two individuals $w_{1},w_{2}\in\mathrm{Pre}(u)$ such that $u$ catches up to either $w_{1}$ or to $w_{2}$.
-Indeed, if $u$ does not catch up to both, then $u$ can not catch up to all but at most one predecessor.
-Hence, for any $u\in\mathcal{P}_{2}$ we need to identify two predecessors and $u$ needs to catch up to either one of them.
-We let c.u.t.
-be short for ‘catches up to’ for ease of writing.
-As a result, we have
-$\mathcal{P}_{2}\subseteq\mathrm{BP}(T){}$ $\cup\bigcup_{u\in\mathcal{A}^{+}}\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\}$ (4.90)
-$\cup\!\!\!\!\!\!\!\!\!\!\bigcup_{\begin{subarray}{c}u\in\mathcal{A}(K)\setminus\{V_{1}\}\\ D^{(u)}=\infty\end{subarray}}\!\!\!\!\!\!\!\!\!\!\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }u\text{ or }V_{2}\}$
-$\cup\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\}.$
-Let us take some time to argue why this inclusion holds.
-First, we observe that we can partition $\mathcal{U}_{\infty}$ into all individuals born up to time $T$, i.e.
-$\mathrm{BP}(T)$, and all individuals born after time $T$.
-The three unions on the right-hand side of (4.90) consider all individuals born after time $T$.
-Furthermore, each set in one of the two big unions, as well as the final set on the right-hand side, identifies two predecessors to which an individual $uv$ or $V_{1}v$ has to catch up.
-We thus conclude that the right-hand side contains $\mathcal{P}_{2}$.
-To show that $\mathcal{P}_{2}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, it thus suffices to show that the right-hand side is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
-To this end, we first observe that $\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, by Lemma 4.7 and the fact that $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
-It then directly follows that $\mathcal{A}(K)\subseteq\mathrm{BP}(T)$ is also finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
-Since each individual in $\mathrm{BP}(T)$ contributes only a finite number of elements to $\mathcal{A}^{+}$, it follows that $\mathcal{A}^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely as well.
-Furthermore, the sets $\mathrm{BP}(T)$, $\mathcal{A}^{+}$, and $\mathcal{A}(K)$ are all $\mathcal{F}_{T}$ measurable, where $\mathcal{F}_{T}$ is the $\sigma$-algebra generated by the process up to time $T$.
-We aim to prove that the right-hand side of (4.90) is finite conditionally on $\mathcal{F}_{T}$.
-As $\mathcal{A}^{+}$ and $\mathcal{A}(K)$ are $\mathcal{F}_{T}$ measurable and finite $\mathbb{P}_{\mathcal{S}}$ almost surely, we are hence left to argue that each of the sets in the unions is finite conditionally on $\mathcal{F}_{T}$.
-These follow from similar arguments, with minor differences only.
-Let us start with the sets
-$\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1}\text{ or}V_{2}\},$ (4.91)
-where $u$ is in $\mathcal{A}^{+}$. We prove that for any $u\in A^{+}$ and $j\in\{1,2\}$,
-[…]
-We obtain this by showing that the conditional expectation of the size of the set if finite. We write
-$\mathbb{E}_{\mathcal{S}}{}$ [|$\Big{[}|\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ catches up to }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.93)
-$=\sum_{r=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{j}\,\Big{|}\,\mathcal{F}_{T}\right).$
-We define
-$K_{j}:=\max\{k\in\mathbb{N}:\mathcal{B}(V_{j}k)\leq T\}\qquad\text{for }j\in\{1,2\},$ (4.94)
-as the last child $V_{j}$ gave birth to before time $T$.
-Note that $K_{1}\geq K$ and $K_{2}=K$ $\mathbb{P}_{\mathcal{S}}$-almost surely by definition of $V_{1}$ and $V_{2}$.
-We then write the conditional probability as
-$\mathbb{P}_{\mathcal{S}}\Bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\mathcal{B}(u)+\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{j}+n}\!\!X(uvi)\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!\!X(V_{j}i),$ (4.95)
-$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in[r]\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
-We now use that $u\in\mathcal{A}^{+}$, so that $u$ is not born by time $T$.
-At the same time, $V_{j}K_{j}$ is born by time $T$.
-That is, $\mathcal{B}(u)>T\geq\mathcal{B}(V_{j}K_{j})$.
-As a result, for $j=2$ we obtain the upper bound
-$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K+1}^{K+n}\!\!\!\!X(V_{2}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right),$ (4.96)
-where we also use that $K_{2}=K$ and that the offspring random variables $D^{(uv)}$ and $D^{(uv_{1}\cdots v_{\ell-1})}$ for $\ell\in[r]$ are independent of $\mathcal{F}_{T}$, as well as independent of the inter-birth random variables.
-Moreover, all inter-birth times in this upper bound are independent of $\mathcal{F}_{T}$ as well, since individual $V_{2}K$ is born exactly at time $T$ by definition.
-As a result, we can omit the conditioning on $\mathcal{F}_{T}$ $\mathbb{P}_{\mathcal{S}}$-almost surely to arrive at
-$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K+1}^{K+n}\!\!\!\!X(V_{2}i)\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$ (4.97)
-When $j=1$, we instead have the upper bound
-$\mathbb{P}_{\mathcal{S}}{}$ $\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{1}+n}\!\!X(uvi)\leq\mathcal{B}(V_{1}K_{1})-T+\!\!\!\!\sum_{i=K_{1}+1}^{K_{1}+n}\!\!\!\!X(V_{1}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}$ (4.98)
-×$\times\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$
-Here, the inter-birth time $X(V_{1}(K_{1}+1))$ (as well as $T$ and $K_{1}$) depends on $\mathcal{F}_{T}$.
-However, by the memoryless property of the exponential distribution, it follows that $\mathcal{B}(V_{1}K_{1})+X(V_{1}(K_{1}+1))-T\sim X(V_{1}(K_{1}+1))^{\prime}$, where $X(V_{1}(K_{1}+1))^{\prime}$ is an i.i.d.
-copy of $X(V_{1}(K_{1}+1))$, which is independent of $\mathcal{F}_{T}$.
-As a result, we arrive at the equivalent upper bound as for $j=2$ $\mathbb{P}_{\mathcal{S}}$-almost surely, namely
-$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{1}+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K_{1}+1}^{K_{1}+n}\!\!\!\!X(V_{1}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$ (4.99)
-Here, only the random variable $K_{1}$ depends on $\mathcal{F}_{T}$, all other random variables are independent of $\mathcal{F}_{T}$. Lemma 4.8 then yields for both $j=1$ and $j=2$ the upper bound
-$\prod_{i=K_{j}+1}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)},$ (4.100)
-where we recall that $K_{2}=K$ is deterministic. However, we can almost surely obtain a deterministic upper bound by using that $K_{j}\geq K$ for both $j\in\{1,2\}$, which yields
-$\prod_{i=K+1}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{K}X(i)\bigg{)}\right]\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)}.$ (4.101)
-Using this in (4.93), we arrive at
-$\mathbb{E}_{\mathcal{S}}{}$ [|$\Big{[}|\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.102)
-[…]
-×$\times\sum_{r=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)}$
-[…]
-[…]
-where we use (4.7) in the last step.
-The upper bound is finite by the conditions in (PHC).
-We thus arrive at (4.92).
-This implies that the first line of the right-hand side of (4.90) is $\mathbb{P}_{\mathcal{S}}$-almost surely finite conditionally on $\mathcal{F}_{T}$.
-We proceed with the sets on the second line.
-We show that for any $u\in\mathcal{A}(K)\setminus\{V_{1}\}$ such that $D^{(u)}=\infty$,
-[…]
-and
-[…]
-The first probability being equal to one $\mathbb{P}_{\mathcal{S}}$-almost surely follows directly from Proposition 4.13.
-Indeed, by the definition of $V_{1}$ and $V_{2}$, it follows that any individual $u\in\mathcal{A}(K)\setminus\{V_{1}\}$ has given birth to fewer than $K$ children by time $T$ if $u\neq V_{2}$ and to exactly $K$ children if $u=V_{2}$.
-Hence, the event in the conditional probability is independent of $\mathcal{F}_{T}$.
-Proposition 4.13 then yields the desired result, by replacing $\varnothing$ with $u$ and ‘$v$ catches up to $\varnothing$’ with ‘$uv$ catches up to $u$’.
-The proof of (4.104) follows in an analogous way when following the proof of Proposition 4.13, since we know that $V_{2}$ has produced $K$ children by time $T$, whereas the individuals $uv_{1}$ with $v_{1}>K$ have all not been born by time $T$.
-To conclude the proof, it remains to show that the set on the final line on of (4.90) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$. That is, for $j\in\{1,2\}$,
-[…]
-Here, individuals $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants may have been born by time $T$, so that a slightly more careful argument is required. Let us define
-$\mathcal{T}_{T}:=\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)\leq T\}$ (4.106)
-to be the sub-tree rooted at $V_{1}$ but restricted to the children $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants that are born by time $T$.
-It is clear that $\mathcal{T}_{T}\subset\mathrm{BP}(T)$, so that it is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and that it is $\mathcal{F}_{T}$ measurable.
-For $w\in\mathcal{T}_{T}$, we define
-$K(w):=\max\{0\leq k\leq D^{(w)}:wk\in\mathcal{T}_{T}\}$ (4.107)
-to be the ‘last’ child of $w$ that is included in $\mathcal{T}_{T}$. We then have the inclusion
-$\{V_{1}v:{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{j}\}$ (4.108)
-$\subseteq\bigcup_{w\in\mathcal{T}_{T}}\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{j}\}.$
-As a result, we have
-$\mathbb{E}_{\mathcal{S}}\Big{[}{}$ |$|\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.109)
-[…]
-Each expected value in the sum on the right-hand side can be written as
-$\sum_{r=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{j}\,\Big{|}\,\mathcal{F}_{T}\right),$ (4.110)
-and each such probability equals
-$\mathbb{P}_{\mathcal{S}}\bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\mathcal{B}(wK(w))+\!\!\!\!\sum_{i=K(w)+1}^{v_{1}}\!\!\!\!X(wi)+\sum_{\ell=2}^{r}\sum_{i=1}^{v_{\ell}}X(wv_{1}\cdots v_{\ell-1}i)+\!\!\!\!\sum_{i=1}^{K_{1}+n}\!\!X(wvi)$ (4.111)
-$\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!X(V_{j}i),D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,r\}\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
-We observe that
-$\mathcal{B}(wK(w))$ $<T<\mathcal{B}(w(K(w)+1))=\mathcal{B}(wK(w))+X(w(K(w)+1)),$
-and
-$\mathcal{B}(V_{j}K_{j})$ $\leq T<\mathcal{B}(V_{j}(K_{j}+1))=\mathcal{B}(V_{j}K_{j})+X(V_{j}(K_{j}+1)),$
-with equality in the second line only for $j=2$. Moreover, again using the memoryless property of the exponential distribution,
-$\mathcal{B}(wK(w))+X(w(K(w)+1))-T$ $\sim X(w(K(w)+1))^{\prime}$
-and
-$\mathcal{B}(V_{j}K_{j})+X(V_{j}(K_{j}+1))-T$ $\sim X(V_{j}(K_{j}+1))^{\prime},$
-where $X(w(K(w)+1))^{\prime}$ and $X(V_{j}(K_{j}+1))^{\prime}$ are i.i.d.
-copies of $X(w(K(w)+1))$ and $X(V_{j}(K_{j}+1))$, respectively, independent of all other random variables and also independent of $\mathcal{F}_{T}$.
-We thus obtain that the conditional probability can be bounded from above by
-$\mathbb{P}_{\mathcal{S}}\bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\sum_{i=K(w)+1}^{v_{1}}\!\!\!\!X(wi)+\sum_{\ell=2}^{r}\sum_{i=1}^{v_{\ell}}X(wv_{1}\cdots v_{\ell-1}i)+\!\!\!\!\sum_{i=1}^{K_{j}+n}\!\!X(wvi)\leq\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!X(V_{j}i),$ (4.112)
-$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in\{2,\ldots,r\}\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
-The dependence on $\mathcal{F}_{T}$ is only in the random variables $K(w)$ and $K_{j}$, all other random variables are independent of $\mathcal{F}_{T}$.
-Also using that the offspring random variables are independent of the inter-birth random variables and applying Lemma 4.8 thus yields the upper bound
-$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K_{j}+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\prod_{i=K(w)+1}^{v_{1}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.113)
-×$\times\prod_{\ell=2}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$
-Using this in (4.110) yields
-$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K_{j}+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.114)
-×$\times\sum_{r=1}^{\infty}{}$ $\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\prod_{\ell=2}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$
-Since $K_{j}\geq K$ by definition, we obtain the upper bound
-$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.115)
-×$\times\sum_{r=0}^{\infty}\Bigg{(}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v}X(i)\bigg{)}\right]\Bigg{)}^{r}$
-[…]
-where we use (4.7) in the final step. By the conditions in (PHC), the first three terms are finite. The final sum is finite when $D^{(w)}$ is finite. In the case that $D^{(w)}=\infty$, it equals
-$\sum_{v_{1}=K(w)+1}^{\infty}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ $\leq\mathbb{P}\!\left(D=\infty\right)^{-1}\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]^{-1}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\prod_{i=1}^{v}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.116)
-$=\mathbb{P}\!\left(D=\infty\right)^{-1}\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]^{-1}\widehat{\mu}(\lambda),$
-where we again use (4.7) to obtain the final equality.
-This upper bound is $\mathbb{P}_{\mathcal{S}}$-almost surely finite for any $w\in\mathcal{T}_{T}$, since $K(w)<\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
-As a result, since $\mathcal{T}_{T}$ is a set wit a $\mathbb{P}_{\mathcal{S}}$-almost surely finite size, the upper bound in (4.109) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, which yields (4.105).
-Combined with (4.104), (4.103), and (4.92), we thus conclude that
-$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{P}_{2}|<\infty\,|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (4.117)
-Since the probability measure is bounded from above, it follows that $\mathcal{P}_{2}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, which completes the proof. ∎
-Adaptations for the proof of Proposition 4.4 with $m>2$.
-We now discuss the adaptations of the proof for $m=2$ to extend to the general case $m\geq 2$.
-To show that $\mathcal{P}_{m}$ has a finite size when $m>2$, we define
-$T:=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}\geq m\},$ (4.118)
-and we let $V_{1},\ldots,V_{m}$ be the first $m$ individuals that produce $K$ children, ordered such that $\mathcal{B}(V_{1}K)<\mathcal{B}(V_{2}K)<\cdots<\mathcal{B}(V_{m}K)=T$.
-The sets $\mathcal{A}(K)$ and $\mathcal{A}^{+}$ have the same definition.
-We then have
-$\mathcal{P}_{m}\subseteq\mathrm{BP}(T){}$ $\cup\bigcup_{u\in\mathcal{A}^{+}}\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1},V_{2},\ldots,\text{ or }V_{m}\}$ (4.119)
-$\cup\bigcup_{u\in\mathcal{A}(K)\setminus\{V_{1},\ldots,V_{m-1}\}}\!\!\!\!\!\!\!\!\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{c.u.t.\ }u,\text{ or }V_{2},\ldots,V_{m}\}$
-$\cup\bigcup_{j=1}^{m-1}\{V_{j}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{j}v)>T,D^{(V_{j}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{1},\ldots,\text{ or }V_{m}\}.$
+$K_{j}\coloneq\sup\{k\in\mathbb{N}\colon\mathcal{B}(V_{j}k)\leq T\}\qquad\text{for }j\in\{1,2\}.$ (4.90)
+Here, we note that $K_{2}=K$ by definition of $T$ and $V_{2}$, and $K_{1}\geq K$. We then write
+$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.91)
+$=\sum_{m=1}^{\infty}\sum_{v_{1}=K+1}^{\infty}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\,|\,\mathcal{F}_{T}\right).$
+Since $K_{2}=K$ by definition of $T$ and $V_{1},V_{2}$, we can write each probability as
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\mathcal{B}(V_{2}K)+\sum_{i=K}^{v_{1}-1}E^{(V_{2})}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(V_{2}v_{1}\cdots v_{\ell-1})}_{i}+\!\!\sum_{i=0}^{K_{1}+j}\!E^{(V_{2}v)}_{i}\leq\mathcal{B}(V_{1}K_{1})+\!\!\sum_{i=K_{1}}^{K_{1}+j}\!E^{(V_{1})}_{i}$, (4.92)
+$D^{(V_{2}v)}=\infty,D^{(V_{2}v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\}\,\bigg|\,\mathcal{F}_{T}\bigg).$
+Since $\mathcal{B}(V_{1}K_{1})\leq\mathcal{B}(V_{2}K)=T\leq\mathcal{B}(V_{1}(K_{1}+1))$ by the definition of $T$ and $K_{1},K_{2}$, and by omitting the event that $D^{(V_{2}v)}=\infty$, we can bound this probability from above by
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\!\sum_{i=K}^{v_{1}-1}\!E^{(V_{2})}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(V_{2}v_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=1}^{K_{1}+j}\!E^{(V_{2}v)}_{i}\leq\mathcal{B}(V_{1}K_{1})-T+\!\sum_{i=K_{1}}^{K_{1}+j}\!E^{(V_{1})}_{i}$, (4.93)
+$D^{(V_{2}v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\}\,\bigg|\,\mathcal{F}_{T}\bigg).$
+Only the random variables $V_{1},V_{2},K_{1},T$, and $E^{(V_{1})}_{K_{1}}$ depend on $\mathcal{F}_{T}$.
+All other random variables are independent of $\mathcal{F}_{T}$.
+By the memoryless property of the exponential distribution and the strong Markov property, we have
+$\mathcal{B}(V_{1}K_{1})-T+E^{(V_{1})}_{K_{1}}\sim E_{K_{1}}^{\prime},$ (4.94)
+where $E_{K_{1}}^{\prime}$ is an independent copy of $E^{(V_{1})}_{K_{1}}$, which is also independent of $\mathcal{F}_{T}$. The probability thus equals
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\sum_{i=K}^{v_{1}-1}\!E^{(V_{2})}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(V_{2}v_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=1}^{K_{1}+j}\!E^{(V_{2}v)}_{i}\leq E_{K_{1}}^{\prime}+\!\sum_{i=K_{1}}^{K_{1}+j}\!E^{(V_{1})}_{i}$, (4.95)
+$D^{(V_{2}v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\}\,\bigg|\,\mathcal{F}_{T}\bigg).$
+We now apply Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6), to obtain the upper bound
+$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}{}$ $\prod_{i=K_{1}}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\bigg(\prod_{i=0}^{K_{1}}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\bigg)\mathbb{P}\!\left(D\geq v_{1}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=K}^{v_{1}-1}E_{i}\bigg)\right]$ (4.96)
+$\times\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
+Since $\mathbb{E}\left[\exp(\lambda(E_{i}^{\prime}-E_{i}))\right]\geq 1$ by Jensen’s inequality and $\mathbb{E}\left[\exp(-\lambda E_{i})\right]\leq 1$ for any $i\in\mathbb{N}$ and as $K_{1}\geq K$ by definition, we obtain an upper bound when replacing $K_{1}$ with $K$ in the first two products.
+Then, by the independence of the inter-birth times, we have
+$\bigg(\prod_{i=0}^{K}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\bigg)\mathbb{P}\!\left(D\geq v_{1}\right)\mathbb{E}\left[\exp\bigg(-\lambda\!\!\!\sum_{i=K}^{v_{1}-1}\!\!\!E_{i}\bigg)\right]=\mathbb{P}\!\left(D\geq v_{1}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{1}-1}E_{i}\bigg)\right].$ (4.97)
+The right-hand side can thus be incorporated in the product on the second line of (4.96), so that this expression (with $K_{1}$ bounded by $K$) simplifies to
+$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}{}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$ (4.98)
+Using this upper bound in (4.91) and writing $c=1/\mathbb{P}\!\left(\mathcal{S}\right)$, we thus obtain
+$\mathbb{E}_{\mathcal{S}}\Big[{}$ $|\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.99)
+$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K+1}^{\infty}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
+Switching summation and products and extending the range of the variable $v_{1}$ to $v_{1}\in\mathbb{N}$ yields the upper bound
+$c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m}.$ (4.100)
+By writing the inner sum as $\widehat{\mu}(\lambda)$ and using that, by Assumption (F-L), we can take $\lambda>0$ large enough so that $\widehat{\mu}(\lambda)<1$, as follows from (4.2), we conclude that the upper bound is finite.
+Using this in (4.69) combined with the earlier conclusion that the set $\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, we obtain
+$\mathbb{P}_{\mathcal{S}}\left(|\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\,\Big|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (4.101)
+Sets in (4.70). This proof follows a similar approach as for the sets in (4.69). By conditioning on $\mathcal{F}_{T}$, we have
+$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.102)
+$=\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\,\Big|\,\mathcal{F}_{T}\right).$
+Recalling $K_{j}$ from (4.90), observing that that $K_{1}\geq K$ and $K_{2}=K$ by definition of $T$ and $V_{1},V_{2}$, we can write each of the probabilities in (4.102) as
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists n\in\mathbb{N}_{0}\colon\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=0}^{K_{j}+n}\!E^{(uv)}_{i}\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\sum_{i=K_{j}}^{K_{j}+n}\!E^{(V_{j})}_{i}$, (4.103)
+$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
+Since $u\in\mathcal{A}_{+}$, we know that $\mathcal{B}(u)>T\geq\mathcal{B}(V_{j}K_{j})$, so that we obtain the upper bound
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists n\in\mathbb{N}_{0}\colon\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=0}^{K_{j}+n}\!E^{(uv)}_{i}\leq\mathcal{B}(V_{j}K_{j})-T+\!\!\!\sum_{i=K_{j}}^{K_{j}+n}\!E^{(V_{j})}_{i}$, (4.104)
+$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
+Only the random variables $K_{j}$, $T$, $V_{j}$, and $E^{(V_{j})}_{K_{j}}$ depend on $\mathcal{F}_{T}$ (the latter only for $j=1$).
+Since the inter-birth times are exponentially distributed and by the strong Markov property, it follows that
+$\mathcal{B}(V_{j}K_{j})+E^{(V_{j})}_{K_{j}}-T\sim E_{K_{j}}^{\prime}\qquad\text{for both }j\in\{1,2\},$ (4.105)
+where $E_{K_{j}}^{\prime}$ is an i.i.d.
+copy of $E^{(V_{j})}_{K_{j}}$ that is independent of $\mathcal{F}_{T}$.
+In fact, for $j=2$, since $K_{2}=K$ we have $\mathcal{B}(V_{2}K_{2})+E^{(V_{2})}_{K_{2}}-T=E^{(V_{2})}_{K_{2}}$, which is independent of $\mathcal{F}_{T}$.
+For both $j=1$ and $j=2$, we thus obtain
+$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists n\in\mathbb{N}_{0}\colon\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=0}^{K_{j}+n}\!E^{(uv)}_{i}\leq E_{K_{j}}^{\prime}+\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!E^{(V_{j})}_{i}$, (4.106)
+$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
+We can then apply Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6), to obtain the upper bound
+$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K_{j}}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{i=0}^{K_{j}-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\prod_{\ell=1}^{m}\Bigg(\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right]\Bigg).$ (4.107)
+Since $\mathbb{E}\left[\exp(\lambda(E_{i}^{\prime}-E_{i}))\right]\geq 1$ by Jensen’s inequality and that $\mathbb{E}\left[\exp(-\lambda E_{i})\right]\leq 1$ for any $i\in\mathbb{N}$ and as $K_{j}\geq K$ for both $j\in\{1,2\}$, we can obtain an upper bound by replacing $K_{j}$ with $K$.
+Substituting this upper bound in (4.102) and writing $c=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}$, we obtain
+$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.
+}V_{j}\}|\,\Big|\,\mathcal{F}_{T}\Big]\leq\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]$ (4.108)
+$c\times\prod_{i=0}^{K-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\prod_{\ell=1}^{m}\bigg(\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right]\bigg).$
+By switching summation and products, this equals
+$c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{i=0}^{K-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m}.$ (4.109)
+By writing the inner sum as $\widehat{\mu}(\lambda)$ and using that, by Assumption (F-L), we can take $\lambda>0$ large enough so that $\widehat{\mu}(\lambda)<1$, as follows from (4.2), we conclude that the upper bound is finite.
+Using this in (4.70), we thus have that
+$\mathbb{P}_{\mathcal{S}}\left(|\{uv{}\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.
+}V_{1}\text{ or }V_{2}\}|<\infty\,\Big|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (4.110)
+We can thus use the inclusion for $\mathcal{P}_{2}$ in (4.67), combined with having shown that all sets on the right-hand side of (4.68), (4.69), and (4.70) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$, to finally arrive at
+$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{P}_{2}|<\infty\,|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (4.111)
+By taking expectation with respect to $\mathcal{F}_{T}$, we thus arrive at the desired result and conclude the proof. ∎
+Adaptations for the proof of Proposition 4.8 with $m=1$ and $m>2$.
+When $m=1$, the proof simplifies somewhat.
+In this case, we instead define the stopping time $T$, as in (4.64), as the time at which there is at least $1$ individual $u$ such that $D^{(u)}=\infty$ that has produced at least $K$ children by time $T$.
+As the inter-birth times are exponentially distributed, it follows that this individual, say $V_{1}$, has produced exactly $K$ children at time $T$.
+The domination of $\mathcal{P}_{1}$ is similar as in (4.67), though now we no longer need the second line (as $V_{2}$ does not exist in this case).
+Furthermore, in the first line we require the individuals $uv$ to catch up to $u$ only, and in the third line to $V_{1}$ only.
+Bounding the set on the first line then directly follows from Proposition 4.8, and bounding the sets on the third line follows in an analogous way to bounding the sets on the right-hand side of (4.70).
+When $m>2$, we define $T$ as the first time that $m$ many individuals $V_{1},\ldots,V_{m}$ such that $D^{(V_{i})}=\infty$ have produced at least $K$ children.
+We bound $\mathcal{P}_{m}$ in the same way as in (4.67), namely
+$\mathcal{P}_{m}\subseteq\mathcal{A}_{-}{}$ $\cup\!\!\!\!\!\!\!\!\!\!\bigcup_{\begin{subarray}{c}u\in\mathcal{A}_{-}\setminus\{V_{2},\ldots,V_{m}\}\\ D^{(u)}=\infty\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }u\text{ or }V_{i}\text{ for some }i\in[m]\setminus\{1\}\}$ (4.112)
+$\cup\bigcup_{j=2}^{m}\{V_{j}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{i}\text{ for some }i\in[m]\}$
+$\cup\bigcup_{u\in\mathcal{A}_{+}}\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{i}\text{ for some }i\in[m]\}.$
 Proving that each of the sets on the right-hand side is finite then follows similar arguments as in the proof for $m=2$.
-4.5. Unique persistent $m$-hubs
-We conclude this section by proving Theorem 4.5.
-Proof of Theorem 4.5.
-It follows from Theorem 4.2 that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when Assumption 4.1 and (PHC) are satisfied.
-From the proof of Theorem 4.2 (using Proposition 4.14 in particular), it follows that the number of persistent $m$-hubs is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
-We assume that there is more than one persistent $m$-hub with positive probability and argue by contradiction.
-Suppose that both $u$ and $v$ are both persistent $m$-hubs.
-It follows that there exists $T>0$ such that their offspring is the same for all time $t\geq T$.
-That is, $\deg_{t}(u)=\deg_{t}(v)$ for all $t\geq T$.
-This is equivalent to there existing $N\in\mathbb{N}$ such that $\mathcal{B}(uk)=\mathcal{B}(vk)$ for all $k\geq N$.
-That is,
-$\mathcal{B}(u)+\sum_{i=1}^{k}X(ui)=\mathcal{B}(v)+\sum_{i=1}^{k}X(vi)\qquad\text{or, equivalently,}\qquad\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{k}(X(ui)-X(vi))=0.$ (4.120)
-Let us set $S_{i}:=X(ui)-X(vi)$.
-It is clear that the series $\sum_{i=1}^{\infty}S_{i}$ is finite almost surely by Assumption 4.1 and the second condition in (PHC).
-Since $\mathcal{B}(v)$ and $\mathcal{B}(u)$ can be represented as finite sums of inter-birth random variables, the equality on the right can only hold with positive probability when the distribution of $\sum_{i=1}^{\infty}S_{i}$ has an atom.
-However, the conditions in Theorem 4.5, combined with Lemma 4.9 and Remark 4.10, yield that the distribution does not have an atom.
-This contradiction thus implies that $u=v$ $\mathbb{P}_{\mathcal{S}}$-almost surely and there thus exists a unique persistent $m$-hub.
-To conclude the proof, we show (4.5).
-We know that there exists a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$.
-Suppose $u$ is the unique persistent $m$-hub and $v$ is the unique persistent $m^{\prime}$-hub, with $1\leq m<m^{\prime}<\infty$ two integers.
-We now assume that
-$\mathbb{P}_{\mathcal{S}}\left(\exists T>0\ \forall t\geq T:\deg_{t}(u)>\deg_{t}(v)\right)=0,$ (4.121)
-and aim to show a contradiction. Since $u$ and $v$ are the unique $m$-hub and $m^{\prime}$-hub, respectively, with $m<m^{\prime}$, we have
-$\mathbb{P}_{\mathcal{S}}\left(\exists T^{\prime}>0\ \forall t\geq T^{\prime}:\deg_{t}(u)\geq\deg_{t}(v)\right)$ $=1,$
-and
-$\mathbb{P}_{\mathcal{S}}\left(\exists T^{\prime\prime}>0\ \forall t\geq T^{\prime\prime}:\deg_{t}(u)=\deg_{t}(v)\right)$ $=0.$
-Combining this with our assumption, it follows that $\mathbb{P}_{\mathcal{S}}$-almost surely,
-$\mathcal{B}(uk)\leq\mathcal{B}(vk)<\mathcal{B}(u(k+1))\text{ holds for infinitely many }k\in\mathbb{N}$ (4.122)
-is satisfied. That is, $u$ and $v$ have the same number of offspring in the non-empty intervals $[\mathcal{B}(vk),\mathcal{B}(u(k+1)))$, for infinitely many $k$. This implies that
-$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{k}(X(vi)-X(ui))<X(u(k+1))\text{holds for infinitely many }k\in\mathbb{N}.$ (4.123)
-We now fix an arbitrary $\varepsilon>0$. Under the condition in (4.4), it follows that $X(u(k+1))\leq\varepsilon$ for all but finitely many $k$. Hence, the inequalities
-$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{k}(X(vi)-X(ui))<\varepsilon$ (4.124)
-hold for infinitely many $k$. Then, since the infinite sum $\sum_{i=1}^{\infty}(X(vi)-X(ui))$ is almost surely convergent under the second condition in (PHC) and $(2)$ in Assumption 4.1,
-$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{\infty}(X(vi)-X(ui))\leq\varepsilon,$ (4.125)
-with positive probability. As $\varepsilon$ is arbitrary, it follows that
-$\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{\infty}(X(vi)-X(ui))=0$ (4.126)
-holds with positive probability.
-However, combining the conditions in (4.3) with Lemma 4.9 yields that the distribution of $\sum_{i=1}^{\infty}(X(vi)-X(ui))$ has no atoms.
-Since $\mathcal{B}(v)$ and $\mathcal{B}(u)$ can be represented as finite sums of inter-birth random variables, it thus follows that (4.126) occurs with probability zero, which leads to a contradiction and concludes the proof.
-∎
 5. Persistence in the CTPAVD branching process
 In this section we prove Theorem 2.6.
-Unlike in the analysis carried out in Section 4, a persistent $m$-hub, as in the sense of (PH-$m$), does not emerge $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$, since individuals live an almost surely finite time (i.e.
+Unlike in the analysis carried out in Section 4, a persistent $m$-hub, as in the sense of (PH-$m$), does not exist $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$, since individuals live an almost surely finite time (i.e.
 $L<\infty$ a.s.).
-This requires a more direct approach where we provide precise asymptotic expansions of the random variables $O_{t}^{\mathrm{ct}}$ and $I_{t}^{(1),\mathrm{ct}}$ to show that their difference is a tight sequence of random variables (with respect to $t$), as in (3.14).
+Indeed, an almost surely finite lifetime implies that $O_{t}^{\mathrm{ct}}$ diverges to infinity with $t$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
+Since $I_{t}^{(1),\mathrm{ct}}\geq O_{t}^{\mathrm{ct}}$ for all $t$, we thus also have $I_{t}^{(1),\mathrm{ct}}\to\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
+To show instead that persistence occurs, in the sense of (P), we use a more quantitative approach where we provide precise asymptotic expansions of the random variables $O_{t}^{\mathrm{ct}}$ and $I_{t}^{(1),\mathrm{ct}}$ to show that their difference is a tight sequence of random variables (with respect to $t$), as in (3.19).
+That is, for a mapping $s\mapsto f(s)$ to be determined, we show that for any $\varepsilon>0$ small and $A>0$ large, there exist $t^{\prime},K,M>0$ such that for all $t\geq t^{\prime}$, the events
+$|O_{s}^{\mathrm{ct}}-f(s)|$ $<K\quad\text{for all }s\in[t-A,t+A],$ (5.1)
+and
+$I_{s}^{\mathrm{ct}}-O_{s}^{\mathrm{ct}}$ $<K\quad\text{for all }s\in[t-A,t+A],$ (5.2)
+occur with probability at least $1-\varepsilon$.
+We first collect several preliminary results in Section 5.1 and prove (5.1) in Section 5.2.
+We then prove several results on the offspring of ‘young’ and ‘old’ individuals in Section 5.3, to finally prove (5.2) in Section 5.4.
 5.1. Preliminary results
 We first state some preliminary results that are used throughout the remainder of this section.
-We recall the definition of the CTPAVD branching process in Section 3, and that we write $D$ for the offspring of an individual, $S_{k}$ for the time an individual needs to give birth to $k$ children, and $L:=S_{D+1}$ for the lifetime of an individual.
+We recall the definition of the CTPAVD branching process in Section 3, and that we write $D$ for the offspring of an individual, $S_{k}$ for the time an individual needs to give birth to $k$ children, and $L\coloneq S_{D+1}$ for the lifetime of an individual.
 Furthermore, we recall the sequences $\varphi_{1},\varphi_{2},$ and $\rho_{1}$ from (2.9) and define the sequence
-$\rho_{2}(k):=\sum_{i=0}^{k-1}\Big{(}\frac{d(i)}{b(i)+d(i)}\Big{)}^{2}\qquad\text{for }k\in\mathbb{N}.$ (5.1)
+$\rho_{2}(k)\coloneq\sum_{i=0}^{k-1}\Big(\frac{d(i)}{b(i)+d(i)}\Big)^{2}\qquad\text{for }k\in\mathbb{N}.$ (5.3)
 We extend the domain of $\rho_{2}$ to $\mathbb{R}_{+}$ by linear interpolation. We then have the following results.
@@
 $\mathbb{P}\!\left(D\geq k\right)=\mathrm{e}^{-\rho_{1}(k)-\mathcal{O}(1)}.$ (5.6)
+The lemma follows from an asymptotic expansion of the tail distribution of $D$, as in (2.6).
 Lemma 5.2 (Probability to survive with many children, Lemma $5.9$ in [23]).
@@
 $\mathbb{P}\!\left(D\geq k,S_{k}\leq t,S_{D+1}>t^{\prime}\right)\geq\mathrm{e}^{-x(t^{\prime}-t)}\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathbbm{1}_{\{S_{k}\leq t\}}\mathrm{e}^{x(S_{k}-t)}\right].$ (5.8)
-We then introduce the following notation. First, recall that $\mathcal{S}$ denotes the event that the branching process survives, and recall that
-$\mathbb{P}_{\mathcal{S}}\left(\cdot\right):=\mathbb{P}\!\left(\cdot\,|\,\mathcal{S}\right)\quad\text{and}\quad\mathbb{E}_{\mathcal{S}}[\cdot]:=\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ (5.7)
+Recall that $\mathcal{S}$ denotes the event that the branching process survives, and that
+$\mathbb{P}_{\mathcal{S}}\left(\cdot\right)\coloneq\mathbb{P}\!\left(\cdot\,|\,\mathcal{S}\right)\quad\text{and}\quad\mathbb{E}_{\mathcal{S}}[\cdot]\coloneq\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ (5.9)
 denote the conditional probability measure and its corresponding expected value, under the event $\mathcal{S}$. For $0\leq s<t<\infty$, we further define
-$\mathcal{B}(s,t):=\{u\in\mathcal{U}_{\infty}:\sigma_{u}\in[s,t]\}$ (5.8)
+$\mathcal{B}(s,t)\coloneq\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\in[s,t]\}$ (5.10)
 as the set of individuals born in the time interval $[s,t]$.
@@
 Moreover, with
-^$\widehat{\mathcal{R}}^{\lambda^{*}}(t):=\int_{0}^{t}\mathrm{e}^{-\lambda^{*}u}\,\mathcal{R}(\mathrm{d}u),$ (5.10)
+$\widehat{\mathcal{R}}^{\lambda^{*}}(t)\coloneq\int_{0}^{t}\mathrm{e}^{-\lambda^{*}u}\,\mathcal{R}(\mathrm{d}u),$ (5.12)
 the following are equivalent:
@@
     $W>0$ almost surely on $\mathcal{S}$.
-Proposition 5.3 is a direct result of [32, Theorem $5.4$] and [18] (see also [32, Proposition $1.1$] for a condensed version).
-We have the following useful corollary, which is a part of Corollary $6.6$ in [23].
+Proposition 5.3 is a direct result of [30, Theorem $5.4$] and [18] (see also [30, Proposition $1.1$] for a condensed version).
 Corollary 5.4 (Corollary $6.6$ in [23]).
-Suppose that $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied, and let $\mathcal{S}:=\{\lim_{t\to\infty}|\mathrm{BP}(t)|=\infty\}$ be the event that the process $\mathrm{BP}$ survives.
+Suppose that $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied, and let $\mathcal{S}\coloneq\{\lim_{t\to\infty}|\mathrm{BP}(t)|=\infty\}$ be the event that the process $\mathrm{BP}$ survives.
 Then,
@@
   2. (b)
-    Suppose that Assumption (N-E) is satisfied and that $\overline{d}:=\limsup_{i\to\infty}d(i)<\infty$.
+    Suppose that Assumption (N-E) is satisfied and that $\overline{d}\coloneq\limsup_{i\to\infty}d(i)<\infty$.
 Then, $\limsup_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\leq\overline{d}$.
 In particular, when $d$ converges to zero, then $\rho_{1}=o(\varphi_{1})$.
-Similarly, if $\underline{d}:=\limsup_{i\to\infty}d(i)>0$, then $\liminf_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\geq\underline{d}$.
+Similarly, if $\underline{d}\coloneq\limsup_{i\to\infty}d(i)>0$, then $\liminf_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\geq\underline{d}$.
 In particular, if $d$ tends to infinity, then $\varphi_{1}=o(\rho_{1})$.
@@
 In this sub-section we study the quantity $O_{t}^{\mathrm{ct}}$; the birth-time of the oldest alive individual in $\mathrm{BP}(t)$ conditionally on survival.
-To this end, we first obtain precise asymptotic distributional properties of the lifetime $L=S_{D+1}$ of individuals in the branching process, to then prove when the oldest individual alive at time $t$ is born.
+To this end, we first obtain precise asymptotic distributional properties of the lifetime $L=S_{D+1}$ of individuals in the branching process, to then prove an asymptotic expansion for $O_{t}^{\mathrm{ct}}$ as $t\to\infty$.
 5.2.1. Lifetime distribution
@@
 That is $\mathbb{P}\!\left(L>t\right)=\exp(-(d^{*}+o(1))t)$ as $t$ tends to infinity.
-The following result provides a higher-order result for the tail distribution when Assumption (C-V) is also satisfied.
+A more precise asymptotic expansion is provided in [23, Lemma $6.3$] when Assumption (C-V) is not satisfied.
+We complement this result by the following lemma, which provides a (different) higher-order expansion when Assumption (C-V) is satisfied.
 Lemma 5.6 (Third-order asymptotic lifetime distribution).
@@
 $\sup_{t\geq 0}\big|\log(\mathbb{P}\!\left(L>t\right))+d^{*}t+\mathcal{K}_{\alpha}(t)\big|<\infty.$ (5.15)
-Remark 5.7.
-This result complements [23, Lemma $6.2$] where a different third-order rescaling is provided for $\log\mathbb{P}\!\left(L>t\right)$ in the case that Assumption (C-V) is not satisfied.
 Proof.
@@
 Since $\varphi_{2}$ converges and $d$ converges, it follows that $\rho_{2}$, as in (5.3), converges.
-Furthermore, as $\varphi_{2}(k)={\rm Var}(S_{k})$ converges and $\varphi_{1}(k)=\mathbb{E}\left[S_{k}\right]$, it follows that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$.
-Lower bound. We set $k:=\lfloor\varphi_{1}^{-1}(t)\rfloor$. Then, using Lemma 5.1 and using the above observations,
+Furthermore, as $\varphi_{2}(k)={\rm Var}(S_{k})={\rm Var}(S_{k}-\varphi_{1}(k))$ converges and $\varphi_{1}(k)=\mathbb{E}\left[S_{k}\right]$, it follows that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$ by Kolmogorov’s two series theorem [19, Theorem 2.5.6].
+Lower bound. Set $k\coloneq\lfloor\varphi_{1}^{-1}(t)\rfloor$. Then, using Lemma 5.1 and the above observations,
 $\mathbb{P}\!\left(L>t\right)$ $\geq\mathbb{P}\!\left(D\geq k\right)\mathbb{P}\!\left(S_{k}>t\right)$ (5.16)
-$=\mathrm{e}^{-\rho_{1}(k)+\mathcal{O}(1)}\mathbb{P}\!\left(S_{k}>t\right)$
-$\geq\mathrm{e}^{-\rho_{1}(\varphi_{1}^{-1}(t))+\mathcal{O}(1)}\mathbb{P}\!\left(S_{k}-\varphi_{1}(k)>0)\right)$
-$=\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)+\mathcal{O}(1)}(\mathbb{P}\!\left(M_{\infty}>0\right)+o(1)),$
+$=\exp(-\rho_{1}(k)+\mathcal{O}(1))\mathbb{P}\!\left(S_{k}>t\right)$
+$\geq\exp(-\rho_{1}(\varphi_{1}^{-1}(t))+\mathcal{O}(1))\mathbb{P}\!\left(S_{k}-\varphi_{1}(k)>0\right)$
+$=\exp(-d^{*}t-\mathcal{K}_{\alpha}(t)+\mathcal{O}(1))(\mathbb{P}\!\left(M_{\infty}>0\right)+o(1)),$
 which yields the desired lower bound.
 Upper bound. We set
-$M:=\lceil(1-\varepsilon)t\rceil\qquad\text{and}\qquad k_{i}:=\lceil\varphi_{1}^{-1}(t-i)\rceil\quad\text{for }i\in\{0,\ldots,M\}.$ (5.15)
+$M=M(t)\coloneq\lceil(1-\varepsilon)t\rceil\qquad\text{and}\qquad k_{i}\coloneq\lceil\varphi_{1}^{-1}(t-i)\rceil\quad\text{for }i\in\{0,\ldots,M\}.$ (5.17)
 Further, we fix $\theta\in(d^{*},R)$ and take $\varepsilon\in(0,1)$ small enough such that $\theta(1-\varepsilon)>d^{*}$. Then, using Lemma 5.1 and Chernoff bounds,
@@
 $\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)}+\mathrm{e}^{-\theta(1-\varepsilon)t}+\sum_{i=0}^{M-1}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}.$ (5.20)
-As $\theta(1-\varepsilon)>d^{*}$ by the choice of $\theta$ and $\varepsilon$, and $\mathcal{K}_{\alpha}=o(t)$ by Lemma 5.5 $(c)$, it follows that the second term is negligible compared to the first term.
+As $\theta(1-\varepsilon)>d^{*}$ by the choice of $\theta$ and $\varepsilon$, and $\mathcal{K}_{\alpha}=o(t)$ by Lemma 5.5(c)(c), it follows that the second term is negligible compared to the first term.
 To deal with the sum, we define
-$\widetilde{\alpha}(k)=\sum_{i=0}^{k-1}\frac{|d(i)-d^{*}|}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}_{0},\quad\text{and}\quad\mathcal{K}_{\widetilde{\alpha}}(t):=\widetilde{\alpha}\big{(}\varphi_{1}^{-1}(t)\big{)}\qquad\text{for }t\geq 0,$ (5.19)
+$\widetilde{\alpha}(k)=\sum_{i=0}^{k-1}\frac{|d(i)-d^{*}|}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}_{0},\quad\text{and}\quad\mathcal{K}_{\widetilde{\alpha}}(t)\coloneq\widetilde{\alpha}\big(\varphi_{1}^{-1}(t)\big)\qquad\text{for }t\geq 0,$ (5.21)
 where we extend the domain of $\widetilde{\alpha}$ to $\mathbb{R}_{+}$ by linear interpolation. We then distinguish between two cases.
@@
 $+\sum_{i=C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t)}^{M-1}\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t-(i+1))-(\theta-d^{*})i+\mathcal{O}(1)}.$
-We note that this is possible since $\mathcal{K}_{\widetilde{\alpha}}(t)=o(t)$ (this follows in the same way as $\mathcal{K}_{\alpha}(t)=o(t)$ in Lemma 5.5 $(c)$) and $M\geq(1-\varepsilon)t$.
-Using the second part of Lemma 5.5 $(c)$ yields
+We note that this is possible since $\mathcal{K}_{\widetilde{\alpha}}(t)=o(t)$ (this follows in the same way as $\mathcal{K}_{\alpha}(t)=o(t)$ in Lemma 5.5(c)(c)) and $M\geq(1-\varepsilon)t$.
+Using the second part of Lemma 5.5(c)(c) yields
 $\mathcal{K}_{\alpha}(t-(i+1))-\mathcal{K}_{\alpha}(t)=o(i)\quad\text{for any }i\leq C^{\prime}\mathcal{K}_{\widetilde{\alpha}}(t).$ (5.24)
@@
 for some constant $C_{2}>0$. For the second sum on the right-hand side of (5.23), we bound for each index $i$ in the sum,
-$-\mathcal{K}_{\alpha}(t-(i+1))\leq\sup_{x\in[\varepsilon t,t]}|\mathcal{K}_{\alpha}(x)|\leq\sup_{x\in[\varepsilon t,t]}\mathcal{K}_{\widetilde{\alpha}}(x)=\mathcal{K}_{\widetilde{\alpha}}(t),$ (5.24)
+$\mathcal{K}_{\alpha}(t-(i+1))\leq\sup_{x\in[\varepsilon t,t]}|\mathcal{K}_{\alpha}(x)|\leq\sup_{x\in[\varepsilon t,t]}\mathcal{K}_{\widetilde{\alpha}}(x)=\mathcal{K}_{\widetilde{\alpha}}(t),$ (5.26)
 where the final step uses that $\mathcal{K}_{\widetilde{\alpha}}$ is increasing. As a result, we obtain
@@
 For $K>0$, we define
-$\ell_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-K,\qquad u_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K.$ (5.28)
+$\ell_{t}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-K,\qquad u_{t}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K.$ (5.30)
 Then, as $O_{s}^{\mathrm{ct}}$ is increasing in $s$,
@@
 For the first probability, we use Lemma 5.6 and the definition of $\ell_{t}$ to further bound this from above by
-$\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\big{(}\lambda^{*}{}$ $\ell_{t}-d^{*}(t-A-\ell_{t})-\mathcal{K}_{\alpha}(t-A-\ell_{t})+C_{1}\big{)}$ (5.31)
-$=\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\Big{(}{}$ $\mathcal{K}_{\alpha}(r(t))-\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big{)}-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\Big{)},$
-where $C_{1}>0$ is a sufficiently large constant. We use Lemma 5.5 $(c)$ to obtain that for any $K,A>0$ there exists $t^{\prime}>0$ such that
+$\exp\big(\lambda^{*}{}$ $\ell_{t}-d^{*}(t-A-\ell_{t})-\mathcal{K}_{\alpha}(t-A-\ell_{t})+C_{1}\big)$ (5.33)
+$=\exp\Big({}$ $\mathcal{K}_{\alpha}(r(t))-\mathcal{K}_{\alpha}\Big(\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big)-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\Big),$
+where $C_{1}>0$ is a sufficiently large constant. We use Lemma 5.5(c)(c) to obtain that for any $K,A>0$ there exists $t^{\prime}=t^{\prime}(K,A)>0$ such that for all $t\geq t^{\prime}$,
 $\Big|\mathcal{K}_{\alpha}\Big(\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big)-\mathcal{K}_{\alpha}\Big(\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big)\Big|<1.$ (5.34)
-That is, the left-hand side is at most $1$ for all $t\geq t^{\prime}=t^{\prime}(K,A)$. Then, Assumption $\mathcal{K}_{\alpha}$ yields that there exists a constant $C_{2}>0$ such that for all $t>0$,
+Then, Assumption $\mathcal{K}_{\alpha}$ yields that there exists a constant $C_{2}>0$ such that for all $t>0$,
 $\Big|\mathcal{K}_{\alpha}\Big(\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big)-\mathcal{K}_{\alpha}(r(t))\Big|\leq C_{2}.$ (5.35)
 We thus arrive for all $t\geq t^{\prime}$ at the upper bound
-$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)\leq\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\big{(}C_{2}+1-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\big{)}+\varepsilon/4\leq\varepsilon/2,$ (5.34)
-where we choose $K$ sufficiently large so that the first term is at most $\varepsilon/4$ in the final step.
+$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)\leq\exp\big(C_{2}+1-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\big)+\varepsilon/4\leq\varepsilon/2,$ (5.36)
+where we choose $K$ sufficiently large so that the exponential term is at most $\varepsilon/4$ in the final step.
 We note that we can first choose such a $K$ and then take $t^{\prime}=t^{\prime}(K,A)$ sufficiently large such that (5.34) holds.
@@
 For ease of writing, we define the events
-$\mathcal{E}_{t}:=\{\forall v\in\mathcal{B}(\widetilde{u}_{t},u_{t}):L^{(v)}\leq t+A-\widetilde{u}_{t}\},\qquad\mathcal{D}_{t}:=\{|\mathcal{B}(\widetilde{u}_{t},u_{t})|>m\}\quad\text{for }m\in\mathbb{N},t\geq 0.$ (5.36)
+$\mathcal{E}_{t}\coloneq\{\forall v\in\mathcal{B}(\widetilde{u}_{t},u_{t}):L^{(v)}\leq t+A-\widetilde{u}_{t}\},\qquad\mathcal{D}_{t}\coloneq\{|\mathcal{B}(\widetilde{u}_{t},u_{t})|>m\}\quad\text{for }m\in\mathbb{N},t\geq 0.$ (5.38)
 We can thus bound
@@
 Conditionally on $|\mathcal{B}(\widetilde{u}_{t},u_{t})|$ and upon the event $\mathcal{D}_{t}$, we can use the independence of the lifetimes of distinct individuals to bound the terms in the expected value from above by
-$\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{P}\!\left(\mathcal{E}_{t}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right)\leq\mathbbm{1}_{\mathcal{D}_{t}}\big{(}1-\mathbb{P}\!\left(L\geq\overline{r}(t)-\widetilde{u}_{t}\right)\big{)}^{m}\leq\exp\big{(}-m\mathbb{P}\!\left(L\geq\overline{r}(t)-\widetilde{u}_{t}\right)\big{)},$ (5.39)
+$\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{P}\!\left(\mathcal{E}_{t}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right)\leq\mathbbm{1}_{\mathcal{D}_{t}}\big(1-\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\big)^{m}\leq\exp\big(-m\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\big),$ (5.41)
 where we have omitted the indicator random variable and used that $1-x\leq\mathrm{e}^{-x}$ for $x\in\mathbb{R}$ in the last step. Combining all of the above we thus arrive at
 $\mathbb{P}_{\mathcal{S}}\left(O_{t+A}^{\mathrm{ct}}\geq u_{t}\right)\leq\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\big(-m\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\big)+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq m\right).$ (5.42)
-We choose $m=M^{-1}\exp(\lambda^{*}u_{t})$ for a large constant $M>0$ to be determined to obtain the upper bound
+We choose $m=M^{-1}\exp(\lambda^{*}u_{t})$ for a large constant $M>0$ to be determined later to obtain the upper bound
 $\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\big(-M^{-1}\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\mathrm{e}^{\lambda^{*}u_{t}}\big)+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq M^{-1}\mathrm{e}^{\lambda^{*}u_{t}}\right).$ (5.43)
@@
 where $C_{3}>0$ is a large constant.
-We again use Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$ to obtain similar bounds as in (5.32) and (5.33).
-We thus obtain, for some constants $C_{3}>0$ and $t^{\prime}=t^{\prime}(K,A)>0$ that for all $t>t^{\prime}$ we have have the upper bound
-$\mathbb{P}(\mathcal{S})^{-1}\exp\Big{(}-M^{-1}\exp\Big{(}C_{3}+(d^{*}/2+\lambda^{*})K-d^{*}A-C_{3}\Big{)}\Big{)}$ (5.43)
-bound the first two terms in the exponential from below by $-C_{4}$, uniformly in $A$ and $K$ and for all $t>t^{\prime}$, where $t^{\prime}>0$ and $C_{4}>0$ are large constants.
+We again use Lemma 5.5(c)(c) and Assumption $\mathcal{K}_{\alpha}$ to obtain similar bounds as in (5.34) and (5.35).
+We thus obtain, for some constants $C_{4}>0$ and $t^{\prime}=t^{\prime}(K,A)>0$ that for all $t>t^{\prime}$ we have have the upper bound
+$\mathbb{P}(\mathcal{S})^{-1}\exp\Big(-M^{-1}\exp\Big(C_{4}+(d^{*}/2+\lambda^{*})K-d^{*}A-C_{3}\Big)\Big).$ (5.45)
+Given $A$, we bound $C_{4}-d^{*}A-C_{3}\geq-C_{5}$ for some large constant $C_{5}=C_{5}(A,d^{*})$¿0.
 Then, we choose $K$ sufficiently large so that the entire term is at most $\varepsilon/4$.
@@
 $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(\widetilde{u}_{t},u_{t})|\leq M^{-1}\mathrm{e}^{\lambda^{*}u_{t}}\right)={}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t})|\mathrm{e}^{-\lambda^{*}u_{t}}-|\mathcal{B}(0,\widetilde{u}_{t})|\mathrm{e}^{-\lambda^{*}\widetilde{u}_{t}-\lambda^{*}K/2}\leq M^{-1}\right)$ (5.46)
-[…]
+$\leq{}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t})|\mathrm{e}^{-\lambda^{*}u_{t}}\leq\frac{2}{M}\right)$
+$+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,\widetilde{u}_{t})|\mathrm{e}^{-\lambda^{*}u_{t}}\geq\frac{1}{M}\mathrm{e}^{\lambda^{*}\frac{K}{2}}\right).$
 We can bound the first term on the right-hand side from above by
@@
 5.3. The richest individual
-In this sub-section we study $I^{\mathrm{ct}}_{t}$, the birth-time of the richest individual in $\mathrm{BP}(t)$.
+In this sub-section we provide an asymptotic expansion of $I^{\mathrm{ct}}_{t}$, the birth-time of the richest individual in $\mathrm{BP}(t)$.
 That is, the individual that is alive at time $t$ and has produced the largest offspring by time $t$.
 To this end, we show that individuals that are ‘too young’ produce ‘small’ offspring, whereas there exist ‘old’ individuals that produce large offspring.
+Combined, these then imply that $I_{t}^{\mathrm{ct}}$ is close to $O_{t}^{\mathrm{ct}}$.
 5.3.1. Young individuals have small offspring and no individual has too large offspring
-In this section we show that individuals that are alive in an interval of time centred at $t$ but born ‘too late’, that is, born too long after the oldest alive individual has been born, have too small offspring to attain the largest offspring among all alive individuals.
+We start by showing that individuals that are alive in an interval of time centred at $t$ but born ‘too late’, that is, born too long after the birth-time $O_{t}^{\mathrm{ct}}$ of the oldest alive individual (at time $t$), have too small offspring to attain the largest offspring among all alive individuals.
 Similarly, we show that there does not exist any individual that attains a ‘too large’ offspring by time $\approx t$.
 Fix $C_{1},C_{1}^{\prime}>0$ and set
-$k$ $:=\lceil\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big{)}\rceil,\qquad u_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t)),$ (5.48)
-$k^{\prime}$ $:=\lceil\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+C_{1}^{\prime}\big{)}\rceil.$
-Then, we have the following lemma.
+$k$ $\coloneq\lceil\varphi_{1}^{-1}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big)\rceil,\qquad u_{t}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t)),$ (5.50)
+$k^{\prime}$ $\coloneq\lceil\varphi_{1}^{-1}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+C_{1}^{\prime}\big)\rceil,$
+where we recall the function $r$ from Assumption $\mathcal{K}_{\alpha}$. Then, we have the following result.
 Lemma 5.8.
@@
 For any $\varepsilon,A,C_{1}>0$ there exist $C_{2},t^{\prime}>0$ such that for all $t>t^{\prime}$,
-$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)<k\right)\geq 1-\varepsilon.$ (5.49)
+$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)<k\right)\geq 1-\varepsilon.$ (5.51)
 Similarly, for any $\varepsilon,A>0$ there exists $C_{1}^{\prime},t^{\prime}>0$ such that for all $t>t^{\prime}$,
-$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)<k^{\prime}\right)\geq 1-\varepsilon.$ (5.50)
+$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)<k^{\prime}\right)\geq 1-\varepsilon.$ (5.52)
+The first result states that all individuals born ‘too late’, namely after time $u_{t}+C_{2}$, that survive up to time $s\in[t-A,t+A]$, have a degree smaller than $k$ with probability close to $1$.
+The second result tells us that no individual that survives up to time $s\in[t-A,t+A]$ has degree larger than $k^{\prime}$ with probability close to one.
 Proof.
 We bound the probability of the complement of the event of interest from above by $\varepsilon$. That is, we bound
-$\mathbb{P}_{\mathcal{S}}{}$ $\big{(}\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k\big{)}$ (5.51)
+$\mathbb{P}_{\mathcal{S}}{}$ $\big(\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)\geq k\big)$ (5.53)
 $={}$ $\mathbb{P}_{\mathcal{S}}\left(\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t):D^{(v)}\geq k,S_{k}^{(v)}\leq s-\mathcal{B}(v),S^{(v)}_{D^{(v)}+1}>s-\mathcal{B}(v)\right).$
@@
 For individuals $v\in\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, we can bound their birth-time $\mathcal{B}(v)$ from below and above by $u_{t}+C_{2}+(j-1)$ and $u_{t}+C_{2}+j$, respectively.
-Similarly we can bound $s$ from below and above by $t-A$ and $t+A$, respectively.
+Similarly we can bound $s$ from below and above by tAt-A and $t+A$, respectively.
 As a result, we bound the above probability from above by
@@
 It thus remains to bound the sum from above by $\varepsilon/2$ to arrive at the desired result.
-For any $j\in[\lceil t-u_{t}\rceil]$, we omit the part of the event that $S_{D+1}>t-A-(u_{t}+C_{2}+j)$, use the independence of $D$ and $S_{k}$ and use a Chernoff bound to arrive at the upper bound
+For any $j\in[\lceil t-u_{t}\rceil]$, we omit the part of the event that $S_{D+1}>t-A-(u_{t}+C_{2}+j)$, use the independence of $D$ and $S_{k}$ and use a Chernoff bound with $\theta>0$ to arrive at the upper bound
 $\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]\exp\big(\theta(t+A-(u_{t}+C_{2}+(j-1)))\big).$ (5.55)
 Applying the inequality $\log(1+x)\geq x-x^{2}/2$ for $x>0$, we can bound
-$-\log\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]=\sum_{i=0}^{k-1}\log\Big{(}1+\frac{\theta}{b(i)+d(i)}\Big{)}\geq\theta\varphi_{1}(k)-\frac{\theta^{2}}{2}\varphi_{2}(k).$ (5.54)
-Since $\varphi_{2}(k)$ is increasing and converges to $\varphi_{2}(\infty):=\lim_{k\to\infty}\varphi_{2}(k)$ as $k$ tends to infinity by Assumption (C-V), and by using Lemma 5.1, we finally arrive at the upper bound
+$\log\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]=\sum_{i=0}^{k-1}\log\Big(1+\frac{\theta}{b(i)+d(i)}\Big)\geq\theta\varphi_{1}(k)-\frac{\theta^{2}}{2}\varphi_{2}(k).$ (5.56)
+Since $\varphi_{2}(k)$ is increasing and converges to $\varphi_{2}(\infty)\coloneq\lim_{k\to\infty}\varphi_{2}(k)$ as $k$ tends to infinity by Assumption (C-V), and by using Lemma 5.1, we finally arrive at the upper bound
 $\mathbb{P}\!\left(D\geq k\right){}$ $\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]\exp\big(\theta(t+A-(u_{t}+C_{2}+(j-1)))\big)$ (5.57)
@@
 $+\theta\big(C_{1}-C_{2}+A-(j-1)\big)+\frac{\theta^{2}}{2}\varphi_{2}(\infty)\Big).$
-We apply Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$ and set $\theta=(C_{2}+(j-1)-(A+C_{1}))/\varphi_{2}(\infty)$ (which is positive for all $j\geq 1$ when $C_{2}$ is large), to bound this for all $t>t^{\prime}$ by
+We apply Lemma 5.5(c)(c) and Assumption $\mathcal{K}_{\alpha}$ and set $\theta=(C_{2}+(j-1)-(A+C_{1}))/\varphi_{2}(\infty)$ (which is positive for all $j\geq 1$ when $C_{2}$ is large), to bound this for all $t>t^{\prime}$ by
 $\exp\Big(-\lambda^{*}\Big(\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big)+C+d^{*}C_{1}-\frac{1}{2}\frac{(C_{2}+(j-1)-(A+C_{1}))^{2}}{\varphi_{2}(\infty)}\Big),$ (5.59)
-where $C>0$ and $t^{\prime}$ are fixed constants, where $C$ is independent of all other constants and parameters and $t^{\prime}=t^{\prime}(C_{1},d^{*},\lambda^{*})$.
+where $C>0$ and $t^{\prime}$ are fixed constants, where $C$ is independent of all other constants and parameters, and $t^{\prime}=t^{\prime}(C_{1},d^{*},\lambda^{*})$.
 By substituting this upper bound in the sum in (5.54), we can bound it from above by
@@
 5.3.2. There exists an old individual with large offspring
-The final ingredient to prove persistence is to show that there exists an ‘old’ individual that survives until time $\approx t$ and has a sufficiently large offspring.
-Combined with the result in Lemma 5.9, this implies that the birth-time of the individual with the largest offspring (i.e.
+The next step is to show that there exists an ‘old’ individual that survives until time $\approx t$ and has a sufficiently large offspring.
+Combined with Proposition 5.7 and Lemma 5.8, this implies that the birth-time of the individual with the largest offspring (i.e.
 $I_{t}^{\mathrm{ct}}$) cannot be ‘too large’ compared to $O_{t}^{\mathrm{ct}}$.
@@
 For any $\varepsilon,A>0$ there exist $C_{1},t^{\prime}>0$ such that for all $t>t^{\prime}$,
-$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k\right)\geq 1-\varepsilon.$ (5.60)
+$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)\geq k\right)\geq 1-\varepsilon.$ (5.62)
 Proof.
@@
 $\mathbb{P}_{\mathcal{S}}{}$ $\big(\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1}):D^{(v)}\geq k,S^{(v)}_{k}\leq s-\mathcal{B}(v),S^{(v)}_{D^{(v)}+1}>s-\mathcal{B}(v)\big)$ (5.63)
-[…]
+$\geq{}$ $\mathbb{P}_{\mathcal{S}}\big(\exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1}):D^{(v)}\geq k,S^{(v)}_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S^{(v)}_{D^{(v)}+1}>t+A-u_{t}\big).$
 Here we bound $s$ and $\mathcal{B}(v)$ from below and above to obtain the inequality. Similar to the steps in (5.37) through (5.42), we bound the right-hand side from below for any $m\in\mathbb{N}$ by
-[…]
+$1{}$ $-\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})|\leq m\right)$ (5.64)
 $-\mathbb{P}\!\left(\mathcal{S}\right)^{-1}\exp\Big(-m\mathbb{P}\!\left(D\geq k,S_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S_{D+1}>t+A-u_{t}\right)\Big).$
@@
 $\mathbb{P}({}$ $D\geq k,S_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S_{D+1}>t+A-u_{t})$ (5.65)
-[…]
-×$\times\exp\big{(}-(d^{*}+\eta)(2A+\tfrac{1}{2})C_{1}\big{)}.$
+$\geq{}$ $\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathbbm{1}\{S_{k}\leq t-u_{t}-A-\tfrac{1}{2}C_{1}\}\exp\big(-(d^{*}+\eta)(t-u_{t}-A-\tfrac{1}{2}C_{1}-S_{k})\big)\right]$
+$\times\exp\big(-(d^{*}+\eta)(2A+\tfrac{1}{2}C_{1})\big).$
 We bound the expected value from below by restricting the range of $S_{k}$ in the indicator to $S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]$, to obtain
 $\mathbb{E}\big[{}$ $\mathbbm{1}\{S_{k}\leq t-u_{t}-A-\tfrac{1}{2}C_{1}\}\exp\big(-(d^{*}+\eta)(t-u_{t}-A-\tfrac{1}{2}C_{1}-S_{k})\big)\big]$ (5.66)
-[…]
-[…]
+$\geq{}$ $\mathbb{E}\big[\mathbbm{1}\{S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]\}\exp\big(-(d^{*}+\eta)(t-u_{t}-A-\tfrac{1}{2}C_{1}-S_{k})\big)\big]$
+$\geq{}$ $\mathbb{P}\!\left(S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]\right)\mathrm{e}^{-\frac{1}{2}(d^{*}+\eta)C_{1}}.$
 We then use that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$ as $k$ tends to infinity, since $\varphi_{2}(k)={\rm Var}(S_{k})$ converges by Assumption (C-V).
@@
 We observe that the right-hand side is strictly positive for any choice of $A$ and $C_{1}>0$ and can thus be bounded from below by
-$c_{\infty}(A,C_{1}):=\frac{1}{2}\mathbb{P}\!\left(M_{\infty}\in[-A,-A+\tfrac{1}{2}C_{1}]\right),$ (5.67)
+$c_{\infty}(A,C_{1})\coloneq\frac{1}{2}\mathbb{P}\!\left(M_{\infty}\in[-A,-A+\tfrac{1}{2}C_{1}]\right),$ (5.69)
 for all $t$ (and thus $k$) sufficiently large. Combining this with (5.66) in (5.65) yields the lower bound
@@
 $\rho_{1}(k)={}$ $\rho_{1}\big(\varphi_{1}^{-1}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big)\big)+o(1)$ (5.71)
-[…]
-Using Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$, we can bound the right-hand side from above for all $t>t^{\prime\prime}$ by
+$={}$ $d^{*}\Big[\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\Big]+\mathcal{K}_{\alpha}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big)+o(1).$
+Using Lemma 5.5(c)(c) and Assumption $\mathcal{K}_{\alpha}$, we can bound the right-hand side from above for all $t>t^{\prime\prime}$ by
 $\lambda^{*}\Big[\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))\Big]-d^{*}C_{1}+C=\lambda^{*}u_{t}-d^{*}C_{1}+C,$ (5.72)
@@
 We use this in (5.64) and set, for some large constant $M>0$ to be determined,
-$m:=M^{-1}\exp\big{(}\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big{)},$ (5.73)
+$m\coloneq M^{-1}\exp\big(\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big),$ (5.75)
 to arrive at
-$\mathbb{P}_{\mathcal{S}}(\forall{}$ $s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k)$ (5.74)
+$\mathbb{P}_{\mathcal{S}}(\forall{}$ $s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)\geq k)$ (5.76)
 $\geq 1{}$ $-\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})|\leq M^{-1}\exp\big(\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big)\right)$
@@
 $\mathbb{P}_{\mathcal{S}}\big({}$ $|\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})|\leq M^{-1}\exp\big(\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big)$ (5.77)
-[…]
-[…]
+$={}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t}+\tfrac{1}{2}C_{1})|-|\mathcal{B}(0,u_{t})|\leq M^{-1}\exp\big(\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big)\right)$
+$\leq{}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t}+\tfrac{1}{2}C_{1})|\exp\big(-\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big)\leq 2M^{-1}\right)$
 $+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t})|\exp\big(-\lambda^{*}u_{t}\big)\geq M^{-1}\exp\big(\tfrac{\lambda^{*}}{2}C_{1}\big)\right)$
-[…]
+$\leq{}$ $\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq 2M^{-1}\right)+\mathbb{P}\!\left(\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\geq M^{-1}\exp\big(\tfrac{\lambda^{*}}{2}C_{1}\big)\right).$
 We first choose $M$ sufficiently large so that, by (5.14) in Corollary 5.4, the first term on the last line is smaller than $\varepsilon/3$.
@@
 Proof.
-Fix $\varepsilon,A>0$, $t^{\prime}>0$ large (according to Proposition 5.8, Lemma 5.9, and Lemma 5.10) and recall $k=k(C_{1})$ and $u_{t}$ from (5.48). Then, for any $t>t^{\prime}$:
+Fix $\varepsilon>0$ small, $A>0$ and $t^{\prime}>0$ large (according to Proposition 5.7, Lemma 5.8, and Lemma 5.9) and recall $k=k(C_{1})$ and $u_{t}$ from (5.50). Then, for any $t>t^{\prime}$:
   1. (a)
-    We take $K>0$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$: $O_{s}^{\mathrm{ct}}>u_{t}-K$.
+    We take $K>0$ sufficiently large so that, with probability at least $1-\varepsilon/3$, for all $s\in[t-A,t+A]$ we have $O_{s}^{\mathrm{ct}}>u_{t}-K$.
   2. (b)
-    We take $C_{1}>0$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$ $\exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}$: $d_{v}(s)\geq k$.
+    We take $C_{1}>0$ sufficiently large so that, with probability at least $1-\varepsilon/3$, for all $s\in[t-A,t+A]$ there exists $v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}$ such that $\deg_{s}(v)\geq k$.
   3. (c)
-    We take $C_{2}>\tfrac{1}{2}C_{1}$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$ $\forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}$: $d_{v}(s)<k$.
+    We take $C_{2}>\tfrac{1}{2}C_{1}$ sufficiently large so that, with probability at least $1-\varepsilon/3$, for all $s\in[t-A,t+A]$ and for all $v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}$ we have $\deg_{s}(v)<k$.
 $(a)$, $(b)$, and $(c)$ follow from Proposition 5.7, Lemma 5.9, and Lemma 5.8, respectively.
-Then, $(b)$ and $(c)$ combined imply that at $s$ the alive individual with the largest offspring has birth-time at most $u_{t}+C_{2}$, for all $s\in[t-A,t+A]$, with probability at least $1-\tfrac{2}{3}\varepsilon$.
-That is, $I_{s}^{\mathrm{ct}}<u_{t}+C_{2}$ for all $s\in[t-A,t+A]$ with probability at least $1-\tfrac{2}{3}\varepsilon$.
+Then, $(b)$ and $(c)$ combined imply that $I_{s}^{\mathrm{ct}}\leq u_{t}+C_{2}$ for all $s\in[t-A,t+A]$, with probability at least $1-\tfrac{2}{3}\varepsilon$.
 Combined with $(a)$, we thus conclude that the result follows for $M=C_{2}+K$.
 ∎
-We finally prove Theorem 2.6.
+We finally prove Theorem 2.6 by translating the results proved in this section for the CTPAVD branching process to the desired results for the discrete time PAVD model, leveraging Proposition 3.1.
 Proof of Theorem 2.6.
@@
 $\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\in\big[\tfrac{1}{\lambda^{*}}\log n-A,\tfrac{1}{\lambda^{*}}\log n+A\big]\right)\geq 1-\varepsilon/3,\qquad\text{for all }n\geq n_{0}.$ (5.80)
-We let $F_{n,A}:=\big{[}\frac{1}{\lambda^{*}}\log n-A,\frac{1}{\lambda^{*}}\log n+A\big{]}$ for ease of writing. Then,
+We let $F_{n,A}\coloneq\big[\frac{1}{\lambda^{*}}\log n-A,\frac{1}{\lambda^{*}}\log n+A\big]$ for ease of writing. Then,
 $\mathbb{P}_{\mathcal{S}}{}$ $\bigg(\Big|O_{\tau_{n}}^{\mathrm{ct}}-\frac{d^{*}}{\lambda^{*}(\lambda^{*}+d^{*})}\log n-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<M\bigg)$ (5.81)
-[…]
+$\geq{}$ $\mathbb{P}_{\mathcal{S}}\left(\forall s\in F_{n,A}:\Big|O_{s}^{\mathrm{ct}}-\frac{d^{*}}{\lambda^{*}(\lambda^{*}+d^{*})}\log n-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<M\right)$
 $-\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\not\in F_{n,A}\right).$
 The second term on the right-hand side is at most $\varepsilon/3$ for all $n\geq n_{0}$ by (5.80).
-Using Proposition 5.8 with $t=\frac{1}{\lambda^{*}}\log n$ we can take $M=M(A,\varepsilon)$ and $n_{1}=n_{1}(A,\varepsilon)\geq n_{0}$ such that the first term on the right-hand side is at least $31-\varepsilon/$ 3 for all $n\geq n_{1}$ as well, so that the entire lower bound is at least $1-\frac{2}{3}\varepsilon$.
+Using Proposition 5.7 with $t=\frac{1}{\lambda^{*}}\log n$ we can take $M=M(A,\varepsilon)$ and $n_{1}=n_{1}(A,\varepsilon)\geq n_{0}$ such that the first term on the right-hand side is at least $1-\varepsilon/3$ for all $n\geq n_{1}$, so that the entire lower bound is at least $1-\frac{2}{3}\varepsilon$.
 We then define
-$\ell_{n}:=\frac{d^{*}}{\lambda^{*}+d^{*}}\log n+\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)},$ (5.80)
+$\ell_{n}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}\log n+\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big),$ (5.82)
 and write
@@
 $\mathbb{P}_{\mathcal{S}}{}$ $\bigg(\Big|\log O_{n}-\frac{d^{*}}{\lambda^{*}+d^{*}}\log n-\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<M^{\prime}\bigg)$ (5.84)
-[…]
-[…]
-$-\mathbb{P}_{\mathcal{S}}\left(N\Big{(}\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\geq\mathrm{e}^{M^{\prime}+\ell_{n}}\right).$
-The first probability is at least $1-\frac{2}{3}\varepsilon$ by taking $M^{\prime}=2\lambda^{*}M$ and applying (5.79) with $M$ large.
+$={}$ $\mathbb{P}_{\mathcal{S}}\left(O_{n}\in\big[\mathrm{e}^{-M^{\prime}+\ell_{n}},\mathrm{e}^{M^{\prime}+\ell_{n}}\big]\right)$
+$\geq{}$ $\mathbb{P}_{\mathcal{S}}\left(O_{\tau_{n}}^{\mathrm{ct}}\in\Big[\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}},\frac{1}{\lambda^{*}}\ell_{n}+\frac{M^{\prime}}{2\lambda^{*}}\Big]\right)-\mathbb{P}_{\mathcal{S}}\left(N\Big(\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}}\Big)\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$
+$-\mathbb{P}_{\mathcal{S}}\left(N\Big(\frac{1}{\lambda^{*}}\ell_{n}+\frac{M^{\prime}}{2\lambda^{*}}\Big)\geq\mathrm{e}^{M^{\prime}+\ell_{n}}\right).$
+The first probability is at least $1-\frac{2}{3}\varepsilon$ by taking $M^{\prime}=2\lambda^{*}M$ and applying (5.81) with $M$ and $n$ sufficiently large.
 Then, since $N(t)\geq|\mathcal{B}((0,t))|$ for all $t\geq 0$ deterministically,
-$\mathbb{P}_{\mathcal{S}}\left(N\Big{(}-\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(\Big{|}\mathcal{B}\Big{(}0,-\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\Big{|}\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ (5.83)
+$\mathbb{P}_{\mathcal{S}}\left(N\Big(\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}}\Big)\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(\Big|\mathcal{B}\Big(0,\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}}\Big)\Big|\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ (5.85)
 $\leq\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq\mathrm{e}^{-M^{\prime}/2}\right),$
@@
 In a similar manner, using Corollary 5.4 and that $N(t)\leq 2|\mathcal{B}(0,t)|$, we can bound the final probability on the right-hand side of (5.84) from above by $\varepsilon/6$ as well when $M$ is large enough, to finally yield the lower bound $1-\varepsilon$, which concludes the proof of (2.20).
-We continue by proving that $I_{n}/O_{n}$ is tight. We apply Proposition 5.11 to conclude that there exist $n_{1}\geq n_{0}$ and $M=M(\varepsilon,A,n_{1})>0$ such that for all $n\geq n_{1}$,
+We continue by proving that $I_{n}/O_{n}$ is tight.
+We apply Proposition 5.10 and (5.80) to conclude that for any $\varepsilon>0$ there exist $n_{1}\geq n_{0}$ and $M=M(\varepsilon,A,n_{1})>0$ such that for all $n\geq n_{1}$,
 $\mathbb{P}_{\mathcal{S}}\left(I^{\mathrm{ct}}_{\tau_{n}}-O^{\mathrm{ct}}_{\tau_{n}}<M\right)\geq\mathbb{P}_{\mathcal{S}}\left(\forall s\in F_{n,A}:I^{\mathrm{ct}}_{s}-O^{\mathrm{ct}}_{s}<M\right)-\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\not\in F_{n,A}\right)\geq 1-\frac{2}{3}\varepsilon.$ (5.86)
@@
 $-\mathbb{P}_{\mathcal{S}}\left(N(O^{\mathrm{ct}}_{\tau_{n}})\leq C^{-1}\exp\big(\lambda^{*}O^{\mathrm{ct}}_{\tau_{n}}\big)\right).$
-We can bound the first term on the right-hand side from below by $31-2\varepsilon/$ 3 for all $n\geq n_{0}$ using (5.84) and setting $M^{\prime}:=C^{2}\mathrm{e}^{\lambda^{*}M}$.
+We can bound the first term on the right-hand side from below by $1-2\varepsilon/3$ for all $n\geq n_{1}$ using (5.86) and setting $M^{\prime}\coloneq C^{2}\mathrm{e}^{\lambda^{*}M}$.
 The second term can be bounded from above by using that $N(t)\leq 2|\mathcal{B}(0,t)|$ deterministically for all $t\geq 0$, so that
@@
 $\mathbb{P}_{\mathcal{S}}{}$ $\bigg(\Big|\varphi_{1}(\max_{v\in\mathcal{A}_{\tau_{n}}^{\mathrm{ct}}}\deg_{\tau_{n}}(v))-\frac{1}{\lambda^{*}+d^{*}}\log n+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<K\bigg)$ (5.92)
-[…]
+$\geq{}$ $\mathbb{P}_{\mathcal{S}}\bigg(\forall s\in F_{n,A}:\Big|\varphi_{1}(\max_{v\in\mathcal{A}_{s}^{\mathrm{ct}}}\deg_{s}(v))-\frac{1}{\lambda^{*}+d^{*}}\log n+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<K\bigg)$
 $-\mathbb{P}_{\mathcal{S}}\left(\tau_{n}\not\in F_{n,A}\right).$
-The final term is at most $\varepsilon/2$ for $A$ sufficiently large by (5.78).
-Then, the first probability on the right-hand side is also at least $21-\varepsilon/$ 2 by setting $C_{1}^{\prime}=K$ in Lemma 5.9 and $C_{1}=K$ in Lemma 5.10 and taking $K$ sufficiently large (with $t=\frac{1}{\lambda^{*}}\log n$ and $n$ sufficiently large).
-We thus arrive at the desired result and conclude the proof.
+The final term is at most $\varepsilon/2$ for $A$ and $n$ sufficiently large by (5.80).
+The first probability on the right-hand side is also at least $1-\varepsilon/2$ by setting $C_{1}^{\prime}=K$ in Lemma 5.8 and $C_{1}=K$ in Lemma 5.9 and taking $K$ sufficiently large (with $t=\frac{1}{\lambda^{*}}\log n$ and $n$ sufficiently large).
+We thus arrive at the desired result, which concludes the proof.
 ∎
@@
 Still, there are several further extensions and open problems that we would like to state here.
-Existence of persistent $m$-hubs for more general CMJ branching processes.
-In Section 4 we studied the existence of persistent $m$-hubs in CMJ branching processes.
-Here, we were able to provide general conditions for the non-existence of persistent $m$-hubs, but could only prove the existence of persistent $1$-hubs (i.e.
-a persistent maximum degree) under general conditions.
-For persistent $m$-hubs with $m\geq 2$ we required the inter-birth times to be exponentially distributed, as the proofs leveraged the memoryless property at a few places.
-This leads to the following open problem.
+Characterising the difference between the richest and oldest individual.
+In the ‘rich are old’ regime, we have showed that the difference between the birth-time of the richest and oldest individual in $(\mathrm{BP}(t))_{t\geq 0}$, that is, $I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}$, is a tight sequence of random variables with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$.
+It remains unclear if more precise results can be proved.
 Open Problem 6.1.
-Can Theorem 4.4 be extended to hold for more general inter-birth random variables that satisfy Assumption 4.1?
-Furthermore, the results in Theorems 4.2 and 4.4 are restricted to the setting in which
-$\sum_{i=1}^{\infty}X(i)=\infty\text{ almost surely and }\widehat{\mu}(\lambda)<1\text{ for some $\lambda>0$}.$ (6.1)
-The existence of persistent $1$-hubs in CMJ branching processes that do not satisfy at least one conditions does exist, see e.g.
-[33, 26, 25, 31], but these works restrict themselves to the case $D=\infty$ almost surely.
-Can such results be proved in more general settings when $\mathbb{P}\!\left(D=\infty\right)\in[0,1)$ as well?
-Open Problem 6.2.
-Let $(\mathrm{BP}(t))_{t\geq 0}$ be a CMJ branching process whose offspring and inter-birth times satisfy Assumption 4.1.
-If either $\sum_{i=1}^{\infty}X(i)<\infty$ with positive probability or $\widehat{\mu}(\lambda)=\infty$ for all $\lambda>0$, under what conditions do persistent $m$-hubs exist, or does persistence occur?
-Additionally, can one prove an equivalent result as in Theorems 4.2 or 4.4 without the assumption that $\widehat{\mu}(\lambda)<1$ for some $\lambda>0$?
-Characterising the difference between the richest and oldest individual.
-In the ‘rich are old’ regime, we show that the difference between the birth-time of the richest and oldest individual in $(\mathrm{BP}(t))_{t\geq 0}$, that is, $I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}$, is a tight sequence of random variables with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$.
-It remains unclear if one can obtain a functional convergence of this random process.
-Open Problem 6.3.
 Under the conditions in Theorem 2.6, characterize the random process $(I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t})_{t\geq 0}$.
@@
 We expect that in this more general setting, persistence can be shown to occur when $\theta<R$ and when Assumption (C-V) is satisfied. We leave this as an open problem.
-Open Problem 6.4.
+Open Problem 6.2.
 Suppose the sequences $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied.
-Assume that $\theta:=\lim_{t\to\infty}-\frac{1}{t}\log\mathbb{P}\!\left(L>t\right)$ exists and recall $R$ from (2.14).
+Assume that $\theta\coloneq\lim_{t\to\infty}-\frac{1}{t}\log\mathbb{P}\!\left(L>t\right)$ exists and recall $R$ from (2.14).
 If $\theta<R$ and Assumption (C-V), does persistence occur in the sense of (P)?
@@
 Other modelling choices.
-In the (embedding of the) PAVD model the size of the offspring of an individual equals the number of children said individual has produced.
+In the (embedding of the) PAVD model the offspring of an individual equals the number of children said individual has produced.
 Similarly, the rate at which an individual produces its next child or dies depends on the total number of children produced.
-Instead, one could count only alive children towards the offspring, or let the birth and death rates depend on the number of alive children.
+Instead, one could let the birth and death rates depend on the number of alive children only.
 Either way of defining the model can be motivated from applications.
-However, the analysis for these alternative choices present more challenges, since now either an individual needs to produce $n$ children out of which $n-k$ should die to have an offspring of size $k$, or the rates depend not only on the parent individual but also on the lifetimes of its children.
+However, analysing this alternative model presents more challenges, since now an individual needs to produce $n$ children out of which nkn-k should die to have an offspring of size $k$.
+The birth and death rates of an individual thus depend the lifetimes of its children.
 Deijfen concludes from simulations presented in [13] that the asymptotic behaviour of the degree distribution should not change significantly, and we expect this to be the case for other properties such as persistence as well.
@@
 He also thanks Tejas Iyer for a discussion regarding the proof of Proposition 4.9.
+Appendix A Persistent hubs in general CMJ branching processes
+In this appendix we discuss how part of Theorem 4.1 can be proved for more general branching processes.
+To this end, we recall the general definition of CMJ branching processes at the start of Section 3.
+We make the following assumptions regarding the offspring random variable $D$ and the inter-birth time random variables $(X(i))_{i\in\mathbb{N}}$.
+Assumption A.1.
+  1. (i)
+    $D$ and $(X(i))_{i\in\mathbb{N}}$ are mutually independent.
+  2. (ii)
+    $X(i)<\infty$ for all $i\in\mathbb{N}$ almost surely.
+  3. (iii)
+    $\sum_{k=1}^{\infty}\mathbb{P}\!\left(D\geq k\right)\prod_{i=0}^{k-1}\mathbb{P}\!\left(X(i)=0\right)<1$.
+  4. (iv)
+    $\mathbb{P}\!\left(D=\infty\right)>0$.
+Part $(i)$ is natural for many classical branching process models (e.g.
+the Bienaymé-Galton-Watson branching process, the Yule branching process, and the Bellman-Harris branching process).
+Part $(ii)$ implies that $D$ corresponds to the number of children born to an individual, since each child will be born a finite amount of time after their parent is born.
+Part $(iii)$ can be viewed as the mean number of immediate children.
+That is, the number of children born at the same time as their parent.
+We assume this number is sub-critical, to avoid the branching process being able to grow infinitely large in finite time with positive probability.
+Finally, Part $(iv)$ states that individual produce infinite offspring with positive probability, which is necessary for the (non-)existence of persistent hubs to be a non-trivial question.
+It is relatively straightforward to check that the CTPAVD branching process defined in Section 3 that we analyse in Sections 4 and 5 is an example of a branching process that satisfies Assumption A.1.
+Let us now discuss the further assumptions used in Section 4. The point process $\mathcal{R}$, as in (2.6), is now defined as
+$\mathcal{R}\coloneq\sum_{k=1}^{D}\delta_{\sum_{i=1}^{k}X(i)}.$ (A.1)
+With $\mu$ the density of $\mathcal{R}$ and $\widehat{\mu}$ its Laplace transform, we have
+$\widehat{\mu}(\lambda)=\mathbb{E}\left[\sum_{k=1}^{D}\exp\bigg(-\lambda\sum_{i=1}^{k}X(i)\bigg)\right].$ (A.2)
+Assumption (F-L) then remains identical, though now for the more general $\widehat{\mu}$ stated here.
+Similar to the argument in (4.2), Assumption A.1(iii) implies that Assumption (F-L) is equivalent to the existence of a $\lambda>0$ such that $\widehat{\mu}(\lambda)<1$.
+Then, Assumption (N-E) is generalised to
+$\sum_{i=1}^{\infty}X(i)=\infty\qquad\text{almost surely},$ (N-E’)
+which implies that no individual can produce infinite offspring in finite time, preventing the branching process from growing infinitely large in finite time with positive probability.
+In Section 4 we worked with Assumption (F-D) not being satisfied, which ensured that individuals can produce infinite offspring with positive probability.
+This is now replaced by Assumption A.1(iv).
+Finally, Assumption (C-V), in light of Lemma 4.5 (and Remark 4.6), is now replaced by the more general
+$\sum_{i=1}^{\infty}(X(i)-X(i)^{\prime})<\infty\qquad\text{almost surely},$ (C-V’)
+where $(X(i)^{\prime})_{i\in\mathbb{N}}$ is an i.i.d. copy of $(X(i))_{i\in\mathbb{N}}$.
+With these more general definitions and assumptions at hand, we have the following result, which generalises part of Theorem 4.1.
+Theorem A.2.
+Consider the CMJ branching process $(\mathrm{BP}(t))_{t\geq 0}$ as defined in Section 3, where the offspring $D$ and inter-birth times $(X(i))_{i\in\mathbb{N}}$ satisfy Assumption A.1.
+  1. (a)
+    If Assumption (C-V’) is not satisfied, then, for any $m\in\mathbb{N}$, the branching process $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely.
+  2. (b)
+    If Assumption (C-V’) is satisfied and there exists $\lambda>0$ and $K\in\mathbb{N}$ such that
+    $\widehat{\mu}(\lambda)<1\qquad\text{and}\qquad\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]<\infty.$ (A.3)
+    Then, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $1$-hub.
+Remark A.3.
+Theorem A.2 does not include the existence of persistent $m$-hubs for $m>1$.
+This is due to the fact that, in the proof of Theorem 4.1(b), we rely on the memoryless property of the exponential distribution when $m>1$.
+[…]
+Iyer [25, Theorem 2.8] present sufficient conditions under which a persistent $1$-hub, when it exists, is unique, in the sense that there exists a unique individual in the branching process such that it attains the largest offspring for all but a finite amount of time.
+This result can directly be extended to the more general class of CMJ branching process as well as to persistent $m$-hubs for any $m\in\mathbb{N}$.
+introduced in this appendix.
+Theorem A.4 (Uniqueness of persistent $m$-hubs, Theorem 2.8 in [25]).
+Consider the CMJ branching process $(\mathrm{BP}(t))_{t\geq 0}$ as defined in Section 3, where the offspring $D$ and inter-birth times $(X(i))_{i\in\mathbb{N}}$ satisfy Assumption A.1.
+Fix $m\in\mathbb{N}$ and suppose that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely.
+Also suppose one of the following conditions is satisfied.
+  1. (1)
+    We have
+    $\sum_{k=1}^{\infty}\mathbb{P}\!\left(0\leq\sum_{i=1}^{k}(X(i)^{\prime}-X(i))-\sum_{i=1}^{k}X(i)<X(k+1)\right)<\infty.$ (A.4)
+  2. (2)
+    Both of the following conditions hold:
+    1. (i)
+      We have $\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)\neq X(i)^{\prime}\right)=\infty$, or there exists $i\in\mathbb{N}$ such that $\mathbb{P}\!\left(X(i)\neq X(i)^{\prime}\right)=1$.
+    2. (ii)
+      For any $\varepsilon>0$ we have $\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)>\varepsilon\right)<\infty$.
+Then, the persistent $m$-hub is $\mathbb{P}_{\mathcal{S}}$-almost surely unique.
+Remark A.5.
+[…]
+We conclude this section with a brief discussion where the proof of Theorem 4.1 requires adjustments to be able to prove Theorem A.2.
+Adapting the proof of Theorem 4.1(a)(a).
+The proof of Part $(a)$ of Theorem 4.1 relies on Part $(c)$ of Lemma 4.7, where we replace Assumption (C-V) with Assumption (C-V’).
+The proof of Lemma 4.7 can be adapted easily by simply replacing the birth times $(E^{(u)}_{i})_{i\in\mathbb{N}}$ and $(E^{(v)}_{i})_{i\in\mathbb{N}}$ with $(X(ui))_{i\in\mathbb{N}}$ and $(X(vi))_{i\in\mathbb{N}}$, respectively.
+Part $(a)$ leverages Lemma 4.2, which holds as we still assume that $\widehat{\mu}(\lambda)<1$ for some $\lambda<1$.
+Part $(b)$, assuming Assumption (C-V’) is satisfied, requires one additional step.
+Namely, we distinguish between whether the distribution of the series $\sum_{j=N}^{\infty}(X(ui)-X(vi))$ has an atom.
+If it does not, the same proof can be applied.
+If it does, then [25, Theorem 1.13] tells us that there exists a sequence $(c_{j})_{j\in\mathbb{N}}$ of real numbers, where $c_{j}=0$ for all but finitely many $j$, such that
+$\mathbb{P}\!\left(X(uj)-X(vj)=c_{j}\right)>0\quad\text{for all }j\in\mathbb{N},\text{ and}\quad\sum_{j=1}^{\infty}\mathbb{P}\!\left(X(uj)-X(vj)\neq c_{j}\right)<\infty.$ (A.5)
+Hence, by the Borel-Cantelli lemma, $X(uj)=X(vj)$ for all but finitely many $j$ (as $c_{j}=0$ for these $j$).
+As a result, we have that either (4.26) or (4.27) is satisfied, or that $\mathbb{P}_{\mathcal{S}}$-almost surely,
+$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=0}^{k-1}(X(uj)-X(vj))=0\qquad\text{for all }k\geq K_{0}.$ (A.6)
+In any of these three cases, the conclusion of the proof, that $\mathcal{B}(uk)\leq\mathcal{B}(vk)$ for all $k\geq K_{0}$ or $\mathcal{B}(vk)\leq\mathcal{B}(uk)$ for all $k\geq K_{0}$, is still valid.
+Finally, for Part $(c)$ of Lemma 4.7, replacing $E^{(v)}_{i}$ with $X(vi)$ for $v=u$ and $v=uj$ and using that Assumption (C-V’) is not sufficient for the proof to hold.
+The proof of Theorem 4.1(a)(a) then follows in the same manner, again replacing $E_{i}^{(v)}$ with $X(vi)$ for the appropriate individuals $v$, and assuming Assumption (C-V’) is not satisfied.
+This yields Theorem A.2(a)(a).
+Adapting the proof of Theorem 4.1(b)(b).
+As in Part $(a)$, we substitute $E^{(v)}_{i}$ with $X(vi)$ for any individual $v\in\mathcal{U}_{\infty}$.
+Instead of Assumptions (N-E), (F-L), (C-V), and (F-D), we instead use Assumptions A.1, (C-V’), and (A.3).
+Then, as the second (A.3) allows us to use Lemma 4.3, Proposition 4.8 follows in the same manner.
+For Proposition 4.9 when $m=1$ requires minor adaptations, as discussed after the proof of Proposition 4.9 already.
+These adaptations are sufficient for the proof to work in the more general setting we are in, here, as well.
+When $m>1$, we were not able to avoid using the memoryless property of the exponential distribution at several places, so that a full generalisation of Theorem 4.1(b)(b) is out of reach, though we expect the result to be true.
+The proof of Theorem 4.1(b)(b) itself then requires no further adaptations, with the exception of the proof of the uniqueness of persistent $1$-hubs.
+Here, by using one of the conditions in A.4, one can prove that the condition in (4.58) is satisfied with probability zero.